A Complete Bibliography of ACM Transactions on Multimedia Computing, Communications and Applications

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

25 October 2025 Version 1.81

Title word cross-reference

(k, n) [LZD^{+}22]. (t, s, k, n) [LZW^{+}21]. 1 [ZCL^{+}24]. 2 [ATMM24, HLC^{+}25, LL23, LZYY24, PJL23, WDJ^{+}21]. 3 [AP10, ARE13, CFN^{+}23, CEE09, CGNG15, Chu15, DP06, DVA21, DÇ07, GZHD12, GLW20, GHQ24, GS11a, HH12a, HLM^{+}22, HSZ^{+}18, HLC^{+}25, JP11, KH13, LLP06, LYZ^{+}18, LLS^{+}21, LCL^{+}23b, LZWC23, LGM^{+}24, LSS^{+}24a, LC12, LCT^{+}12, LOJZ18, LZC^{+}19, LKW^{+}22, LCZ^{+}24b, LZYY24, LDZ^{+}24, LGF^{+}14, LWWZ20, MC11, NLW^{+}21, NALM23, PB19, PB14, QXG^{+}24, RS16, RHS12, SAZ^{+}15, SHOG12, SGY^{+}23, SZZ^{+}23, SHF^{+}24, SLNL20, SHWC19, SWK^{+}22, TLZ^{+}20, WXW^{+}22, WAK^{+}12, WHY19, XFSZ20, YWN^{+}10a, YWN^{+}10b, YI14, YGL^{+}23, YH14, YLHL22, YCFX23, YHZ19, ZSYW23, ZDE16, ZFSX21, ZLC^{+}25, ZYC^{+}24, vRPPP23]. 360 [LLA^{+}21, QJ23, ZLOL18]. 4 [CRL20, ZFY^{+}25]. 5 [CXW^{+}19]. ^{2} [BLW^{+}24, BPB^{+}22, HQF^{+}19, XWDC23, ZSYW23, ZHD^{+}23, ZWJ^{+}24]. ^{3} [WWWL23]. ∘ [CE23, FHH22, YM24]. δ(q_{k})/F_{2} [DDK24]. l_{1/2} [FBG22]. P [DG17b]. U [HRL^{+}24].

-Degree [LLA^{+}21, QJ23, ZLOL18]. -Encoded [DG17b]. -Net [HRL^{+}24]. -Score-Based [RTR21]. -Threshold [LZW^{+}21].

1 [MPSR05]. 19 [ASI^{+}21, CLZ^{+}21, LWZW21, LCL22, RHAG21]. 1993 [JPS05]. 1Mbps [RXC14]. 1s [Ano14, Ano20].

2 [WLW^{+}23]. 2-based [BLS^{+}19]. 2000 [ZC12]. 2003 [PSS05]. 2007 [BSH08]. 2008 [CDGJ09]. 2009 [XSEZ10]. 201 [WZ20]. 2010 [SLYE11]. 2011 [GPHOH12]. 2013 [ZZMS14]. 2014 [CCO15, HT15]. 2016 [TB17]. 2017 [CHHH18]. 2018 [LKM^{+}19]. 2020 [LVM^{+}21]. 2023 [GWZ25]. 20th [NLS13]. 2BiVQA [TFHM24]. 2s [TOM12].

3.0 [GHS^{+}24]. 360 [LLC^{+}25]. 360-Degree [VTD22, LLC^{+}25, NTT20, ZZMZ21, dCAGF^{+}25]. 3D [HSS^{+}25, LKL25a, LCH^{+}25, RBLJ25, YXK^{+}25]. 3K [WMH^{+}22]. 3s [Dao17, Sha21, TOM12]. 3V3D [ZCXL23].

4D [YZD^{+}25].

5G [KJJ^{+}21].

6DOF [CKRB23]. 6G [AAT^{+}22, LWH^{+}22, MSG22]. 6G-based [AAT^{+}22].

7S [Ano11].

802.11 [LCK09]. 802.11e [PBS12]. 8K [GGML22, YGMY24]. 8S [Ano12].

AABLSTM [ZSLW23]. AB-LSTM [LZL20b]. ABR [BRZS18, PHS^{+}20, SWZ25, YDJ24]. Abstraction [LDZ^{+}24, JSL07, TV07]. Accelerating [GGML22, ZT22]. Accelerator [MGG17]. accelerometer [RHE10]. acceptable [KS09]. Access [APP^{+}22, LZW^{+}21, WFC^{+}25b, CW10, PZ08, YH13, ZO13]. access-pattern-driven [ZO13]. accessibility [HWY^{+}11]. accessible [AP13]. Accessing [CFP15]. accuracy [HH08b]. Accurate [APC^{+}25, DBSL24, FYY^{+}21, GWWC25, HD19, HSG23, LYL^{+}21b, QLW^{+}24, ZLZ22b, ZWZ^{+}23a, LJP08, LML^{+}13]. achieve [LGX^{+}08]. Achieving [CCG^{+}08]. ACM [JPS05, TB17, BSH08, BLJX10, CDGJ09, CHHH18, Dao17, GPHOH12, Geo05, HT15, KZHC13, NLS13, PSS05, RJ05, SLYE11, XSEZ10, ZZMS14, ZTB20, HGKS24]. ACMNet [CDL^{+}20]. Acoustic [WJ15b, YSSF24, FYH10]. Across [YKW^{+}22]. ACTE [BTBZ20]. Acted [THR^{+}22]. Action [BUS^{+}21, BLW^{+}24, CLZ^{+}23a, CZZ^{+}24, DVA21, DZWH23, FXMZ24, HYSL20, HZL^{+}21, HSZ^{+}18, HZL^{+}25, HP17, LLZ^{+}25, LMX24, LZZ^{+}25, LSL^{+}20, LLC^{+}24b, LLL^{+}24b, LHC^{+}25, MZZ^{+}20, SZL^{+}22, TLY^{+}22, TSH^{+}25, WWY^{+}24, WCG^{+}25, XW25, XZZL23, XLW23a, XW17, XLZ^{+}21, XJW^{+}22, YHQH17, ZHL19, ZFSX21, ZML^{+}24, ZLT^{+}25a, ZLL20, ZSS^{+}23, ZYY^{+}20, ZZW^{+}22, WLHL13]. Action-Aware [HZL^{+}25, CZZ^{+}24, WLHL13]. Actions [ICZ^{+}22, LLCH17]. Activated [XZL^{+}22a]. Activation [WSX^{+}23]. Active [CLZ^{+}21, FKCW20, JYZC12, KIT^{+}24, KUH^{+}22, QCW^{+}25, TCS^{+}25, XWZ^{+}24b, ZZP^{+}20]. Activities [GCL^{+}25]. Activity [CWTG22, HYG^{+}21, LXL^{+}18, MHCG19, NRUT20, YSY^{+}22, ZCD15]. Actor [CLZ^{+}23a]. Actor-Action [CLZ^{+}23a]. Actual [CHHH25]. Ad [LB15, WWL13]. Adaptation [DWS^{+}24, HXZ^{+}20, JLL^{+}25, JXLD25, KBB21, LLZ^{+}22, PWY^{+}25, PFC^{+}16, SYS17, WBRZ17, WWQ^{+}25b, WLSZ22, XSD^{+}22, YJM^{+}19, YHZ23, ZLY^{+}24, ZCYL24, CXS^{+}08, WBL09]. Adaptative [dMGRT24]. Adapted [Ala21]. Adapters [WZRD25]. Adapting [YHG^{+}24, ZDD^{+}24, ZHG^{+}25]. Adaptive [AH20, APM21, BBZ18, BYOZ20, BQBLN18, CPSH14, CDL^{+}20, CHHH25, CDZ^{+}17, DFXY20, DLC^{+}22, GHAM25, GZLW18, GCZ^{+}24, GHY^{+}24, HLW^{+}21, HZC^{+}16, HJWW19, HXZ^{+}20, HSL^{+}20, HJZ^{+}23, HLS^{+}22, KPL^{+}22, KFP^{+}25, KABB20, LSCH25, LLZ^{+}21, LLZ^{+}22, LZZ23b, LYC^{+}25, Lin15, LX21, LHS^{+}21, LYW^{+}24, LTHT24, LDZ^{+}20, MLJ^{+}22, MAGT23, MN16, NTT20, PLZW18, PFC^{+}16, PVWD18, QHFX21, SLG^{+}24, SZC24, SYS17, TAT^{+}25, TAPP^{+}15, TNH^{+}21, VTP^{+}20, WJS^{+}21, WMM^{+}23, WXS^{+}25, WBB^{+}17, WY23, WSX^{+}24, XWL^{+}25a, BLS^{+}19, YYX^{+}24, YZJ^{+}25, YWLZ23, YCLH22, YBTX23, ZQRS18, ZT22, Zha19, ZYW^{+}24, ZML^{+}24, ZWZ^{+}23a, ZZL^{+}24c, ZWZ^{+}23b, Bag11, BAK13, GKW08, LC12, LYJ^{+}13, LMLC14, MB08, XMST07]. Adaptively [YWG^{+}20]. AdGPT [HQLW25]. adjacency [QLSQ12]. Adjusting [WCK05]. Adjustment [BWP^{+}24, CXL^{+}22]. admission [GHP^{+}06]. Adult [TESU16, HZW^{+}11]. Advanced [CH15, Hon19, LLL^{+}25, YM24, SWS21]. Advances [Cho13, GTLG14a, TAT^{+}25, Ano13]. Advancing [ZZL^{+}25b]. AdvErsarial [GTP^{+}22, CFG^{+}24, CZW^{+}25, CSŞ^{+}25, DHP23, DYL^{+}25, DLD^{+}22, FDZ^{+}25, FNH22, FBGD23, GYY^{+}25, HCZ^{+}22, JMS23, LZL^{+}21a, LLJC21, LJL^{+}24, LZH^{+}24b, LJW^{+}25, LKW^{+}22, LNH23, LCZ^{+}24b, LZW^{+}24, LYH24, MPTD22, NWL^{+}20, PQ19, PWY^{+}25, SSK20, SZZY20, SGS21, SL22, TZLZ21, WDX^{+}23, WYGC24, WWC^{+}24, WLMZ25, WWZ^{+}25, WSX^{+}24, XTL^{+}21, XWDC23, YZS^{+}23, YLLL23, YNY25, ZY21, ZCL^{+}22, ZCXL23, ZSS^{+}23, WSX^{+}23]. Adversarially [LWH^{+}23]. Advertising [HQLW25, MLHL12, SCFL14]. Advisor [RUD23]. AED [PWY^{+}25]. AED-PADA [PWY^{+}25]. Aerial [YHH^{+}22, ZLL^{+}24a, ZWZ^{+}23a]. Aesthetic [Abd18, JLL^{+}22, XLN^{+}21, BSS11, FPH^{+}08]. Aesthetics [CLN^{+}21, ZZZ14]. Aesthetics-Guided [ZZZ14]. Affect [MGP19]. Affective [APRS^{+}19, HXZ^{+}20, KBI^{+}23, LGT^{+}24, WHY18, XZL^{+}22b, ZTR^{+}22, ZHLJ22, ZZS^{+}25, ZWS^{+}20, ZJSJ20]. Affine [YZY^{+}25, GZHD12]. affine-transformation-invariant [GZHD12]. Affinity [AQL^{+}20, FXF^{+}23, LYL^{+}21b]. Affordance [CZS^{+}25]. Affordance-Aware [CZS^{+}25]. after [JPS05]. Against [FBGD23, CYZ^{+}24, DYL^{+}25, LZW^{+}24, WWC^{+}24, XLF^{+}25]. AGAR [dMGRT24]. Age [DLO^{+}20, YML^{+}22, YLCC18]. Age-Invariant [YML^{+}22]. Agent [DJWZ25, GCL^{+}25, TSZ^{+}23, ZZMZ21, MEA^{+}21]. Aggregate [SEK12]. Aggregating [XXG^{+}21, YTL^{+}24b]. Aggregation [GLZW23, LLZ^{+}25, PB19, WLTL23, WXS^{+}25, YMS^{+}24]. agile [SLKS12]. Agnostic [VTD22, XSD^{+}22]. Agreement [SSP21]. Ahead [NCL^{+}23]. AI [CS22, CJHH21, CLS^{+}21, EAK^{+}25, GHAM25, HCM^{+}22, MHF24, USM^{+}25, WSLZ25, YNLZ22]. AI-empowered [HCM^{+}22]. AI-Manipulated [YNLZ22]. Aided [LZL^{+}24c, RPE^{+}17, WCF^{+}17, PSN^{+}24, YYX^{+}24]. AIGC [ZSX^{+}25]. Algorithm [AAS^{+}20, ABR17, APM21, BLMP18, CH15, GKSB17, GGML22, LLL^{+}21, LSS^{+}24a, PLZW18, PFC^{+}24, SMK^{+}24, SAZ^{+}15, SSSA23, SB23, TZLZ21, YXJ^{+}24, YJM^{+}19, YDJ24, ZWL^{+}21, Bag11, LBD08, LWL08, WT10]. Algorithms [GL12, WBB^{+}17, NC13, BDV08]. Aligned [LTD^{+}21, SXY^{+}23, WHL^{+}21]. Aligning [ZXD^{+}23a]. Alignment [CWW^{+}25, FZZ^{+}25, GJY^{+}25, HLZ^{+}21, HDH^{+}24, HJZ^{+}23, HZL^{+}25, JCB25, LZT^{+}20, LDL^{+}25, LYW^{+}25, MAE^{+}21, SZC24, SZW^{+}24, TY23, WDR^{+}24, WDC^{+}25, WWQ^{+}25b, WGHX24, XLX^{+}24, YLK^{+}20, YWL^{+}24, ZH18, ZHS20, ZYJZ25, ZLJ^{+}25, MSL10]. All-in-one [JXTC21]. Alleviating [QCW^{+}25]. Allocation [DMF17, LZG^{+}25, PYZ^{+}20, PHS^{+}20, QZXH14, XZL^{+}22b, YGNT19, ZZL21, TC08]. ALOHA [ZHG^{+}25]. Alone [MGP19]. ALP [CPSH14]. Alternate [XSL^{+}23]. Alzheimer [KN21b, LMS^{+}24, TRK^{+}20]. Am [BUS^{+}21]. AMC [ZWZ^{+}23b]. AMIL [SZZY20]. among [CY11]. Amphitheater [CGNG15]. Amplitude [WHL^{+}21]. Amplitude-Angle [WHL^{+}21]. AMSA [WMM^{+}23]. ANAGL [MBXZ24]. Analysis [ADMM25, ABR17, BTBZ20, BA20, BZDX^{+}18, CZC15, CDL^{+}20, CHHH25, CF22, DZW^{+}21, DMF^{+}20, FZW^{+}25b, GS19, HDW^{+}18, HWWL20, KD18, LCL14, LLC^{+}24a, LKM17, LKM^{+}19, LWZW21, LQS21, MXH^{+}23, MKSB17, MZGY17, PCH^{+}20, QRTM24, RSE16, RHS^{+}20, SS17, SZM^{+}21, SHSR24, SS20, SAL^{+}22, TNZ^{+}25, TESU16, VVSV17, WBRZ17, WLCG21, WMM^{+}23, WZZ^{+}23, YV23, YTR^{+}25, YXZQ22, ZZX^{+}20, ZSLW23, ZWL^{+}24, ZYW^{+}24, ZLJ^{+}25, ZGWZ23, ADCB07, CY11, Eff13, EGEM06, GIL^{+}10, GFB^{+}14, HCW^{+}07, HM10, KK08, MTTH13, SX13, SWH06, XMST07, XXD^{+}08, YC08, ZI13]. Analytic [CS17, CSSZ23, WKST08]. Analytic-Based [CSSZ23]. Analytical [GS18, TSHP05]. Analytics [GLWK19, HCM^{+}22, LLC^{+}25, LSL^{+}20, WAD^{+}18, WFZ^{+}21a, Wan21, YNC18, ZYZE19a, ZYZE19b]. Analyzing [CYM^{+}25, SKW^{+}15]. Anatomy [LW23]. Anchor [LKL^{+}25b]. Ancient [XZY^{+}25]. ANet [BLW^{+}24]. Angle [WHL^{+}21]. Animating [LYLL23]. animation [KP08, SY09]. anniversary [NLS13]. Annotation [HOSS13, LLZ^{+}21, RPE^{+}17, SS20, WKE16, YSZZ14, DCC^{+}13, LWL^{+}12, LCS09, QHR^{+}08]. Annulus [YLLZ25]. Anomaly [LQG25, PWHP24, XWW^{+}25, ZYW^{+}25]. Anonymization [HMB^{+}25]. Answer [LGY^{+}22]. Answering [BBB^{+}23, EL25, HWZL24, LXB^{+}22, LJL^{+}24, LXL^{+}18, LYW^{+}25, PN16, PSSZ24, WLS^{+}24b, WLSN24, YTOH22, ZSZ^{+}25, ZXY^{+}20]. Anti [FDZ^{+}25, MBXZ24]. Anti-Forensics [FDZ^{+}25]. Anti-Sparse [MBXZ24]. Anticipation [ZML^{+}24]. Anxiety [LZH^{+}24a]. Aortic [HLZ^{+}20]. API [QSZ^{+}21]. Appearance [CHLW19, FKCW20, GZW^{+}24b, SR25, XZY^{+}24b, CCD07]. Appearance-consistent [CHLW19]. Application [AAA^{+}21, HLZ^{+}20, LLS^{+}21, LYW^{+}22, LWWZ20, Zho16, EGEM06, SX13]. Applications [ASLA18, APM21, dCAGF^{+}25, Ber18, CPSH14, FZW^{+}25a, GZT21, GTLG14a, JLA^{+}23, KUH^{+}22, KJJ^{+}21, LLYH14, MAX^{+}24, NHP^{+}16, NALM23, PCH^{+}20, REP^{+}19, RVTC25, SCXC15, SAZ^{+}15, SRPH16, WAC^{+}25, XLH^{+}24, YCGM14, Ano13, BDV08, DY09, GDGC07, GA12a, HNL08, HCKL13, LML^{+}13, MGCH13, MMW10, MTTH13, OMP07, PZ08, PS05, SOC^{+}13, BLJX10, Geo05, SWS21, SLK21, HGKS24]. Applied [DDK24, YJM^{+}24]. Applying [HAAM19, NHP^{+}16]. Approach [BYOZ20, BZDX^{+}18, CGPCR18, CTBC22, DZW^{+}21, FPA24, GDM^{+}23, GHS^{+}24, GWM^{+}14, GGB24, HLM^{+}22, HJMY15, HLW^{+}21, HSN^{+}14, HLD18, HDZ^{+}15, HWW25, JW21, KUH^{+}22, KZGH15, LQH18, LZXY20, LZH^{+}20, LLZ^{+}21, LWZ^{+}21a, LMF^{+}14, LGY^{+}22, LTHT24, MN16, MC19, RVTC25, SED25, DMSRL18, SAL^{+}22, TVZ^{+}19, UMM24, WWHW14, WBRZ17, WDX^{+}23, XGZ^{+}24, XFSZ20, XZH^{+}21, YHQH17, ZWL^{+}17, ZMH^{+}20, ZDZ^{+}22, ZLL^{+}24a, BAK13, BSS11, DCM13, MC11, NC13, PZ08, RHE10, WK10, WAK^{+}12]. Approaches [FAA18, SWS21, WHX^{+}25, WKN25, LK07]. Approximate [CLP17]. Approximation [HJ23, YWG^{+}20, ZZB24]. Approximation-Guided [HJ23].

Approximator [AMT25]. Apps [MNPOF22]. APRICOD [ZO13]. AQM [KABB20]. Arbitrary [FYW^{+}24, FXF^{+}23, LZZ^{+}24c, MZH^{+}24, WWY^{+}25]. Arbitrary-Scale [FYW^{+}24]. Arbitrary-Shaped [FXF^{+}23]. Archaeological [SRDJ18]. Architectural [BLH^{+}25]. ArchiTecture [GTP^{+}22, DWC^{+}21, FAA18, KCC17, PA20, SHIE15, WMW^{+}22, CVV06, OMP07, ZCAP08]. Architectures [SCB^{+}24]. archives [SWH06]. Area [LCC^{+}14a, SHIE15]. Areas [Hua23]. arrangement [HCS12]. Array [ZZCZ21]. Art [CS22, CDA^{+}24, GTLG14b, LCL23a, WSLZ25, WDE25, WFZ25, WF25, YLHL22, LSDJ06, XLN^{+}21]. ART-UP [XLN^{+}21]. Articles [LLO^{+}20, SSBT20]. Artifact [LZL^{+}24a, XZLZ23]. Artifacts [CWL24, SKW^{+}15]. Artificial [DZW^{+}21, GEL^{+}15, PDD22, TMB^{+}22]. Artistic [CTBC22, SW18, WSLZ25]. Artwork [SBU^{+}17]. ASAP [ZQRS18]. ASIFusion [LYW^{+}24]. Assembling [XTL^{+}21]. Assessing [AdSGMS25, HAE11]. Assessment [CZY^{+}21, CLN^{+}21, FPA24, HLY^{+}23, JLL^{+}22, KN21a, LZH^{+}24a, LXJ^{+}24, LGLZ20, LYXY23, SDJ17, TCL^{+}23, TFHM24, TSP^{+}24, WLL^{+}19, YTR^{+}25, YWC^{+}25, YDCZ23, YWW^{+}24, ZSMZ21, ZGL^{+}23, ZSZ^{+}24a, ZSX^{+}25, ZZL^{+}25b, ZZL^{+}25c, ZYC^{+}24, MVW08, MRS^{+}07, RHS12]. Assessment-driven [ZGL^{+}23]. assigned [AAT^{+}22]. Assistance [FKW22, SAL21b]. Assisted [BRZS18, HYT^{+}25, KCC17, NY24, OE19, SP21, SK23, TCS^{+}25, WCY^{+}18, XJFS25, YXZQ22, ZLW17, CDZ^{+}17, MZL^{+}18]. Association [LZZ^{+}23a, WCX^{+}14, WHL^{+}23, YYBX24, GA12b]. Associations [YZRW22]. Asymmetric [CLP17, FHF^{+}24, LLH^{+}24, WZL^{+}24]. Asynchronous [SKA24, XGZ^{+}24, ZWLZ24]. Atom [HZC^{+}16]. atoms [JCC^{+}10]. Atrial [HZW22]. Attack [CYZ^{+}24, DYL^{+}25, LYH^{+}25, WWC^{+}24, ZZC^{+}23, FLZ^{+}12]. Attack-Defending [LYH^{+}25]. Attacks [CFG^{+}24, CQA^{+}25, DLD^{+}22, FBGD23, SED25, XLF^{+}25, XWDC23, YNY25]. Attention [APC^{+}25, CFN^{+}23, CBSC18, DVA21, DNL^{+}23, DWL^{+}23, DMF^{+}20, EFF^{+}25, FZYW20, GLZW23, dMGRT24, GSY^{+}24, HH19, HYG^{+}25, HQF^{+}19, HDH^{+}24, HCZ^{+}25, HWWL20, HWW^{+}23, JWH21, LPY^{+}19, LLS^{+}22, LDQ^{+}25, LZZS23, LSN^{+}20, LZL20b, LX21, LGY^{+}22, LYW^{+}25, LYD^{+}21, LHC^{+}25, PKL25, RLAM25, SHSR24, SLC25, SGY^{+}23, SO22, SJC^{+}25, SZB^{+}22, TWY24, TLZ^{+}21, TCL^{+}23, WHY18, WDPX23, WLTL23, WKS^{+}23, WWS^{+}24, WSLC24, WQTX25, WHW18, WHY19, WSX^{+}23, XFZ^{+}19, XJW^{+}22, YRE^{+}20, YHXL20, YZG^{+}20, YCZ^{+}23, YZ23, ZHL19, ZHS20, ZSS20, ZHD^{+}23, ZZJ^{+}23, ZLX24, ZZW^{+}24, ZZW^{+}19, ZYLN21, ZLF^{+}24, ZHC^{+}23, ZXD^{+}23b, ZMZ^{+}23, ZXY^{+}20, AQL^{+}20, BGP11, FPH^{+}08, HQF^{+}20, JC08, YML^{+}22, YWW^{+}24, ZZY^{+}23]. Attention-Augmented [ZHC^{+}23]. Attention-Aware [HQF^{+}19]. Attention-Based [HWWL20, LSN^{+}20, LZL20b, LX21, LYD^{+}21, RLAM25, TLZ^{+}21]. attention-centered [FPH^{+}08]. Attention-Consistency [XJW^{+}22]. Attention-Directed [WLTL23]. Attention-Guided [PKL25, WWS^{+}24, ZLF^{+}24]. Attentional [WWY^{+}24, YCL25]. Attentions [LLJ^{+}20]. Attentive [HRL^{+}24, YV23, ZJZC20, ZTR^{+}22, ZLL^{+}22]. Attraction [DNPG^{+}17]. attractiveness [NLN^{+}13]. Attribute [Ala21, CYW24, DLL^{+}20, HC25, HLX^{+}14, JLL^{+}22, LZD^{+}21, LLZ^{+}24, TGC23, XCM^{+}23, YSF^{+}21, YZY^{+}25, ZZY^{+}14, ZH18, ZYZ^{+}25b]. Attribute-Augmented [ZZY^{+}14].

Attribute-Aware [HLX^{+}14, ZYZ^{+}25b]. Attribute-consistent [XCM^{+}23]. Attribute-Controlled [CYW24]. Attribute-wise [YSF^{+}21]. Attributed [SBS23]. Attributes [APRS^{+}19, CWL24, CTBC22, EFF^{+}25, FSX14, GLT^{+}20, XQG^{+}23]. Attribution [APC^{+}25]. Auction [WLQL12, ZXZ^{+}24a]. Auction-based [WLQL12]. Audio [AMMG16, BSSNF^{+}20, Ber18, BPFA24, CHHH18, HLTH24, JCC^{+}10, LLJW15, LXZ^{+}25, LLYH14, MBK^{+}23, NPG^{+}22, SBU^{+}17, SRDJ18, TB17, XXD^{+}08, YZYX24, YTRC19, ZYO20, ZWH^{+}23, ZYT^{+}23, ZLX24, ZXZ^{+}25, ZZC^{+}25a, ZHG^{+}25, ZSZ^{+}23, LSQ11, MVW07, SOC^{+}13, YQH12]. Audio-Tactile [MBK^{+}23]. Audio-Visual [HLTH24, YZYX24, ZYO20, ZYT^{+}23, ZLX24, ZXZ^{+}25, ZZC^{+}25a, ZHG^{+}25, BPFA24, JCC^{+}10, LXZ^{+}25, ZWH^{+}23, ZSZ^{+}23]. Audiovisual [DG17a, KN21a, SDJ17, Tas20, Tas22, Tas24, ZCHD23]. Augmentation [BYM^{+}18, CJHH21, EL25, IAJ^{+}24, QCW^{+}25, TSH^{+}25, WSLC24, XW25, YSSF24]. Augmentation-based [IAJ^{+}24]. Augmentations [HPZ25]. Augmented [AWG^{+}15, CPL^{+}23, HQF^{+}20, LL15, LDZ^{+}24, REP^{+}19, SCB^{+}24, SRDJ18, SAL21b, YLL^{+}24, ZRZX23, ZZY^{+}14, ZHC^{+}23, ZMZ^{+}23, GYN12]. Augmenting [CSŞ^{+}25, ZCX^{+}24]. Authentic [PYZ25]. Authentication [GAH25, HM10, NHZ^{+}24, SSP21, WWC^{+}24, WH22, XLF^{+}25, ZCL^{+}24, DY09, SG07]. Authenticity [LCH^{+}25]. Author [VMP20]. Authoring [AdSGMS25, SGW08, SLKS12, BB11, BH05, BCG13, KP08]. Authorization [XWZ^{+}24b]. Autism [DMF^{+}20]. Auto [GHQ24, HYLG23, LLS^{+}21, LZZ^{+}24a, PLB^{+}24]. Auto-Encoder [LZZ^{+}24a, LLS^{+}21]. Auto-encoding [GHQ24]. Auto-Knee [PLB^{+}24]. Auto-Regressive [HYLG23]. Autoencoder [ARA^{+}23, BSSNF^{+}20, FFCM24, LWZH19, XLZ^{+}24, ZYT^{+}23, XTL^{+}21]. Autoencoders [FWLA15]. Automated [CLZ^{+}21, CZZ21, LLO^{+}20, VCO15, ZRCH08]. Automatic [AdSGMS25, BYM^{+}18, BZ05, FYY^{+}21, GDM^{+}23, KD18, LLZ^{+}21, LZH^{+}24a, MGP19, NN21, RSB11, SBU^{+}17, SXM^{+}06, SB23, VM12, YMX^{+}16, YMY^{+}21, YSZZ14, HCS12, JWL06, MSL10]. Automaton [DDK24]. Autonomous [GCL^{+}25, PSN^{+}24, SZZZ25, TJZ^{+}25, YWG^{+}20]. Autoregressive [AAPLV25, LLSX20, CZZ^{+}24]. Auxiliary [CZW^{+}24, CZW15, GYY^{+}25, YWW^{+}24]. AV [XC06]. Avatar [SHIE15, ZYW^{+}25, ZPR^{+}25]. award [Ste12c]. Aware [ASI^{+}21, CZS^{+}25, CXY^{+}25, CPP^{+}14, CAJ19, CZW^{+}25, DYSX14, FJF^{+}25, FAA18, FSK^{+}15, GNC17, GGSV25, GWWC25, HC25, HDZ^{+}15, HQF^{+}19, HZPL21, HCZ^{+}25, HLX^{+}14, HZL^{+}25, KKGE18, LPS15, LYJ^{+}15, LYZ^{+}24, LZWL25, LLJW15, LL15, LZL^{+}24b, MOL^{+}22, OE19, PLB^{+}24, QSZ^{+}21, RK15, SBU^{+}17, SGY^{+}23, TFY^{+}25, WBRZ17, WAD^{+}18, WCLC18, WDC^{+}25, WZZ25, WCY^{+}18, WLZ^{+}24b, YRE^{+}20, YDJ24, YYY^{+}24, YCLH22, ZQRS18, Zha19, ZHL19, ZDD^{+}24, ZYZ^{+}25a, ZGD^{+}19, ZHL^{+}25, ZJL^{+}21, ZZCZ21, ZZL25e, AH20, BCG13, CPP^{+}13, CXL^{+}22, CZZ^{+}24, CSZ^{+}25, EGEM06, HL24, HXZ^{+}20, JD23, LCL23a, LWH^{+}22, LPW^{+}22, LWL08, LTX^{+}23, LWP22, MGY22, MYX^{+}23b, NSM^{+}24, QHFX21, RXY^{+}25, SYX23, WZC^{+}13, WLHL13, WTZ^{+}20, WDJ^{+}21, WHY19, WXQC20, WGHX24, XYL^{+}23, YHP25, ZJT^{+}23, ZYZ^{+}25b, YJM^{+}19]. Awareness [ZLT^{+}25a]. Axial [FQJ^{+}25]. Axial-shunted [FQJ^{+}25].

back [Eff13]. Backdoor [ZYH^{+}24]. Backdoors [HYW^{+}25]. Background [NSK^{+}21, WSLM18, MVW07].

Backpropagation [JLL^{+}25]. Backpropagation-Free [JLL^{+}25]. Backtracing [CLP^{+}23]. Bag [WZNM14, XJG^{+}22]. Bag-of-Visual [WZNM14]. Bag-of-words [XJG^{+}22]. bags [HZW^{+}11]. Baking [LOJZ18]. Balanced [CHHH25, XWL^{+}25a, ZLZ22b]. Balancing [ZZP^{+}20, DL14]. Balloon [YMY^{+}21]. Bandwidth [BTBZ20, HLD18, SSP^{+}24, HB08, MMW10, REV^{+}12]. bandwidth-lookup [REV^{+}12]. Banks [MKS20]. Barcode [JGY^{+}24]. Based [Abd18, ARA^{+}23, AR15, BWP^{+}24, CGPCR18, CHLW19, CLWW24, CWL24, CXX^{+}24, CZX^{+}24, CQA^{+}25, CZW^{+}25, CSSZ23, VTD22, DMF17, DG17a, DZW^{+}21, DSL^{+}25, DWC^{+}21, EWSZ15, EFF^{+}25, FZW^{+}25b, FF23, FSK^{+}15, FFCM24, FZZ^{+}25, GLW20, GLT^{+}20, GZLW18, GCZ^{+}24, GZZ^{+}24, GS18, GS19, GSDT21, HJMY15, HMB^{+}25, HLD18, HYSL20, HMOS17, HSL^{+}20, HWW25, HWWL20, HSX22, Hua23, HLS^{+}22, JXL^{+}25, JTZ^{+}16, JLK^{+}24, KN21a, KUH^{+}22, LYZ^{+}18, LPY^{+}19, LYL^{+}21a, LLZ^{+}21, LZL^{+}23, LLG^{+}23, LCL^{+}24a, LLL^{+}25, LZWL25, LKL^{+}25b, LLC^{+}25, LLL^{+}21, Lin15, LZLJ22, LMF^{+}14, LSN^{+}20, LLA^{+}21, LZW^{+}21, LWZ21b, LWP22, LYW^{+}24, LCZ^{+}24b, LSS^{+}24b, LZG^{+}25, LZXP25, LWHC24, LGC^{+}24, LDZ^{+}20, MY15, MKY^{+}24, MXZ^{+}24, MATW17, MKC21, MNPOF22, MGG17, NCMM21, NPG^{+}22, NHP^{+}16, PYZ^{+}20, PDD16, PWX^{+}24, RTR21, RS16, SSBT20, SED25, SCXC15, SLG^{+}24, SWM^{+}24, SLC25, SRPH16, SRAA17, SJC^{+}25, SB23, SKA24, SZB^{+}22, SLZ24, TYY^{+}18, TWFW24, TZ24, TFY^{+}25]. Based [TLZ^{+}20, TZLZ21, UJLS22, VVSV17, WBRZ17, WLHT19, WZ20, WHL^{+}21, WJS^{+}21, WCL23, WSW^{+}24, WCG^{+}25, WFZ25, WWQ^{+}25b, WBB^{+}17, WLL^{+}19, WDS21, WLW^{+}23, WLL^{+}24, WZZL25, XWW^{+}25, XW25, XLZ^{+}23, XLF^{+}25, XHYX23, XLZ^{+}21, XDL^{+}21, XHL^{+}24b, XJB25, XZY^{+}25, YTL^{+}21, YBW^{+}25, YXYB21, YZS^{+}22, YLLL23, YDCZ23, YH14, YWY^{+}23, YJTN18, YHZ23, ZCL^{+}22, ZZY^{+}14, ZZL^{+}17, ZYZE19a, ZZP^{+}20, ZGR21, ZSLW23, ZZJ^{+}23, ZDQ^{+}23, ZZLL17, ZXD^{+}23b, ZHL^{+}25, ZZC^{+}15, ZZCZ21, ZZW^{+}22, ZLC^{+}23, ZZL^{+}23, ZZL^{+}24d, ADMM25, ADCB07, Ala21, AAT^{+}22, ASVE13, AA23, AS22b, AMT25, BBUD24, BWA13, BAK13, BTWZ22, BPFA24, BCNP24, CWTG22, CJP^{+}21, CL07, CML^{+}13, CXL^{+}22, CLZ^{+}23a, CSZ^{+}25, CSW^{+}25b, CZZ21, DWS^{+}24, DP06, DHP23, DVA21, DBSL24, FKCW20, FLM^{+}06, GLC05, GZW^{+}24a, HYLD20, HC22, IAJ^{+}24, JLZ^{+}21, JC10, JXTC21, JSL07, KVB^{+}24, KIT^{+}24, KK08, LSDJ06, LLKL11, LLHS12, LMXJ21, LZL^{+}21a, LYZX21, LXB^{+}22, LML^{+}24, LJW^{+}25, LWZ^{+}21a]. based [LWL08, LSQ11, LZL20b, LX21, LHS^{+}21, LYXY23, LYW^{+}25, LYD^{+}21, LYXA22, MGCH13, MAGT23, MB08, MHT^{+}13, NR24, NSK^{+}21, NW08, NWL^{+}20, PZ08, QRTM24, QLSQ12, QJ23, RHE10, RLAM25, RLXW24, REV^{+}12, SSP21, SG07, SMTR22, SCX^{+}25, SZHY19, SLZ^{+}21, SBS23, SDK^{+}21, SSSA23, SZZ^{+}23, SZFR23, TLZ^{+}21, TLY^{+}22, TFHM24, TMB^{+}22, TP22, TCJ08, UFJ21, VTP^{+}20, VMP20, VPSS^{+}13, WZC^{+}13, WDPX23, WLL23, WLS^{+}24b, WYGC24, WXS^{+}25, WXJL25, WLQL12, WJQ^{+}22a, WY23, XZCG23, XQG^{+}24, XGZ^{+}24, XMST07, XHZ^{+}21, XZL^{+}22b, BLS^{+}19, YWF^{+}24, YLLZ25, YZL^{+}14, YGMY24, ZSJL22, ZWM12b, ZMH^{+}20, ZHLJ22, ZTQX23, ZWLZ24, ZSO13, ZLL^{+}24b, ZG08, ZLF^{+}24, ZLF^{+}23, ZGY^{+}24b, ZWF^{+}20, GSB^{+}22]. Baseline [DSL^{+}22, XHL^{+}24b]. Basis [HZC^{+}16]. Bat [AAS^{+}20]. Bayesian [DLC^{+}22, Tas20, WHJ20, ZFSX21]. Be [GEL^{+}15, XWL^{+}25b]. Beat [ML11].

Beautiful [LXL^{+}14]. BEC [ZHD^{+}23]. Become [WWW^{+}22]. Becomes [JSEI16]. before [YZL^{+}14]. Behave [ROST20]. Behavior [CE23, CDL^{+}20, HDW^{+}18, MYGX21, NSK^{+}21, QRTM24, ZZX^{+}20, ZMZ^{+}23, BGP11, NH10]. Behavior-Oriented [NSK^{+}21]. Behaviors [LCC^{+}14b, MZL^{+}18, ZZB^{+}21]. Benchmark [BLW^{+}24, LSL^{+}20, SSZ^{+}23, SW18, WJ15a]. Benchmarking [ZSX^{+}25, ZGWZ23]. benefits [WWGT09]. Berkeley [MPSR05]. BERT [WZZ^{+}23]. Best [CHHH18, HT15, TB17, ZZMS14, ZTB20, CDGJ09, GPHOH12, JYZC12, KZHC13, SLYE11, Ste12c, XSEZ10]. Between [ACVM20, CTBC22, LZZ^{+}24c, XZY^{+}24a, XSL^{+}23]. Beyond [CYM^{+}25, GCTG24, GCC^{+}25, HTT^{+}11, HXH^{+}24, HJZ^{+}23, RXC14, TGC23, WCLC18, WHX^{+}25]. Bi [JWH21, LHY^{+}24, SR22, TFHM24, XZC^{+}25, LYC^{+}12, CJP^{+}21]. Bi-dimensional [SR22]. Bi-Directional [JWH21, XZC^{+}25, LHY^{+}24]. bi-layer [LYC^{+}12]. Bi-LSTM-based [TFHM24]. Bi-manual [CJP^{+}21]. Bias [GNC^{+}23, ZSZ^{+}25]. biased [LYC^{+}12]. Bibliometric [HGKS24]. BiC [HZS^{+}24]. BiC-Net [HZS^{+}24]. Bidirection [LSXZ23]. Bidirectional [HLW^{+}25, JWF18, LZL20b, WYM18, ZHD^{+}23, ZLC^{+}25, ZTX^{+}23]. Big [LCL22, WFZ^{+}21a, KS09]. Bilateral [LLL^{+}21, WCX^{+}14]. bilingual [ENHN09]. billions [ZR13]. BiLSTM [LYD^{+}21, SCX^{+}25]. BiLSTM-based [SCX^{+}25]. BIM [LCL22]. Bin [SB23]. Binary [CLP17, DHT^{+}19, LQH18, WCL^{+}22, XLZ^{+}22]. Binocular [PSL^{+}24]. Biological [CISF^{+}25]. Biomedical [RTR21, SSBT20, TKPL20, WWL^{+}23]. Biometric [BCNP24, DWS^{+}24, CW10]. biometrics [CFG^{+}24]. Biquadratic [TYY^{+}22]. BiRe [XLZ^{+}22]. BiRe-ID [XLZ^{+}22]. birth [Geo05]. Bit [HXH^{+}24, LZG^{+}25, PYZ^{+}20, SSP^{+}24, HH11b, LJP08].

Bitcoin [YWF^{+}24]. Bitcoin-based [YWF^{+}24]. Bitrate [APM21, MAGT23, SYS17, THA^{+}25, REV^{+}12]. Bitstreams [CQA^{+}24]. Black [CYZ^{+}24, WWY^{+}21]. Black-Box [WWY^{+}21, CYZ^{+}24]. Blend [JWW^{+}24]. Blind [AP10, GSB^{+}22, LCZ^{+}24a, LGLZ20, MBK^{+}23, RS16, YTR^{+}25, YWC^{+}25, ZYC^{+}24]. blink [WT10]. Block [CQA^{+}25, EC16, LHY^{+}24, TWKK21, YMS^{+}24]. Block-Level [TWKK21]. Block-Permutation [CQA^{+}25]. Blockchain [LHJ^{+}24, NPG^{+}22, RLXW24, WDS21, XZCG23, XGZ^{+}24, ZXZ^{+}24a, LWZ^{+}21a]. Blockchain-Based [NPG^{+}22, RLXW24, XZCG23]. Blockchain-driven [ZXZ^{+}24a]. Blur [JMS23]. BMIF [XZCG23]. Body [CC17, GWWC25, GMWW25, LZZ^{+}24c, STSW23, YCFX23, CJZ^{+}24]. book [ARE13]. Books [SVA^{+}21, RSB11]. Bool [XJB25]. Boolean [WY23]. Boolean-based [WY23]. Boosted [QZXH14, YZXY15]. Boosting [CLP^{+}23, DPL^{+}23, LGX^{+}24, LWLC24, SYY^{+}24, SHT^{+}24, WBH^{+}23, WLMZ25, XSL^{+}23, ZPY^{+}23, LGX^{+}08, WK10]. Bootstrap [WZD^{+}20]. Botanists [ZWJ^{+}24]. Both [Hon19]. Bottleneck [HYG^{+}25, KFP^{+}25]. Bottlenecks [ZMX^{+}23]. Bottom [KBB21, PLYM23]. Bottom-up [KBB21, PLYM23]. Boundary [HSR^{+}23, LZXY20, LLZ^{+}25, NXL^{+}25, YZ23, ZLT^{+}25a, ZXZ^{+}24b]. Bounded [WJS^{+}21]. Box [WWY^{+}21, CYZ^{+}24]. Brain [HAAM19, KMK^{+}21, LDL^{+}25, XYJ^{+}20]. Brain-Machine [LDL^{+}25]. Branch [CW25, CZS^{+}25, LYH24, SJC^{+}25, CH21, TH22]. Branched [ATMM24]. Brand [GUH^{+}20]. Breaking [ZMX^{+}23]. Bridge [SHF^{+}24]. Bridging [MKY^{+}24, ZZTL24]. broadcast [HH11c]. Broadcasting [BLW^{+}24, CGNG15, PCB^{+}21, SkFM18, ZRCH08]. Brownfield [TP22]. Browse [SX11].

Browser [LKM^{+}19, LVM^{+}21, RW12]. Browsing [LKM^{+}19, LLW^{+}13]. Brushstroke [FYY^{+}21]. Brushstrokes [WLZF22]. Buffer [BZDX^{+}18, CPCM21, LCK09]. builder [BB11]. Building [LLSC12].

CABAC [CI23]. Cache [KZGH15]. Cache-Centric [KZGH15]. Caches [KZGH15]. Caching [ABR17, CPP^{+}14, CE23, HJWW19, CPP^{+}13, GS11b, ILL08, LS05, ZO13]. Calibration [FZZ^{+}25, LWL24, SO22, WWY^{+}21, AP13, WGL^{+}23]. Call [Ano13, BLJX10, GS11a, KDC08, Ste12c]. Calorie [SYX23]. Calorie-aware [SYX23]. Camcorder [HLS^{+}22, SKW^{+}15]. Camera [GEL^{+}15, HZSC20, LCC^{+}14a, LZL^{+}24b, NAK15, ZDZ^{+}22, ZZCZ21, XHL^{+}24a]. Camera-Aware [LZL^{+}24b]. Cameraman [GEL^{+}15]. camerawork [GL08]. CamMark [SKW^{+}15]. Camouflaged [CW25, LTX^{+}23, NXL^{+}25]. Campaign [HJMY15]. campus [CY11]. campus-wide [CY11]. Can [GEL^{+}15, SkFM18, WDR^{+}24, XWL^{+}25b, YNY25, LW07]. Cancer [WZD^{+}20]. candidates [SLYE11]. Canonical [LKM17]. Capacity [MSS^{+}24, GZHD12]. CapST [APC^{+}25]. Capsule [APC^{+}25]. CAPTAIN [FKW22]. Caption [LX21, YBTX23, AYH^{+}23, HMUC21]. Captioned [AMG23]. Captioning [AYH^{+}23, CYW24, CPL^{+}23, CZZ^{+}24, CBSC18, DNL^{+}23, HH19, JW21, JWH21, LML^{+}24, LLL^{+}24a, LYL^{+}25, LYD^{+}21, MOL^{+}22, MXZ^{+}24, NDC^{+}23, NCL^{+}23, PLYM23, SCB^{+}24, SXY^{+}23, TWL19, WHY18, WYM18, WLH^{+}21, WHW18, WHY19, WLCH22, WZZ^{+}22, XLZ^{+}24, YHXL20, YZG^{+}20, YTL^{+}24b, ZGY^{+}24a]. Capture [HSR18, LZL20a, ZRCH08]. capturing [ZYM^{+}10]. CAQoE [JD23]. Car [BCNP24]. Cardinalities [AR15]. Care [RHAG21, Whi13]. Caricature [HLC^{+}25]. Carrying [YWC24]. Carving [Hon19, ZHG^{+}21]. Cascade [ZHD^{+}23]. Cascaded [FNH22, JK25, LYC^{+}25, STSW23, ZT22]. Cascades [CSJC17]. Cascading [JWW^{+}24]. Case [HHH22, RXC14]. Cast [HLD18]. casual [GL08]. Categories [ZLL20]. Categorization [HLZ^{+}21, SSSK18, WTD^{+}21]. Category [BSQ^{+}24, FJF^{+}25, LXL^{+}23, SWK^{+}22, ZCX^{+}24]. Category-Level [BSQ^{+}24]. Category-Region-Aware [FJF^{+}25]. Category-Stitch [LXL^{+}23]. Causal [DWS^{+}24, JLK^{+}24, PLZT22, Tas20, Tas22, YLYB24, ZCT^{+}07]. Cause [CHHH25]. Cause-Guided [CHHH25]. CCA [ZYO20, ZYT^{+}23]. CD [ZSYW23]. CDN [AH20, YLZ^{+}10]. Celebrity [LSK^{+}15]. CelebrityNet [LSK^{+}15]. Cell [PA20, TS20]. Cellular [YZL^{+}24, ZQRS18]. center [VPSS^{+}13]. centered [FPH^{+}08, WBL09, CBS08]. Centric [CWXW25, GCC^{+}25, KZGH15, PVWD18, ZLT^{+}25b, DCO10, NLKB24, PKLK23]. cepstral [ML11]. certainty [NT08]. Ch [YLD^{+}25]. Ch-Transformer [YLD^{+}25]. Chaff [DKJ^{+}21]. Chaff-less [DKJ^{+}21]. Chain [HCY^{+}23, LLC^{+}24a, RLXW24, XGZ^{+}24, XGZ^{+}24]. Chain-of-Thought [LLC^{+}24a]. Challenges [BLH^{+}25, GTLG14b, PVWD18, RSAH25, Sin21, SKVHC18b, TAT^{+}25, USM^{+}25, LSDJ06, UTK^{+}08]. Chamfer [ZSYW23]. Change [FQJ^{+}25, YJTN18, YTL^{+}24b]. Change-Resistant [YJTN18]. Changing [CHS^{+}23]. Channel [GQSG24, HCJ^{+}24, KLS^{+}18, LYL^{+}21a, LZLJ22, LHC^{+}25, TLZ^{+}21, WSW^{+}24, XLF^{+}25, XYL^{+}23, ZHD^{+}23, HH11c]. Channel-aware [XYL^{+}23]. Channel-Level [HCJ^{+}24]. Channel-wise [GQSG24]. Channels [YP15, CLC05]. Character [TWY24, XYC^{+}23, SY09]. Characteristic [CSZ^{+}25]. Characteristics [LGLZ20, LZZ^{+}24c, PYZ^{+}20, VGNL10]. Characterizing [HPH^{+}20, MZL^{+}18]. Characters [TWY24]. Charts [MKC21]. Chat [WH22]. ChatGPT [HQLW25]. Cheating [SZFR23]. Chebyshev [HD19]. Chief [dB16]. Child [QRTM24]. Child-computer [QRTM24]. Children [AMC^{+}25, DMF^{+}20, ONH07]. China [Hua23]. Chinese [LYD^{+}21, WXQC20]. choice [NT08]. Chrominance [ZCQ^{+}23]. Chunk [OE19]. Chunked [BTBZ20]. Chunklets [KABB20]. chunks [SX11]. Cinematographic [WPRC18]. cinematography [GL08]. Circle [HLX^{+}14]. Cities [GCF^{+}21]. City [LCL22]. CL [BPB^{+}22, ZWJ^{+}24]. CL-LeafNet [ZWJ^{+}24]. Class [KIT^{+}24, LZH^{+}24b, LYL^{+}23b, SYY^{+}24, WCF^{+}17]. Classes [SS24]. Classification [Ala21, BRG24, CWTG22, CIPE18, CAJ19, CSW^{+}25a, FZW^{+}25a, FJF^{+}25, FBGD23, FLG^{+}21, HHL^{+}22, HXH^{+}24, JCSL19, KIT^{+}24, KN21b, LZH^{+}20, LZH^{+}24b, LZZ^{+}24b, LHS^{+}21, LDL^{+}25, LXZ^{+}25, LJG^{+}24, MEA^{+}21, NSK^{+}21, NDX^{+}21, NLW^{+}21, QZXH14, RUD23, RHS^{+}20, TYY^{+}22, TDW^{+}23, TS20, WCLC18, WLHT19, WZ20, WBH^{+}23, WZWY23, WLFL23, WJ15b, XHL^{+}21, YBW^{+}25, YGL^{+}23, YCZ^{+}23, ZRZX23, ZSJL22, ZZP^{+}20, ZTR^{+}22, ZLZ22b, ZHC^{+}23, ZXD^{+}23a, ZXD^{+}23b, FLM^{+}06, JCC^{+}10, LYC11, TV07, ZI13]. Classifier [ZH18, ZYQM24, LGX^{+}08]. Classifying [TWB^{+}23]. Click [LPY^{+}19, TYY^{+}18]. Click-Based [LPY^{+}19]. Client [WBL09]. Client-centered [WBL09]. Clientless [yHcCzH^{+}21]. Clients [YJM^{+}19, GSM^{+}08, WLHL13]. Clinical [YRE^{+}20]. Clip [CLZ^{+}23a, PN16, BBUD24, JCB25, YZZZ25]. CLIP-based [BBUD24]. Clip-level [CLZ^{+}23a]. Clips [HCM^{+}25, LMF^{+}14, WWW^{+}22]. close [HCKL13]. Cloth [RTM^{+}24, YZD^{+}25]. Clothed [ZXS^{+}25]. Clothes [CHS^{+}23]. Clothing [JWF18, LSN^{+}20, LZZ^{+}24c, LZL^{+}24c, SLG^{+}24]. Cloud [AGC^{+}18, ACGH18, AR15, CWL24, CZZ21, HLD18, HDZ^{+}15, HH25, HSS^{+}25, IVS^{+}20, JLL^{+}25, LA15, LSCH25, LDT^{+}18, LYXY23, NCMM21, QLGL24, RTR21, RBLJ25, RSR25, SRAA17, SSSK18, TRRB20, WCY^{+}18, XHZ^{+}21, XZY^{+}24b, YZY^{+}25, ZLW17, ZZL^{+}25b, ZZL21, ZXZ^{+}24b, ZZY^{+}23, LOJZ18]. Cloud-Assisted [WCY^{+}18, ZLW17]. Cloud-Based [HLD18, SRAA17, CZZ21]. Cloud-Device [JLL^{+}25]. Cloud-Edge [XHZ^{+}21]. Clouds [dMGRT24, LSS^{+}24a, WHX^{+}25, WXS^{+}25, YGL^{+}23, ZYC^{+}24]. CloudVR [MSKYJ21]. Cluster [LGM^{+}24, ZYO20]. Cluster-CCA [ZYO20]. Clustering [CZW15, FZYY18, FMG20, FBG22, GLL^{+}24, HCW^{+}07, KPK^{+}24, RVTC25, SMK^{+}24, TGSF21, TKK^{+}17, WYGC24, XYJ^{+}20, XWDC23, YGNT19, YZ23, ZWL15, ZLW^{+}21, ZXS^{+}24, ZZZ^{+}22, CFGW05, hHLC10, ULIS07]. CM [PQ19]. CM-GANs [PQ19]. CMAF [YSSF24]. CMHNE [HQF^{+}19]. CNN [DVA21, LZH^{+}20, LCL^{+}23b, LJZ^{+}22, LWZ21b, SG22b, TLZ^{+}20, WFC^{+}25a, WHZ^{+}24, ZSLW23, ZZY18]. CNN-based [DVA21]. CNN-RNN [LZH^{+}20, ZSLW23]. CNNs [GSDT21, HGGZ23, NRUT20, SSY20]. Co [HQF^{+}20, JWH21, LYW^{+}25, SHSR24, ZCY^{+}19, ZTB20, ZX14, CQA^{+}25]. Co-Attention [JWH21, SHSR24, HQF^{+}20]. Co-Located [ZTB20]. Co-Occurrence [ZCY^{+}19]. Co-occurring [LYW^{+}25]. CO-PMHT [ZX14]. Co-XOR [CQA^{+}25]. Coarse [CYZ23, HSR^{+}23, HJZ^{+}23, LPW^{+}22, LKL^{+}25b, LYZZ24, SHWC19]. Coarse-to-Fine [CYZ23, HJZ^{+}23, LKL^{+}25b, LYZZ24, SHWC19]. COBIRAS [SSP^{+}24]. coconstruction [VNC^{+}11].

Code [ZLS^{+}25]. Codebook [JCSL19]. Codec [LTHT24, ZLL^{+}24b]. Codecs [SK23]. Coded [LSS^{+}24a]. Codes [CLP17, KMSW18, LQH18, LZLJ22, MWL^{+}24, PRH14, XLN^{+}21, ZZL^{+}24b]. Codesign [HDZ^{+}15]. Coding [CI23, DG17b, GZL^{+}20, GGML22, HHH22, HYLG23, LYZ^{+}18, LLL^{+}25, LJ25, LJZ^{+}22, LZG^{+}25, SAF19, XFSZ20, ZLK^{+}19, ZZL^{+}23, ADCB07, GL12, HH11c, MC11, RHS12, ZLLT13]. Cognition [WBR25]. Cognition-Inspired [WBR25]. Cognitive [WJXN24, ZYL^{+}24, MC11]. Coliseum [BBT^{+}05]. Collaboration [AH20, CZZ21, FAA18, JLL^{+}25, JXTC21, Wan21]. Collaborations [KLS^{+}18]. Collaborative [HQF^{+}19, LOJZ18, SWM^{+}24, WC12, WKE16, ZWZ^{+}23b, CZ24, hHLC10, LT14]. Collage [BC15]. Collapse [WWY^{+}21]. Collection [TFY^{+}25]. Collections [FHG^{+}17, CFGW05]. Collective [Yan17]. Collectiveness [LCW16]. Collision [BOZ17]. Collocated [LZL^{+}24c]. collusion [FLZ^{+}12]. Color [HD19, LZD^{+}22, LWP22, LZG^{+}24, MSS^{+}24, PKLK23, QDX^{+}24, SGY^{+}24, WSW^{+}24, XLH^{+}23, YWNW15, YDCZ23, YZ23]. Color-gray-difference [YDCZ23]. Color-Unrelated [XLH^{+}23]. Colour [ARA^{+}23]. Combined [ZQKH19, QLSQ12]. Combining [CC17, WHZ^{+}24, WJ15b, XZY^{+}24b, YLCC18]. come [Cha13]. Comfort [Chu15]. Comic [YMY^{+}21]. Comics [TKK^{+}17]. comment [CTGP08]. Commenting [FFCM24]. Commerce [VVSV17]. Common [PQ19, ZWH^{+}23]. Commonsense [MYX23a, WGHX24]. Commonsense-aware [WGHX24]. Communicate [PRH14]. Communication [AAT^{+}22, AS22b, CIPE18, CCF^{+}24, GNC17, HWLC19, MON21, MSG22, NLKB24, WDE25, ZYL^{+}24, ZZL^{+}25d, Cho13]. Communication-Efficient [ZYL^{+}24]. Communications [BLJX10, DG17a, GZ20, Geo05, HGKS24, JTZ^{+}16, Tas20, Tas22, Tas24, ZLT^{+}25b, SSTK07]. Communitarian [PRPO23]. communities [ZCY^{+}13]. Community [CSW^{+}25a, LXL^{+}18, YWG^{+}20, SSS13]. Commuting [ACVM20]. Compact [DHT^{+}19, MCM^{+}09, SZTL16]. Comparison [LSL^{+}20, THA^{+}25, ZWM12a, ZDZ^{+}23, HM10]. Comparisons [yHcCzH^{+}21]. Compatibility [QSZ^{+}21, YSF^{+}21]. Compatibility-Aware [QSZ^{+}21]. Compatible [BPB^{+}22, PFC^{+}24, XJG^{+}22]. Compensated [GKSB17]. Compensation [LPS15, PYBY25]. Complementarily [FJF^{+}25]. Complementarity [XZY^{+}24a]. Complementary [CYZ23, LSN^{+}20, TWL19, XPP^{+}23, ZYY^{+}20]. Completely [LCZ^{+}24a]. Completion [HZC^{+}16, LZZS23, LLCH17, LLWX24, MHW^{+}19, XLX^{+}24, YLL^{+}23, ZZC^{+}25b, LYCJ12]. Complex [CW25, CGGC20, CSW^{+}25b, GHZW25, NSM^{+}24, TWFW24, ZJL^{+}21, MVW07]. Complexity [JYZ^{+}25, LPS15, SAF19, ZZLL17, HH08b, HXL24, IB10]. Component [AMMG16, VMP20, ZZB24]. Component-based [VMP20]. Components [LXRL23, ZCQ^{+}23]. Composed [BBUD24, DDL^{+}25]. Composing [HSS^{+}25, ZXX22]. Composite [WZ20]. Composition [FKW22, WWY^{+}24, ZYZ^{+}25b, EGEM06]. Compositional [LJW^{+}25, ZML^{+}24]. Comprehension [UFJ21, WKS^{+}23]. Comprehensive [AS22b, CCD24, CLWW24, DOD23, FKW22, SSZ^{+}23, SZF^{+}25, ZZS^{+}25]. Compress [FBGD23]. Compressed [DG17b, HZL^{+}21, HSG23, LZL^{+}24a, QJ23, WXZ^{+}23, XZY^{+}24b, YKQ^{+}21, ZRZX23, LCSX11, LMC^{+}22, NH10, RBBS25]. Compressing [WFC^{+}25a]. Compression [APP^{+}22, CWL24, HLW^{+}21, HSL^{+}20, JYZ^{+}25, JLLL25, LLS^{+}21, LLYH14, LDZ^{+}20, PKLK23, PFC^{+}24, QLGL24, SK23, SS22, TZ24, TZZ25, TWKK21, TLZL24, WHL^{+}21, WSW^{+}24, WHX^{+}25, XZC^{+}25, YYX^{+}24, YZJ^{+}25, YNY25, YZY^{+}25, YGMY24, ZZB24, ZZT24, FDKB11, KK08, KMP05]. Compressive [EC16, HXL24]. Computation [CLS17, GYF^{+}21, SZHY19, YH14, ZYL^{+}17, San11]. Computational [LK07, TKPL20, YTR^{+}25]. Computed [PDD16, RLAM25]. Computer [HLW^{+}21, KPK^{+}24, LPD^{+}24, QXG^{+}24, RPE^{+}17, WCF^{+}17, ZSZ^{+}24a, And13, CCD07, QRTM24, WYM07]. Computer-Aided [RPE^{+}17, WCF^{+}17]. Computing [AGC^{+}18, ACGH18, BLJX10, CCF^{+}24, CLS^{+}21, GLGG25, Geo05, HC22, HGKS24, MZL^{+}18, MSKYJ21, NCMM21, NLKB24, WJXN24, WFZ^{+}21b, WCL23, XFQ^{+}21, XZL^{+}22b, YXZQ22, ZZC^{+}23, ZGD21, ZZS^{+}25, ZWS^{+}20, ZJSJ20, LWL08]. Computing-Assisted [YXZQ22, MZL^{+}18]. Computing-Based [WCL23]. Concealment [HSS^{+}25]. Concept [MGS18, PDD16, SWK^{+}22, WCF^{+}17, YZG^{+}20, BAK13, JCC^{+}10, KIT^{+}24, LGX^{+}08]. Concept-Based [PDD16]. concept-oriented [LGX^{+}08]. Concepts [SKVHC18b, ZXD^{+}23a]. Concurrent [UFJ21]. Conditional [LLL^{+}24a, LYL^{+}25, PKL25, TXW^{+}25, WWL20, WWL^{+}24, XLZ^{+}24, YSC21, YH13]. Conditions [HCM^{+}25]. conductor [RWP07]. Conference [BSH08, CHHH18, JPS05, NLS13, PSS05, TB17]. Confidence [CDL^{+}20, GLC05]. confidence-based [GLC05]. confidentiality [YC08]. Configuration [SSSK18]. Configurations [HHH22]. Confusion [IKTD25]. Congestion [GNC17]. Connected [FZW^{+}25a, SG22b, YXYB21]. Connection [AAS^{+}20, CLS17, CSSZ19, LCS17, LCL^{+}21]. Connections [LSYM19, YW23, TNEcC08]. connectivity [CKRB23]. conquer [YGMY24]. Consensus [WHW18, XGZ^{+}24].

Consensus-based [XGZ^{+}24].

Consideration [YHZ19]. considerations [PZ08]. Considering [Hon19, YBO14]. Consistency [LQD^{+}24, LWHC24, WC23, WKE16, XJW^{+}22, XLS^{+}25, YPSC22, ZWYS20, MMW10]. Consistent [LXT^{+}24, MSFR25, WLZ^{+}24a, XW17, ZWY21, CHLW19, XCM^{+}23, ZWZL22]. Constant [CPSH14]. Constrained [GZZ^{+}24, PB14, WXQC20, YHXL20, FDKB11]. Constraint [NDX^{+}21, SGY^{+}24]. Constraints [CF22, LZDL25, LGC^{+}24]. Constructed [LSK^{+}15]. Constructing [ZG08]. Construction [HCJ^{+}24, LYXY23, LCL22, THA^{+}25, CML^{+}13]. Constructs [Tas22]. Consumption [SS17]. contained [AMG23]. Content [BRZS18, BCP14, CZY^{+}21, VTD22, DP06, EAK^{+}25, FF23, FSK^{+}15, HMOS17, HXZ^{+}20, KKGE18, KABB20, LSDJ06, MY15, MHT^{+}13, PYZ^{+}20, SVA^{+}21, TP22, WXZ^{+}23, WHY19, XMST07, YSZ15, ZZY^{+}14, ZDD^{+}24, ZZW^{+}19, ZYZ^{+}25b, DY09, Eff13, EGEM06, GA12b, GFB^{+}14, GL12, HMVI13, LLC11, SG07, WWL13, WZC^{+}13, WBL09, Yan10, ZO13]. Content-adaptive [XMST07]. Content-Agnostic [VTD22]. Content-Aware [FSK^{+}15, KKGE18, ZDD^{+}24, HXZ^{+}20, WHY19]. Content-Based [HMOS17, ZZY^{+}14, DP06, LSDJ06, MHT^{+}13, TP22, SG07]. Contents [Ano11, Ano12, Ano14, Ano20, Dao17, Sha21, TOM12, ZCY^{+}13]. Context [APM21, CLZ^{+}21, CC17, CBSC18, GHY^{+}24, HL24, HLS^{+}22, JD23, LWH^{+}22, LL15, RK15, SBU^{+}17, SGY^{+}23, SB23, TCS^{+}25, WDPX23, WLMZ25, WWL^{+}23, XWL^{+}25a, YSZ15, YCLH22, ZHG^{+}25, APV08, ADCB07, CCK06, FLM^{+}06, KTM^{+}06, KS13, LWL08, RW12, YGHH12, ZI13]. Context-Assisted [TCS^{+}25]. Context-Aware [LL15, RK15, SBU^{+}17, SGY^{+}23, YCLH22, JD23, LWH^{+}22, LWL08].

Context-Based [SB23, FLM^{+}06]. Context-detail-aware [HL24]. Context-Feature [GHY^{+}24]. context-specific [ADCB07]. Contextual [DYN^{+}25, HQF^{+}20, LJW^{+}25, MPE^{+}11, MYX^{+}23b, SHT^{+}24, ZCXL23, LYCJ12, MLHL12, SGS21]. contingent [DÇ07, LW07]. Continual [XHL^{+}24b, YWLZ23]. continuity [LYC^{+}12]. continuity-biased [LYC^{+}12]. Continuous [FDBP23, SSP^{+}24, ZYCD24, SKSZ13]. Contour [LZT^{+}20]. contract [ZWM12a]. Contrast [LZD^{+}22, SB23, WCG^{+}25, WQL18, YDCZ23, ZWX24]. Contrast-distorted [YDCZ23]. Contrast-Enhanced [LZD^{+}22]. Contrasted [MYX^{+}23b]. Contrastive [BBUD24, CYZ^{+}24, CZX^{+}24, GZW^{+}25, HWZL24, LZZ^{+}24a, LYH^{+}25, LSX^{+}25, LGW^{+}25, LZXP25, MXZ^{+}24, PZJL22, RE24, SHF^{+}24, WLZ^{+}24b, XLZ^{+}24, ZLL^{+}24a, ZLX24, ZCYL24, ZZC^{+}25a, ZSZ^{+}25, ZLF^{+}24]. Control [CDZ^{+}17, GNC17, MY15, MN16, NAK15, WJS^{+}21, XWZ^{+}24b, YXJ^{+}24, ZZLL17, CCG^{+}08, CW10, GHP^{+}06, HNL08, HCKL13, HH11c, IB10, LJP08, PZ08, PBS12]. Controllable [LYL^{+}25, SLL^{+}24, WWY^{+}25]. Controlled [BFAS15, CYW24, SLZ25, AYH^{+}23]. Controller [Cla18]. Controlling [FPA24, WWX^{+}21]. convenient [LLW^{+}13]. Convergence [SCFL14]. Conversation [FQJ^{+}25, LWY22, SO22]. Conversational [BGP11]. Conversations [BPFA24]. conversion [HYLD20]. Convex [THA^{+}25]. ConvNet [TS20]. Convo [BV25]. Convolution [JLL^{+}21, LZJ^{+}20, LHZ^{+}23, LD25, MXH^{+}23, XLZ^{+}21]. Convolutional [ARA^{+}23, AM24, CZZ^{+}23, CPL^{+}23, DHT^{+}19, DLL^{+}20, FLG^{+}21, HZH24, KMK^{+}21, LWZH19, LZC^{+}19, LWP22, LYW^{+}22, MAZ22, QHFX21, SL23, WCCN23, WFC^{+}25a, WYMX23, XFZ^{+}19, XHL^{+}21, XLW^{+}23b, YLCC18, ZZW^{+}24, ZZG^{+}20, ZZL25e]. CookGAN [LNH23]. Cooperation [SRPH16]. cooperative [GS11b, ILL08]. Coordinated [LS05]. coordinates [AH12]. Coordination [LMXJ21, NAK15, OMP07]. Copies [SKW^{+}15, CWC10, KH13]. Copy [CZZ^{+}23, LYC^{+}25, LTL^{+}24, LH20, YP15, YTL^{+}24a, MHT^{+}13]. Copyright [GCTG24, MSS^{+}24, NPG^{+}22, WWZ^{+}25, WFZ25, ZNYZ25]. corners [VPSS^{+}13]. Corpus [Hua23]. Corrected [HYLG23]. Correcting [HWLC19]. Correction [YWC24, ZDZ^{+}22, WCK05]. correlated [AKO07]. Correlates [KYVE14]. Correlation [DZW^{+}21, DLL^{+}20, HCJ^{+}24, LYZY21, LZZ^{+}25, LKM17, LWZW21, PRGA18, PSL^{+}24, RLY^{+}21, SZM^{+}21, SZL^{+}24, WLSZ22, YLK^{+}20, YZZZ25, YP20, YTRC19, YCLH22, ZZJ^{+}23, ZLC^{+}25, ZXS^{+}24, ZZLL17, ZZL^{+}24c, ZWF^{+}20, CL07, CL12]. Correlation-Based [ZZLL17]. Correlation-CLIP [YZZZ25]. Correlative [QHR^{+}08]. Correspondence [FWLA15, JLL24, WCX^{+}14, WLCG21, XLT^{+}23]. Correspondences [MHCG19]. Corresponding [SSK20]. Corruption [XSD^{+}22]. Corruption-Agnostic [XSD^{+}22]. Cosine [LLC^{+}21]. Cost [CCF^{+}24, LDT^{+}18, YF22]. Cost-effective [CCF^{+}24]. Cost-Efficient [LDT^{+}18]. Counterfactual [HWZL24, LYX23]. Counterfeit [CSSZ19, CSSZ23]. Counterfeit-goods [CSSZ19]. Counterfeiting [CQA^{+}25]. Counterparts [SZLL17]. Counting [LZZ^{+}23a, WMW^{+}22, ZZL^{+}17, ZJZC20, ZZL^{+}24a, ZLT^{+}25a]. Coupled [SSZ^{+}23]. Course [ZYLN21]. Courses [ZYLN21]. Covariance [GZL^{+}20, ZFSX21]. Covert [AS22b, JTZ^{+}16, ZZL^{+}25d]. COVID [ASI^{+}21, CLZ^{+}21, LWZW21, LCL22, RHAG21]. COVID-19 [ASI^{+}21, CLZ^{+}21, LWZW21, LCL22, RHAG21]. CovLets [ZGM^{+}20]. CRAR [ZT22]. Creating [CS22, TKK^{+}17]. Creation [FF23, HCM^{+}25, MGG17, QSZ^{+}21, WXQC20, AVJ05, MB08, RSB11]. Creative [MHF24]. Creativity [EAK^{+}25, PDD22]. Creator [RABC24]. Creators [JSEI16]. Credit [ZGR21]. criterion [PZ08]. criterion-based [PZ08]. Critical [ZCT^{+}07]. Cropping [ZYZ^{+}25b, FDKB11]. Cross [BXMH15, CW25, CFN^{+}23, CE23, CDL^{+}20, CZX^{+}24, CZQ^{+}22, DWL^{+}23, EWSZ15, FWLA15, GJY^{+}25, HCZ^{+}22, HLW^{+}24, HLY21, HSN^{+}14, HPW20, JWF18, JCB25, KLX^{+}24, LZL20a, LYL^{+}21a, LLG^{+}23, LLZ^{+}25, LD25, LWZ^{+}21a, LYDY24, LKM17, LWZH19, LLN^{+}24, LSS^{+}24b, LDL^{+}25, MAE^{+}21, MSFR25, PQ19, RLXW24, RBLJ25, SZM^{+}21, SZL^{+}24, SHF^{+}24, SJC^{+}25, SLZ25, TLY^{+}22, WKS^{+}23, WLS^{+}24b, WSW^{+}24, WXH^{+}25, WCG^{+}25, WJQ^{+}22b, XYJ^{+}20, XGZ^{+}24, XWH^{+}21, XTL^{+}21, XZL^{+}25, XLS^{+}25, XZC^{+}25, YSXH16, YBW^{+}25, YTOH22, YZXY15, YLK^{+}20, YSSF24, YLL^{+}24, YZZZ25, YP20, YTRC19, ZYO20, ZCXL23, ZWH^{+}23, ZX14, ZWY21, ZZJ^{+}23, ZYT^{+}23, ZYQM24, ZCL^{+}24, ZGY^{+}24a, ZLLH25, ZLJ^{+}25, ZZL^{+}24c, ZFC^{+}24, ZZL^{+}24d, HH11a, PS05]. Cross-Attention [SJC^{+}25]. Cross-blockchain [LWZ^{+}21a]. Cross-chain [RLXW24, XGZ^{+}24]. Cross-Channel [LYL^{+}21a, WSW^{+}24]. Cross-Consistent [MSFR25]. Cross-Dataset [TLY^{+}22, WCG^{+}25]. Cross-Domain [SZM^{+}21, WXH^{+}25, XYJ^{+}20, YZXY15, ZX14, ZFC^{+}24, LSS^{+}24b, ZZJ^{+}23, ZYQM24, PS05]. Cross-interaction [YBW^{+}25]. Cross-Layer [CW25, EWSZ15, HSN^{+}14, HH11a]. Cross-level [LLZ^{+}25]. Cross-lingual [ZGY^{+}24a]. Cross-media [HPW20].

Cross-Modal [CZX^{+}24, FWLA15, GJY^{+}25, HCZ^{+}22, JCB25, LYDY24, LLN^{+}24, LDL^{+}25, MAE^{+}21, SZL^{+}24, SHF^{+}24, SLZ25, XWH^{+}21, XZC^{+}25, YTOH22, YLK^{+}20, YSSF24, YZZZ25, YP20, YTRC19, ZYO20, ZLLH25, ZLJ^{+}25, ZZL^{+}24d, CDL^{+}20, CZQ^{+}22, DWL^{+}23, HLW^{+}24, KLX^{+}24, LD25, PQ19, WKS^{+}23, WJQ^{+}22b, XTL^{+}21, XZL^{+}25, XLS^{+}25, YLL^{+}24, ZWH^{+}23, ZWY21, ZYT^{+}23]. Cross-Modality [LLG^{+}23, LWZH19, WLS^{+}24b, ZZL^{+}24c]. Cross-Network [YSXH16]. Cross-Platform [BXMH15]. Cross-Resolution [CFN^{+}23]. Cross-session [ZCL^{+}24]. Cross-slice [ZCXL23]. Cross-space [RBLJ25]. Cross-Triplet [JWF18]. Cross-User [CE23]. Cross-View [HLY21, LKM17]. CrossFormer [LYDY24]. Crossmodal [MHCG19]. Crowd [GF17, HDZ^{+}15, LZZ^{+}23a, MC19, WMW^{+}22, YHH^{+}22, ZJZC20, ZZL^{+}24a]. Crowd-Cloud [HDZ^{+}15]. Crowd-sourced [MC19]. Crowded [ZZB^{+}21]. CrowdGraph [ZZL^{+}24a]. Crowdsensing [WWZ24]. Crowdsourced [ZLW17, ZMH^{+}20]. Crucial [LLL^{+}22]. Cryptography [LZD^{+}22, SA16, SZFR23, WFC^{+}25b, YWC24, WYK12]. CryptoLesion [TRRB20]. Cryptosystem [DDK24, SZHY19, TWM^{+}25]. CSG [CXX^{+}24]. CT [XYJ^{+}20, XZLZ23]. CTU [ZZLL17]. CTU-Level [ZZLL17]. CU [YGMY24]. Cubism [TLY^{+}22]. Cue [ZWL^{+}17]. Cues [GCC^{+}25, XXG^{+}21, XFSZ20, NLN^{+}13]. Cultural [BBB^{+}23]. Cumulative [LZL21b, TNH^{+}21]. CUR [XLW^{+}23b]. Current [Sin21, TAT^{+}25]. Curriculum [LJG^{+}24, PLZT22]. Curve [PLB^{+}24]. Cutting [GGB24]. Cutting-edge [GGB24]. Cyber [TFY^{+}25]. Cyber-Physical [TFY^{+}25]. Cycle [SL22, YTL^{+}24b, ZWZL22]. Cycle-consistent [ZWZL22]. Cyclic [ZZY^{+}23]. Cyclone [SYM^{+}25]. CZLoD [WAK^{+}12].

D [SAZ^{+}15, Ste10, Ste12c, AP10, ARE13, ATMM24, BSQ^{+}24, CFN^{+}23, CEE09, CRL20, CCH^{+}24, Chu15, DP06, DVA21, DÇ07, GZHD12, GLW20, GHQ24, GS11a, HH12a, HLM^{+}22, HSZ^{+}18, HLC^{+}25, JP11, KH13, LLP06, LYZ^{+}18, LLS^{+}21, LCL^{+}23b, LZWC23, LGM^{+}24, LSS^{+}24a, LC12, LCT^{+}12, LOJZ18, LZC^{+}19, LKW^{+}22, LL23, LCZ^{+}24b, LZYY24, LDZ^{+}24, LWWZ20, MC11, NLW^{+}21, NALM23, PB19, PB14, PJL23, QXG^{+}24, RS16, RHS12, SHOG12, SGY^{+}23, SZZ^{+}23, SHF^{+}24, SLNL20, SHWC19, SWK^{+}22, TLZ^{+}20, WCLC18, WXW^{+}22, WWS^{+}24, WAK^{+}12, WHY19, WJQ^{+}22b, WWWL23, XFSZ20, XZC^{+}25, YWN^{+}10a, YWN^{+}10b, YI14, YGL^{+}23, YH14, YLHL22, YCFX23, YHZ19, ZSYW23, ZDE16, ZFSX21, ZCL^{+}24, ZHY^{+}25, ZLC^{+}25, ZYC^{+}24, ZFY^{+}25, vRPPP23]. D-CNN [TLZ^{+}20]. D-Convolutional [LZC^{+}19]. D-Ghost-Free [LL23]. D-HEVC [LYZ^{+}18]. DaaS [ZWZL21]. DAG [GHY^{+}24]. DAG-YOLO [GHY^{+}24]. Daily [GCL^{+}25]. Dance [WWW^{+}22, WSLM18, ZWC^{+}22]. DanceNet [ZWC^{+}22]. Dancer [HYSL20]. Dark [HRL^{+}24]. Darkness [ZXT^{+}25]. DASH [AS20, BYOZ20, KCC17, OE19, SSP^{+}24, WBRZ17, YJM^{+}19, YDJ24, ZQRS18]. DashReStreamer [HCM^{+}25]. Data [ASI^{+}21, BYM^{+}18, CYM^{+}25, CSBJ24, FSEW25, FYZ^{+}21, FMG20, GLL^{+}24, GCL^{+}25, HCM^{+}22, IPMD25, IAJ^{+}24, KBI^{+}23, LZY^{+}24, LJP08, LHY^{+}24, LHJ^{+}24, LCL22, NRUT20, PJL23, RLAM25, SMK^{+}24, Sin21, SLK21, SAL^{+}21a, SKC24, SKA24, TYY^{+}18, TFY^{+}25, TKPL20, TKK^{+}17, TJZ^{+}25, TNP^{+}18, VF25, WFZ^{+}21a, Wan21, WCL23, WSLC24, WF25, WWWL23, XQG^{+}23, XW25, XLZ^{+}23, XHYX23, XZL^{+}25, XZC^{+}25, YXZQ22, YHX^{+}23, YHQH17, YCZ^{+}24, ZZP^{+}20, ZLW^{+}21, ZCY^{+}19, ZWS^{+}20, ZJSJ20, ZLH16, ZLT^{+}25b, vRPPP23, AP10, ATM06, COM^{+}11, KS13, LT14, LH12, RHE10, WDCX07, YC08]. Data-Dependent [WWWL23]. Database [THR^{+}22]. databases [HOSS13]. Dataset [BBB^{+}23, CISF^{+}25, HLY^{+}23, HZZ^{+}24, LXJ^{+}24, LSL^{+}20, LYXY23, MSFR25, SSZ^{+}23, TLY^{+}22, WMH^{+}22, WCG^{+}25, XHL^{+}24b, ZDZ^{+}23, ZSX^{+}25]. Datasets [Che24, JLL^{+}22]. DBDGAN [JMS23]. DBGAN [LYH24]. DCNNs [MGG17]. DCVS [FBSGR25]. DDIFN [LLZ^{+}23]. DDoS [SP21]. DDPM [WZZ^{+}24]. De-raining [SGS21]. De-redundancy [WGL^{+}23]. dear [Ste10]. Debiasing [ZML^{+}24]. Deblurring [LSXZ23]. decades [LWLZ13]. Decentralization [WAC^{+}25]. Decentralized [GHS^{+}24, GAH25, KT21]. Decision [GLT^{+}20, LYZ^{+}18, SAZ^{+}15, WWZ24, ZLK^{+}19, ZCS^{+}20, CXS^{+}08, IB10]. Declarative [DMSRL18]. Decodable [ZZL^{+}24b]. Decoded [YYX^{+}24]. Decoder [FHF^{+}24, LPS15, LFP^{+}22, TY23, ZJZC20, CPL^{+}23, MPSR05]. Decoder-Complexity-Aware [LPS15]. Decoding [EC16, JLLL25, LGT^{+}24, LZLJ22, YWC24, ZLS^{+}25]. Decolorization [LWZ21b]. decomposability [LYCJ12]. Decomposition [HZC^{+}16, LYW^{+}22, SR22, WLL23, WFC^{+}25a, WDLW23, XJB25, YML^{+}22, ZLF^{+}23, BWA13]. Decoupled [CZL^{+}23a, FJF^{+}25, HHGW22, HPZ25]. Decoupling [DSL^{+}25, LZXP25, PHH^{+}24, ZSZ^{+}24b]. decrypting [NLN^{+}13]. Deep [AMC^{+}18, AC19, APC^{+}25, AS22b, BYM^{+}18, CAJ19, CLZ^{+}21, CZZ^{+}23, DBSL24, DHT^{+}19, GLW20, GYF^{+}21, GLT^{+}20, GCF^{+}21, GHR^{+}22, GMWW25, GSDT21, HVC^{+}20, HAAM19, HLM^{+}24, HHH22, HYLG23, IAJ^{+}24, JLA^{+}23, JWF18, JLW^{+}18, JCSL19, KPK^{+}24, LKL25a, LPY^{+}19, LZXY20, LLJ^{+}20, LMXJ21, LZZ23b, LGM^{+}24, LCZ^{+}24a, LHS^{+}21, LH22, LYXA22, LPD^{+}24, MHF24, MPM^{+}25, MDAE19, NSM^{+}24, NALM23, ODMD17a, ODMD17b, PHH^{+}24, PWX^{+}24, PS17, QRTM24, QWH^{+}21, QJ23, RHAG21, SMTR22, SBU^{+}17, SBS23, SHZ^{+}20, SHSR24, SZL^{+}24, SHE21, SLBS20, TLZ^{+}21, TDW^{+}23, TZ24, TZZ25, TP22, VVSV17, WYM18, WZTL19, WZ20, WTZ^{+}20, WFZ^{+}21a, WCZ^{+}21, Wan21, WCCN23, WDX^{+}23, WTL^{+}25, WSLM18, WLL^{+}19, WDLW23, XWY21, XZL^{+}25, XLS^{+}25, YV23, YNY25, YDJ24, YTRC19, YTL^{+}24a, ZYO20, ZYL^{+}17, ZYZE19a, ZYZE19b, ZZP^{+}20, ZRZ^{+}21, ZGD21, ZSLW23, ZTQX23, ZYQM24, ZZT24, ZZS^{+}25, ZCY^{+}19, ZLL20, ZLH16, ZLL^{+}22, ZZMZ21, ZLC^{+}23, ZZL^{+}23, ZQKH19, YXZQ22]. Deep-based [LMXJ21]. Deep-fake [APC^{+}25]. Deep-learning [TZ24]. Deep-Learning-Based [WLL^{+}19]. Deep-Neural [TP22]. Deeper [GLT^{+}20, LCY^{+}23]. Deepfake [BBL^{+}24, BHD24, CGW^{+}24, GLL^{+}22, GGB24, KFA^{+}24, LHL^{+}24, SED25, SLC25, TYL^{+}25, UMM24, WCCN23, YLD^{+}25, ZLX24, ZNNH25, ZZL^{+}25a, NR24, YJM^{+}24]. Deepfaked [IAJ^{+}24]. Deepfakes [AMC^{+}25, FDZ^{+}25]. Deeply [XZL^{+}22a]. DeepProduct [JLW^{+}18]. DeepSearch [WHF^{+}18]. Defending [LYH^{+}25]. Defense [FBGD23, SP21, WFZ25, WSX^{+}23]. Defined [ZLT^{+}25b]. Defining [GG06, HC22]. Defocus [JMS23]. Defogging [DWC^{+}21, GZZ^{+}24, LLL^{+}21]. Deformable [dMGRT24, WZL^{+}24]. Deformation [MLJ^{+}22, ZZLW25]. DEGAN [LCZ^{+}24b]. Degradation [YTR^{+}25]. Degraded [JXTC21]. Degree [VTD22, LLC^{+}25, LLA^{+}21, NTT20, QJ23, ZLOL18, ZZMZ21, ZMZ^{+}23, dCAGF^{+}25]. Dehazing [SG22a, SZB^{+}22]. Delay [AGC^{+}18, ACGH18, Cla18, CCG20, MFL^{+}16, MATW17, WCY^{+}18, GHP^{+}06, HNL08, HH11c]. Delay-Aware [WCY^{+}18]. Delay-Moving [Cla18]. Delay-Sensitive [AGC^{+}18, ACGH18]. Delays [XW17]. delivering [KS09, SLKS12]. Delivery [BWP^{+}24, HDZ^{+}15, SGYX22, ZWZL21, CCG^{+}08, DY09, GL12, Hua13, LH12, PS05, QS10]. Delving [GLT^{+}20]. Demand [CPP^{+}14, EL25, HSN^{+}14, CPP^{+}13, CE10, GSM^{+}08, LLKL11, QS10, SAAH10, TC08]. Demonstrative [DJWZ25]. Denoising [CLWW24, HZL^{+}16, HH25, LCL^{+}23b, XLW^{+}23b, XSL^{+}23, YLZ^{+}21, YBTX23]. Dense [BLW^{+}24, CXW^{+}19, LCL^{+}21, LZC^{+}19, ZYC^{+}24]. Densely [LWY22, SG22b, YXYB21]. DenseNet [LYD^{+}21, WZ20]. DenseNet-201-Based [WZ20]. DenseNet-BiLSTM [LYD^{+}21]. Density [YHH^{+}22]. Dependencies [SLBS20, ZPL^{+}23]. Dependency [PWHP24, COM^{+}11]. Dependent [LYZ^{+}18, WWWL23, JC08]. deployments [TC08]. Depression [LZH^{+}24a, ZLL^{+}24c]. Depth [BCNP24, CED^{+}16, DYN^{+}25, HLY21, LKW^{+}22, LL23, LYZZ24, LCZ^{+}24b, LLWX24, LGC^{+}24, PSN^{+}24, PB19, PSL^{+}24, QXG^{+}24, RS16, SK23, SGY^{+}24, YLZ^{+}21, ZLY^{+}24]. Depth-Based [LCZ^{+}24b, RS16, BCNP24]. Depth-of-Field [LL23]. Deraining [HL24]. Derivation [LYL^{+}21b, ZZL^{+}23]. Derivative [LSQ11, ZQKH19]. Derivative-based [LSQ11]. derive [RHS12]. Deriving [SSSK18]. Description [QSZ^{+}21]. Descriptions [RYZ^{+}23]. Descriptive [TWB^{+}23]. Descriptor [AR15, CFN^{+}23, HHL^{+}22, ZGM^{+}20]. Descriptors [GZL^{+}20]. Design [AAA^{+}21, CDZ^{+}17, FF23, GHW^{+}24, HH12b, HSX22, KMSW18, NHP^{+}16, REP^{+}19, ROST20, SKR09, SLMJ24, SS20, THA^{+}25, WBRZ17, YZS^{+}23, JC08, YC08]. Designed [HZZ^{+}24]. Designing [AAA^{+}21, BLMP18, PRH14, RABC24]. Designs [BFAS15, UFJ21]. Detail [LCZ^{+}24b, PYBY25, STSW23, XZY^{+}24a, YQZP24, HL24]. Detail-Enhanced [LCZ^{+}24b]. Detail-preserving [YQZP24]. Detailed [LKM^{+}19, RPE^{+}17]. Details [XZY^{+}24b]. Detect [MAZ22, WYM07]. Detecting [CYMW07, CSSZ19, JC10, JJML24, LHL^{+}24, ULIS07, WHL^{+}21, ZCY^{+}13, CWC10]. DeTection [GTP^{+}22, Abd18, BXMH15, BHD24, BOZ17, BLW^{+}24, CW25, CZW^{+}24, CZC15, CZZ^{+}23, CCH^{+}24, CLQQ25, CSSZ23, CSBJ24, CSŞ^{+}25, DPL^{+}23, FZW^{+}25a, FNH22, FQJ^{+}25, FXF^{+}23, GLL^{+}22, GGB24, GZH17, GHY^{+}24, GGA^{+}20, HYW^{+}25, HAAM19, HLM^{+}24, HWHL18, HZW22, HZZ^{+}24, HGGZ23, IAJ^{+}24, JLZ^{+}21, JMS23, KFA^{+}24, KBB21, KFP^{+}25, KEYY22, KLS^{+}18, KWL^{+}25, LQZH14, Li23, LDL24, LCL^{+}24a, LYC^{+}25, LLZ^{+}25, LQG25, LTL^{+}24, LG24, LZD^{+}21, LTX^{+}23, LH20, LZL20b, LZL21b, LMC^{+}22, LDZ^{+}24, LWH20, LWZ^{+}25, MKS20, NR24, NXL^{+}25, OL25, PWY^{+}25, PWHP24, QHFX21, QSY^{+}23, QLW^{+}24, RNR^{+}22, SED25, SBU^{+}17, SYY^{+}24, SZC24, SLC25, SGY^{+}23, SRAA17, SG22b, STSW23, SR25, TWM^{+}25, TLZ^{+}20, UMM24, WMH^{+}22, WXW^{+}22, WCCN23, WHL^{+}23, WWS^{+}24, WLSX24, WTL^{+}25, WWX^{+}21, WF25, XWW^{+}25, XZY^{+}24a, XPP^{+}23, XG24, XWW^{+}21, YWG^{+}20, YZS^{+}22, YHX^{+}23, YJM^{+}24, YLD^{+}25, YNLZ22, YTL^{+}24a, YZ23, ZZC^{+}23, ZYW^{+}25, ZWL^{+}17, ZWL^{+}21, ZCQ^{+}23, ZLZX23]. Detection [ZCHD23, ZWLZ24, ZLX24, ZNNH25, ZZL^{+}25a, ZZC^{+}25a, ZWZ^{+}23a, ZWR^{+}20, ZLL^{+}24c, ZYY^{+}20, ZJL^{+}21, ZZL25e, BAK13, KO11, LYZY21, LW07, MHT^{+}13, SEK12, VPSS^{+}13, XC06]. Detector [GGSV25, QDX^{+}24, YPSC22, ZWLZ24]. Detectors [BBL^{+}24, DYL^{+}25]. determination [MVW07]. Development [CISF^{+}25]. Developments [MHF24]. Device [HJMY15, JLL^{+}25, Zho16, SSTK07]. Device-Based [HJMY15]. Device-to-Device [Zho16]. Devices [CFP15, Chu15, HSN^{+}14, LYJ^{+}15, LHF^{+}14, MKSB17, NSJB17, PB14, RK15, SCX^{+}25, SLZ^{+}21, VLE25, WHF^{+}18, HH12a, HH11b, LLSC12, SCFL14]. df [RVTC25]. Diabetes [SDK^{+}21]. Diabetic [MAZ22, SHE21]. Diagnosing [WLZ12]. Diagnosis [CLZ^{+}21, JGJ^{+}20, KMK^{+}21, KN21b, LWZW21, MEA^{+}21, RHAG21, RPE^{+}17, SHE21, TRK^{+}20, WCF^{+}17, WWY^{+}21]. Dialocalization [FYH10]. Dialog [CSZ^{+}25, FZYW20]. Dialogue [WZZL25, ZJT^{+}23]. diarization [FYH10]. Dice [LLJW15]. Dictionaries [ZRZX23]. Dictionary [NDX^{+}21]. Did [ICZ^{+}22]. DIEN [ZCL^{+}24]. DiFace [SLZ25]. Difference [TWKK21, ZCXL23, YDCZ23]. Differentiable [WZWY23]. Differential [XZL^{+}25, ZQKH19]. Differentially [GYF^{+}21]. Differentiated [GQG^{+}24]. Diffusion [GGB24, HWW25, LKL^{+}25b, LYL^{+}25, SLZ25, TXW^{+}25, TJZ^{+}25, WLZ^{+}24a, WZZ^{+}24, WWY^{+}25, ZBYZ25, ZFY^{+}25, ZXS^{+}25, DL14]. Diffusion-model [GGB24]. Digging [SGY^{+}24]. Digital [AE22, BCP14, CHHH18, Dao17, LCL22, NY24, RD17, SVA^{+}21, TB17, WCL23, WQL18, CBS08, CFGW05, FLZ^{+}12, LLC11, SEK12, XMST07]. Dilated [HRL^{+}24]. Dilemma [KMK^{+}21]. dimensional [JP11, NWL^{+}20, NWNL21, SR22, ZCXL23, ZSX^{+}25]. Dimensions [AR15]. DIP [GGB14]. DIPS [BBRS23]. DiRaC [SS24]. DiRaC-I [SS24]. Direct [ZDZ^{+}22]. Directed [MOL^{+}22, WLTL23]. Directing [PCB^{+}21]. Direction [CLP^{+}23, ZZLL17, ZLC^{+}23]. Directional [JWH21, XZC^{+}25, LHY^{+}24]. Directions [MHF24, WYGC24, ZGL^{+}18, CBS08, RJ05]. Director [WPRC18, PCB^{+}21]. DIRformer [HWW25]. Dirichlet [QZXH14, YGNT19]. DISA [CZS^{+}25]. Discarding [BLS^{+}19]. disciplinary [ZCS^{+}20]. Discomfort [YHZ19]. Discovering [FSX14, LSDK12, YYY^{+}24, YGNT19]. Discovery [AMG23, CLS17, CSSZ19, FDBP23, GUH^{+}20, JXLD25, LCS17, LWZ^{+}21a, SCX^{+}25, ZRZ^{+}21, AH12, GLC05, JP11, ZO13]. Discrepancy [RLY^{+}21]. Discrete [BZDX^{+}18, WDLW23, ZWL15, ZLLH25]. Discrete-Time [BZDX^{+}18]. Discriminant [LWZW21, ZLC^{+}23]. Discrimination [WJQ^{+}22a, XWW^{+}25, XHL^{+}24a]. Discriminative [LLL^{+}22, MWL^{+}24, SSY20, SYY^{+}24, XG24, YI14, ZCD15, ZWZL22, ZNNH25]. Discriminatively [ZZY18]. discriminator [LLZ^{+}23]. Disease [LMS^{+}24, MEA^{+}21, SLL^{+}21, TRK^{+}20, ZSJL22]. Diseases [KN21b]. Disentangle [STSW23]. Disentangled [CZS^{+}25, CHS^{+}23, FDBP23, LXT^{+}24, YZS^{+}23]. Disentanglement [CZL23b, GQG^{+}24, JLK^{+}24, TNZ^{+}25]. Disentangling [DBBD23, LSL^{+}24]. Dish [WDJ^{+}21]. Dishonest [YLS^{+}23]. Disk [RWW05]. Disorder [DMF^{+}20, YRE^{+}20]. Disorder-Aware [YRE^{+}20]. Display [LL23]. DisplayCast [CBR14]. Displays [FHH22, DÇ07, LW07, SCFL14]. Disproportionate [AGCC25]. Disrupt [YNY25]. Disruptor [TYL^{+}25]. Dissecting [TCP^{+}20, LCC^{+}14b]. Dissemination [LB15]. Dissimilarity [MKC21]. Dissimilarity-Based [MKC21]. Distance [ATS19, CXX^{+}24, LH20, WLHT19, ZSYW23, ZLC^{+}23, hHLC10]. Distances [YP15]. Distill [JMS23]. Distill-DBDGAN [JMS23]. Distillation [GSY^{+}24, JMS23, LCL^{+}21, YZS^{+}22, ZYJZ25, ZSS^{+}23]. Distilled [LYL^{+}23b]. Distilling [PLZT22]. Distinction [WHJ20].

Distinctive [MXZ^{+}24]. Distinguish [GGB24]. Distinguishing [AMC^{+}25]. Distorted [ZGL^{+}23, YDCZ23]. Distortion [BA20, PJL23, TSP^{+}24, YTL^{+}21, HH08a, HH08b]. Distributed [BYOZ20, BOZ17, CNG22, DMF17, EG17, KD18, LCS17, MN16, XHYX23, ZCAP08, DL14, GG06, MMW10, OMP07, Yan10, ZO13, ZCT^{+}07, GGB14]. Distribution [BRZS18, CSZ^{+}25, LTD^{+}21, LZW^{+}23, NPG^{+}22, QLW^{+}24, SLZ24, Zho16, CCG^{+}08, GHP^{+}06, WZC^{+}13, YHX^{+}23, YH13]. Distribution-based [CSZ^{+}25]. Distributions [ZWL15]. Dive [WTL^{+}25]. Diverse [LLL^{+}24a, SS24, XLZ^{+}24, YMJ^{+}25]. Diversely [TS22]. Diversely-Supervised [TS22]. Diversification [DNPG^{+}17]. Diversity [CWW^{+}25, LGX^{+}24]. Diversity-Representativeness [CWW^{+}25]. Divide [YGMY24]. Divide-and-conquer-based [YGMY24]. DLRF [ZRZ^{+}21]. DLRF-Net [ZRZ^{+}21]. DNA [ARA^{+}23, NCMM21, NLKB24, WCL23, ZZC^{+}23]. DNA-Based [NCMM21]. DNA-centric [NLKB24]. DNN [TSP^{+}24, XWZ^{+}24b]. Do [ROST20]. Doctor [KMK^{+}21]. Document [SJC^{+}25]. Documents [CFP15, DMSRL18, BB11]. Does [COM^{+}11]. Domain [CSZ^{+}25, CTBC22, DWS^{+}24, DG17b, GQSG24, GS19, HYG^{+}25, JLK^{+}24, JXLD25, KBB21, LA15, LLZ^{+}22, LCL^{+}24a, LXL^{+}23, PWY^{+}25, SZHY19, SZM^{+}21, SZC24, SZLL17, WTD^{+}21, WXH^{+}25, WDC^{+}25, WCG^{+}25, WZRD25, WWQ^{+}25b, WF25, WLSZ22, WWWL23, XYJ^{+}20, XSD^{+}22, YZXY15, YHG^{+}24, YHZ23, ZX14, ZH18, ZZTL24, ZLY^{+}24, ZCYL24, ZNNH25, ZFC^{+}24, vRPPP23, BA20, CLQQ25, FYW^{+}24, LTD^{+}21, LSS^{+}24b, PS05, QDX^{+}24, ZZJ^{+}23, ZYQM24]. Domain-Aware [WDC^{+}25, CSZ^{+}25]. Domain-Incremental [YHG^{+}24].

Domain-invariant [LLZ^{+}22, ZNNH25]. Domain-Separated [HYG^{+}25]. Domain-Specific [ZH18]. Domical [GS11b]. Dominated [LGX^{+}24]. Done [BUS^{+}21]. Double [GCZ^{+}24, HH19, HLC^{+}25, Lin15, TFHM24, WHL^{+}21, ZXS^{+}24, ZXD^{+}23b]. Double-Layer [GCZ^{+}24]. DoubleAUG [QDX^{+}24]. Down [GWM^{+}14, PLYM23]. DPDFormer [LYZZ24]. DQ [BYOZ20]. DQ-DASH [BYOZ20]. DQFormer [HPZ25]. Dr [Ste10]. Dress [FCL^{+}22]. Drift [GKSB17]. Drift-Compensated [GKSB17]. DriveDiTFit [TJZ^{+}25]. Driven [BRZS18, DVA21, LMS^{+}24, PCB^{+}21, PFC^{+}16, SSSK18, SCZJ24, SLL^{+}24, WLL^{+}24, XFQ^{+}21, YXZQ22, ZWC^{+}22, HTT^{+}11, HDH^{+}24, HYG^{+}21, RBLJ25, SY09, TXW^{+}25, YMY^{+}21, ZGL^{+}23, ZXZ^{+}24a, ZO13]. Drivers [HDZ^{+}15]. Driving [TJZ^{+}25, ZYZ^{+}13]. DRL [XZL^{+}22b]. Drone [GLT^{+}20]. Drone-Based [GLT^{+}20]. Dropout [GQSG24, XZZL23, GQSG24]. dropouts [DCM13]. Dropping [FSK^{+}15]. DSI [Yan10]. DTI [CGNG15]. DTSD [XWW^{+}25]. DTV [LGF^{+}14]. Dual [CW25, CZS^{+}25, CKRB23, CH21, FHF^{+}24, HDH^{+}24, JW21, LCL^{+}23b, LDQ^{+}25, LWLC24, LCC^{+}14a, LLZ^{+}23, LLN^{+}24, LWW^{+}25, LYH24, QDX^{+}24, QLGL24, RYZ^{+}23, SZLL17, TCL^{+}23, WBH^{+}23, WCG^{+}25, WYMX23, WHW^{+}21, XWW^{+}25, XLZ^{+}21, XLZ^{+}24, YW23, ZWH^{+}23, ZZG^{+}20, ZHL^{+}25]. Dual-Attention [LDQ^{+}25, TCL^{+}23]. Dual-Branch [CW25, CZS^{+}25, CH21]. Dual-Camera [LCC^{+}14a]. Dual-connectivity [CKRB23]. Dual-Decoder [FHF^{+}24]. Dual-discriminator [LLZ^{+}23]. Dual-Domain [SZLL17, WCG^{+}25]. Dual-Lens [LCL^{+}23b]. Dual-Model [QLGL24]. Dual-path [HDH^{+}24, ZZG^{+}20]. Dual-Priors-Guided [LWW^{+}25].

Dual-Stream

[WHW^{+}21, XLZ^{+}21, ZHL^{+}25]. Dual-style [QDX^{+}24]. Dual-Threshold [LWLC24]. Dummy [BHD24]. duo [ONH07]. Duplicate [LQZH14, ZWYS20, DCO10, ZLLT13, ZHLY11]. Duration [OE19, ICZ^{+}22]. Duration-Aware [OE19]. Duration-informed [ICZ^{+}22]. During [KP15]. Dyadic [BBRS23, ONAGP19]. Dynamic [Ala21, dCAGF^{+}25, BTWZ22, CCH^{+}24, DL14, FLG^{+}21, HH25, HWHL18, HSS^{+}25, KLX^{+}24, LZH^{+}24b, LLS^{+}24, LCH^{+}25, LOJZ18, LCL^{+}24b, LTHT24, LDZ^{+}20, MVW08, SCZ^{+}25, SR25, TY23, UJLS22, WWQ^{+}25a, WFC^{+}25b, XLH^{+}24, YZRW22, YM24, Zha19, ZDD^{+}24, CXS^{+}08, IB10, ZWM12b]. dynamically [Bag11]. Dynamics [KP15, TVZ^{+}19, ZXS^{+}25].

e-book [ARE13]. e-Commerce [VVSV17]. E-detector [ZWLZ24]. e-Health [ACC^{+}21, SLK21]. E-Healthcare [SSP21]. E-learning [ASLA18]. Early [GZLW18, KN21b, LYZ^{+}18, ZLK^{+}19, ZSS^{+}23]. Early-Late [GZLW18]. Earthquake [YZYX24]. Ease [AdSGMS25]. ECCNAS [WMW^{+}22]. EDCA [PBS12]. EDEN [LFP^{+}22]. Edge [GHAM25, HLW^{+}21, LWH^{+}22, LZWC23, LDL24, LZWL25, MZL^{+}18, MSKYJ21, XHZ^{+}21, XFQ^{+}21, YXZQ22, ZZW^{+}24, GGB24]. Edge-Based [LZWL25]. Edge-guided [LZWC23]. eDiaPredict [SDK^{+}21]. Edit [WLZ^{+}24a]. Editing [ATMM24, dCAGF^{+}25, JJML24, WPRC18, YBTX23, CTGP08]. Editor [dB16]. Editor-In-Chief [dB16]. Editorial [SKVHC18a, Ste11, Ste12a, Ste12b, Ste12c, Ste13a, Ste13b, Ste14, WJXN24, YNC18, ZYZE19b, JPS05, Geo05]. Education [FMIS17]. Educational [CZZ21]. EEG [CIPE18, JGJ^{+}20, KYVE14, LGT^{+}24, XHL^{+}21]. Effect [AMMG16, LXRL23, NH10, ZLH16, ZSO13]. Effective [DWC^{+}21, KVB^{+}24, KEYY22, LYL23a, LJG^{+}24, RTR21, SZF^{+}25, TYL^{+}25, YWC^{+}25, YYS13, CCF^{+}24, LGX^{+}08, MC11, SCFL14, ZG08]. Effectiveness [ZLOL18]. Effects [LL23, TKK^{+}17, WBR25]. Efficiency [DG17b, KZGH15, LZWL25, MPM^{+}25, SAF19, TZ24]. Efficient [APP^{+}22, AMT25, BPM15, BHD24, CLS17, DBSL24, DG17b, FXMZ24, GYN12, GZH17, GGML22, HZS^{+}24, HJMY15, KVB^{+}24, KD18, LGT^{+}24, LDT^{+}18, LSXZ23, LZT^{+}20, LH20, LHC^{+}25, MAGT23, NNK25, NY24, PLZW18, QS10, RNR^{+}22, RW12, San11, SS20, SR25, TDW^{+}23, TYL^{+}25, WDCX07, WMW^{+}22, WZWY23, WGL^{+}23, WFC^{+}25a, WWL^{+}24, WB16, WGM25, XLZ^{+}22, YXZQ22, YGL^{+}23, YW23, Zha19, ZWY21, ZLLH25, ZYL^{+}24, ZBYZ25, ADCB07, CML^{+}13, HH12a, LLSC12, LXZ^{+}25, PBS12, WZRD25, YSG^{+}06, YH13, ZG08]. efficiently [HH11c]. EfficientNet [ZCL^{+}24]. Ego [LGC^{+}24, QXG^{+}24]. Ego-Motion [QXG^{+}24, LGC^{+}24]. Egocentric [HWHL18, HYG^{+}21, ZML^{+}24, ZSS^{+}23]. EGroupNet [DLO^{+}20]. eHealth [RTR21]. eigenfaces [Tur13]. Eigenvector [WLHT19]. Eigenvector-Based [WLHT19]. EiMOL [SSSA23]. Elaboration [BXMH15]. elderly [CYMW07]. Elective [ZLLH25]. Electric [LK18]. Electrical [RD17]. Electrocardiogram [ZCL^{+}24]. Electrodermal [YSY^{+}22]. Elements [GUH^{+}20]. Elevating [EL25]. elicited [LGT^{+}24]. ELVIS [MA10]. Embedded [HZW22, ZJT^{+}23]. Embedding [BA20, CXY^{+}25, CGGC20, CLP17, DNL^{+}23, FHP23, HMUC21, HCZ^{+}22, HQF^{+}19, HPLL23, JWF18, LZW^{+}19, NDX^{+}21, WZNM14, WYGC24, WTL24, WFZ25, XGX^{+}25, ZZY18, ZWR^{+}20, MSS^{+}24]. Embedding-and-retrieval [HMUC21]. Embeddings [LPY^{+}19, ZYLN21, ZZG^{+}20]. Embezzle [ZNYZ25]. Emerging [BXMH15, SKVHC18b]. EMES [MAGT23].

Emoji [ZJT^{+}23]. Emoji-embedded [ZJT^{+}23]. Emotion [ASLA18, BTWZ22, CIPE18, CZ24, HYG^{+}25, HYSL20, KP15, LZH^{+}20, LYX23, YMY^{+}21, YSY^{+}22, ZGD^{+}19, ZWF^{+}20, LYC11]. Emotion-Based [HYSL20]. Emotion-driven [YMY^{+}21]. Emotional [HMB^{+}25, KTK^{+}17, LWP22]. Emotions [SWS^{+}22, THR^{+}22]. Empirical [LGT^{+}24, SR22, Tas22, ZFC^{+}24, CY11]. Employing [GLT^{+}20]. Empowered [LHZ^{+}23, ZYL^{+}24, HCM^{+}22]. Enabled [BBZ18, KJJ^{+}21, MSG22, ZZC^{+}25b, BCNP24, DY09, LYXA22, SMTR22]. Enabling [LL15, WAD^{+}18, YWN^{+}10a, YWN^{+}10b]. Encoded [DG17b, SWZ25]. Encoder [CPL^{+}23, LFP^{+}22, LZZ^{+}24a, MSC^{+}23, ZJZC20, LLS^{+}21]. Encoder-Decoder [LFP^{+}22, CPL^{+}23]. Encoders [MAE^{+}21]. Encoding [AS22a, AGCC25, IVS^{+}20, KD18, LPS15, LZH^{+}24a, LHY^{+}24, SAF19, SSSK18, WGM25, XG24, YMS^{+}24, ZHD^{+}23, GHQ24, IB10, MAGT23]. Encrypted [CQA^{+}24, GZH17, GGA^{+}20, LA15, LHY^{+}24, SZHY19, WHZ^{+}24, WLW^{+}23, XHYX23, vRPPP23]. Encrypted-Image [XHYX23]. Encryption [ARA^{+}23, HQF^{+}25, NCMM21, PFC^{+}24, SS22, SSSA23, XJG^{+}22]. End [HNL08, HPZ25, LZL^{+}23, LSN^{+}20, LZG^{+}25, WJQ^{+}22b, WZZL25, XYL^{+}23, ZXX22, CVV06, ZRCH08]. End-to-End [HPZ25, LSN^{+}20, LZG^{+}25, WZZL25, HNL08, LZL^{+}23, WJQ^{+}22b, XYL^{+}23, ZXX22, ZRCH08]. Endoscopy [HLZ^{+}20, WCF^{+}17]. Energy [HH12a, HSN^{+}14, LXZ^{+}25, SLP15]. Energy-efficient [HH12a, LXZ^{+}25]. Enforcing [AQL^{+}20, DXY^{+}21]. Engagement [RSAH25]. engine [LLKL11, JWL06]. Engineering [PJL23]. Engines [MKSB17]. Enhance [ZHG^{+}25]. Enhanced [AMMG16, APP^{+}22, CSW^{+}25b, DLL^{+}20, FHP23, GHR^{+}22, LWY22, LML^{+}24, LZWL25, LZL21b, LZD^{+}22, LCZ^{+}24b, MLQM14, NDC^{+}23, PHH^{+}24, PSSZ24, PWHP24, QSY^{+}23, SWK^{+}22, TXWK24, WKS^{+}23, WB16, WZZL25, XZC^{+}25, YHX^{+}23, ZYL^{+}24, ZZL^{+}24d, AG13, DLO^{+}20, GA12b, LCS09]. Enhancement [GLZW23, HHGW22, HLY21, HLZ^{+}21, HZPL21, HSG23, HRL^{+}24, LA15, LZXP25, LWW^{+}25, PLB^{+}24, QJ23, SWZ25, SB23, WCZ^{+}21, WQL18, WLZ^{+}24b, XZY^{+}24a, XWW^{+}21, XJB25, YWC24, YYX^{+}24, ZSMZ21, ZCYL24, ZXT^{+}25, ZLF^{+}23, ZHL^{+}25, BSS11, HWY^{+}11]. Enhancing [BOZ17, BCNP24, DYN^{+}25, GA12a, LLW^{+}13, LCH^{+}25, MPM^{+}25, WYGC24, WZZ^{+}24, WWZ^{+}25, XZY^{+}25, YWH^{+}17, ZLL^{+}24d, CBJ^{+}09, YLZ^{+}10]. EnIter [DYN^{+}25]. enjoy [Dao17]. enrich [CBJ^{+}09]. Enriched [VCO15, LYX23]. enroute [LS05]. Ensemble [AMC^{+}18, SHE21, SDK^{+}21, ZZJ^{+}23, ZXS^{+}24, GLC05]. Ensemble-based [SDK^{+}21]. Entailment [HPW20, SSBT20]. Entailment-Based [SSBT20]. Entertainment [MA10]. Entertainment-Led [MA10]. Entity [AMG23]. Entropy [JYZ^{+}25, KN21b, XZC^{+}25, YTOH22]. Entropy-Guided [YTOH22]. Environment [AS22a, GTP^{+}22, NCMM21, PB14, XHZ^{+}21, YMA17, YYY^{+}24, AVJ05, CYMW07, MVW08]. Environments [CZZ21, EG17, HEA14, SHIE15, YZRW22, ARE13, DL14, MVW07, MCM^{+}09, MRS^{+}07, YWN^{+}10a, YWN^{+}10b, ZWM12b, ZCT^{+}07]. EOGT [PWHP24]. Epilepsy [HAAM19]. Epileptic [XHL^{+}21]. Epipolar [ZPL^{+}23]. Equipment [TC08]. Equivariant [LQD^{+}24, ZY21]. Era [NY24, ZZL^{+}25c]. Errata [CPP^{+}14]. Error [HYLG23, HSS^{+}25, YWC24, WCK05]. Error-Corrected [HYLG23]. errors [YSG^{+}06]. ESRNet [RNR^{+}22]. Essential [RHS^{+}20]. Establishing [GHS^{+}24]. Estimation [BSQ^{+}24, BBL^{+}24, BPFA24, BCNP24, CZW15, CLZ^{+}21, CLQQ25, CZW^{+}25, DSL^{+}22, DYN^{+}25, DLO^{+}20, DLL^{+}20, FH20, JCF^{+}22, KRKK14, LKL^{+}25b, LK18, LYZZ24, LZXP25, LGC^{+}24, NT08, PLZW18, PSL^{+}24, RSAH25, SG22a, SZZY20, WTZ^{+}20, WQL18, XFSZ20, YLCC18, YHH^{+}22, YWY^{+}23, ZZTL24, ZLY^{+}24, ZWR^{+}20, BDV08]. Estimators [CI23]. Eternal [LVM^{+}21]. Evaluate [WDX^{+}23]. Evaluating [BLMP18, GCL^{+}25, KFP^{+}25, KMK^{+}21]. Evaluation [AGCC25, CXL^{+}22, CDZ^{+}17, HLY^{+}23, IPMD25, LCZ^{+}24a, LYW^{+}23b, LYXA22, MNPOF22, NTT20, PHS^{+}20, RPE^{+}17, SMN^{+}22, THA^{+}25, UFJ21, WKN25, YTL^{+}21, YJM^{+}24, FPH^{+}08, HH12b, VM12, WT10, GSB^{+}22]. EVASR [LZWL25]. EVCS [WFC^{+}25b]. EVCS-DAS [WFC^{+}25b]. Event [AMC^{+}18, AC19, CHLW19, DLZ^{+}17, HTT^{+}11, LLJ^{+}20, MKS20, PCB^{+}21, QZXH14, TVK18, YRE^{+}20, YZX16, YLK^{+}20, ZX14, ZWLZ24, CFGW05, PS05, TCJ08, XC06]. Event-based [ZWLZ24]. Event-Driven [PCB^{+}21, HTT^{+}11]. Events [AE22, MC19, ZCT^{+}07]. Ever [WWW^{+}22]. Evolution [HNS13, MAX^{+}24]. Evolutionary [ZZL21]. Evolving [KRKK14, SZEST21, WRK14, WFC^{+}25b, LCSX11]. Examine [LMS^{+}24, MHCG19]. Examining [MHT^{+}08]. Example [DLZ^{+}17, PWY^{+}25]. Examples [LZW^{+}24, WLMZ25]. Exemplar [BTWZ22]. Exercising [FMIS17]. EXP [TCP^{+}20]. Expanded [WSLZ25]. Expanding [ZZL^{+}24b]. Expanding-Window [ZZL^{+}24b]. Expansion [HMOS17, WTD^{+}21]. Experience [BPT^{+}15, DLC^{+}22, FHH22, FPA24, HEA14, PVWD18, RT14, SkFM18, SVA^{+}21, YCGM14, ZQRS18, MHT^{+}08]. Experience-Centric [PVWD18]. Experiential [Sun13]. Experimentation [TCP^{+}20]. expert [ZWZ^{+}23b]. Explainable [AAA^{+}21, CJHH21, CLS^{+}21, KMK^{+}21, LLC^{+}21, LWZW21, SHE21, UMM24, XHL^{+}21, YSF^{+}21, ZGD21]. Explaining [ZYQM24]. Explanation [LMS^{+}24, XWY21]. Explanation-Driven [LMS^{+}24]. Explanations [YJM^{+}24]. Explicit [DNL^{+}23, YWL^{+}24, ZWH^{+}23]. Explicit-Implicit [YWL^{+}24]. Exploiting [And13, HYW^{+}25, LYC11, WWL13, XJW^{+}22, YGL^{+}23, YZW^{+}24, ZLZX23, ZZW^{+}19]. Exploration [BRG24, DFXY20, FSX14, PB14, USM^{+}25, WWHW14]. Explore [HQLW25]. Exploring [CL07, CL12, CSŞ^{+}25, GLW20, GCC^{+}25, HMUC21, JLL24, LXRL23, LML^{+}24, LZF^{+}22, WLZ08, XLW23a, XWW^{+}21, YRE^{+}20, ZZS^{+}25]. Exposing [YQH12]. Exposure [HSR18, LCL^{+}23b, LNC^{+}25]. Exposure-Induced [LNC^{+}25]. Expression [HC25, LYL23a, MDAE19, SCZ^{+}25, SHWC19, WHJ20, WWL20, WKS^{+}23, YLCC18, YTY^{+}24, ZFY^{+}25, GZW^{+}24a, LWP22]. Expressions [HMB^{+}25, RE24]. Extended [HT15, QLGL24, RVTC25, WDE25]. extensible [TCJ08]. Extension [SAF19]. Extensive [NN21]. external [XC06]. Extraction [FYY^{+}21, KN21b, LGM^{+}24, TDW^{+}23, TYL^{+}25, WLH^{+}25, YRE^{+}20, ZLL^{+}24a, ENHN09, MB08]. extrapolation [ZWM12b]. Extremely [HRL^{+}24]. Eye [HWLC19, MHCG19, SLZ^{+}21, CCD07, FPH^{+}08, GDGC07, JC08, KK08, MRS^{+}07, NT08, ULIS07, VPSS^{+}13, WT10]. Eye-based [SLZ^{+}21]. eye-gaze [MRS^{+}07]. eye-gaze-position [KK08]. eye-movement-dependent [JC08]. eye-tracking [GDGC07]. Eyebrow [HWW^{+}23]. eyes [WYM07].

Face [CFN^{+}23, CZX^{+}24, CGW^{+}24, CC17, DKJ^{+}21, FHK^{+}23, GHR^{+}22, HYW^{+}25, HMB^{+}25, HWW^{+}23, KP15, LZXY20, LMXJ21, LLJC21, LLH^{+}24, LLC^{+}21, LXT^{+}24, LZT^{+}20, LW23, LLC^{+}24b, MHW^{+}19, MKY^{+}24, PB19, PCH^{+}20, SLC25, SLZ25, SR25, TM25, TSZ^{+}23, WWC^{+}24, WTL^{+}25, XLT^{+}23, YML^{+}22, YPSC22, YHH24, ZNYZ25, ZH18, ZHS20, ZZC^{+}25a, KO11]. Face-Mask [FHK^{+}23]. Face-top [LMXJ21]. FaceDefend [ZNYZ25]. Faces [CTBC22, DMF^{+}20, YNLZ22]. FaceSigns [NHZ^{+}24]. Facial [BTWZ22, BPFA24, CFG^{+}24, FNH22, HPH^{+}20, HC25, LKL25a, LZXY20, LYL23a, LWP22, LW23, LLC^{+}24b, LWWZ20, MDAE19, QRTM24, RE24, SSK20, SHWC19, WWL20, YLD^{+}25, ZFY^{+}25]. Facial-expression-aware [LWP22]. Factor [WZ20]. Factorization [LLN^{+}24, WSLM18, ZWY21, ZJL^{+}21, WC12]. Factors [FDBP23]. Fair [AS20, PFC^{+}16]. Fake [CZC15, KWL^{+}25, QHFX21, YLS^{+}23, YNLZ22, APC^{+}25]. Fakes [WTL^{+}25]. False [PRPO23]. False-Free [PRPO23]. far [Cha13]. Fashion [CYW24, DBBD23, GHW^{+}24, WB16, YZS^{+}23, YSF^{+}21, ZSLW23]. Fast [Che24, CWC10, FYY^{+}21, GZT21, HZL^{+}21, HSG23, LQZH14, LLL^{+}21, LZT^{+}20, LLN^{+}24, SAZ^{+}15, WHF^{+}18, WGHX24, XW17, YF22, ZLK^{+}19, ZLC^{+}25, CLC05, IB10, ZO13, AGCC25, WWY^{+}25]. Fast-Paced [XW17]. FastCNN [HSG23]. Faster [DSL^{+}22, DYSX14, TLZ^{+}20]. FasterPose [DSL^{+}22]. FCM [XYJ^{+}20]. Feature [AQL^{+}20, Ber18, CZW^{+}24, CZL23b, CZ24, DHP23, DSL^{+}25, DNL^{+}23, DLO^{+}20, GHY^{+}24, GQG^{+}24, HZH24, HC25, HWW^{+}23, LBD08, LGM^{+}24, LXC^{+}24, LLZ^{+}25, LG24, LWZH19, LZL21b, LZG^{+}25, LZXP25, LWH20, MYX^{+}23b, MEA^{+}21, NWNL21, QSY^{+}23, SLZ24, SZW^{+}24, TYL^{+}25, TSH^{+}25, WCLC18, WHL^{+}21, WLTL23, WGL^{+}23, WLB^{+}25, WJQ^{+}22a, XWY21, XPP^{+}23, XG24, XHL^{+}21, XWH^{+}21, XZLZ23, XYL^{+}23, Yan17, YML^{+}22, YBW^{+}25, YZX16, YGL^{+}23, YMS^{+}24, YLD^{+}25, YXK^{+}25, YFLF19, YLYB24, ZYL^{+}17, ZNNH25, ZLP^{+}14, ZZW^{+}19, ZWZ^{+}23a, ZZZ^{+}22, ZLF^{+}24, ZZW^{+}22, CXL^{+}22, DCC^{+}13, GLC05, LGX^{+}08, ML11, MHT^{+}13, NDC^{+}23, San11, WXW^{+}22]. Feature-Aware [HC25]. Feature-enhanced [DLO^{+}20]. Feature-fusion [XZLZ23]. Feature-level [TSH^{+}25]. Features [AC19, BBUD24, BPFA24, CHS^{+}23, CJHH21, DBBD23, DHT^{+}19, DLL^{+}20, GQG^{+}24, GWWC25, HZH24, HMOS17, JLW^{+}18, KWL^{+}25, LTL^{+}24, LPC^{+}18, LYW^{+}23b, LLWX24, MHW^{+}19, ONAGP19, SLL^{+}21, SR25, WWZ24, WJ15b, WLW^{+}23, YI14, YNLZ22, YTL^{+}24a, ZGM^{+}20, ZCQ^{+}23, ZZJ^{+}23, ZLY^{+}24, QLSQ12, XC06]. Features-Enhanced [DLL^{+}20]. Federated [CCF^{+}24, LZL^{+}24b, SMTR22, ZYH^{+}24, ZYL^{+}24, ZGWZ23, ZZL^{+}24d]. Feedback [CIPE18, CJP^{+}21, DNPG^{+}17, FHG^{+}17, HCY^{+}23, KBI^{+}23, LHZ^{+}23, LS21, YBO14, YWY^{+}23]. Feedforward [YWY^{+}23]. FER [LYL23a]. Few [AM24, FJF^{+}25, HHL^{+}22, JMLL20, LLH^{+}24, SYY^{+}24, TY23, WLB^{+}25, WXJL25, YBW^{+}25, ZLL^{+}24d, ZFC^{+}24]. Few-Shot [AM24, FJF^{+}25, HHL^{+}22, LLH^{+}24, TY23, ZLL^{+}24d, ZFC^{+}24, JMLL20, SYY^{+}24, WLB^{+}25, WXJL25, YBW^{+}25]. FG [ADMM25]. Fibrillation [HZW22]. Fidelity [BCP14, LZYY24, MFL^{+}16, WC23, XLT^{+}23, ZXS^{+}25, LLC11]. Field [APP^{+}22, CRL20, GZT21, HSL^{+}20, LDL24, LZL^{+}24a, LG24, LL23, WHZ^{+}24, WYY^{+}25, ZWL^{+}17, ZZLW25, ZLH16]. Fields [CXX^{+}24, HH25, PYBY25]. Fifteen [Whi13]. Figure [SSBT20]. File [PFC^{+}24, JKKL08]. Files [WH22]. Filling [STSW23]. Film [ADMM25, GAB^{+}17, USM^{+}25, WPRC18].

Films [KTK^{+}17, VNC^{+}11]. Filter [HLY21, JLL^{+}21, LZZ23b, WHZ^{+}24]. Filtering [GS18, GS19, LLL^{+}25, LLC^{+}25, LLL^{+}21, LLCH17, LJZ^{+}22, RSE16]. Filtering-Based [LLC^{+}25]. Filters [LYZY21, YCLH22, ZZL^{+}24c]. Filtration [YLL^{+}23]. FIN [LWH20]. Financial [DŞB^{+}22]. Find [WLC^{+}20, Dao17]. Finding [TAS16, YZL^{+}14, KH13]. Findings [LVM^{+}21]. Fine [CYZ23, CF22, FZYY18, GJY^{+}25, HSR^{+}23, HLZ^{+}21, HJZ^{+}23, JJML24, LPW^{+}22, LKL^{+}25b, LYZZ24, MAE^{+}21, MPTD22, RBLJ25, SCZ^{+}25, SHWC19, TYY^{+}18, TYY^{+}22, TNZ^{+}25, TJZ^{+}25, WFZ^{+}21b, XQG^{+}23, XLH^{+}23, ZSYW23, ZLZ^{+}22a, ZCX^{+}24, ZLL^{+}24c, HH08a, HH08b]. Fine-Grained [GJY^{+}25, HSR^{+}23, MAE^{+}21, MPTD22, SCZ^{+}25, TYY^{+}18, XQG^{+}23, XLH^{+}23, ZLZ^{+}22a, ZLL^{+}24c, CF22, HLZ^{+}21, RBLJ25, TYY^{+}22, TNZ^{+}25, WFZ^{+}21b, ZSYW23, ZCX^{+}24, HH08a, HH08b]. Fine-tuning [FZYY18, TJZ^{+}25]. Finger [CJP^{+}21]. Fingerprint [XWZ^{+}24b]. Fingerprinting [BSSNF^{+}20, FLZ^{+}12]. Fingerprints [LCL^{+}24a, YP15]. Fire [GGSV25, KEYY22]. first [AH12, She13]. Fisheye [YLLZ25]. FishFormer [YLLZ25]. Fitting [LZT^{+}20, WZZ^{+}24]. Fixation [MZGY17]. fixations [ULIS07]. Fixed [MKSB17]. flash [JKKL08]. Flexible [CJZ^{+}24, PS05, YWC^{+}25, ZWY21]. Flexibly [WWY^{+}25]. flickr [JGZ^{+}11]. Flow [CW25, GZW^{+}24b, WLL^{+}24, XFQ^{+}21, ZZTL24]. Focal [TLZ^{+}20]. focus [RW12]. Focused [ZZL25e]. Fog [LWZ^{+}21a]. Fog-based [LWZ^{+}21a]. Following [HKYW14]. Food [JMLL20, LNH23, MHL^{+}25, PS17, SYX23, SLMJ24, WDJ^{+}21, YMS^{+}24, YMJ^{+}25]. Fool [LZX^{+}21]. Foot [LHF^{+}14]. Force [WJS^{+}21]. Foreground [NSK^{+}21, ZHL19, ZJL^{+}21].

Foreground-Aware [ZHL19]. Foreground-Background [NSK^{+}21]. Forensic [RSE16]. Forensics [Ber18, FDZ^{+}25, GS18, GS19, VNDB24, WQL18]. Forest [GLT^{+}20, KEYY22, WZD^{+}20]. Foreword [Hae10]. forgeries [YQH12]. Forgery [CLQQ25, LLJC21, LYC^{+}25, SJC^{+}25, SR25, WWX^{+}21, XZY^{+}24a, ZTQX23, ZZC^{+}25a]. Form [HLTH24, MZZ^{+}20]. Formal [GDM^{+}23]. Format [XJG^{+}22]. Format-compatible [XJG^{+}22]. Formation [DBPM25]. Former [BV25]. forward [Row13, WCK05]. Fountain [ZZL^{+}24b]. Four [CI23]. Fourier [SL23]. FoV [LHZ^{+}23]. Foveated [DÇ07, IVS^{+}20, YWY^{+}23, DE12, NH10]. FPGA [DWC^{+}21, MGG17]. FPGA-Based [MGG17]. Fractal [JTZ^{+}16]. Fractional [PLZW18]. Fractional-Pixel [PLZW18]. fragile [NHZ^{+}24]. Fragment [CBJ^{+}09]. Frame [BWP^{+}24, DOD23, HYLD20, HZH24, JLL24, LJ25, PYZ^{+}20, SGS21, TDW^{+}23, TSH^{+}25, BLS^{+}19, ZHD^{+}23, ZLC^{+}25, ZZW^{+}22, YQH12]. Frame- [TSH^{+}25]. Frame-level [PYZ^{+}20]. Frames [DG17b, LLCH17, PB19]. Framework [ADMM25, AE22, CZS^{+}25, CPP^{+}14, CLS17, DLL^{+}24, DG17b, FMG20, GLWK19, GCF^{+}21, HYG^{+}25, HMB^{+}25, HCM^{+}25, HQF^{+}25, HP17, JXTC21, JK25, JMS23, KT21, KEYY22, KJJ^{+}21, LCS17, LLC^{+}25, LLL23, LZL^{+}24c, LWZW21, MGP19, QRTM24, QWH^{+}21, RHAG21, SP21, SSZ^{+}23, SWZ25, SDK^{+}21, SS20, SZLL17, TZZ25, USM^{+}25, VMP20, WHF^{+}18, WAD^{+}18, WZL^{+}24, XLX^{+}24, XWZ^{+}24b, YJM^{+}24, YSY^{+}22, ZZY^{+}14, ZZT24, ZSZ^{+}22, ZLL^{+}24b, CPP^{+}13, CXS^{+}08, HH11a, HCS12, JSL07]. Free [CKRB23, GZW^{+}24b, JLL^{+}25, LMX24, LL23, LWZ^{+}25, PRPO23, TLZL24, YGMY24]. Free-space-optics [CKRB23]. Frequency [FYW^{+}24, GS19, LK18, LTX^{+}23, YLLL23, ZLF^{+}23, LXRL23]. Frequency-aware [LTX^{+}23]. Frequency-based [ZLF^{+}23]. Frequency-domain [FYW^{+}24]. Fresh [XHL^{+}24b]. friend [Ste10]. Friendly [KCC17, SZ12, WCK05]. Full [CZY^{+}21, CJZ^{+}24, HLY21, TCL^{+}23]. Full-body [CJZ^{+}24]. Full-Reference [TCL^{+}23, CZY^{+}21]. Fully [FZW^{+}25a, PZJL22, ZZZ^{+}22]. Function [HLZ^{+}20, MAZ22, SLMJ24]. Functions [HZSC20, JK25]. fungible [GCTG24]. Fused [LSXZ23]. Fusing [CZY^{+}21, CXX^{+}24, SVF12, YHH^{+}22, YSY^{+}22]. Fusion [CZW^{+}24, DNL^{+}23, DWC^{+}21, DLL^{+}20, GZW^{+}24a, GLT^{+}20, GZLW18, GCZ^{+}24, HZC22, HYG^{+}25, HCJ^{+}24, KWL^{+}25, LLZ^{+}21, Li23, LLZ^{+}23, LYW^{+}24, LLC^{+}24b, NLW^{+}21, NDC^{+}23, OL25, RLAM25, SZF^{+}25, SYM^{+}25, TY23, TXWK24, WRK14, WCLC18, Wan21, WXW^{+}22, WWZ24, WLSX24, WWQ^{+}25a, WZZ25, XZCG23, XC06, XWH^{+}21, YSSF24, YF22, YZYX24, ZRZ^{+}21, ZTQX23, ZWL^{+}24, ZLJ^{+}25, ZZL^{+}25a, ZZLW25, ZWZ^{+}23a, ZLL^{+}24c, ZZW^{+}22, CDA12, CW10, GHY^{+}24, Hae10, MGY22, NRUT20, RHE10, WK10, XZLZ23, WRK14]. Fusion-Based [DWC^{+}21]. Fusionand [YML^{+}22]. Future [CKRB23, GZ20, MHF24, TAT^{+}25, USM^{+}25, ZGL^{+}18, BCG13, Hua13, RJ05]. Fuzzy [DKJ^{+}21, HZW22, JGJ^{+}20, LYD^{+}21, Bag11].

G [CXW^{+}19]. GAHE [CXY^{+}25]. Gait [ZDZ^{+}23]. Gaits [YKW^{+}22]. Game [Cla18, CCG20, FMIS17, JSEI16, LLKL11, LLJW15, MKSB17, SSSK18, AH12, JC08]. Game-on-demand [LLKL11]. Gamepad [BFAS15]. Gamer [LSCH25]. Games [BFAS15, FF23, RSAH25, XW17, ZZL21, HMVI13]. Gaming [IVS^{+}20, LSCH25, LDT^{+}18, REP^{+}19, SSSK18]. GAN [AAPLV25, DHP23, GDM^{+}23, GGB24, HMB^{+}25, HYT^{+}25, LCL^{+}24a, NR24, WTL^{+}25, WWY^{+}21, WWWL23, WZW24, YSC21, ZSZ^{+}22, ZTX^{+}23]. GAN-Assisted [HYT^{+}25]. GAN-Based [HMB^{+}25, NR24]. GAN-Generated [WTL^{+}25]. GANK [SR25]. GANonymization [HMB^{+}25]. GANs [LXRL23, LZL^{+}24c, PQ19, VF25, WWL20, WWWL23, XZLZ23, ZLZX23]. Gap [ZZTL24]. GAPG [GDM^{+}23]. Garment [PKL25]. Garment-Person [PKL25]. Gated [HWWL20, LDL24, ZZW^{+}24]. Gateway [CNG22]. Gating [YTL^{+}24b]. Gaussian [YGNT19, ZLW^{+}21]. Gaze [HWLC19, SO22, TKK^{+}17, ADCB07, DÇ07, KK08, LW07, MRS^{+}07]. gaze-contingent [DÇ07, LW07]. gaze-tracking [ADCB07]. GCN [ZZJ^{+}23]. GD [PYBY25]. GD-NeRF [PYBY25]. Ge [WGM25]. Gender [CS17]. General [LQH18, ZLX24, ZNNH25]. Generalizability [PWY^{+}25]. Generalizable [LCZ^{+}24c, PYBY25, WZRD25, XGX^{+}25]. Generalization [GQSG24, JLK^{+}24, LCL^{+}24a, LXL^{+}23, SLMJ24, WXH^{+}25]. Generalization-Oriented [SLMJ24]. Generalized [QDX^{+}24, SLC25, YNLZ22, NC13]. Generated [WTL^{+}25, ZZZ14, ZCY^{+}13]. Generating [FDZ^{+}25, GCL^{+}25, LLJW15, LZW^{+}24, SZLL17, XLN^{+}21, YLHL22, ZXS^{+}25, SGW08]. Generation [AAPLV25, ATMM24, BC15, CISF^{+}25, CZX^{+}24, DHP23, DBPM25, DŞB^{+}22, DLD^{+}22, DYZ^{+}23, GDM^{+}23, HVC^{+}20, HYSL20, HLC^{+}25, IPMD25, LSL^{+}24, LWL24, LJ25, LNH23, LW23, LLC^{+}24b, MKY^{+}24, SWS^{+}22, SHT^{+}24, SLL^{+}24, TJZ^{+}25, VF25, VCO15, WLB^{+}25, WWWL23, XJFS25, XXG^{+}21, XCC^{+}23, XLW23a, XWL^{+}25a, YMX^{+}16, YMY^{+}21, YHP25, YSC21, YCFX23, YMJ^{+}25, YZD^{+}25, ZPY^{+}23, ZJT^{+}23, ZSZ^{+}25, ZZL^{+}25d, ZWC^{+}22, CDA12, HMVI13, NLKB24, SSTK07, XXD^{+}08]. Generative [CSŞ^{+}25, DHP23, EAK^{+}25, GYY^{+}25, LZL^{+}21a, LLJC21, LWH^{+}23, LNH23, LCZ^{+}24b, LYH24, NWL^{+}20, PYBY25, PQ19, SSK20, SL22, SLZ24, TZLZ21, WGM25, XTL^{+}21, YZS^{+}23, YLLL23, ZZL^{+}25d]. Generator [FZYW20]. Generic [BZDX^{+}18, RT14, SZTL16, JCC^{+}10]. Genes [RHS^{+}20]. genetic [ASVE13]. Genre [BRG24, WJ15b]. Geo [FSX14, MC19]. Geo-Informative [FSX14]. Geo-localization [MC19]. Geographic [YSZ15, AP13]. Geometric [CLWW24, LZDL25, LH20, PSL^{+}24, SR25, WLCG21, YWNW15, GZHD12, RHE10, WYM07, ZLLT13]. Geometrical [WDX^{+}23]. Geometries [SSK20]. Geometry [CXY^{+}25, QLGL24, WXS^{+}25, ZLS^{+}25, ZPL^{+}23, LLKL11]. Geometry-Aware [CXY^{+}25]. Georeferenced [TAS16]. Georganas [Ste10, Ste12c]. Geospatial [YSZZ14]. geotagged [PJ13]. Gesture [LHF^{+}14, ZHY^{+}25, ZLP^{+}14]. Gestures [KTK^{+}17, XLH^{+}24]. GHOSM [AA23]. Ghost [LL23]. GJFusion [HCJ^{+}24]. Glands [LWZY22]. Global [GS18, GS19, HMOS17, HZPL21, JCF^{+}22, LGM^{+}24, LLH^{+}24, LTL^{+}24, MYGX21, NCL^{+}23, PWHP24, RXC14, WZZ^{+}24, WQTX25, XZY^{+}24b, ZFSX21, ZYZ^{+}25b]. Global-Local [HZPL21, JCF^{+}22, LLH^{+}24]. Gloss [TXW^{+}25]. Gloss-driven [TXW^{+}25]. GLPose [JCF^{+}22]. GNN [HSL^{+}20]. GNN-Based [HSL^{+}20]. Goal [AKO07, MOL^{+}22]. Goal-Directed [MOL^{+}22]. Goal-oriented [AKO07]. Going [ZZB^{+}21]. Good [GEL^{+}15]. goods [CSSZ19]. GP [LZDL25]. GP-HSI [LZDL25]. GPSView [ZYZ^{+}13]. GPU [DBSL24]. GPU-based [DBSL24]. Gradient [CHHH25, GZZ^{+}24, HH25, LLS^{+}24, SWM^{+}24]. Grading [SHE21, ZGR21]. Grain [ADMM25]. Grained [GJY^{+}25, HCZ^{+}22, HSR^{+}23, LGW^{+}25, MAE^{+}21, MPTD22, QLGL24, SCZ^{+}25, TYY^{+}18, XQG^{+}23, XLH^{+}23, YTL^{+}24b, ZLZ^{+}22a, ZYJZ25, ZLL^{+}24c, CF22, HH08a, HH08b, HLZ^{+}21, RBLJ25, TYY^{+}22, TNZ^{+}25, WFZ^{+}21b, ZSYW23, ZCX^{+}24]. Granular [DPL^{+}23, LFP^{+}22]. Granularity [OL25, SZC24, WLZF22]. Graph [Ala21, AA23, CWTG22, CXY^{+}25, CZW^{+}25, CZQ^{+}22, FBG22, FLG^{+}21, dMGRT24, HCWM14, JLZ^{+}21, KPK^{+}24, LZJ^{+}20, LZL^{+}21a, LLZ^{+}22, LLS^{+}22, LWL24, LYC^{+}25, LZZS23, LYDY24, LSS^{+}24b, MYGX21, MYX23a, MBXZ24, MXH^{+}23, QHFX21, RBLJ25, SGYX22, SYX23, SLNL20, SWK^{+}22, SHT^{+}24, TGC23, WKS^{+}23, WSW^{+}24, WSLC24, WXS^{+}25, WYMX23, WLSZ22, XHL^{+}21, XLZ^{+}21, XLX^{+}24, XWL^{+}25a, YLZ^{+}21, YCZ^{+}23, ZPY^{+}23, ZZL^{+}24a, ZZW^{+}24, ZXD^{+}23b, ZGY^{+}24b, ZZL25e, CML^{+}13]. Graph-Based [CZW^{+}25, AA23, JLZ^{+}21, LZL^{+}21a, WXS^{+}25]. Graph-RNN [dMGRT24]. Graphics [MBK^{+}23, PB14, PJL23, ZSZ^{+}24a, CVV06, SNC12]. Graphics-Constrained [PB14]. Graphs [LML^{+}24, HOSS13]. GraSP [BR22]. Grassmannian [BR22]. Gray [YDCZ23]. Grayscale [DXY^{+}21]. Green [CXW^{+}19, GLGG25]. grey [QWH^{+}21]. GreyReID [QWH^{+}21]. GRHFMR [WYY^{+}25]. Grid [SZFR23]. Grid-based [SZFR23]. GridCast [CSJ^{+}08]. Grocery [AWG^{+}15]. Grounding [HSR^{+}23, LYW^{+}23a, LLG^{+}23, LQD^{+}24, WSLC24, WGHX24, YTY^{+}24]. Group [BBRS23, DLO^{+}20, MON21, ONAGP19, LLL23, MGP19]. Grouping [WY23]. Groups [YYY^{+}24]. Growing [HWHL18]. GSyncCode [ZLS^{+}25]. Guaranteed [LB15]. GuessUNeed [ZYLN21]. Guest [JPS05, SKVHC18a, YNC18]. Guidance [HGGZ23, LZD^{+}21, SLZ24, WHW18]. Guide [NWNL21]. Guided [CZL^{+}23a, CWL24, CHHH25, DNL^{+}23, GGSV25, HJ23, HPLL23, HLC^{+}25, LZXY20, LLS^{+}22, LCL^{+}23b, LXC^{+}24, LDQ^{+}25, LJZ^{+}22, LLYL22, LYW^{+}25, LWW^{+}25, MLJ^{+}22, NXL^{+}25, PKL25, TM25, TY23, WLSX24, WHW^{+}21, WWL^{+}23, WZW24, YTOH22, YWW^{+}24, YYY^{+}24, ZZZ14, ZYC^{+}24, FCL^{+}22, HVC^{+}20, JLL^{+}21, LZWC23, LQG25, LLL23, LLL^{+}24b, LSX^{+}25, SZF^{+}25, WWS^{+}24, WSLC24, XG24, YZD^{+}25, ZSZ^{+}25, ZLF^{+}24, ZWZ^{+}23b]. Guided-Learning [WHW^{+}21]. Guides [SBU^{+}17]. Guiding [CLP^{+}23, WHY18].

H.264 [IB10, LLHS12]. H.265 [GKSB17]. H.265/HEVC [GKSB17]. Habit [YHZ19]. Hallucination [LZXY20, QCW^{+}25]. Hamming [WZNM14]. Hand [GWWC25, HWHL18, LHF^{+}14, XLH^{+}24, YXK^{+}25, ZHY^{+}25]. Handheld [LHF^{+}14]. Handling [ATM06, MDMK06]. handoff [LWL08]. Hands [FPA24]. handwritten [ASVE13]. Happen [ICZ^{+}22]. Haptic [CIPE18, CJP^{+}21, HEA14, Tas20, Tas22, Tas24, ARE13, ASVE13, KTM^{+}06, SOC^{+}13]. Haptic-Audiovisual [Tas20, Tas22, Tas24]. Haptic-based [CJP^{+}21]. haptic-to-audio [SOC^{+}13]. haptic-to-video [SOC^{+}13]. Harder [TH22]. Hardware [NY24, SK23]. Hardware-Assisted [NY24]. Harmonic [YWNW15]. Harmonious [TH22]. Harnessing [WLSN24]. HARR [MWL^{+}24]. Harvesting [WLL^{+}19]. Hash [MWL^{+}24]. Hashing [CZZ^{+}23, DWL^{+}23, HLW^{+}24, KLX^{+}24, KVB^{+}24, LTL^{+}24, LH20, LLN^{+}24, NSM^{+}24, SHZ^{+}20, WCL^{+}22, WDC^{+}25, WDLW23, XZL^{+}25, XLS^{+}25, YP20, YTL^{+}24a, ZWY21, ZWZL22, ZLLH25, ZLX^{+}25, ZZL^{+}24d, CWC10, LYJ^{+}13, LMLC14]. hatching [WYK12]. Haze [FHF^{+}24, SG22a]. Hazy [CH15]. HCCL [HWZL24]. HCI [DWS^{+}24, LZH^{+}24a, LMS^{+}24]. HCMS [WWL^{+}24]. HCNCT [RLXW24]. HDR [CZL^{+}23a, LCL^{+}23b, LNC^{+}25]. Head [AAPLV25, BBL^{+}24, BCNP24, FHH22, LCH^{+}25, NDX^{+}21, SLL^{+}24, XLH^{+}23, ZH18, ZZMZ21]. Head-Mounted [FHH22]. Head-Shoulder [XLH^{+}23]. Health [ACC^{+}21, MYGX21, SLK21, WDJ^{+}21]. Health-aware [WDJ^{+}21]. Healthcare [AAA^{+}21, FYZ^{+}21, HCM^{+}22, KT21, LYW^{+}22, LYXA22, SSP21, SAL^{+}21a, SKA24, VLE25]. Healthy [AWG^{+}15]. hearing [HWY^{+}11]. hearing-impaired [HWY^{+}11]. Heart [MHCG19]. Heart-Rate [MHCG19]. heat [DL14]. Helpers [RXC14]. Heritage [BBB^{+}23]. Heterogeneous [HQF^{+}19, HPW20, KEYY22, LPS15, LYDY24, MYGX21, SYX23, WWZ24, WJQ^{+}22b, ZCY^{+}19, ZWS^{+}20, ZJSJ20, GSM^{+}08, LLSC12, LC12, QS10]. Heuristic [PFC^{+}16]. HEVC [CQA^{+}24, DG17b, GKSB17, HSG23, LYZ^{+}18, LZLJ22, LMC^{+}22, MAGT23, PLZW18, PYZ^{+}20, RBBS25, SAZ^{+}15, YXJ^{+}24, ZLK^{+}19, ZZLL17]. HEVC-Based [LZLJ22, MAGT23]. HEVC-compressed [RBBS25, LMC^{+}22]. HGAN [NWL^{+}20]. Hidden [JGY^{+}24, XYJ^{+}20, ZLS^{+}25]. Hiding [LHY^{+}24, PJL23, SL22, Sin21, WH22, XLZ^{+}23, XHYX23, YCZ^{+}24, vRPPP23, LT14]. Hierarchical [BHD24, BV25, CSBJ24, GSY^{+}24, HLW^{+}24, HWZL24, HB08, LZH^{+}20, LLJ^{+}20, LPW^{+}22, LYZ^{+}24, LXC^{+}24, LZW^{+}23, PSSZ24, QLW^{+}23, SMK^{+}24, SGYX22, TYY^{+}22, WBH^{+}23, WWL^{+}24, XLX^{+}24, YCL25, ZRZ^{+}21, ZZTL24, ZCL^{+}12]. Hierarchy [TJN14, ZZY^{+}14, LGX^{+}08]. HIFGAN [YLLL23]. High [CBR14, DG17b, FHK^{+}23, LXRL23, LCC^{+}14a, LX21, LZYY24, MWL^{+}24, MSS^{+}24, MFL^{+}16, SGS21, SAF19, TZ24, WC23, XLT^{+}23, YLLL23, YMJ^{+}25, ZGD^{+}19, ZXS^{+}24, ZZL^{+}25d, ZWF^{+}20, ZXS^{+}25, CVV06, GZHD12]. high-capacity [GZHD12]. High-Efficiency [DG17b, SAF19]. High-embedding [MSS^{+}24]. high-end [CVV06]. High-Fidelity [MFL^{+}16, WC23, XLT^{+}23, ZXS^{+}25]. High-Frequency [YLLL23, LXRL23]. High-level [LX21]. High-Order [ZGD^{+}19, ZXS^{+}24, ZWF^{+}20]. High-Quality [MWL^{+}24, YMJ^{+}25, FHK^{+}23, SGS21]. High-Resolution [LCC^{+}14a, ZZL^{+}25d]. Higher [FDZ^{+}25, KABB20]. Higher-Performance [KABB20]. Higher-Quality [FDZ^{+}25]. Highlight [Abd18]. HIL [HPW20]. Hints [DYN^{+}25]. Hippocampus [HCY^{+}23]. Histogram [SB23, XLZ^{+}23]. Histograms [GS18]. Histopathological [TS20]. Historical [ZYZ^{+}25a, HNS13, SAL^{+}22]. History [LLYH14, WSLZ25, WH22, LT14]. Hitchcock [SGW08]. HKA [XLX^{+}24]. HMM [GZLW18]. Hoc [LB15]. Holistic [NWL^{+}20, SAL^{+}22, BSS11]. Home [GCL^{+}25, CYMW07, GS11b]. Homogeneity [CGW^{+}24]. Homogenization [YZZZ25]. Homomorphic [vRPPP23]. Horizon [WSLZ25]. Hourglass [ZHS20]. house [MVW08]. Household [BV25]. HSI [LZDL25]. HTTP [BBZ18, CDZ^{+}17, HXZ^{+}20, LTHT24, SYS17, TAT^{+}25, TNH^{+}21, WLW^{+}23, BLS^{+}19]. HTTP/2 [WLW^{+}23, BLS^{+}19]. HTTP/2-based [BLS^{+}19]. Hull [THA^{+}25]. Human [CWTG22, CZS^{+}25, CDL^{+}20, CJZ^{+}24, CZW^{+}25, CF22, DSL^{+}22, DVA21, DJWZ25, DMF^{+}20, FH20, GEL^{+}15, GMWW25, HYSL20, HSZ^{+}18, JCF^{+}22, KPK^{+}24, KRKK14, LZL20a, LZDL25, LKL^{+}25b, LLCH17, LSL^{+}20, LPD^{+}24, MHF24, MLJ^{+}22, NRUT20, QXG^{+}24, RD17, SZZY20, SOC^{+}13, WWY^{+}24, WBR25, XFSZ20, YHQH17, YCFX23, ZCD15, ZHL19, ZZX^{+}20, ZZB^{+}21, ZZT24, ZLP^{+}14, ZTX^{+}23, ZLL20, ZLT^{+}25b, ZWJ^{+}24, ZXS^{+}25, And13, DCO10, JP11, LZL^{+}21a, TCW^{+}13, WYM07, CBS08]. Human-Agent [DJWZ25]. Human-centered [CBS08]. Human-Centric [ZLT^{+}25b, DCO10]. Human-Computer [KPK^{+}24, And13, WYM07]. Human-like [YCFX23]. Human-Scene [LZDL25]. Hybrid [AAS^{+}20, AA23, BQBLN18, CZ24, HLD18, LLL^{+}21, LZZ^{+}24b, LZY^{+}24, LJG^{+}24, LGF^{+}14, MEA^{+}21, NY24, DMSRL18, VPSS^{+}13, XWH^{+}21, YXK^{+}25, ZGZ^{+}22, TS20]. Hybrid-Cast [HLD18]. Hyper [CXY^{+}25, LZZS23, SGW08]. Hyper-Hitchcock [SGW08]. Hyper-node [LZZS23]. Hyper-Relational [CXY^{+}25]. Hypercube [WTL24]. Hypergraph [GZW^{+}25, LZZ^{+}23a, LD25, TBC^{+}11]. Hyperspectral [HZSC20, HZC22, LH22, WBH^{+}23, ZYZ^{+}25a]. hypervideo [SGW08]. Hypomimia [SLL^{+}21]. hypotheses [JLL24]. Hypothesis [CI23].

I/O [RWW05]. I2P [SZW^{+}24]. ICA [ZWL^{+}21]. ICN [HJWW19]. Id [GLT^{+}20, LSL^{+}24, LSX^{+}25, XLZ^{+}22, ZZJ^{+}23]. ID-guided [LSX^{+}25]. iDAM [LZZ23b]. Identification [CJHH21, CZL23b, CS17, DSL^{+}25, DLL^{+}18, DKJ^{+}21, FZZ^{+}25, GGB14, HCZ^{+}25, JGJ^{+}20, LZC^{+}19, LWH^{+}23, LZL^{+}24b, LXC^{+}25, PZJL22, PJ13, PWX^{+}24, SLZ^{+}21, WXCW24, WWQ^{+}25a, WKN25, WWH17, WLW^{+}23, XQG^{+}23, XQG^{+}24, XGX^{+}25, XWH^{+}25, YTL^{+}21, YP15, YJTN18, YLYB24, ZGZ^{+}22, ZLZ^{+}22a, ZSZ^{+}22, AQL^{+}20, CWW^{+}25, CHS^{+}23, DFXY20, FZYY18, GHZW25, HLM^{+}22, HZH24, LZJ^{+}20, LYZX21, LLS^{+}24, LWHC24, LCZ^{+}24c, QWH^{+}21, RLY^{+}21, SJC^{+}19, TH22, TGC23, WDPX23, WZRD25, WHW^{+}21, WJQ^{+}22b, WZW24, XHL^{+}24a, XYW^{+}22, XLH^{+}23, YWH^{+}17, ZZZ^{+}22, ZGWZ23]. Identifier [WLS^{+}24a]. Identify [PA20]. Identifying [LLYH14, SS24]. Identity [ASVE13, CZL23b, LZYY24, MHW^{+}19, WWL20, WZW24, XHL^{+}24a, XLT^{+}23, CW10]. Identity-Matched [XLT^{+}23]. Identity-Preserving [MHW^{+}19, WWL20]. IEEE [LCK09, PBS12]. Illumination [GHR^{+}22, GHZW25, LOJZ18, PLB^{+}24, YJTN18, ZLZX23]. Illumination-Aware [PLB^{+}24]. Illumination-Enhanced [GHR^{+}22]. illustration [JWL06]. Image [AYH^{+}23, ARA^{+}23, APP^{+}22, AGCC25, AR15, BBUD24, BSQ^{+}24, CYW24, CZL^{+}23a, CCD24, CGGC20, CAJ19, CRL20, CZY^{+}21, CLQQ25, CZQ^{+}22, CBSC18, CLN^{+}21, CSW^{+}25a, DLL^{+}24, DDK24, DHP23, DBPM25, DHT^{+}19, DE12, DWC^{+}21, DDL^{+}25, EC16, FYW^{+}24, FHP23, FBGD23, FBG22, GSB^{+}22, GYY^{+}25, GLZW23, GAH25, GCZ^{+}24, GZZ^{+}24, GS19, HMUC21, HZC22, HHGW22, HH19, HD19, HLY^{+}23, HJ23, HL24, HXH^{+}24, HWW25, HWWL20, HSX22, HPLL23, HRL^{+}24, HZZ^{+}24, HQF^{+}25, IKTD25, JXL^{+}25, JWH21, JYZ^{+}25, JCB25, JXTC21, JWW^{+}24, JLLL25, KVB^{+}24, KIT^{+}24, KKGE18, KBI^{+}23, LA15, LQZH14, LSYM19, LLSX20, LCL^{+}21, LLZ^{+}21, LCZ^{+}24a, LML^{+}24, LLH^{+}24, LXC^{+}24, LXJ^{+}24, LYC^{+}25, LDQ^{+}25, LLL^{+}21, LZZ^{+}24b, LTL^{+}24, LTD^{+}21, LW24, LYCJ12, LYJ^{+}13, LZW^{+}19, LGLZ20, LH20, LLO^{+}20, LX21, LHS^{+}21, LGY^{+}22, LH22, LLYL22, LLZ^{+}23, LNH23, LLP^{+}23, LZW^{+}23, LYW^{+}23b, LYW^{+}24, LLL^{+}24a, LYH^{+}25, LNC^{+}25, LYL^{+}25, LYD^{+}21, MGY22, MWL^{+}24, MXZ^{+}24, NSK^{+}21, NWL^{+}20]. Image [NSM^{+}24, PLYM23, PLB^{+}24, PYZ25, PDD16, PKLK23, PHH^{+}24, PFC^{+}24, QLW^{+}23, RS16, RSE16, RNR^{+}22, RUD23, SCB^{+}24, SZHY19, SBS23, SZM^{+}19, SRAA17, SG22b, SS22, SSSA23, SZZ^{+}23, SHF^{+}24, SLMJ24, SJC^{+}25, SYM^{+}25, SS20, SZTL16, SZZT18, SZB^{+}22, SLZ24, TWM^{+}25, TYY^{+}18, TYY^{+}22, TCL^{+}23, TWKK21, TSP^{+}24, TP22, TZLZ21, TLZL24, WHF^{+}18, WHY18, WYM18, WLHT19, WCZ^{+}21, WMH^{+}22, WBH^{+}23, WXZ^{+}23, WLTL23, WGL^{+}23, WYGC24, WLZ^{+}24a, WSW^{+}24, WDC^{+}25, WQTX25, WFZ25, WWX^{+}21, WLH^{+}21, WZZ25, WYK12, WQL18, WYY^{+}25, WB16, WHW18, WHY19, WXQC20, WLCH22, WY23, WWL^{+}23, WWWL23, WLZ^{+}24b, XYJ^{+}20, XJFS25, XZCG23, XLF^{+}25, XFSZ20, XHYX23, XWH^{+}21, XWW^{+}21, XDL^{+}21, XLW^{+}23b, XCM^{+}23, XSL^{+}23, XLZ^{+}24, YLZ^{+}21, YTL^{+}21, YLLL21, YLS^{+}23, YTR^{+}25, YBW^{+}25, YWC^{+}25, YTOH22, YWNW15, YHXL20, YXYB21, YQC^{+}21, YLLL23, YLL^{+}23, YWW^{+}24, YSSF24, YQZP24, YXK^{+}25, YLL^{+}24, YHZ23, YTL^{+}24a, YMJ^{+}25, YFLF19, YZG^{+}20, YBTX23, YCZ^{+}23, ZY21, ZCXL23, ZRZX23]. Image [ZSMZ21, ZSJL22, ZR13, ZZY^{+}14, ZLN^{+}16, ZSS20, ZGD21, ZXX22, ZCQ^{+}23, ZGL^{+}23, ZTQX23, ZZB24, ZCYL24, ZZT24, ZGY^{+}24a, ZXT^{+}25, ZLZ22b, ZLL^{+}24b, ZZLW25, ZZG^{+}20, ZSZ^{+}24b, ZLH16, ZWYS20, ZHC^{+}23, ZXD^{+}23a, ZXD^{+}23b, ZLF^{+}23, ZZL^{+}25d, ZBYZ25, ZHL^{+}25, ZWZ^{+}23b, ZYZ^{+}25b, ZHS^{+}25, BWA13, CLC05, DCC^{+}13, hHLC10, HOSS13, LBD08, LMLC14, MLHL12, MB08, QLSQ12, TCW^{+}13, WLH^{+}25, YGHH12, YYS13, YW23, YHP25, ZYM^{+}10, ZI13, ZG08, ZLLT13, ZHLY11]. Image-Based [HSX22, NWL^{+}20, SZZ^{+}23, ZSJL22]. Image-caption [HMUC21]. image-keyword [LMLC14]. Image-Sentence [LZW^{+}19, XWH^{+}21]. Image-Text [HWWL20, LZW^{+}19, PYZ25, YLL^{+}23, ZZG^{+}20, ZSZ^{+}24b, CZQ^{+}22, FHP23, YLL^{+}24]. Image-to-Image [LLH^{+}24, ZY21]. Image/Video [LLSX20]. ImageNet [MKS20]. Imagery [BCP14, CSBJ24, HYT^{+}25, WLFL23, LLC11].

Images [AMC^{+}25, BYM^{+}18, BV25, CFN^{+}23, CH15, CS17, CLS17, CSSZ23, CC17, DNPG^{+}17, DPL^{+}23, GUH^{+}20, GGB24, GHR^{+}22, GHY^{+}24, HZL^{+}16, IAJ^{+}24, JGY^{+}24, JLL^{+}22, JJML24, KBB21, LYLL23, LSK^{+}15, LWZY22, LHY^{+}24, LZG^{+}24, MSS^{+}24, NNK25, PSN^{+}24, PA20, TM25, TS20, VCO15, WHZ^{+}24, WLL^{+}19, XJG^{+}22, XWZ24a, XWW^{+}21, Yan17, YHH^{+}22, YDCZ23, YWY^{+}23, YCZ^{+}24, YHZ19, ZHG^{+}21, ZLL^{+}24a, ZSZ^{+}24a, ZCX^{+}24, ZWZ^{+}23a, ZLL20, ZZC^{+}15, ZZMZ21, ZWJ^{+}24, HCW^{+}07, HZW^{+}11, LMLC14, PJ13]. ImageSense [MLHL12]. Imagine [ZXX22]. Imaging [FMG20, TKPL20, YKQ^{+}21, DÇ07]. Imbalance [KIT^{+}24]. Imbalanced [ZZP^{+}20]. IMCE [AVJ05]. Img2Doc [LXB^{+}22]. Immersive [HHH22, KTK^{+}17, WDE25, ARE13, BBT^{+}05, Cho13, MRS^{+}07, WAK^{+}12, YWN^{+}10a, YWN^{+}10b]. Immune [YLS^{+}23]. Impact [CCG20, KLS^{+}18, PDD22, SVA^{+}21, WAC^{+}25, AG13, Cha13]. Impaired [HCM^{+}25, HWY^{+}11]. Impersonation [SED25]. Implementation [GZT21, GGML22, SS20, PZ08]. Implications [CE23, KWK25, RABC24]. Implicit [CIPE18, YWL^{+}24, ZWH^{+}23, ZXZ^{+}25]. Importance [LZG^{+}25]. Importance-Based [LZG^{+}25]. Important [REP^{+}19, XWL^{+}25b, ZCL^{+}22]. Impression [BBRS23]. Impressions [ONAGP19]. Improve [KZGH15, WDR^{+}24]. Improved [AC19, Ber18, BCP14, CPCM21, HHL^{+}22, SMK^{+}24, SZFR23, TLZ^{+}20, ZQRS18]. Improvements [TWM^{+}25]. Improving [ENHN09, IKTD25, MXZ^{+}24, PYZ25, PDD16, PWY^{+}25, WTZ^{+}20, WJQ^{+}22a, WZZ^{+}22, YHH^{+}22, YHX^{+}23, YZL^{+}24, ZFFX24, CSJ^{+}08, GL08]. IMUs [CZW^{+}25]. In-a-group [MGP19]. In-Loop [LLL^{+}25, JLL^{+}21, LZZ23b, LJZ^{+}22]. in-the-Wild [ZDZ^{+}23, ZSZ^{+}24a]. In-video [LWL^{+}12]. Incentive [SRPH16, ILL08]. Incentive-Based [SRPH16]. Incentives [WLQL12]. Inception [PA20]. Incomplete [FMG20, GLL^{+}24, LZZ^{+}24a, SZL^{+}24, ZLW^{+}21, ZLLH25, ZLH16]. Inconsistency [ZZL^{+}25a]. Inconspicuous [SZLL17]. Incorporating [LGX^{+}08]. Increasing [SZM^{+}19]. Incremental [LYL^{+}23b, PB19, YZS^{+}22, YHG^{+}24, YZYX24, BB11]. independent [MON21]. Index [WLC^{+}20, WHZ^{+}24, XZY^{+}24b, LLHS12]. index-based [LLHS12]. Indexing [WZNM14, CWC10, VV11, Yan10]. Individual [TMB^{+}22, TSP^{+}24, ZGL^{+}18, ZQKH19]. Indoor [LL15, SSZ^{+}23, SZZZ25, ZLY^{+}24]. Induced [LNC^{+}25]. Industrial [KUH^{+}22]. inefficiencies [WLZ12]. Inertial [SZZZ25]. Inference [MYX23a, PLZT22, PLYM23, TWFW24, Tas22, ZGY^{+}24b, MPE^{+}11]. Influence [ROST20, WZZL25, SX13]. Information [AG13, BRG24, CW25, CC17, FYW^{+}24, LZL^{+}24c, MBK^{+}23, NLW^{+}21, OL25, PWHP24, QWH^{+}21, SZF^{+}25, SHT^{+}24, UFJ21, WLSN24, WLH^{+}25, WFZ25, WDS21, WH22, XJFS25, XCC^{+}23, YLLL23, YWW^{+}24, YWC24, YSZ15, ZYCD24, ZZC^{+}15, ASVE13, CCK06, Cha13, HAE11, LSDJ06, LWL^{+}12, TBC^{+}11, XC06, ZWM12a, ZWM12b]. Information-Assisted [XJFS25]. Information-Based [YLLL23, ZWM12b]. Informative [FSX14, WZNM14, ZFSX21]. informed [ICZ^{+}22]. Infrared [CZL23b, DSL^{+}25, HZH24, HCZ^{+}25, LWZY22, LYW^{+}24, WWQ^{+}25a, WZL^{+}24, YLYB24, ZGZ^{+}22, LLS^{+}24]. Injection [LYW^{+}24]. Injection-Based [LYW^{+}24]. Inner [LXB^{+}22]. Innovation [Sin21]. Innovations [USM^{+}25]. Innovative [RABC24]. Inpainting [JWW^{+}24, LDQ^{+}25, LW24]. Input [Cla18]. Insertion [CZS^{+}25, MB08]. insertion-extraction [MB08]. Insight [RLY^{+}21]. Inspired [GLGG25, WBR25]. Instance [LLZ^{+}20, LYL^{+}21b, SWS21, TLZ^{+}21, WJ15a, WLSZ22, XZL^{+}22a, XHL^{+}24b, YZW^{+}24, ZZG^{+}20, CLZ^{+}23a, JW21, JLZ^{+}21, SZZY20]. Instance-Based [XHL^{+}24b]. Instance-Level [WJ15a, YZW^{+}24]. INSTRE [WJ15a]. Instruction [HZL^{+}25]. Instrument [TWB^{+}23]. Integrated [CCF^{+}24, LWH20, XHZ^{+}21, YXK^{+}25, ZCL^{+}24, WXCW24, AVJ05]. Integrating [SS22, WLH^{+}21]. Integration [LZL^{+}25, LGF^{+}14, XZY^{+}25, AP13, Yan10]. Integrity [BCP14, CCD24, LHJ^{+}24, SKC24, LLC11]. Intelligence [PDD22, TKPL20, Yan17]. Intelligent [DZW^{+}21, HHLY19, JXL^{+}25, LPD^{+}24, RHS^{+}20, SWM^{+}24, WJXN24, YZS^{+}23, ZZC^{+}23, ZYZE19b, ZWLZ24, ZYL^{+}24, ZWZL21, ZCS^{+}20]. Intensifying [KTK^{+}17]. intent [ZYM^{+}10]. Inter [AMMG16, HYLG23, LJ25, XHL^{+}24a]. Inter-camera [XHL^{+}24a]. Inter-Coding [LJ25]. Inter-Component [AMMG16]. Interact [NSJB17]. Interaction [BBRS23, KPK^{+}24, LLG^{+}23, LJW^{+}25, LZDL25, LLL23, LCL^{+}24b, LGF^{+}14, LHF^{+}14, LPD^{+}24, LQX^{+}24, MON21, MFL^{+}16, QRTM24, QXG^{+}24, RLXW24, RBLJ25, WWS^{+}24, WZZ25, YYY^{+}24, ZYW^{+}25, ZTR^{+}22, And13, RWP07, RHE10, WYM07, YBW^{+}25]. Interaction-process-guided [LLL23]. Interactions [TVZ^{+}19, XWY21, BRA^{+}09, CYMW07, HCKL13, ZSO13]. Interactive [CPSH14, GCF^{+}21, GAB^{+}17, HCKL13, HPW20, HLC^{+}25, LMF^{+}14, LXZ^{+}25, LKM^{+}19, LVM^{+}21, LYXA22, MBK^{+}23, MSG22, NN21, OBBW12, PB14, RTM^{+}24, SCXC15, UTK^{+}08, VNC^{+}11, WWHW14, WBR25, YTOH22, ZZL^{+}25a, ZHS^{+}25, CTGP08, CVV06, CBS08, LJP08, MMW10, NW08, SCFL14, SNC12, TCW^{+}13, TTR12]. InteractNet [LQX^{+}24]. Interest [CWXW25, KRKK14, SED25, SGY^{+}23, TAS16, CL12, FDKB11]. interfaces [FPH^{+}08]. interleaving [CLC05]. Intermodal [ONAGP19]. International [JPS05, NLS13]. Internet [AAA^{+}21, CCG^{+}08, GYF^{+}21, HSL^{+}22, JTZ^{+}16, LWLZ13, LWZ^{+}21a, LQS21, SP21, SMTR22, SZEST21, SKC24, WWL13, WLL^{+}19, YWF^{+}24, ZYM^{+}10]. Interperson [ONAGP19]. Interpolation [DOD23, JTZ^{+}16, LWH^{+}22, ZHD^{+}23, ZLC^{+}25]. Interpretability [DYZ^{+}23, PHH^{+}24, ZZS^{+}25]. Interpretable [DYL^{+}25, GLGG25, JXLD25, ZYQM24]. Intersectional [LZH^{+}24b]. Intervention [MNPOF22, YLYB24]. Interventional [WLB^{+}25]. Intra [GGML22, LYL^{+}21a, LLL^{+}25, WWY^{+}21, XLW23a, ZZL^{+}23, ZLK^{+}19]. Intra-Mode [WWY^{+}21]. intranet [CBR14]. Introduction [ACGH18, BJLX11, CCK06, CDGJ09, CBS08, CCO15, CLS^{+}21, FSEW25, GZ20, GPHOH12, GDGC07, GTLG14a, GWZ25, GMWW25, HT15, IPMD25, Kan12, KZHC13, LX21, MSG22, NLS13, NLKB24, ODMD17b, PCH^{+}20, SWS21, SZEST21, SCXC15, SLYE11, SHOG12, SLK21, SKC24, TKPL20, VNDB24, WFZ^{+}21a, WFZ^{+}21b, XSEZ10, YZRW22, ZZMS14, ZZX^{+}20, ZGD21, ZLZ^{+}22a, ZHLJ22, ZPR^{+}25, ZJSJ20, ZTB20]. Intrusion [CSŞ^{+}25]. Invariance [XZY^{+}24a]. Invariant [LTL^{+}24, LZW^{+}19, LH20, RS16, YML^{+}22, YWNW15, ZLY^{+}24, GZHD12, LLZ^{+}22, RE24, SZL^{+}22, ZNNH25]. Invertible [DXY^{+}21, WWQ^{+}25b]. Investigating [BFAS15]. Investigation [LGT^{+}24, ROST20]. Invisible [JGY^{+}24, MB08, WWZ^{+}25]. IOD [YZS^{+}22]. IoMT [TP22, XHZ^{+}21]. IoT [BWP^{+}24, BCNP24, GLL^{+}24, GTP^{+}22, LHJ^{+}24, SSP21]. IoT-based [SSP21].

IoT-enabled [BCNP24]. IoV

[XFQ^{+}21, YXZQ22]. IP

[BDV08, CNG22, SSTK07, XWZ^{+}24b]. IPT [Hua23]. IPTV [LCC^{+}14b, VGNL10]. IR [RHE10]. iris [VPSS^{+}13]. Irrelevant [LCZ^{+}24c]. IRTS [ZWZL21]. Ischemic [AM24]. ISODATA [TLZ^{+}20]. ISODATA-Improved [TLZ^{+}20]. ISP [AH20, FAA18, SZ12]. ISP-friendly [SZ12]. Issue [ACGH18, CCO15, CLS^{+}21, FSEW25, GZ20, GTLG14a, GWZ25, GMWW25, HT15, IPMD25, LPD^{+}24, MSG22, NLKB24, ODMD17b, PCH^{+}20, SWS21, SZEST21, SCXC15, SLK21, SKC24, SKVHC18a, TKPL20, VNDB24, WJXN24, WFZ^{+}21b, ZZMS14, ZYZE19b, ZZX^{+}20, ZGD21, ZLZ^{+}22a, ZHLJ22, ZPR^{+}25, ZJSJ20, BLJX10, BJLX11, CCK06, CBS08, Dao17, GS11a, GDGC07, Hae10, Kan12]. ISTN [XZL^{+}22b]. Iterative [BSQ^{+}24, DLL^{+}18, DYN^{+}25, LGM^{+}24, ZLL^{+}24a]. Iteratively [LZZ23b].

JAUNE [HP17]. JDAN [WHL^{+}23]. Jitter [CPCM21, YZL^{+}24]. JND [YWY^{+}23]. Joint [CH21, CJZ^{+}24, EC16, FH20, FHF^{+}24, GWWC25, HP17, LZLJ22, LWHC24, MSFR25, PFC^{+}24, WHL^{+}23, WXJL25, XW25, XZL^{+}22b, YLCC18, YQZP24, YF22, YZY^{+}25, YZG^{+}20, ZRZX23, ZH18, ZHS20, ZLX24, ZYW^{+}24, ZSZ^{+}24b, FYH10, ZC12]. Joint-Dataset [MSFR25]. Joint-Level [GWWC25]. Jointly [SWS^{+}22, WCL^{+}22, ZSZ^{+}22]. JoT [ZSZ^{+}22]. JoT-GAN [ZSZ^{+}22]. JPEG [LZL^{+}24a, PFC^{+}24, WHL^{+}21, XJG^{+}22, YCZ^{+}24, ZC12]. JPEG-2000 [ZC12]. JPEG-compatible [PFC^{+}24]. Just [TWKK21].

KAN [LWW^{+}25]. KANformer [LWW^{+}25]. Kernel [ASLA18, CZL^{+}23a, DLL^{+}18, GZL^{+}20, LKM17, WZNM14, ZSS20, QLSQ12]. kernels [QHR^{+}08]. Key [SSP21, TDW^{+}23, WKN25, XCC^{+}23, ZHL19, YH13]. keyframe [ZHLY11]. Keypoint [XG24]. Keypoints [WLL^{+}24]. Keypoints-Driven [WLL^{+}24]. Keystrokes [THR^{+}22]. keyword [LMLC14]. keywords [XXD^{+}08]. KF [WLL^{+}24]. KF-VTON [WLL^{+}24]. Kinematic [GWWC25]. Kinematic-Aware [GWWC25]. Kinship [ZLC^{+}23]. Knee [PLB^{+}24]. kNN [PRGA18]. Knowledge [AAA^{+}21, CWW^{+}25, CXY^{+}25, CGGC20, DJWZ25, FHP23, GSY^{+}24, HYG^{+}21, JXL^{+}25, JP11, JMS23, LXB^{+}22, LZZS23, LYX23, LSS^{+}24b, LYW^{+}25, MYX23a, PLZT22, QHFX21, SWK^{+}22, TLZL24, WDR^{+}24, WLS^{+}24b, WXCW24, WLH^{+}21, XLX^{+}24, ZYJZ25, ZSS^{+}23, ZZL25e]. Knowledge-Aware [ZZL25e, QHFX21]. Knowledge-based [LXB^{+}22, LYW^{+}25, WLS^{+}24b]. Knowledge-driven [HYG^{+}21]. Knowledge-enriched [LYX23]. Knowledge-integrated [WXCW24]. KPI [WKN25].

L [TS20, ZHD^{+}23]. Label [LQH18, LYC^{+}12, LNH23, NDX^{+}21, ZWY21, ZHC^{+}23, ZXD^{+}23a, ZXD^{+}23b, JW21, LYCJ12, RXY^{+}25, WXJL25, YBW^{+}25, YCZ^{+}23]. Label-to-region [LYC^{+}12]. Labeled [ASI^{+}21]. labeling [LJG^{+}24, TCW^{+}13]. Labels [RUD23, RXY^{+}25, ZSJL22, ZLN^{+}16, ZLLH25, ZLZ22b]. Ladder [THA^{+}25]. LAN [BPM15]. LANBIQUE [GSB^{+}22]. Landmark [FNH22, YSZ15]. Landmarking [SHWC19]. landmarks [JGZ^{+}11]. Landslide [TLZ^{+}21]. Language [CLP^{+}23, FSEW25, GCC^{+}25, GGSV25, GZLW18, GMWW25, HZPL21, HZL^{+}25, LFP^{+}22, LWZ^{+}25, NNK25, QCW^{+}25, RXY^{+}25, SZF^{+}25, TWL19, TXW^{+}25, WDR^{+}24, WSLC24, WLH^{+}25, ZCL^{+}22, ZSZ^{+}25, ZDQ^{+}23, ZHS^{+}25, CL07, ML11, MGCH13, MSC^{+}23, ZMX^{+}23, GSB^{+}22]. Language-Based [ZCL^{+}22, ZDQ^{+}23, GSB^{+}22]. Language-guided [SZF^{+}25, WSLC24, ZSZ^{+}25]. Language-substitutable [WLH^{+}25]. Laplacian [RBBS25]. Large [ASI^{+}21, BBB^{+}23, BLW^{+}24, CML^{+}13, FHG^{+}17, HDW^{+}18, LSK^{+}15, LCC^{+}14a, QCW^{+}25, REP^{+}19, SW18, WMH^{+}22, WDC^{+}25, XWZ24a, YSY^{+}22, ZDZ^{+}23, ZZL^{+}25c, ZWS^{+}20, ZJSJ20, ZWYS20, ZHS^{+}25, ZFC^{+}24, CE10, DCM13, VGNL10, WDCX07, WLZ08, ZLLT13]. Large-Area [LCC^{+}14a]. Large-Scale [BLW^{+}24, HDW^{+}18, LSK^{+}15, REP^{+}19, WDC^{+}25, YSY^{+}22, ZDZ^{+}23, ZWYS20, BBB^{+}23, CML^{+}13, SW18, WMH^{+}22, XWZ24a, ZWS^{+}20, ZJSJ20, CE10, VGNL10, WLZ08, ZLLT13]. Late [GZLW18, LW07]. Latency [HLD18, LLC^{+}25, BLS^{+}19, BTBZ20, GKW08]. Latent [JK25, LXT^{+}24, MGS18, QZXH14, WCF^{+}17, WHJ20, WWY^{+}25, YGNT19, ZRZ^{+}21]. Latin [SZFR23]. Launching [HJMY15]. Layer [CW25, EWSZ15, FSK^{+}15, GCZ^{+}24, GQSG24, HSN^{+}14, HSL^{+}20, AAS^{+}20, HH11a, LYC^{+}12, NDX^{+}21]. Layer-wise [GQSG24]. Layered [CLC05, GL11, ZC12]. Layers [FZW^{+}25a]. Layerwise [KBB21]. Layout [CSW^{+}25b, WLH^{+}25, YMX^{+}16, BZ05, HCW^{+}07]. Layout-image [WLH^{+}25]. LayoutEnc [CSW^{+}25b]. Layouts [YMY^{+}21]. LBSNs [ZCY^{+}13]. LCSNet [XYL^{+}23]. Lead [ZLT^{+}25b, ONH07]. Leaf [ZWJ^{+}24]. LeafNet [ZWJ^{+}24]. Learn [MNPOF22, NNK25]. Learnable [GYY^{+}25, LSX^{+}25, YHP25]. Learned [JYZ^{+}25, YYX^{+}24, YZJ^{+}25, ZDD^{+}24, ZZY18]. Learning [AQL^{+}20, ASLA18, AM24, AS20, AS22b, BBUD24, BWP^{+}24, BPT^{+}15, BBZ18, BHD24, BC15, BPB^{+}22, CFN^{+}23, CHS^{+}23, CAJ19, CLZ^{+}23a, CLWW24, CZX^{+}24, Che24, CSW^{+}25a, DWS^{+}24, VTD22, DG17a, DPL^{+}23, DBSL24, DLL^{+}18, DJWZ25, FJF^{+}25, FH20, FNH22, FLG^{+}21, FXF^{+}23, GLW20, GHQ24, GLGG25, GLL^{+}22, GUH^{+}20, GCF^{+}21, GMWW25, GZW^{+}25, GSDT21, HZC22, HZS^{+}24, HWZL24, HLW^{+}21, HAAM19, HLM^{+}24, HHH22, HZL^{+}21, HLW^{+}25, HPW20, HJZ^{+}23, IPMD25, IAJ^{+}24, JLA^{+}23, JW21, JLZ^{+}21, JMLL20, JCF^{+}22, JXTC21, JLK^{+}24, JJML24, JMS23, KN21a, KPL^{+}22, KT21, KIT^{+}24, KUH^{+}22, KN21b, LCY^{+}23, LQH18, LPY^{+}19, LZXY20, LZX^{+}21, LYZX21, LLZ^{+}21, LLS^{+}22, LZWC23, LZH^{+}24b, LZL^{+}24a, LXC^{+}24, LZZ^{+}24a, LYC^{+}25, LJW^{+}25, LCH^{+}25, LWLC24, LYDY24, LZY^{+}24, LXT^{+}24, LWZH19, LHS^{+}21, LXL^{+}23, LWH^{+}23, LLP^{+}23, LYXY23, LZW^{+}23, LLZ^{+}24, LZL^{+}24b, LHJ^{+}24, LQD^{+}24, LYH^{+}25, LDL^{+}25, LSX^{+}25, LGW^{+}25, LZXP25, LWHC24, LWZW21, LZF^{+}22, LYXA22, LPD^{+}24, LCZ^{+}24c, MBXZ24]. Learning [MWL^{+}24, MHF24, MPM^{+}25, MSC^{+}23, MAX^{+}24, MGS18, MYX^{+}23b, MSFR25, MDAE19, MHL^{+}25, MKC21, MPTD22, NDX^{+}21, NWNL21, NCL^{+}23, ODMD17a, ODMD17b, PLZT22, PYZ25, PZJL22, PRGA18, PQ19, PHH^{+}24, PWX^{+}24, PSL^{+}24, PS17, QXG^{+}24, QLW^{+}24, RHAG21, RYZ^{+}23, RK15, RE24, RXY^{+}25, SMTR22, SS24, SZZY20, SWS21, SSY20, SXY^{+}23, SHE21, SWH06, SGYX22, SHF^{+}24, SZW^{+}24, THR^{+}22, TSZ^{+}23, TCS^{+}25, TYY^{+}18, TLZ^{+}21, TLY^{+}22, TGC23, TY23, TRK^{+}20, VVSV17, WWHW14, WYM18, WLHT19, WZ20, WTZ^{+}20, WZD^{+}20, WMM^{+}23, WBH^{+}23, WLFL23, WLS^{+}24b, WLB^{+}25, WXJL25, WWX^{+}21, WLL^{+}19, WHW^{+}21, WJQ^{+}22a, WLCH22, WLZ^{+}24b, WGHX24, XQG^{+}23, XGX^{+}25, XLZ^{+}24, XHL^{+}24b, XLT^{+}23, Yan17, YLZ^{+}21, YZXY15, YLK^{+}20, YZRW22, YPSC22, YW23, YHG^{+}24, YHH24, YNY25, YP20, YDJ24, YTRC19, YCLH22, YLYB24, ZCL^{+}22, ZWH^{+}23, ZWX24, ZYL^{+}17, ZYZE19a, ZYZE19b, ZZP^{+}20, ZZX^{+}20, ZMH^{+}20, ZGD21, ZTR^{+}22, ZGZ^{+}22, ZHLJ22, ZXX22, ZGL^{+}23, ZLX24]. Learning [ZLY^{+}24, ZXZ^{+}24a, ZCYL24, ZYW^{+}24, ZNNH25, ZZS^{+}25, ZYZ^{+}25a, ZZL^{+}25b, ZZL^{+}25a, ZHY^{+}25, ZCY^{+}19, ZGD^{+}19, ZZW^{+}19, ZJT^{+}23, ZYH^{+}24, ZYL^{+}24, ZLL^{+}24b, ZLL20, ZLL^{+}24d, ZZL^{+}24c, ZWF^{+}20, ZZMZ21, ZLC^{+}23, ZZL^{+}23, ZFC^{+}24, CL07, GZW^{+}24a, hHLC10, LYL^{+}23b, MRS11, RUD23, TZ24, TTR12, ZSJL22, ZI13, ZHLY11, WWW^{+}22]. Learning-Based [BWP^{+}24, CLWW24, DG17a, KN21a, VVSV17, ZYZE19a, ZZP^{+}20, ZZL^{+}23, AS22b, JXTC21, LYXY23, LYXA22, ZMH^{+}20, ZLL^{+}24b]. Least [XW17]. Lecture [WKN25, ZRCH08]. Led [MA10]. Length [KMSW18]. Lens [LCL^{+}23b]. LEO [FZW^{+}25b]. Lesion [RTR21, RLAM25, TRRB20]. Less [XQG^{+}23, DKJ^{+}21]. lessons [She13]. Level [BSQ^{+}24, CWXW25, GWWC25, HCJ^{+}24, LXL^{+}18, TXWK24, TWKK21, WJ15a, WC23, YZS^{+}22, ZZLL17, ZWYS20, CLZ^{+}23a, KDC08, LLZ^{+}25, LX21, LZL21b, LCL^{+}24b, MSL10, OMP07, PYZ^{+}20, SJC^{+}19, TSH^{+}25, YV23, YZW^{+}24, ZWZ^{+}23a, ZLF^{+}24]. Leveraging [APC^{+}25, CSW^{+}25b, LZL^{+}25, MSKYJ21, TSH^{+}25, WCZ^{+}21]. LFGAN [CRL20]. LFIZW [WYY^{+}25]. LFIZW-GRHFMR [WYY^{+}25]. LFR [DHP23]. LFR-GAN [DHP23]. Library [Dao17]. licenses [SEK12]. LIDAR [YF22, WHX^{+}25, ZXZ^{+}24b]. Life [JSEI16]. Lifelog [XDL^{+}21]. Lifelong [BPB^{+}22, CWW^{+}25, XZL^{+}25, XLS^{+}25]. Light [APP^{+}22, CRL20, GZT21, GHR^{+}22, GHZW25, HHGW22, HSL^{+}20, LDL24, LG24, LWW^{+}25, PLB^{+}24, WHZ^{+}24, WYY^{+}25, WLZ^{+}24b, XWW^{+}21, ZWL^{+}17, ZXT^{+}25, ZHL^{+}25, ZSMZ21, ZLF^{+}23]. Light-Aware [WLZ^{+}24b]. Lightweight [DLL^{+}24, HZZ^{+}24, KMK^{+}21, LCL^{+}21, LQG25, SSP21, WLTL23, WGL^{+}23, YMS^{+}24, ZHD^{+}23, ZSZ^{+}23, MHT^{+}08]. Like [WPRC18, ZWJ^{+}24, YCFX23]. likelihood [MC19]. LiLTv2 [WLH^{+}25]. Limited [ASI^{+}21, LHZ^{+}23, WWWL23]. LINE [WH22]. Linear [JYZ^{+}25, LK18, RSE16]. lingual [ZGY^{+}24a]. Linguistic [CZZ^{+}24, SWS^{+}22, WDR^{+}24, ZGY^{+}24a]. Link [LYJ^{+}15, TWY24, VCO15, HCW^{+}07]. Link-Aware [LYJ^{+}15]. links [HNL08]. Lip [LCH^{+}25]. Lipreading [XYL^{+}23]. Listen [WLC^{+}20]. Literature [HC22]. Live [BQBLN18, FFCM24, GGB14, HWLC19, JSEI16, LCC^{+}14b, MN16, MATW17, PCB^{+}21, RXC14, SS17, SkFM18, WAD^{+}18, YNY25, YH14, ZMH^{+}20, KW11, SZ12, WWGT09, WLZ08, WLZ12]. Livecast [MZL^{+}18, ZLW17]. Liver [JXL^{+}25]. LiveSky [YLZ^{+}10]. Load [AMT25, DL14]. Local [BR22, DHP23, DYZ^{+}23, HHL^{+}22, HMOS17, HZPL21, HCZ^{+}25, HWW^{+}23, JCF^{+}22, LSXZ23, LLH^{+}24, LTL^{+}24, MYGX21, NDX^{+}21, WXS^{+}25, XZY^{+}24b, YBW^{+}25, ZHD^{+}23, ZZJ^{+}23, ZHG^{+}25, QLSQ12]. Local-Aware [HCZ^{+}25]. Local-Global [MYGX21]. Localization [DZWH23, ICZ^{+}22, LZL^{+}24a, LZZ^{+}25, LNT^{+}21, PN16, SRAA17, SJC^{+}25, XZZL23, XWZ24a, ZTQX23, ZDQ^{+}23, FYH10, LML^{+}13, MC19, NC13, WT10]. Localizing [CLZ^{+}23b]. Located [ZTB20]. Locating [JGY^{+}24]. Location [FSX14, FF23, NHP^{+}16, REV^{+}12, SG07]. Location-Based [FF23, NHP^{+}16, REV^{+}12]. locations [PJ13]. LOD [DÇ07]. log [JKKL08]. log-structured [JKKL08]. Logic [YH14]. Logits [WSX^{+}24]. Logo [HLM^{+}24, WMH^{+}22]. LogoDet [WMH^{+}22]. LogoDet-3K [WMH^{+}22]. Lollipop [RD17]. Long [ATS19, CLZ^{+}23b, HLTH24, HZL^{+}25, LZZ^{+}24b, RUD23, SLBS20, TSZ^{+}23, WWY^{+}24, ZTX^{+}23, ZPL^{+}23, ZZC^{+}25b]. Long-Distance [ATS19]. Long-Instruction [HZL^{+}25]. Long-range [ZPL^{+}23]. Long-Tail [TSZ^{+}23, RUD23]. Long-Tailed [WWY^{+}24, LZZ^{+}24b]. Long-Term [CLZ^{+}23b, SLBS20, ZZC^{+}25b, ZTX^{+}23]. Look [HWLC19, WLC^{+}20, Eff13]. Looking [DCO10, NCL^{+}23, Row13]. lookup [REV^{+}12, ZCL^{+}12]. Loop [LLL^{+}25, JLL^{+}21, LZZ23b, LJZ^{+}22]. Lorenz [SSSA23]. Loss [BHD24, CPSH14, CYZ^{+}24, JTZ^{+}16, JK25, LZT^{+}20, MAZ22, SLMJ24, TH22, TGSF21, TLZ^{+}20, XJW^{+}22, XYW^{+}22, ZGR21, ZLL^{+}24a, ZZG^{+}20, CLC05]. Loss-Based [ZGR21]. lossless [MTTH13]. lossy [LLP06]. Low [BTBZ20, GKW08, GHR^{+}22, HHGW22, HXL24, LLSX20, LLC^{+}25, LWW^{+}25, MFL^{+}16, MATW17, PLB^{+}24, PJL23, SAF19, SZM^{+}21, WFC^{+}25a, WZZ^{+}22, WLZ^{+}24b, XWW^{+}21, BLS^{+}19, YWG^{+}20, ZSMZ21, ZRZ^{+}21, ZXT^{+}25, ZLF^{+}23, ZHL^{+}25, FKFB05]. Low-Complexity [SAF19, HXL24]. Low-Delay [MATW17]. Low-Interaction-Delay [MFL^{+}16]. Low-Latency [LLC^{+}25, BLS^{+}19, BTBZ20, GKW08]. Low-Light [GHR^{+}22, HHGW22, LWW^{+}25, PLB^{+}24, WLZ^{+}24b, ZXT^{+}25, ZHL^{+}25, ZSMZ21, ZLF^{+}23]. low-power [FKFB05]. Low-Rank [LLSX20, SZM^{+}21, ZRZ^{+}21, WFC^{+}25a, YWG^{+}20]. Low-Resource [WZZ^{+}22]. LPIC [LSX^{+}25]. LRTA [SG22a]. LSH [XHZ^{+}21]. LSH-based [XHZ^{+}21]. LSTM [CWTG22, LZL20b, LWZ21b, TFHM24, YHXL20, YSC21, ZYL^{+}24]. LSTM-GAN [YSC21]. LSTMs [WYM18]. LTE [FSK^{+}15]. Lung [WZD^{+}20]. Lyrics [DYZ^{+}23, YTRC19, YSC21, MSL10].

MAC [BPM15]. Machine [BWP^{+}24, VTD22, DG17a, GEL^{+}15, KN21a, LHJ^{+}24, LDL^{+}25, TRK^{+}20, TLZL24, XYC^{+}23, ZZX^{+}20, ZZT24]. Machines [LZG^{+}25]. MAINet [WZZ25]. Major [TESU16]. Make [HLY21]. Makeup [LZYY24]. malicious [JC10]. Malware [GTP^{+}22]. managed [PS05]. Management [CPCM21, FAA18, GCTG24, LCL22, MY15, PVWD18, SWM^{+}24, SKVHC18a, SKVHC18b, WLS^{+}24a, DÇ07, ZCAP08]. Managing [AS22a]. MANETs [SKSZ13]. Manifold [HPH^{+}20, YGL^{+}23, WYM07]. Manipulated [YNLZ22]. Manipulation [FDBP23, LZL^{+}24a, MKY^{+}24, RNR^{+}22, SO22, SG22b]. manual [CJP^{+}21]. many [CBR14, CL07]. many-to-many [CBR14]. many-to-one [CL07]. Map [CWL24, HZH24, LCL23a, SK23, WZTL19, WZW24, ZZW^{+}19]. Mapping [HAAM19, NALM23, SLG^{+}24, SGY^{+}24, WWQ^{+}25b, XDL^{+}21, ZCY^{+}19]. Margin [SYY^{+}24]. Marker [JGY^{+}24]. Markerless [LZL20a]. Market [WDJ^{+}21]. MARS [EFF^{+}25, WLFL23]. Mashup [QSZ^{+}21]. Mask [CZL^{+}23a, FDZ^{+}25, FHK^{+}23, MLJ^{+}22, WTZ^{+}20, YPSC22]. Mask-aware [WTZ^{+}20]. Mask-Guided [MLJ^{+}22]. Masked [HWW^{+}23, TM25]. Masking [AMMG16, YZL^{+}24]. Massive [BYM^{+}18, EG17, ZWL15]. Massively [ZZL21]. Mastering [GGB24]. Match [BLW^{+}24, ZLLT13]. Matched [XLT^{+}23]. Matching [CDL^{+}20, CYZ23, CZQ^{+}22, CTBC22, HLM^{+}22, JLL24, LLS^{+}22, Lin15, LZW^{+}19, LSN^{+}20, LNT^{+}21, SLNL20, XWH^{+}21, YF22, YLL^{+}24, ZT22, ZMX^{+}23, CWC10, MHT^{+}13, ZHLY11]. matchings [GYN12]. Mathematics [GLGG25]. Mathematics-Inspired [GLGG25]. Matrix [HZC^{+}16, WSLM18, WDLW23, ZWY21, ZJL^{+}21, QLSQ12, WC12]. matter [COM^{+}11]. Matters [KN21a, TYL^{+}25, ZZZ^{+}22, Han13].

Matting [QLW^{+}23, YQC^{+}21]. Max [HZW22]. Maximize [SSP^{+}24]. Maximizing [ZZC^{+}25b, MMW10]. MD [ZWL^{+}21]. Me [ATS19, HWLC19]. Mean [SZHY19]. Meaningful [HQLW25, SA16]. Means [FSEW25]. Measure [SMN^{+}22]. Measurement [FZW^{+}25b, LCL14, LWWZ20, PSN^{+}24, YF22]. Measurement-Based [FZW^{+}25b]. Measurements [SkFM18, TESU16, GHP^{+}06]. Measures [HMUC21]. Measuring [LCW16, ZGL^{+}18]. MEC [ZZC^{+}25b]. Mechanism [CXL^{+}22, LDQ^{+}25, Lin15, LS21, SHIE15, SGY^{+}23, SYS17, SZB^{+}22, WWZ^{+}25, YZG^{+}20, YZ23, ZZP^{+}20, ZXZ^{+}24a]. Mechanisms [DZWH23, XWL^{+}25b, BZ05, ZWM12a]. Media [BPT^{+}15, CSJC17, FZW^{+}25b, FPA24, FHG^{+}17, GTLG14a, LWY22, NSK^{+}21, NHZ^{+}24, OL25, TMB^{+}22, TVK18, XWL^{+}25b, ZLT^{+}25b, ZGL^{+}18, AVJ05, Ano13, BLJX10, BJLX11, CY11, CCG^{+}08, CXS^{+}08, GS11b, GA12b, GSM^{+}08, HPW20, ILL08, LJP08, LSDK12, RSB11, SX11, Sun13, TBC^{+}11, TSHP05, ZO13]. Median [GS18, GS19]. Medical [FYZ^{+}21, JXL^{+}25, LHS^{+}21, LLZ^{+}23, LYH^{+}25, RPE^{+}17, SLK21, SSSA23, TP22, WZZ25, WDS21, XZCG23, XZL^{+}25, YXYB21, ZCXL23, ZWX24, ZZP^{+}20, ZGD21, ZZLW25, ZCS^{+}20]. Medley [LLJW15]. MEDUSA [LTHT24]. Meet [WZZ^{+}22]. Meets [ZXT^{+}25, ZHS^{+}25]. Meibomian [LWZY22]. Meixner [YTL^{+}24a]. Melody [DYZ^{+}23, YSC21]. Memorability [SZM^{+}19]. Memory [QDX^{+}24, YWLZ23, YZL^{+}14, ZHC^{+}23]. Mental [CXL^{+}22, LMS^{+}24]. Merge [LYZ^{+}18, LSXZ23]. Merging [FXMZ24]. Mesh [BLMP18, CLWW24, GWWC25, LDZ^{+}20, YI14, ZSYW23, COM^{+}11, ZSO13]. mesh-based [ZSO13]. Meshes [NALM23, LLP06]. Message [CCH^{+}24, SL22, WXJL25]. messaging [MHT^{+}08]. Meta [GZW^{+}24a, HQF^{+}20, LYL^{+}23b, MGY22, RUD23, VLE25, ZSJL22, GZW^{+}24a, ZLT^{+}25b]. Meta-fusion [MGY22]. Meta-learning [LYL^{+}23b, RUD23, ZSJL22]. Meta-learning-based [GZW^{+}24a]. Meta-MMFNet [GZW^{+}24a]. Meta-path [HQF^{+}20]. Meta-Review [VLE25]. Metadata [MDMK06, MC19]. Metal [XZLZ23]. Metaverse [EAK^{+}25, GCTG24, GHS^{+}24, RLXW24, SCX^{+}25, TFY^{+}25, VLE25, WLS^{+}24a, XGZ^{+}24, ZXZ^{+}24a, ZPR^{+}25, ZLT^{+}25b, CDA^{+}24, SJC^{+}25]. Method [Ala21, AAT^{+}22, ASI^{+}21, BTBZ20, DLD^{+}22, FBG22, GZT21, HCJ^{+}24, IAJ^{+}24, LXJ^{+}24, NSK^{+}21, SMN^{+}22, SZB^{+}22, Tas22, WRK14, XLN^{+}21, XHZ^{+}21, LT14, RWP07, VPSS^{+}13]. Methodological [USM^{+}25]. Methodology [ACC^{+}21, RHS12]. Methods [CCD24, JLA^{+}23, NTT20, RHS^{+}20, THR^{+}22, THA^{+}25, VF25, ZWL^{+}24]. Metric [GGA^{+}20, JD23, LLC^{+}21, LWH^{+}23, LYXY23, TYY^{+}18, WLHT19, WJQ^{+}22a, ZGZ^{+}22, ZLC^{+}23, hHLC10]. Metrics [CISF^{+}25, DLL^{+}18, HLY^{+}23, REP^{+}19, RHS12]. MFECN [LZL21b]. MFGAN [XZLZ23]. Micro [GZW^{+}24a, MBXZ24, SCZ^{+}25]. Micro-Expression [SCZ^{+}25, GZW^{+}24a]. Micro-Video [MBXZ24]. Microblogging [PRPO23]. Microblogs [WCX^{+}14]. Microscopy [PA20]. Middle [ZFFX24, Swa13]. Middleware [LLP06, ZO13]. MILL [TLZ^{+}21]. Millimeter [CKRB23]. Mimicking [TMB^{+}22]. Min [HZW22, CWC10]. min-hashing [CWC10]. Min-Max [HZW22]. Mini [LMS^{+}24]. Mini-Mental [LMS^{+}24]. Mining [CWXW25, GLL^{+}24, JGZ^{+}11, LG24, QSZ^{+}21, YZX16, YP20, YSZZ14, YNLZ22, ZHL19, LWL^{+}12, SX11]. Mirror [MYX^{+}23b]. MIS [WLS^{+}24a].

Mismatch [LZH^{+}24b]. Missing [JK25]. Mix [LXC^{+}25, WZZ^{+}24]. Mix-DDPM [WZZ^{+}24]. Mix-Modality [LXC^{+}25]. Mixed [JLL^{+}22, LMLC14, LHC^{+}25, SHF^{+}24, VMP20, ZGL^{+}23]. Mixing [XW25, ZCX^{+}24]. MixOOD [YHX^{+}23]. Mixtape [FHG^{+}17]. Mixture [WZZ^{+}24, WZRD25, ZLW^{+}21]. Mixup [YHX^{+}23]. MKVSE [FHP23]. ML [AMT25, LNH23]. ML-based [AMT25]. ML-CookGAN [LNH23]. MLIC [JYZ^{+}25]. MM [ZZL^{+}25b]. MM-PCQA [ZZL^{+}25b]. MMFN [NLW^{+}21]. MMFNet [GZW^{+}24a]. MMHCL [GZW^{+}25]. MMS [YK07]. MMSUM [AE22]. MMSys [CHHH18, TB17, ZTB20, CCO15, GWZ25]. MobiCoop [SRPH16]. Mobile [AWG^{+}15, ARE13, CFP15, CKRB23, Chu15, FMIS17, HJMY15, HSN^{+}14, JLW^{+}18, KPL^{+}22, LYJ^{+}15, MZL^{+}18, MKSB17, MNPOF22, NHP^{+}16, ODMD17a, ODMD17b, PS17, RK15, SLZ^{+}21, SkFM18, SRPH16, WHF^{+}18, WWZ24, YJM^{+}19, ZYL^{+}17, Zho16, GS11a, GYN12, HH12a, HH12b, HH11b, HH11c, JYZC12, JKKL08, KS09, LCT^{+}12, LML^{+}13, RHE10, SCFL14, SNC12, SHOG12, YH13, ZC12]. MoDA [WZRD25]. Modal [CZX^{+}24, FWLA15, GJY^{+}25, GZW^{+}25, HCZ^{+}22, JCB25, LDL24, LXC^{+}24, LYDY24, LZL^{+}24c, LLN^{+}24, LDL^{+}25, LYH24, MYX23a, MAE^{+}21, SZL^{+}24, SHF^{+}24, SLL^{+}24, SLZ25, TVZ^{+}19, UJLS22, XWH^{+}21, XZC^{+}25, YTOH22, YLK^{+}20, YSSF24, YZZZ25, YP20, YTRC19, ZYO20, ZTR^{+}22, ZYW^{+}24, ZLLH25, ZZL^{+}25b, ZLJ^{+}25, ZXZ^{+}24b, ZZL25e, ZZL^{+}24d, Ano13, CDL^{+}20, CZQ^{+}22, DWL^{+}23, HLW^{+}24, JLL^{+}25, KLX^{+}24, LCY^{+}23, LD25, LZZS23, LSL^{+}20, LYW^{+}22, LYX23, LLZ^{+}23, LLL^{+}24b, MGP19, OL25, PQ19, QHFX21, Wan21, WKS^{+}23, WXCW24, WWZ24, WJQ^{+}22b, XTL^{+}21, XZLZ23, XZL^{+}25, XLS^{+}25, YBW^{+}25, YLL^{+}24, ZWH^{+}23, ZWX24, ZWY21, ZYT^{+}23]. Modalities [JK25, MZZ^{+}20]. Modality [CZ24, GNC^{+}23, HWWL20, LLG^{+}23, LYZ^{+}24, LWZH19, LZW^{+}19, LXC^{+}25, MKY^{+}24, WZZ25, WWL^{+}24, ZGZ^{+}22, ZZL^{+}24c, ZZW^{+}22, ZFFX24, AE22, LYZY21, NLN^{+}13, WWS^{+}24, WLS^{+}24b]. Modality-Aware [WZZ25]. Modality-collaborative [CZ24]. Modality-Gated [HWWL20]. Modality-Invariant [LZW^{+}19]. Mode [Hua23, LYZ^{+}18, SR22, SAZ^{+}15, WWY^{+}21, ZLK^{+}19, ZZL^{+}23, IB10]. Model [APC^{+}25, BSSNF^{+}20, BTWZ22, CWTG22, CHLW19, CZZ^{+}23, DBPM25, DBSL24, DLC^{+}22, FKCW20, FCL^{+}22, GLW20, HEA14, HZW22, IKTD25, JLL^{+}25, KUH^{+}22, LYZ^{+}18, LLSX20, LZL^{+}21a, LZZ23b, LKL^{+}25b, LQG25, LLL^{+}22, LSN^{+}20, LZL20b, LHS^{+}21, LYZZ24, LMS^{+}24, LWZ^{+}25, MAZ22, MGS18, MEA^{+}21, NLW^{+}21, QLGL24, RT14, RTR21, SG22a, SLC25, SHE21, SZZ^{+}23, SLNL20, TRRB20, WCX^{+}14, WCL23, WLZ^{+}24a, WXQC20, WZZL25, XWW^{+}25, XWL^{+}25b, YLZ^{+}21, ZYL^{+}17, ZLW^{+}21, ZSLW23, ZXZ^{+}24a, ZYW^{+}24, ZXZ^{+}25, ZSX^{+}25, ZSZ^{+}22, ZJT^{+}23, ZLF^{+}23, ZSZ^{+}23, ZFC^{+}24, ZFY^{+}25, GZW^{+}24a, GGB24, JC08, TBC^{+}11, Yan10, YGHH12, ZWM12b, DY09]. Model-Based [LYZ^{+}18, WZZL25]. Modeling [BC15, BZDX^{+}18, CXX^{+}24, CSZ^{+}25, COM^{+}11, FHH22, GWWC25, HAE11, JYZ^{+}25, KTM^{+}06, LLZ^{+}25, MON21, NLKB24, ONAGP19, SKW^{+}15, SS17, SLBS20, STSW23, SWS^{+}22, Tas20, TSP^{+}24, TNH^{+}21, WWL^{+}23, XZC^{+}25, XWL^{+}25a, YSXH16, YTR^{+}25, YBW^{+}25, YSF^{+}21, YHQH17, ZGM^{+}20, ZSO13, ZPL^{+}23, JGZ^{+}11]. Modelling [AA23]. Models [AMC^{+}18, CYZ^{+}24, DG17a, GLGG25, GGSV25, HWW25, LSCH25, NNK25, PB19, QCW^{+}25, RXY^{+}25, TXW^{+}25, WZZ^{+}24, WWY^{+}25, YLHL22, ZLK^{+}19, ZZP^{+}20, ZDD^{+}24, ZZL^{+}25c, ZHY^{+}25, ZYH^{+}24, ZBYZ25, ZHS^{+}25, DP06, GZHD12, HH08b, MCM^{+}09]. Modern [WXQC20]. Modification [WYGC24]. Modified [LL23]. Modular [SHSR24]. Modulations [YWY^{+}23]. Module [GYY^{+}25, HZZ^{+}24, LSXZ23]. MoHGCN [LD25]. Moment [GS18, LNT^{+}21, WGHX24, ZCL^{+}22, ZDQ^{+}23]. Moment-Text [WGHX24]. Moments [HD19, YTL^{+}24a]. Momentum [LD25]. Monitoring [LCC^{+}14a, HAE11, JC10, LCSX11]. Monocular [BSQ^{+}24, LYZZ24, LCZ^{+}24b, LGC^{+}24, PSN^{+}24, PSL^{+}24]. Moore [DDK24]. Mopsi [FMIS17]. MOSAIC [GHAM25]. most [Dao17]. Motion [AAPLV25, CZW15, CJZ^{+}24, Chu15, CCG20, dMGRT24, GZH17, HYLD20, HZL^{+}21, LPS15, LZL20a, LZL^{+}23, LYZ^{+}24, LXT^{+}24, LMC^{+}22, MZZ^{+}20, MSFR25, PLZW18, QXG^{+}24, WLL23, YYX^{+}24, ZWL^{+}21, ZYCD24, ZTX^{+}23, ZJL^{+}21, AP10, JP11, LZP07, LGC^{+}24]. Motion-aided [YYX^{+}24]. Motion-Aware [LYZ^{+}24, ZJL^{+}21]. Motivation [WYMX23]. Mounted [FHH22]. Mouse [CCG20]. Move [LYC^{+}25]. Movement [WSLZ25, ZZMZ21, JC08]. Movements [HPH^{+}20, NT08]. Movie [BRG24, TWY24, WXCW24]. Movies [Abd18]. Moving [Cla18, Li23, ZHL19]. MOWL [MGCH13]. MP3 [YQH12]. MPEG [MPSR05, SG07, WCK05]. MPEG-1 [MPSR05]. MPEG-4 [SG07]. MRFGNet [LJ25]. MRFs [WB16]. MRI [LYH24]. MSSA [LDZ^{+}24]. MTI [Hua23]. MulSeMedia [Ano13, GTLG14a, AdSGMS25, GTLG14b, YCGM14, MHCG19]. Multi [AYH^{+}23, AAS^{+}20, AH20, AE22, CWTG22, CKRB23, CXL^{+}22, CLZ^{+}23a, CLQQ25, CWXW25, DPL^{+}23, DYN^{+}25, DLL^{+}20, DLD^{+}22, EG17, GLWK19, GHQ24, GZW^{+}24a, GSY^{+}24, GCF^{+}21, GQG^{+}24, GZW^{+}25, HCZ^{+}22, HZC22, HLY21, HJ23, HC25, HPZ25, HGGZ23, JW21, JLZ^{+}21, JLL^{+}25, JMLL20, JYZ^{+}25, JXTC21, JLL^{+}22, LYZY21, LCY^{+}23, LZL^{+}21a, LFP^{+}22, LZWC23, LZL^{+}23, LDL24, LYZ^{+}24, LXC^{+}24, LZZ^{+}24a, LKL^{+}25b, LZZS23, LTD^{+}21, LPC^{+}18, LSL^{+}20, LZL21b, LYW^{+}22, LYX23, LLZ^{+}23, LYL^{+}23b, LNH23, LYW^{+}23b, LZL^{+}24c, LDZ^{+}24, LSS^{+}24b, LLL^{+}24b, LCL^{+}24b, LGW^{+}25, LZXP25, LYH24, MYX23a, MGS18, MSKYJ21, MAGT23, MEA^{+}21, MGP19, NAK15, NDX^{+}21, NDC^{+}23, NRUT20, OL25, PS17, QHFX21, QLGL24, RXY^{+}25, SSP21, SZZY20, SGS21, SHIE15, SJC^{+}19, SG22b, SLNL20, SLL^{+}24, TSZ^{+}23, TYY^{+}22, TH22, TXWK24, TVZ^{+}19, UJLS22, WCF^{+}17, WYM18, Wan21, WXW^{+}22, WLZF22, WCL23, WHL^{+}23, WWS^{+}24, WXCW24, WLS^{+}24a, WWC^{+}24, WWZ24, WWQ^{+}25a, WQTX25, WC23, WZZ^{+}22, XXG^{+}21]. Multi [XZLZ23, XLS^{+}25, YV23, YML^{+}22, YBW^{+}25, YMY^{+}21, YP20, YYBX24, YCFX23, YCZ^{+}23, YTL^{+}24b, ZWX24, ZX14, ZWL^{+}17, ZJZC20, ZTR^{+}22, ZSLW23, ZTQX23, ZYCD24, ZYW^{+}24, ZSX^{+}25, ZYJZ25, ZZL^{+}25b, ZZW^{+}19, ZJT^{+}23, ZXS^{+}24, ZWZ^{+}23a, ZLF^{+}24, ZWR^{+}20, ZHC^{+}23, ZXD^{+}23a, ZXD^{+}23b, ZLL^{+}24c, ZXZ^{+}24b, ZLT^{+}25b, ZCS^{+}20, ZZW^{+}22, ZWZ^{+}23b, ZPL^{+}23, ZZL25e, Ano13, NLN^{+}13, YWN^{+}10a]. Multi-Agent [TSZ^{+}23, MEA^{+}21]. Multi-Anchor [LKL^{+}25b]. Multi-Attention [GSY^{+}24]. Multi-Attribute [DLL^{+}20, HC25, JLL^{+}22]. Multi-Avatar [SHIE15]. Multi-branch [TH22]. Multi-Camera [NAK15]. Multi-caption [AYH^{+}23]. Multi-CDN [AH20]. Multi-Class [LYL^{+}23b, WCF^{+}17]. Multi-Concept [MGS18]. Multi-contextual [SGS21]. Multi-Cue [ZWL^{+}17]. Multi-dimensional [ZSX^{+}25]. Multi-disciplinary [ZCS^{+}20].

Multi-domain [CLQQ25, LTD^{+}21].

Multi-encoding [MAGT23]. Multi-Event [ZX14]. Multi-expert [ZWZ^{+}23b]. Multi-Feature [YML^{+}22, CXL^{+}22, NDC^{+}23, WXW^{+}22]. Multi-Food [PS17]. Multi-fusion [NRUT20]. Multi-Grained [HCZ^{+}22, LGW^{+}25, QLGL24, YTL^{+}24b, ZYJZ25]. Multi-Granular [DPL^{+}23, LFP^{+}22]. Multi-Granularity [OL25, WLZF22]. Multi-Guidance [HGGZ23]. Multi-human [LZL^{+}21a]. Multi-Identifier [WLS^{+}24a]. Multi-instance [CLZ^{+}23a, JW21, JLZ^{+}21, SZZY20]. Multi-Label [LNH23, ZHC^{+}23, ZXD^{+}23a, ZXD^{+}23b, JW21, RXY^{+}25, YBW^{+}25, YCZ^{+}23]. Multi-layer [AAS^{+}20, NDX^{+}21]. Multi-Lead [ZLT^{+}25b]. Multi-Level [CWXW25, TXWK24, WC23, LZL21b, LCL^{+}24b, SJC^{+}19, YV23, ZWZ^{+}23a, ZLF^{+}24]. Multi-Media [ZLT^{+}25b]. Multi-Modal [GZW^{+}25, LDL24, LXC^{+}24, LZL^{+}24c, LYH24, MYX23a, SLL^{+}24, TVZ^{+}19, UJLS22, ZTR^{+}22, ZYW^{+}24, ZZL^{+}25b, ZXZ^{+}24b, ZZL25e, JLL^{+}25, LCY^{+}23, LZZS23, LSL^{+}20, LYW^{+}22, LYX23, LLZ^{+}23, LLL^{+}24b, MGP19, OL25, QHFX21, Wan21, WXCW24, WWZ24, XZLZ23, YBW^{+}25, ZWX24, Ano13]. Multi-Modality [LYZ^{+}24, ZZW^{+}22, AE22, LYZY21, WWS^{+}24, NLN^{+}13]. Multi-model [GZW^{+}24a]. Multi-Object [GQG^{+}24, HPZ25, WHL^{+}23, XXG^{+}21, LZL^{+}23, YYBX24]. Multi-Objective [WZZ^{+}22]. Multi-page [YMY^{+}21]. Multi-party [SSP21, YWN^{+}10a]. Multi-peak [JLZ^{+}21]. Multi-perceptual [LYW^{+}23b]. Multi-Person [LZXP25]. Multi-Purpose [SG22b]. Multi-Reference [JYZ^{+}25]. Multi-Representation [LDZ^{+}24]. Multi-Sample [DLD^{+}22]. Multi-Scale [WWQ^{+}25a, ZWR^{+}20, HZC22, HJ23, LZWC23, TYY^{+}22, WQTX25, XLS^{+}25, YP20, ZJZC20, ZTQX23]. Multi-sensor [GLWK19]. Multi-Source [MYX23a, WCL23, LSS^{+}24b]. Multi-Stage [ZYCD24]. Multi-Task [WYM18, WWC^{+}24, CWTG22, GCF^{+}21, JXTC21, ZSLW23, ZZW^{+}19, ZJT^{+}23]. Multi-Tier [MSKYJ21]. Multi-User [EG17, CKRB23]. Multi-Variate [LPC^{+}18]. Multi-View [DYN^{+}25, HLY21, LZZ^{+}24a, SLNL20, TVZ^{+}19, ZXS^{+}24, AE22, GHQ24, JMLL20, WQTX25, YCFX23, ZPL^{+}23]. Multicast [CNG22, LHZ^{+}23, WCO15, GL11, GL12, YC08]. multicasting [HH12a]. Multichannel [LKM17, ZXY^{+}20]. Multichannel-Kernel [LKM17]. multidepth [KMP05]. Multidimensional [Tas20, Tas24]. Multifeature [HSZ^{+}18, YZXY15, ZI13]. MultiFusion [WK10]. Multigranularity [LLZ^{+}25, MRS11]. multihop [HNL08]. multilabel [CML^{+}13, LMLC14, QHR^{+}08]. Multilayer [AS22a, PZ08]. Multilayered [LCC^{+}14a]. Multilevel [CZY^{+}21, WJ15b]. Multilingual [BBB^{+}23]. Multimedia [AMMG16, AC19, AMT25, BSH08, BXMH15, BLJX10, CDGJ09, CXW^{+}19, CHHH18, CLS^{+}21, CW10, CF22, DBSL24, FSEW25, GLL^{+}24, GLGG25, GYF^{+}21, GPHOH12, Geo05, GJAF18, Han13, HSN^{+}14, HCM^{+}22, HT15, HGKS24, JPS05, JJML24, KEYY22, KJJ^{+}21, LQH18, LWZ^{+}21a, LQS21, LS21, LYXA22, MSG22, MLQM14, NCMM21, NPG^{+}22, ODMD17a, ODMD17b, PSS05, RHAG21, RPE^{+}17, SMK^{+}24, SP21, DMSRL18, SWS21, SZEST21, She13, SLYE11, SKC24, SKVHC18a, SKVHC18b, SKA24, TB17, TNEcC08, VNDB24, WJXN24, WKST08, WCX^{+}14, WRK14, WFZ^{+}21a, WDE25, WAC^{+}25, XSEZ10, XFQ^{+}21, YK07, YNC18, YXZQ22, YZRW22, YWF^{+}24, ZZMS14, ZYL^{+}17, ZYZE19a, ZYZE19b, ZCY^{+}19, ZWS^{+}20, ZJSJ20, ZZL^{+}24b, ZLX^{+}25, ZTB20, AG13, ATM06, AKO07, Bag11, BH05, BCG13, CCK06, Cha13, DY09, Eff13, ETF06, GS11a, GDGC07, GA12a, GA12b, GFB^{+}14, GG06, Hae10, HNL08, HM10]. multimedia [HAE11, HNS13, JKKL08, Kan12, KZHC13, LSDJ06, LS05, LWL08, MGCH13, OMP07, PZ08, PS05, RWW05, RJ05, Row13, SKSZ13, SG07, SX13, SSTK07, SHOG12, SOC^{+}13, SKR09, SLKS12, WDCX07, WK10, WZC^{+}13, WBL09, Yan10, YC08, ZCY^{+}13, NLS13]. Multimedia-enabled [LYXA22]. Multimedia-Rich [WCX^{+}14]. Multimodal [APRS^{+}19, AMC^{+}25, CCF^{+}24, CHHH25, CYM^{+}25, CSZ^{+}25, CXC^{+}17, DNPG^{+}17, EAK^{+}25, FHP23, GTLG14a, HYG^{+}25, HQF^{+}19, HYG^{+}21, HCJ^{+}24, JLA^{+}23, JK25, KWL^{+}25, KBI^{+}23, LLC^{+}24a, LTD^{+}21, LLC^{+}24b, LHJ^{+}24, LSX^{+}25, LXZ^{+}25, LGF^{+}14, LHF^{+}14, MBK^{+}23, MXH^{+}23, MZGY17, MHL^{+}25, NLW^{+}21, PSSZ24, QZXH14, RHAG21, SHSR24, SKC24, SZF^{+}25, TNZ^{+}25, TVK18, UMM24, WCLC18, WFZ^{+}21a, WMM^{+}23, WDR^{+}24, WZZ^{+}23, XZH^{+}21, XLX^{+}24, XZY^{+}25, YV23, YNC18, YCL25, YHQH17, YSY^{+}22, ZZX^{+}20, ZLJ^{+}25, ZXT^{+}25, And13, MVW08, MHT^{+}13]. Multimodality [MAX^{+}24, SYM^{+}25, WLC^{+}20]. Multinomial [CHLW19]. Multiparty [DMF17, YWN^{+}10b]. Multipath [KFP^{+}25, WWGT09]. Multiperson [WTZ^{+}20]. Multiphase [EC16]. Multiplanar [LLSX20]. Multiplatform [TVK18]. Multiplayer [XW17, ZZL21]. Multiple [ASLA18, DZW^{+}21, DLL^{+}18, DZWH23, GTLG14a, GZL^{+}20, JLL24, JGJ^{+}20, KPK^{+}24, KP15, LPS15, MGS18, MLQM14, NRUT20, SWS21, TLZ^{+}21, TSP^{+}24, WZ20, WLS^{+}24b, WWX^{+}21, YW23, ZWX24, ZZZ14, ZGM^{+}20, ATM06, NT08, SVF12, TNEcC08, Ano13]. multiple-choice [NT08]. Multiple-hypotheses [JLL24].

Multiple-Scent [MLQM14]. Multiple-Source [JGJ^{+}20]. Multiplexed [GHAM25]. Multiplexing [CXC^{+}17, HH11b]. Multiplicative [SLBS20]. multirelational [LCSX11, LSDK12]. multiresolutional [LW07]. Multiscale [LJ25, LZG^{+}25]. multisensor [HAE11]. Multisensorial [LGF^{+}14]. Multisensory [SVA^{+}21]. Multisource [BQBLN18]. Multispectral [CJHH21]. Multitarget [AY21]. Multiview [EC16, GLL^{+}24, HH12a, RHS12]. Mural [XZY^{+}25]. MuralAgent [XZY^{+}25]. Museum [NHP^{+}16]. Music [BRG24, CFP15, CYM^{+}25, HDZ^{+}15, SY09, TWB^{+}23, WWHW14, WJ15b, YSY^{+}22, YTRC19, ZWC^{+}22, DCM13, FLM^{+}06, HCS12, KS13, LYC11, MSL10, ML11, SXM^{+}06, TBC^{+}11, Whi13, XMST07]. Music-Driven [ZWC^{+}22, SY09]. Music2Dance [ZWC^{+}22]. Musical [LLJW15, ONH07, ZCAP08]. Musicalization [TNP^{+}18]. Mustang [YZL^{+}24]. Mutual [LZL^{+}24b, ZHY^{+}25]. Mutually [LXC^{+}24]. Mutually-Guided [LXC^{+}24]. MV2Flow [HZL^{+}21]. My [GEL^{+}15].

N [FBGD23]. Names [YMJ^{+}25]. Narrated [ICZ^{+}22]. narrative [JSL07]. narrative-based [JSL07]. narratives [UTK^{+}08]. NAS [Che24]. Native [yHcCzH^{+}21]. Natural [AMC^{+}25, FSEW25, HYW^{+}25, LYLL23, LGLZ20, XYC^{+}23]. Navigate [FHG^{+}17]. Navigating [SSS13]. Navigation [CLP^{+}23, HZL^{+}25, PSN^{+}24, TJN14]. Near [LQZH14, MTTH13, ZHLY11, DCO10]. Near-Duplicate [LQZH14, ZHLY11, DCO10]. Near-lossless [MTTH13]. Nearest [KVB^{+}24]. Needs [GQG^{+}24]. Negative [PYZ25, YBO14]. Neighbor [JLL24, KVB^{+}24]. NeRF [PYBY25]. Nested [LJZ^{+}22].

Nested-Residual [LJZ^{+}22]. Net [BV25, HRL^{+}24, LZYY24, CH21, CLQQ25, FHF^{+}24, HZS^{+}24, PA20, WWL20, YXYB21, ZRZ^{+}21]. Nets [WFC^{+}25a]. Network [AYH^{+}23, AM24, BRZS18, CZL^{+}23a, CW25, CPP^{+}14, CFN^{+}23, CGPCR18, CHHH18, CDL^{+}20, CH21, CZZ^{+}24, CCH^{+}24, CLQQ25, CZQ^{+}22, CSSZ23, CDZ^{+}17, CSŞ^{+}25, DHP23, DLO^{+}20, DLL^{+}20, EC16, FZYW20, FYW^{+}24, FNH22, FFCM24, GJY^{+}25, GNC17, GZW^{+}24a, GHR^{+}22, GHY^{+}24, GQG^{+}24, HCZ^{+}22, HCM^{+}25, HJ23, HL24, HC25, HQF^{+}19, HLZ^{+}21, HZPL21, HYLG23, HDH^{+}24, HQF^{+}20, HSG23, HWW^{+}23, HCY^{+}23, JLL^{+}21, JWH21, JWW^{+}24, KPK^{+}24, KCC17, KWL^{+}25, KMK^{+}21, LSK^{+}15, LLJ^{+}20, LCL^{+}21, LYL^{+}21a, LWY22, LFP^{+}22, Li23, LSXZ23, LZL^{+}23, LLS^{+}24, LDL24, LZL^{+}24a, LLL^{+}25, LD25, LJ25, LZZS23, LZZ^{+}24b, LK18, LZD^{+}21, LHL^{+}24, LZZ^{+}25, LZC^{+}19, LZL21b, LNT^{+}21, LWZ21b, LWP22, LKW^{+}22, LLZ^{+}23, LNH23, LYW^{+}24, LZZ^{+}24c, LCZ^{+}24b, LNC^{+}25, LXZ^{+}25, LWH20, LS21, LYH24, MYX23a, MLJ^{+}22, MAZ22, NWNL21, NXL^{+}25, PSSZ24, RNR^{+}22, RBLJ25, RBBS25, RLY^{+}21, SED25, SR22, SGS21, SBS23, SJC^{+}19, SL23, SL22, SZL^{+}22, SHSR24, SK23, SGY^{+}24, TWY24, TWFW24]. Network [TH22, TCL^{+}23, TNZ^{+}25, TB17, TMB^{+}22, TP22, UJLS22, WZ20, WLZF22, WDPX23, WLTL23, WHL^{+}23, WGL^{+}23, WWS^{+}24, WSLC24, WLSX24, WYMX23, WWX^{+}21, WC23, WZZ25, WHY19, WJQ^{+}22b, WLL^{+}24, XJFS25, XXG^{+}21, XPP^{+}23, XHL^{+}21, XLZ^{+}21, XTL^{+}21, XLZ^{+}22, YV23, YSXH16, YZS^{+}23, YXYB21, YXZQ22, YZS^{+}22, YLLL23, YHH24, YCL25, YLL^{+}24, YYY^{+}24, YFLF19, YBTX23, YCZ^{+}23, ZY21, ZCXL23, ZWX24, ZHS20, ZSS20, ZJZC20, ZRZ^{+}21, ZZL^{+}24a, ZZB24, ZZW^{+}24, ZLJ^{+}25, ZLT^{+}25a, ZYLN21, ZSZ^{+}24b, ZLL^{+}22, ZHC^{+}23, ZXD^{+}23b, ZGY^{+}24b, ZLL^{+}24c, ZLT^{+}25b, ZHL^{+}25, ZZW^{+}22, ZWZ^{+}23b, AH12, CPP^{+}13, CXS^{+}08, GL12, LWL08, TC08, WLZ12]. Network-Assisted [BRZS18, KCC17, CDZ^{+}17]. Network-Aware [GNC17]. Network-Based [CGPCR18, LLL^{+}25, SED25, UJLS22]. Network-Driven [YXZQ22]. network-wide [WLZ12]. networked [MCM^{+}09, ZWM12b]. Networking [FZW^{+}25b, GZ20, MY15, FKFB05]. Networks [APC^{+}25, AY21, BOZ17, CXW^{+}19, CZZ^{+}23, CPL^{+}23, CZW^{+}25, CSŞ^{+}25, FSK^{+}15, FLG^{+}21, GYY^{+}25, HWWL20, HZZ^{+}24, KPL^{+}22, KEYY22, KJJ^{+}21, LCL14, LB15, LZJ^{+}20, LLJC21, LYW^{+}22, LZW^{+}24, LYL^{+}25, NPG^{+}22, NWL^{+}20, NLW^{+}21, PLYM23, PQ19, PRPO23, QHFX21, QSY^{+}23, QJ23, RUD23, SMTR22, SSK20, SLBS20, SGYX22, SYX23, SJC^{+}25, SZZT18, TZLZ21, WHJ20, WDX^{+}23, WKS^{+}23, WWY^{+}24, WF25, WGM25, XWY21, XFZ^{+}19, XLH^{+}23, YLCC18, YZL^{+}24, ZQRS18, ZYO20, ZH18, ZWZL22, ZDQ^{+}23, ZLH16, ZQKH19, ZZL25e, BDV08, BRA^{+}09, GS11b, HH12b, HH11a, HH11c, LS05, LLP06, LCK09, LCSX11]. Neural [AM24, AY21, CZZ^{+}23, JLL^{+}21, JLLL25, KMK^{+}21, LYL^{+}21a, LLL^{+}25, LZC^{+}19, LSN^{+}20, LWZ21b, LWP22, LYW^{+}22, LXZ^{+}25, MAZ22, PYBY25, QJ23, SR22, SBS23, SWZ25, SLBS20, SZM^{+}19, SK23, TMB^{+}22, TP22, WZ20, WMW^{+}22, WDX^{+}23, WFC^{+}25a, WWX^{+}21, WF25, XWY21, XHL^{+}21, XLZ^{+}22, ZYO20, ZZL^{+}24a, ZXZ^{+}25, ZYLN21, ZQKH19, LWL08]. Neural-Network-Based [LYL^{+}21a, LWL08]. Neurosymbolic [UMM24]. News [KWL^{+}25, LLO^{+}20, QHFX21, LLW^{+}13, OBBW12, SWH06]. Newsroom [LLO^{+}20]. Next [NLKB24, SSTK07, She13]. Next-generation [NLKB24]. NextSlidePlease [SLKS12]. NFT [WFZ25, ZXZ^{+}24a]. Nicolas [Ste10, Ste12c]. NIR [YHH24]. NIR-VIS [YHH24]. NMF [HZPL21]. NMF-Aware [HZPL21]. NMS [ZMX^{+}23]. No [HLY^{+}23, JD23, LYXY23, LYW^{+}23b, YTL^{+}21, YDCZ23]. No-Reference [HLY^{+}23, JD23, YTL^{+}21, LYXY23, LYW^{+}23b, YDCZ23]. Node [LJG^{+}24, LZZS23]. nodes [JC10]. Nodule [TLZ^{+}20]. Noise [CZW^{+}24, LYL23a, LZY^{+}24, MBXZ24, RSE16, WZZ^{+}24, YMA17, YLLL21]. Noise-Resistant [MBXZ24]. Noise-Tolerant [LZY^{+}24, LYL23a]. Noisy [LZY^{+}24, RUD23, ZSJL22, ZLN^{+}16, WDCX07]. nominations [Ste12c]. Non [CZZ^{+}24, FBSGR25, GCTG24, LGX^{+}24, LGM^{+}24, THR^{+}22, WHL^{+}21, YPSC22]. Non-Acted [THR^{+}22]. Non-Aligned [WHL^{+}21]. Non-autoregressive [CZZ^{+}24]. Non-Dominated [LGX^{+}24]. Non-fungible [GCTG24]. Non-Iterative [LGM^{+}24]. Non-Mask [YPSC22]. Non-Uniform [FBSGR25]. noncontinuous [ZO13]. Nonlinear [RSE16]. Nonlocal [HZL^{+}16, LZZ^{+}24b]. nonrigid [ZHLY11]. Norm [YLZ^{+}21]. Normal [LGC^{+}24]. Normalization [LZT^{+}20]. Normalized [AAT^{+}22]. NOSSDAV [CHHH18, TB17, CCO15, GWZ25]. Note [Ste14, Ste12b, Ste12c, Ste13a]. notice [Ste11]. Noticeable [TWKK21, XW17]. Novel [BQBLN18, DLO^{+}20, DLD^{+}22, GDM^{+}23, HLD18, HWW25, IAJ^{+}24, JD23, JXLD25, LZW^{+}21, LLP^{+}23, QWH^{+}21, SWZ25, SB23, UJLS22, WWZ^{+}25, XLZ^{+}23, XLN^{+}21, ZSLW23, ZSZ^{+}23, LC12]. NR [LJZ^{+}22]. NR-CNN [LJZ^{+}22]. Nuclear [YLZ^{+}21]. Nuclei [PA20, TS20]. Nuisance [KDC08]. Number [AYH^{+}23, Ano11, Ano12, Ano14, Ano20, Sha21, SRAA17, TOM12].

Number-controlled [AYH^{+}23]. NumCap [AYH^{+}23]. Numerically [JLL^{+}22]. nursing [CYMW07].

O [TS20, FMIS17, RWW05]. O-Mopsi [FMIS17]. Obfuscation [CFG^{+}24, TFY^{+}25]. Obfuscation-Based [TFY^{+}25]. Obituary [Ste10]. Object [CW25, CHLW19, CYZ23, CYZ^{+}24, CCH^{+}24, CSBJ24, GGSV25, GQG^{+}24, HZZ^{+}24, HPZ25, HGGZ23, JLZ^{+}21, Li23, LDL24, LSL^{+}24, LG24, LTX^{+}23, LDZ^{+}24, LWH20, NWL^{+}20, NXL^{+}25, PLYM23, PWHP24, QLSQ12, QDX^{+}24, QSY^{+}23, QLW^{+}24, SZM^{+}21, SMN^{+}22, SYY^{+}24, SZC24, SZTL16, WJ15a, WXW^{+}22, WHL^{+}23, WWS^{+}24, WLSX24, WF25, WJQ^{+}22a, WZL^{+}24, XXG^{+}21, XPP^{+}23, XWW^{+}21, YTOH22, YZS^{+}22, ZWLZ24, ZWL^{+}24, ZZW^{+}24, ZYZ^{+}25a, ZWZ^{+}23a, vRPPP23, LS05, LZL^{+}23, TCJ08, YYBX24, ZG08]. Object-Aware [ZYZ^{+}25a]. Object-based [QLSQ12, ZG08]. Objective [LXJ^{+}24, SMN^{+}22, WZZ^{+}22, ZSZ^{+}24a, RHS12]. Objectness [WLL^{+}19]. Objects [dMGRT24, LMX24, WLL^{+}19, VV11]. Observation [LLCH17]. Observers [TMB^{+}22]. Occluded [FHK^{+}23, PWX^{+}24]. Occluders [CLZ^{+}23b]. Occlusion [GZW^{+}24b]. Occlusion-Free [GZW^{+}24b]. Occlusions [CF22, Hon19]. Occupancy [CWL24]. Occupation [CC17]. Occurrence [ZCY^{+}19]. occurring [LYW^{+}25]. OCR [ZZC^{+}15]. octave [ML11]. Octree [QLGL24, YH14]. Octree-Based [YH14]. Odor [KYVE14]. Offering [SSP^{+}24, YC08]. offline [NH10]. Offloading [XHZ^{+}21, YXZQ22]. Offset [LKL^{+}25b, QLW^{+}24, WZZ^{+}24, ZLOL18]. offsets [YQH12]. Olfaction [AMMG16, AG13, GA12a, GA12b]. Olfaction-Enhanced [AMMG16, AG13, GA12b]. Olfactory [HC22]. Olfactory-based [HC22].

OmniArt [SW18]. Omnidirectional [LYW^{+}23b, YWC^{+}25]. Omniscient [YWL^{+}24]. On-Demand [EL25, HSN^{+}14, GSM^{+}08, QS10, SAAH10]. On-Device [JLL^{+}25]. On-Screen [BFAS15]. On-the-Spot [VVSV17]. One [HXH^{+}24, PYBY25, SGYX22, WLC^{+}20, WY23, ZLS^{+}25, CL07, JXTC21]. One-Bit [HXH^{+}24]. One-shot [PYBY25]. One-step [ZLS^{+}25]. One-Stop [SGYX22]. Online [ABR17, Ano11, Ano12, Ano14, Ano20, BBZ18, BTWZ22, CZC15, CSSZ19, CSSZ23, CZZ21, GZLW18, HLW^{+}24, HLW^{+}21, Hua13, JSEI16, JWF18, KLX^{+}24, KPL^{+}22, KRKK14, LSK^{+}15, LZL^{+}23, LZW^{+}24, MVW07, Sha21, SZZT18, SAL21b, TOM12, TAS16, WHL^{+}23, XW17, ZDZ^{+}22, ZLP^{+}14, ZZL21, BRA^{+}09, CXS^{+}08, Dao17, LLKL11, LYC11]. online-only [Dao17]. Only [YMJ^{+}25, Dao17]. ontology [BAK13, FLM^{+}06, MGCH13]. Open [JXLD25, LWH^{+}23, LGW^{+}25, OMP07]. Open-Set [JXLD25]. Open-Vocabulary [LGW^{+}25]. Open-world [LWH^{+}23]. Operating [CHHH18, DDK24, GEL^{+}15, TB17]. Operations [CXX^{+}24, SZHY19, Zha19]. Opinion [PN16]. Opportunistic [HSL^{+}22]. Opportunities [BLH^{+}25, RSAH25, UTK^{+}08]. optics [CKRB23]. Optimal [GLL^{+}24, GL11, TAPP^{+}15, AKO07, HB08, WLQL12]. Optimization [AAS^{+}20, BWP^{+}24, CXW^{+}19, CZZ^{+}24, EWSZ15, FZW^{+}25b, GHAM25, HXZ^{+}20, LLSX20, NW08, PYZ^{+}20, SSSA23, TRRB20, WCY^{+}18, WHW^{+}21, WZZ^{+}22, YWF^{+}24, ZZL21, ZLL^{+}24b, FLZ^{+}12, FYH10, GL12, HH11a, LCK09]. Optimized [HLM^{+}22, HLY21, LSCH25, TP22, WZZ^{+}23, HH08a, YK07]. Optimizing [HHH22, MMW10, WBB^{+}17, ZGWZ23, TCW^{+}13, AH12]. Optional [WLSX24]. Orchestrating [HJWW19]. Orchestration [CZZ21]. Order [XW17, YHQH17, YZY^{+}25, ZGM^{+}20, ZGD^{+}19, ZXS^{+}24, ZCT^{+}07, ZWF^{+}20]. Ordering [YCL25]. Orders [ZQKH19]. Ordinal [ZGR21]. Organism [HZZ^{+}24]. Organization [LMF^{+}14, LWWZ20, LLW^{+}13]. Organizing [WZTL19, SX11]. Oriented [FFCM24, NSK^{+}21, PRPO23, SLMJ24, TAS16, YXJ^{+}24, ZWZ^{+}23a, AKO07, BBUD24, JSL07, LCT^{+}12, LGX^{+}08]. Orienteering [FMIS17]. ORL [BBZ18]. Orthopedic [CGPCR18]. OSNs [LZW^{+}24, SZZT18]. OTT [FAA18]. OTT-ISP [FAA18]. our [Ste10]. Out-of-distribution [YHX^{+}23]. Outlier [LLCH17, DCC^{+}13]. Outline [AA23, PKL25]. Outlook [CS22, SLC25]. Outpainting [XZY^{+}25]. Output [WJS^{+}21, ETF06]. Output-Bounded [WJS^{+}21]. Outsourced [HZL^{+}16, YMA17]. Outsourcing [YWF^{+}24]. Over- [HSR18]. Overhead [CPSH14, CSBJ24]. overlapping [ZCY^{+}13]. Overlay [RXC14, LLSC12, VGNL10, YC08]. Overview [HGKS24, SGW08].

P [BLW^{+}24]. P2P [BQBLN18, CSJ^{+}08, GGB14, LLSC12, MN16, SZ12, WLQL12, WLZ12, YLZ^{+}10, ZCL^{+}12, ZSO13]. P2P/ [BQBLN18]. Paced [WZD^{+}20, XW17]. Pacing [ZQRS18]. Packet [JTZ^{+}16, CLC05]. packet-loss [CLC05]. Packing [SK23]. PADA [PWY^{+}25]. page [HCW^{+}07, HZW^{+}11, YMY^{+}21]. Paillier [SZHY19]. PAINT [GHW^{+}24]. Pair [Lin15]. Paired [YP15]. Pairing [WSX^{+}24]. Pairs [WZZ^{+}22]. Palm [CJHH21]. Palm-Vein [CJHH21]. PANet [LZL^{+}23]. Panoptes [FKFB05]. Panoptic [LLL^{+}24a]. Panorama [PB14, VF25]. Panoramic [WQTX25, XLF^{+}25, YTR^{+}25, ZCHD23, ZZCZ21, DE12]. Panoramic-Image-Based [XLF^{+}25]. Pansharpening [SR22]. paper [SLYE11, Ste12c]. Papers [CHHH18, HT15, TB17, ZZMS14, ZTB20, Ano13, BLJX10, CDGJ09, GPHOH12, GS11a, KZHC13, PSS05, XSEZ10]. PAPooling [WXS^{+}25]. Paradigm [GLGG25, LXC^{+}25, MZL^{+}18, YWC^{+}25, CTGP08]. Parallel [WZWY23, ZWL15, EGEM06, GFB^{+}14]. Parameter [WZRD25, ZBYZ25]. Parameter-efficient [WZRD25]. Parameters [SG22a]. Parametric [DG17a]. Paraphrase [XCC^{+}23]. Parents [AMC^{+}25, ONH07]. Pareto [LGX^{+}24]. Parking [SSZ^{+}23, SZZZ25]. Parkinson [MEA^{+}21, SLL^{+}21]. Parsing [LZL^{+}21a, MHW^{+}19, MLJ^{+}22, WB16]. Part [DVA21, LYZX21]. Part-based [LYZX21]. Part-wise [DVA21]. Partial [KP08, RXY^{+}25, ZWYS20, ZLLT13]. Partial-Duplicate [ZWYS20, ZLLT13]. Partially [NNK25, ZYJZ25]. Participant [YLS^{+}23]. Participatory [HJMY15]. Partition [LZW^{+}21]. Partitioning [YGMY24]. partner [HCKL13]. Parts [HJZ^{+}23, LLL^{+}22]. party [SSP21, YWN^{+}10a]. Passing [WXJL25]. passive [CCD07]. Past [Che24, dB16, BCG13, Hua13]. Patch [JCSL19, YNY25, ZNNH25]. Patch-discriminative [ZNNH25]. Patches [ZCD15]. patching [HB08]. path [HDH^{+}24, HQF^{+}20, YW23, ZZG^{+}20]. Patient [WDS21]. pattern [WT10, ZO13]. Patterns [BBL^{+}24, BR22, WPRC18]. PAV [ZCHD23]. PAV-SOD [ZCHD23]. pay [YH13]. pay-TV [YH13]. Paying [CBSC18, EFF^{+}25]. Payoff [Ala21]. Payoff-based [Ala21]. PCC [LSS^{+}24a, RSR25]. PCQA [ZZL^{+}25b]. peak [JLZ^{+}21]. Peaks [YP15]. PEAR [LLC^{+}24a]. Pedestrian [KBB21, LLZ^{+}24, ZZCZ21]. Pedestrian-Aware [ZZCZ21]. PEDM [ZJT^{+}23]. peer [CSJ^{+}08, CL12, JC10, LH12, TSHP05, VGNL10, WLZ08]. peer-based [JC10]. peer-to-peer [CL12, JC10, LH12, TSHP05, VGNL10, WLZ08]. People [MBK^{+}23, ZZL^{+}17]. Perception [CCF^{+}24, GCF^{+}21, KYVE14, MBK^{+}23, QXG^{+}24, SCZ^{+}25, SJC^{+}19, TMB^{+}22, ZXZ^{+}24b, GA12b, GG06, KTM^{+}06, SOC^{+}13]. Perceptron [AAS^{+}20]. Perceptual [CZZ^{+}23, LKL25a, LGLZ20, SZLL17, TWKK21, YXJ^{+}24, ZSMZ21, ZLX^{+}25, KK08, LYW^{+}23b]. Perceptually [WHZ^{+}24]. Performance [AdSGMS25, BTBZ20, CDZ^{+}17, TCP^{+}20, yHcCzH^{+}21, IKTD25, KABB20, LLL23, MKSB17, WKN25, ZGWZ23, BBT^{+}05, CBR14, WKST08, WWGT09]. Performances [WSLM18, ONH07, ZCAP08]. Perfusion [RLAM25]. Periodontal [CJP^{+}21]. peripheral [DÇ07]. Permutation [CQA^{+}25]. Person [AQL^{+}20, CHS^{+}23, CZL23b, DFXY20, DSL^{+}25, DLL^{+}18, EFF^{+}25, FZYY18, FZZ^{+}25, GLT^{+}20, GHZW25, HLM^{+}22, HZH24, HCZ^{+}25, HJZ^{+}23, LZJ^{+}20, LYZX21, LLS^{+}24, LKM17, LZC^{+}19, LWH^{+}23, LZL^{+}24b, LXC^{+}25, LZXP25, LWHC24, LCZ^{+}24c, PZJL22, PKL25, PWX^{+}24, QWH^{+}21, SED25, SJC^{+}19, TCS^{+}25, TH22, TGC23, WLC^{+}20, WDPX23, WZRD25, WWQ^{+}25a, WHW^{+}21, WJQ^{+}22b, WZW24, XJFS25, XQG^{+}23, XQG^{+}24, XGX^{+}25, XWH^{+}25, XHL^{+}24a, XLZ^{+}22, XYW^{+}22, XLH^{+}23, XCM^{+}23, YWH^{+}17, YZYX24, YJTN18, YLYB24, ZGZ^{+}22, ZZJ^{+}23, ZSZ^{+}22, ZZY18, ZZZ^{+}22, ZGWZ23, AH12, LCL^{+}24b]. Persona [ZJT^{+}23]. Persona-aware [ZJT^{+}23]. Personal [GLWK19, MZL^{+}18, Eff13]. Personality [HSX22, ZGD^{+}19]. Personality-Aware [ZGD^{+}19]. Personalization [CED^{+}16, SSTK07]. Personalized [DYSX14, HCWM14, LXL^{+}18, LZL^{+}24b, SWS^{+}22, WWHW14, WCL^{+}22, YBO14, ZGD^{+}19, WWL13]. PersonArt [CTBC22]. Perspective [BLH^{+}25, CF22, GDM^{+}23, SAL^{+}21a, XGX^{+}25, XXG^{+}21, ZWL^{+}24, DCO10, FPH^{+}08, MDMK06, ZWM12a]. Perspectives [GTLG14b, RABC24]. Perturbation [QDX^{+}24, SMTR22, TZLZ21]. Perturbation-enabled [SMTR22]. Perturbations [WLCH22]. Pervasive [MY15, NSJB17, WGM25]. PGNet [NWNL21]. phase [LYL23a]. PhD [Ste10]. Photo [BC15, FKW22, GHW^{+}24, LLH^{+}24, WWL20, ZLN^{+}16, CFGW05, RSB11]. Photo-Realistic [WWL20, GHW^{+}24]. Photograph [YBO14]. Photographs [TAS16, BSS11]. Photography [RK15, ZLS^{+}25]. Photometries [SSK20]. Photorealistic [MHW^{+}19]. Photos [SAL^{+}22, SSS13]. Phrases [PRH14, ZG08]. Physical [LZDL25, TFY^{+}25, YNY25, RHE10]. Physiological [HCJ^{+}24, ZGD^{+}19, ZWF^{+}20]. piano [ONH07]. picture [KS09]. Pictures [WCX^{+}14]. Picturing [JWL06]. Pieces [TWB^{+}23]. Pilot [WWY^{+}21]. Pinball [TGSF21]. Pipelines [HSS^{+}25]. Pithos [EG17]. Pivot [XLF^{+}25]. Pixel [FXF^{+}23, Lin15, PLZW18, SZW^{+}24, WY23, XLZ^{+}23, ZLL^{+}24a]. Pixel-Residual [XLZ^{+}23]. Pixel-to-Point [SZW^{+}24]. Pixel-wise [ZLL^{+}24a]. PLACE [GQSG24]. placement [LJP08]. PLACID [MGG17]. Placing [TJN14]. planner [ZYZ^{+}13]. planning [REV^{+}12]. Plant [ZSJL22, MCM^{+}09]. Platform [BXMH15, TCP^{+}20, MGG17, SAL21b]. platforms [GFB^{+}14, LT14]. Play [LGM^{+}24]. Player [dCAGF^{+}25, FPA24, RSAH25]. Playlists [ACVM20, CYM^{+}25]. Playout [LCK09]. Pleasant [KYVE14]. Please [WSX^{+}23, Dao17]. Pleasing [BC15]. Plenoptic [HSR18]. Plug [LGM^{+}24].

Plug-and-Play [LGM^{+}24]. PMAL [KUH^{+}22]. PMHT [ZX14]. PMS [BQBLN18]. Pneumonia [LWZW21]. POB [SRAA17]. Pod [RABC24]. Podcasting [RABC24]. Poetry [WXQC20]. Point [CWL24, dMGRT24, HH25, HSS^{+}25, LYJ^{+}15, LSS^{+}24a, LYXY23, LCL^{+}24b, QLGL24, RHAG21, RBLJ25, RSR25, SZW^{+}24, WXCW24, WHX^{+}25, WXS^{+}25, WKN25, XZY^{+}24b, YGL^{+}23, YZY^{+}25, ZZL^{+}25b, ZYC^{+}24, ZXZ^{+}24b, ZZY^{+}23]. Point-of-Care [RHAG21]. Point-to-Point [LYJ^{+}15]. Points [SGY^{+}23, TAS16, YLD^{+}25, ATM06]. Polar [LZLJ22, YWNW15]. Polluters [GGB14]. Pooling [DZWH23, GCZ^{+}24, MXH^{+}23, SMN^{+}22, TYY^{+}22, WCF^{+}17, WCCN23, WTL24, ZGY^{+}24b]. pop [ML11]. Popular [WWW^{+}22]. Popularity [HDW^{+}18, KLS^{+}18, XWL^{+}25b]. Portable [JLW^{+}18]. Portal [TESU16]. Portrait [HVC^{+}20]. Portraitist [HHLY19]. Pose [BSQ^{+}24, BBL^{+}24, BCNP24, BR22, CZW^{+}25, DSL^{+}22, FH20, JCF^{+}22, LKL^{+}25b, LZXP25, SZZY20, SLL^{+}24, WTZ^{+}20, XFSZ20, XCM^{+}23, YJTN18]. Pose- [XCM^{+}23]. Pose-Controllable [SLL^{+}24]. Posed [HHLY19, WHJ20]. Poses [ZDZ^{+}22]. Position [WJS^{+}21, WXS^{+}25, KK08]. Positive [YBO14]. Post [JJML24, NY24, ZDD^{+}24]. Post-Processing [ZDD^{+}24]. Post-Quantum [NY24]. Posts [GUH^{+}20]. Potential [KWL^{+}25, Sin21, MRS^{+}07]. Power [DZW^{+}21, LZWL25, SS17, FKFB05]. Powered [ZZC^{+}25b]. PPNet [AH20]. Pr [WGM25]. Pr-Ge-Ne [WGM25]. Practical [LXC^{+}25, XLH^{+}24, ZMH^{+}20]. Practice [NLKB24, ONH07]. Pre [LFP^{+}22, LWZ^{+}25, ZZC^{+}25a, CLC05]. pre-interleaving [CLC05]. Pre-train [ZZC^{+}25a]. Pre-trained [LWZ^{+}25]. Pre-training [LFP^{+}22]. Precise [LZD^{+}21, LWZ^{+}25, YTL^{+}21].

Precise-Detection-Free [LWZ^{+}25].

Precomputation [WZ20].

Preconditioning [HXL24]. Predict [WWW^{+}22, ZZMZ21]. Predicting [CC17, BUS^{+}21, ZZB^{+}21]. Prediction [BBRS23, BTBZ20, BV25, CZL^{+}23a, CSJC17, DLL^{+}24, VTD22, DG17a, DMF^{+}20, dMGRT24, HYLG23, HDW^{+}18, JD23, LYL^{+}21a, LHZ^{+}23, LLL^{+}25, LZL^{+}25, LK18, LLL23, LXL^{+}18, MYGX21, MZGY17, QLGL24, SDK^{+}21, SYS17, SCZJ24, Tas24, THA^{+}25, TWKK21, WYMX23, XFQ^{+}21, XWL^{+}25b, XWL^{+}25a, YHG^{+}24, YZJ^{+}25, ZXZ^{+}25, ZTX^{+}23, ZSS^{+}23, KW11, SAAH10]. Predictions [MATW17]. Predictive [GLL^{+}22, KK08, LYXA22, YM24, ZWM12a]. Predictors [KN21a]. Preference [LLJW15]. Preferences [BC15, LCY^{+}23]. prefetching [LJP08]. Prerequisite [ZYLN21]. presence [MHT^{+}08, YSG^{+}06]. present [BCG13, Hua13]. Presentation [GLWK19, MBK^{+}23, WBR25, YMX^{+}16]. Presentations [AMMG16, SLKS12]. Preservation [BCP14, LZZ^{+}24c, LZYY24, PFC^{+}24, WSX^{+}23, LLC11]. Preserved [LZJ^{+}20, ZXS^{+}24]. Preserving [ATS19, HMB^{+}25, LQH18, LPY^{+}19, LKW^{+}22, MHW^{+}19, SMTR22, SHZ^{+}20, SKA24, SZZT18, WWL20, YFLF19, KT21, LMC^{+}22, LSS^{+}24b, SKSZ13, TRRB20, XZCG23, YQZP24, ZXZ^{+}24a]. PRest [CLQQ25]. PRest-Net [CLQQ25]. Pretrained [ZFC^{+}24]. Pretraining [MAX^{+}24, WDR^{+}24, XQG^{+}24]. Prevent [ZNYZ25]. Prevention [MNPOF22, SZFR23]. Primitive [LJW^{+}25]. Primitive-based [LJW^{+}25]. Principal [PWY^{+}25, ZZB24]. Principles [FF23, HNS13]. Printing [ZDE16]. Prior [GZZ^{+}24, HXL24, WWL^{+}23, XZC^{+}25, YYX^{+}24, ZFSX21]. Priori [WHW^{+}21]. Priorities [FSK^{+}15]. prioritization [ETF06]. Prioritized [YM24]. Priors [DXY^{+}21, HLY21, HPLL23, LWW^{+}25, SO22, LYC^{+}12]. Privacy [AH20, AAA^{+}21, ACC^{+}21, ATS19, FYZ^{+}21, KWK25, KT21, KCC17, LMC^{+}22, LHJ^{+}24, LSS^{+}24b, PRPO23, SKSZ13, SZEST21, SKA24, SZZT18, TRRB20, TZLZ21, WAD^{+}18, WDS21, XZCG23, ZXZ^{+}24a, ZPR^{+}25, ZYL^{+}24, ZZL^{+}24d, CCG^{+}08, MVW08]. Privacy-Aware [WAD^{+}18]. Privacy-Enhanced [ZYL^{+}24, ZZL^{+}24d]. Privacy-Friendly [KCC17]. Privacy-Oriented [PRPO23]. Privacy-Preserving [ATS19, SKA24, SZZT18, KT21, LMC^{+}22, LSS^{+}24b, TRRB20, XZCG23, ZXZ^{+}24a]. Private [GYF^{+}21]. PRNU [ZTQX23]. PRNU-based [ZTQX23]. ProActive [NR24]. Probability [CI23, CLQQ25, LYZ^{+}18, QLW^{+}24, WBB^{+}17, ZWR^{+}20]. Problem [HXH^{+}24, YWF^{+}24, YHZ23, FYH10]. Procedural [HMVI13]. Process [ETF06, LLL23]. Processing [AMT25, SZHY19, YH14, ZDD^{+}24, EGEM06]. Product [CAJ19, JLW^{+}18, TS22, LWL^{+}12]. Production [TXW^{+}25, ETF06]. Professor [Ste10]. profilable [ZCY^{+}13]. Prognosis [WZD^{+}20]. Program [MNPOF22]. programming [ASVE13]. Progress [BUS^{+}21, UTK^{+}08]. Progressive [CZW15, GQSG24, NWNL21, PB14, QLW^{+}23, SA16, SHWC19, TSH^{+}25, XYC^{+}23, YHG^{+}24, ZRZ^{+}21, ZDQ^{+}23, ZZL^{+}25d, CLC05, COM^{+}11, LLP06, MCM^{+}09]. Projection [LLN^{+}24, YKQ^{+}21, ULIS07]. Projections [ZLOL18]. Projective [RYZ^{+}23]. Prompt [MKY^{+}24, WWQ^{+}25b, XJB25, YHP25, ZYZ^{+}25a]. Prompt-Based [MKY^{+}24, WWQ^{+}25b]. Prompting [JCB25]. Prompts [LSX^{+}25]. Propagation [CCH^{+}24, HCWM14, LCL14, WZC^{+}13, WXJL25, CML^{+}13, HOSS13]. Propagation-based [WZC^{+}13]. PROPANE [PB14]. Properties [KMSW18, WYM07]. Property [GDM^{+}23].

Proposal [XXG^{+}21, XLW23a, ZMX^{+}23, ZGL^{+}18, ZYY^{+}20]. Proposal-Intra [XLW23a]. ProposalVLAD [XLW23a]. Proposed [ABR17]. Protected [AH20]. Protecting [BCP14, LLC11]. Protection [CPSH14, FYZ^{+}21, LHJ^{+}24, MSS^{+}24, NPG^{+}22, TWM^{+}25, TZLZ21, WWZ^{+}25, WFZ25, XWZ^{+}24b, ZNYZ25, CLC05, CCG^{+}08]. Protocol [JTZ^{+}16, LB15, PB14, SSP21, OMP07]. ProtoRefine [ZLL^{+}24d]. Prototype [CZX^{+}24, FJF^{+}25, KLX^{+}24, TY23, WXJL25, ZZL^{+}24d]. Prototype-Based [ZZL^{+}24d, WXJL25]. Prototypes [ZLL^{+}24d]. Prototypical [LZY^{+}24]. Prototyping [GJAF18]. Providing [ZCL^{+}12, ILL08]. Provisioning [LDT^{+}18]. proxies [LS05]. Proxy [KUH^{+}22, RXC14, DY09, ILL08]. proxy-enabled [DY09]. Pruned [WGM25]. Pruning [YHG^{+}24]. Pseudo [LWH^{+}22, LJG^{+}24, MC19, WHY19, ZWX24, ZLZ22b]. Pseudo- [WHY19]. Pseudo-labeling [LJG^{+}24]. Pseudo-Labels [ZLZ22b]. Pseudo-likelihood [MC19]. Pseudo-Siamese [ZWX24]. Pseudo-true [LWH^{+}22]. PSNController [MY15]. Psychology [YYY^{+}24]. Psychology-Guided [YYY^{+}24]. psychophysical [WAK^{+}12]. Pulmonary [TLZ^{+}20]. punishment [BAK13]. Pupil [LZH^{+}24a]. Pupil-wave [LZH^{+}24a]. Pure [ZZL^{+}24a]. Purpose [SG22b]. pursuing [LYCJ12]. pursuit [DCC^{+}13]. PVLMs [XJB25]. Pyramid [FXF^{+}23, QSY^{+}23, XPP^{+}23, YFLF19, MHT^{+}13]. Python [GFB^{+}14].

Q [YXZQ22]. QoE

[SKVHC18a, AS20, BRZS18, BPFA24, EWSZ15, FAA18, FSK^{+}15, GGA^{+}20, HXZ^{+}20, KJJ^{+}21, LSCH25, LCT^{+}12, MHCG19, MATW17, PFC^{+}16, SKVHC18b, SSSK18, Tas20, Tas22, Tas24, WBRZ17, YJM^{+}19, YZL^{+}24, Zha19, ZGL^{+}18]. QoE-Aware [FAA18, WBRZ17, Zha19]. QoE-Based [EWSZ15, FSK^{+}15, MATW17]. QoE-Driven [BRZS18, PFC^{+}16, SSSK18]. QoE-Fair [AS20]. QoE-oriented [LCT^{+}12]. QoS [AH20, AS22a, EGEM06, KW11, Tas24]. QoS-aware [AH20, EGEM06]. QR [XLN^{+}21]. QRNet [YTY^{+}24]. QUality [GSB^{+}22, BQBLN18, CZY^{+}21, DG17a, DLC^{+}22, FDZ^{+}25, FBSGR25, FPA24, HEA14, HLY^{+}23, JD23, KN21a, LB15, LCZ^{+}24a, LXJ^{+}24, LZWL25, LGLZ20, LYXY23, LYW^{+}23b, LH12, MWL^{+}24, PVWD18, QJ23, RT14, SMN^{+}22, SWZ25, SVA^{+}21, TCL^{+}23, TFY^{+}25, TFHM24, TMB^{+}22, TSP^{+}24, TNH^{+}21, WBRZ17, WCY^{+}18, XZY^{+}24b, YTL^{+}21, YXJ^{+}24, YTR^{+}25, YWC^{+}25, YDCZ23, YWW^{+}24, YYX^{+}24, YMJ^{+}25, YCGM14, ZSMZ21, Zha19, ZGL^{+}23, ZSZ^{+}24a, ZSX^{+}25, ZZL^{+}25b, ZZL^{+}25c, ZYC^{+}24, BDV08, FHK^{+}23, GG06, HAE11, KS09, KO11, NH10, PS05, RHS12, SGS21, SkFM18]. Quality-Adaptive [BQBLN18]. Quality-Aware [TFY^{+}25]. quality-managed [PS05]. Quality-of-Experience [DLC^{+}22, SkFM18]. Quantitative [RHS^{+}20, HM10]. Quantization [PKLK23, RSE16, WZTL19]. Quantum [NY24, SL23]. Quaternion [HD19, YWNW15]. Queries [CGGC20, San11]. Query [HMOS17, HPZ25, TY23, YH14, ZHS^{+}25, JYZC12, KP08, LYJ^{+}13, LMLC14, ZYM^{+}10]. query-adaptive [LYJ^{+}13]. Query-Guided [TY23]. Question [BBB^{+}23, EL25, HWZL24, LXB^{+}22, LJL^{+}24, LXL^{+}18, LYW^{+}25, PN16, PSSZ24, WLS^{+}24b, WLSN24, YTOH22, ZSZ^{+}25, ZXY^{+}20]. Questionnaire [HSX22]. questionnaires [NT08]. Questions [LGY^{+}22]. Queuing [BYOZ20]. QuGu [LB15]. QUIC [KFP^{+}25]. Quick [CXX^{+}24]. QuickCSGModeling [CXX^{+}24].

R [BPB^{+}22]. RA [GHAM25]. RA-MOSAIC [GHAM25]. RAC [XGZ^{+}24]. RAC-Chain [XGZ^{+}24]. Radial [HD19]. Radiance [PYBY25]. Radio [YHH^{+}22]. RAG [XZY^{+}25]. RAG-Based [XZY^{+}25]. Rain [FHF^{+}24]. raining [SGS21]. Random [ACVM20, APP^{+}22, SZFR23, WZD^{+}20, YKQ^{+}21]. Randomisation [WF25]. Randomized [LQZH14, BWA13]. range [ZPL^{+}23]. Rank [FZW^{+}25a, KVB^{+}24, LLSX20, LK18, PRGA18, SZM^{+}21, XYW^{+}22, ZRZ^{+}21, WFC^{+}25a, YWG^{+}20]. Rank-based [KVB^{+}24]. Rank-in-Rank [XYW^{+}22]. Ranking [LGX^{+}24, YBO14, YTOH22]. Rare [SS24]. RAST [MZH^{+}24]. RaT2IGen [YHP25]. Rate [DMF17, HYLD20, HH08a, MN16, MHCG19, PFC^{+}16, SSP^{+}24, YXJ^{+}24, YJM^{+}19, ZZLL17, HH08b, HH11b, LCK09, LJP08]. Rate-distortion [HH08a, HH08b]. Ratio [ZZC^{+}25b]. Rationale [WZZL25]. Rationale-Enhanced [WZZL25]. RBFNN [WJS^{+}21]. RBFNN-Based [WJS^{+}21]. RCE [HPW20]. RCE-HIL [HPW20]. RCNN [TLZ^{+}20]. RD [YZS^{+}22]. RD-IOD [YZS^{+}22]. RDH [CQA^{+}25, XHYX23]. RDH-DES [XHYX23]. RDH-EI [CQA^{+}25]. RDIAS [GAH25]. RDO [YGMY24]. RDO-free [YGMY24]. Re [AQL^{+}20, CWW^{+}25, CHS^{+}23, CZL23b, DFXY20, DSL^{+}25, DLL^{+}18, FZYY18, FZZ^{+}25, GL08, GLT^{+}20, GHZW25, HLM^{+}22, HZH24, HCZ^{+}25, LYZY21, LGX^{+}24, LZJ^{+}20, LYZX21, LLS^{+}24, LZC^{+}19, LWH^{+}23, LZL^{+}24b, LXC^{+}25, LWHC24, LCZ^{+}24c, PZJL22, PWX^{+}24, QWH^{+}21, RLY^{+}21, SJC^{+}19, TH22, TGC23, WDPX23, WZRD25, WWQ^{+}25a, WHW^{+}21, WJQ^{+}22b, WZW24, XQG^{+}23, XQG^{+}24, XGX^{+}25, XWH^{+}25, XHL^{+}24a, XLZ^{+}22, XYW^{+}22, XLH^{+}23, YTOH22, YWH^{+}17, YJTN18, YLYB24, ZGZ^{+}22, ZLZ^{+}22a, ZZJ^{+}23, ZSZ^{+}22, ZZZ^{+}22, ZGWZ23, LSL^{+}24]. Re-cinematography [GL08]. Re-detection [LYZY21]. Re-Id [GLT^{+}20, XLZ^{+}22, ZZJ^{+}23, LSL^{+}24]. Re-Identification [CZL23b, DSL^{+}25, DLL^{+}18, FZZ^{+}25, HCZ^{+}25, LZC^{+}19, LWH^{+}23, LZL^{+}24b, LXC^{+}25, PZJL22, PWX^{+}24, WWQ^{+}25a, XQG^{+}23, XQG^{+}24, XGX^{+}25, XWH^{+}25, YJTN18, YLYB24, ZGZ^{+}22, ZLZ^{+}22a, ZSZ^{+}22, AQL^{+}20, CWW^{+}25, CHS^{+}23, DFXY20, FZYY18, GHZW25, HLM^{+}22, HZH24, LZJ^{+}20, LYZX21, LLS^{+}24, LWHC24, LCZ^{+}24c, QWH^{+}21, RLY^{+}21, SJC^{+}19, TH22, TGC23, WDPX23, WZRD25, WHW^{+}21, WJQ^{+}22b, WZW24, XHL^{+}24a, XYW^{+}22, XLH^{+}23, YWH^{+}17, ZZZ^{+}22, ZGWZ23]. Re-Ranking [LGX^{+}24, YTOH22]. Reactions [Abd18, KTK^{+}17]. Reading [SVA^{+}21, ARE13, Dao17]. Real [BPM15, CTBC22, DG17a, DWC^{+}21, EGEM06, FHG^{+}17, GLZW23, GGA^{+}20, yHcCzH^{+}21, HRL^{+}24, IB10, KMP05, LXJ^{+}24, MPM^{+}25, MDAE19, RSR25, SAZ^{+}15, TWM^{+}25, WHL^{+}23, XHL^{+}24b, YZL^{+}24, ZYW^{+}25, ZCYL24, HCKL13, KK08, KW11, SNC12, YSG^{+}06, PYZ25]. Real-Time [BPM15, DG17a, DWC^{+}21, FHG^{+}17, GGA^{+}20, yHcCzH^{+}21, HRL^{+}24, MDAE19, SAZ^{+}15, TWM^{+}25, WHL^{+}23, YZL^{+}24, ZYW^{+}25, EGEM06, GLZW23, IB10, KMP05, MPM^{+}25, RSR25, HCKL13, KK08, KW11, SNC12, YSG^{+}06]. Real-World [XHL^{+}24b, ZCYL24, LXJ^{+}24]. Realistic [DBPM25, DŞB^{+}22, HCM^{+}25, IPMD25, WWL20, GHW^{+}24]. Reality [AWG^{+}15, BLMP18, CGPCR18, LL15, MSKYJ21, REP^{+}19, RVTC25, SCX^{+}25, SRDJ18, SAL21b, VTP^{+}20, VMP20, WDE25, GYN12]. realization [SZ12]. Realtime [XLH^{+}24]. Reasoning [DLZ^{+}17, LLC^{+}24a, LYX23, MYX23a, TWFW24, ZYZ^{+}25b]. Rebuffering [WBB^{+}17]. recall [AG13, YZL^{+}14]. Receivers [LPS15, QS10]. Receptive [LZL^{+}24a]. Recognition [AdSGMS25, AMC^{+}18, AC19, ASLA18, AA23, BYM^{+}18, BTWZ22, BR22, CWTG22, CFN^{+}23, CZ24, DVA21, FSX14, FXMZ24, GZW^{+}24a, GGSV25, GZLW18, GNC^{+}23, GSDT21, HYG^{+}25, HC25, HZW^{+}11, HZPL21, HZL^{+}21, HDH^{+}24, HSZ^{+}18, HYG^{+}21, HWW^{+}23, HP17, JMLL20, KP15, LLJ^{+}20, LZX^{+}21, LYL23a, LLL^{+}21, LMX24, LLZ^{+}24, LCL^{+}24b, LHC^{+}25, LQX^{+}24, MON21, MZZ^{+}20, MGP19, NWNL21, NRUT20, PHH^{+}24, PS17, RNR^{+}22, RE24, RXY^{+}25, SLZ^{+}21, SCZ^{+}25, SZL^{+}22, SLL^{+}21, SLZ25, TYY^{+}18, TLZ^{+}21, TLY^{+}22, TSH^{+}25, WJ15a, WWY^{+}24, WCG^{+}25, WWL^{+}24, XW25, XLH^{+}24, XFZ^{+}19, XLZ^{+}21, XZH^{+}21, XJW^{+}22, XHL^{+}24b, YML^{+}22, YSSF24, YMS^{+}24, YHH24, YWLZ23, YZZZ25, YHQH17, YSY^{+}22, YZYX24, ZHL19, ZFSX21, ZLZ^{+}22a, ZYQM24, ZHY^{+}25, ZLP^{+}14, ZGD^{+}19, ZZW^{+}19, ZLH16, ZWF^{+}20, ZZW^{+}22, ZZY^{+}23, KO11, LZP07, ML11, WT10]. Recognize [THR^{+}22]. Recognizer [LYL23a]. Recognizing [HPW20, LLCH17, MDAE19, YKW^{+}22, ZWJ^{+}24]. Recolored [ZCQ^{+}23]. Recombination [GAB^{+}17]. Recommendation [CWXW25, DBBD23, DYSX14, DBSL24, GZW^{+}25, HCWM14, HQF^{+}20, KZGH15, LCY^{+}23, LSS^{+}24b, LSX^{+}25, MBXZ24, QSZ^{+}21, SYX23, WWHW14, WDJ^{+}21, WWZ24, WSLM18, YSXH16, ZCS^{+}20, KS13, TBC^{+}11, WWL13]. Recommending [ZYLN21]. Reconfigurable [LYJ^{+}15]. Reconstructed [WWH17]. Reconstructing [PB19]. Reconstruction [CZL^{+}23a, CJZ^{+}24, CZ24, FBSGR25, FMG20, GWWC25, HZSC20, HZC22, HSR18, HSL^{+}20, HXL24, LZWC23, LH22, LKW^{+}22, LCZ^{+}24b, LNC^{+}25, SLC25, SWK^{+}22, TLZL24, YKQ^{+}21, YXK^{+}25, ZSYW23, JGZ^{+}11]. Reconstruction-Based [SLC25]. Reconstruction-Free [TLZL24]. Reconstructive [DWL^{+}23]. Recording [HLS^{+}22]. records [LT14]. Recreated [SRDJ18]. Rectification [WSX^{+}23, YLLZ25]. Rectified [ZSJL22]. Recurrent [FZYW20, GZW^{+}24b, HJ23, LSXZ23, MOL^{+}22, SGS21, ZQKH19]. Recycle [PKL25]. Reduce [HLD18]. Reduced [FZW^{+}25a]. Reduced-Rank [FZW^{+}25a]. Reduction [GNC^{+}23, LWL24, LK18, XZLZ23, YMA17, HCS12, JP11]. redundancy [WGL^{+}23]. Redundant [HZSC20]. REED [AH12]. Reenactment [LXT^{+}24, XYC^{+}23, XLT^{+}23]. Reference [GCZ^{+}24, HLY^{+}23, HLC^{+}25, JD23, JYZ^{+}25, LLHS12, LJ25, MXZ^{+}24, TCL^{+}23, YTL^{+}21, ZLF^{+}24, CZY^{+}21, LYXY23, LYW^{+}23b, YDCZ23]. Reference-Based [GCZ^{+}24, MXZ^{+}24, ZLF^{+}24]. Referring [LXC^{+}24, WKS^{+}23, YTY^{+}24]. Refined [LCW16, ZWL^{+}21, LMXJ21]. Refinement [BSQ^{+}24, DHP23, DSL^{+}25, HYLD20, LZL20a, LDL24, LZD^{+}21, PB19, ZT22, ZH18, ZHS20, ZZTL24, ZXZ^{+}24b, ZXY^{+}20]. Reflection [LLYL22]. Region [FJF^{+}25, HLM^{+}22, HWHL18, WLHL13, XZL^{+}22a, YHZ19, ZHS20, ZWYS20, FDKB11, LYC^{+}12]. Region- [WLHL13]. Region-Level [ZWYS20]. region-of-interest [FDKB11]. Regional [GZZ^{+}24, JLLL25, XLW^{+}23b]. Regions [KP15, LGY^{+}22, SZTL16]. Registration [SZW^{+}24, ZZLW25, ZLF^{+}24, DE12]. Regression [WHJ20, ZT22, ZZL^{+}17, ZGR21]. Regressive [HYLG23]. Regularization [HXL24, LGY^{+}22, MSFR25]. Regularized [FBG22, LYZY21, MKC21, SG22a]. Regulation [FZZ^{+}25]. Regulations [KWK25]. rehabilitation [CXS^{+}08]. Reidentification [LKM17, ZZY18]. Reign [LVM^{+}21]. reindexing [WC12]. Reinforced [FZYW20]. Reinforcement [AS20, BBZ18, HLW^{+}21, HLW^{+}25, WWHW14, WTZ^{+}20, YDJ24, ZGL^{+}23]. Related [LLC^{+}24b]. Relation [HZS^{+}24, LMX24, XZH^{+}21, XWL^{+}25a, YHP25, ZPY^{+}23, ZYLN21, ZYZ^{+}25b]. Relation-aware [YHP25]. Relational [CXY^{+}25, LZZS23, LDL^{+}25, PSSZ24, TWFW24]. Relations [LZF^{+}22, WLS^{+}24b]. Relationship [DPL^{+}23, LKW^{+}22, LLZ^{+}24]. Relationships [LML^{+}24, YH14, YZW^{+}24]. Relative [LZL^{+}23]. Relevance [DNPG^{+}17, XDL^{+}21, WWL13]. Relevant [NNK25, ZYJZ25, LYX23]. Reliable [ZWZL21, ZC12]. Relying [WKE16]. remapping [CCD07]. Remote [GHY^{+}24, MSKYJ21, ZSZ^{+}24b, SNC12]. Removal [CWL24, FHF^{+}24, LLYL22, ZLZX23]. Rendered [MSKYJ21]. RenderGAN [MPM^{+}25]. Rendering [MPM^{+}25, KTM^{+}06, SNC12]. Reorganization [ZYCD24]. Repair [SAL21b]. repairing [DCM13]. RePC [SWZ25]. Repetitive [ZLT^{+}25a]. Replay [CWW^{+}25]. Replays [JSEI16]. Replication [Ala21]. report [RJ05]. Repository [PYZ25]. Representation [ATMM24, CXY^{+}25, CSW^{+}25a, GLL^{+}22, GSDT21, HZL^{+}21, HJZ^{+}23, JMLL20, JCF^{+}22, JLK^{+}24, LYZX21, LZZ^{+}24a, LYC^{+}25, LKL^{+}25b, LLL^{+}22, LYDY24, LDZ^{+}24, MSC^{+}23, MAX^{+}24, PQ19, RXY^{+}25, SSY20, SZM^{+}21, SYY^{+}24, SZTL16, TWL19, TGC23, TXWK24, WCL^{+}22, YPSC22, YGL^{+}23, YTL^{+}24b, ZZL^{+}17, ZFSX21, ZTR^{+}22, ZHLJ22, ZYW^{+}24, ZXZ^{+}25, MGCH13, MHT^{+}13]. Representation-Based [ZZL^{+}17]. Representations [BPB^{+}22, CSW^{+}25b, DHT^{+}19, HLW^{+}25, JCSL19, LKL25a, LCH^{+}25, LXT^{+}24, NRUT20, RE24, SXY^{+}23, TAPP^{+}15, YZRW22, ZLL20].

Representative [WLSN24]. Representativeness [CWW^{+}25]. Representing [RT14]. Request [HJWW19]. Requet [GGA^{+}20]. reranking [TTR12]. Resampling [ZYC^{+}24]. research [CBS08, Eff13, RJ05, She13]. Reservation [XFQ^{+}21]. Residual [HCZ^{+}25, JLL^{+}21, LSYM19, LLJ^{+}20, LZD^{+}21, LJZ^{+}22, RBBS25, SLC25, WSLC24, XLZ^{+}23, YHXL20, YZS^{+}22, ZT22, ZLZX23]. Residual-Distillation-Based [YZS^{+}22]. Residual-guided [JLL^{+}21]. Residuals [GS19]. Resilient [HD19]. RESIST [WZZL25]. Resistant [MBXZ24, YLLL21, YJTN18, PJL23]. Resolution [BWP^{+}24, CFN^{+}23, FYW^{+}24, GHR^{+}22, GCZ^{+}24, HJ23, HPLL23, LSYM19, LZWL25, LCC^{+}14a, LZW^{+}23, SYM^{+}25, SGY^{+}24, WXZ^{+}23, WLTL23, WLS^{+}24a, WLW^{+}23, YLLL23, YWL^{+}24, ZSS20, ZYCD24, ZZL^{+}25d, JXTC21, KP08, LCL^{+}21, MGY22, WGL^{+}23, YW23, ZLF^{+}24]. Resource [GHAM25, PHS^{+}20, WZZ^{+}22, XFQ^{+}21, XZL^{+}22b]. resources [FLM^{+}06, MRS11]. Response [SWS^{+}22]. Responses [LZXY20, NT08]. Restorable [MZH^{+}24]. Restoration [CH15, FHK^{+}23, HWW25, LLSX20, RBBS25, SBS23, ZGL^{+}23]. Restore [FBGD23]. Restraint [WJQ^{+}22b]. Result [TJN14]. Retained [XCC^{+}23]. Retargeting [KKGE18, ZHG^{+}21, ZLN^{+}16]. Rethinking [HZH24, LG24, XQG^{+}24, ZQKH19]. Retinopathy [MAZ22, SHE21]. Retreat [RJ05]. Retrieval [BBUD24, CGGC20, CDL^{+}20, CPL^{+}23, CZQ^{+}22, DNPG^{+}17, DLZ^{+}17, DHT^{+}19, DDL^{+}25, FSEW25, FWLA15, FHP23, GLW20, GJY^{+}25, GJAF18, HCZ^{+}22, HZS^{+}24, HLW^{+}24, HLX^{+}14, KVB^{+}24, LQH18, LD25, MWL^{+}24, MGS18, MAE^{+}21, MSFR25, NN21, NWL^{+}20, NLW^{+}21, NSM^{+}24, NNK25, PYZ25, PDD16, QCW^{+}25, SCB^{+}24, SBU^{+}17, SHZ^{+}20, SZL^{+}24, SZZ^{+}23, SHF^{+}24, SLMJ24, SLNL20, SZTL16, TP22, WJ15a, WLHT19, WDC^{+}25, WBR25, XDL^{+}21, XTL^{+}21, XZL^{+}25, YTOH22, YLK^{+}20, YLL^{+}23, YSZ15, YZW^{+}24, YTRC19, YHZ23, ZYO20, ZCL^{+}22, ZWH^{+}23, ZWX24, ZZY^{+}14, ZWY21, ZXX22, ZYT^{+}23, ZLLH25, ZYJZ25, ZSZ^{+}24b, ZWZ^{+}23b, ZFFX24, ZHS^{+}25, ZZL^{+}24d, Cha13, DP06, FLM^{+}06, HMUC21, Han13, hHLC10, LSDJ06, LBD08, LLW^{+}13, LYJ^{+}13, QLSQ12, TCJ08, Whi13, YGHH12, ZG08, ZHLY11]. Retrieval-Augmented [SCB^{+}24]. Retrieving [FSEW25, CW10]. retrospective [LWLZ13]. reuse [MRS11]. revenue [ILL08]. revenue-rewarding [ILL08]. Reversal [LLS^{+}24]. Reversible [LHY^{+}24, PJL23, XLZ^{+}23, XHYX23, YCZ^{+}24, NC13]. Review [CS22, HC22, JLA^{+}23, RSAH25, TRK^{+}20, TAT^{+}25, VLE25, ZYZE19a, ZWL^{+}24, ZZS^{+}25, TV07]. Reviewers [Ste13b]. revision [LT14]. revisited [BZ05]. Revisiting [HHL^{+}22]. revocation [YH13]. revolution [Swa13]. Reward [WZZL25, ZGY^{+}24a, BAK13]. reward-and-punishment-based [BAK13]. rewarding [ILL08]. Rewards [MXZ^{+}24]. Reweighted [WB16]. Rewriting [ZHS^{+}25]. RGB [BSQ^{+}24, CRL20, CCH^{+}24, PSN^{+}24, QWH^{+}21, WCLC18, WWS^{+}24, WLSX24, WJQ^{+}22b, XZC^{+}25, YXK^{+}25, ZHY^{+}25, ZZL^{+}24c, ZZW^{+}22]. RGB-D [BSQ^{+}24, CCH^{+}24, WCLC18, WWS^{+}24, WJQ^{+}22b, XZC^{+}25, ZHY^{+}25]. RGB-grey [QWH^{+}21]. RGB-T [WLSX24, ZZL^{+}24c]. RGBT [LYZ^{+}24, TXWK24, ZWL^{+}24]. RICA [ZWL^{+}21]. RICA-MD [ZWL^{+}21]. Rich [TWL19, WCX^{+}14, LQX^{+}24, SWH06, TBC^{+}11]. Riemannian [HPH^{+}20]. Right [LGY^{+}22]. rights [SEK12]. Rivalry [Wan21]. RNN [dMGRT24, LZH^{+}20, ZSLW23]. Road [HYT^{+}25, ZLL^{+}24a]. roadmap [SZ12]. Robust [CZ24, CLQQ25, CLZ^{+}23b, CCD07, DCC^{+}13, FNH22, FBGD23, FBG22, GKSB17, GAH25, GZL^{+}20, GMWW25, HWZL24, HLW^{+}25, HQF^{+}25, JK25, LTL^{+}24, LK18, LML^{+}13, LPC^{+}18, LWH^{+}23, LLN^{+}24, LZW^{+}24, MSS^{+}24, SG07, SZZZ25, SZM^{+}21, SWM^{+}24, SHE21, SO22, SZZT18, SHWC19, SR25, TXWK24, WLC^{+}20, WLL23, WYY^{+}25, XG24, XLN^{+}21, XSD^{+}22, YLLL21, YWNW15, YTL^{+}24a, ZRZX23, ZSJL22, ZHS20, ZGR21, ZXS^{+}24, ZZL^{+}24c, AP10, CWC10, GZHD12, MB08, VPSS^{+}13, YPSC22]. Robustness [WDX^{+}23]. Rotating [GHY^{+}24]. route [ZYZ^{+}13]. Routing [HJWW19]. RSUIGM [DBPM25]. rules [BAK13]. Running [YKW^{+}22].

S [YHZ19, ZWJ^{+}24]. S3Mix [ZCX^{+}24]. SABR [BRZS18]. SADnet [SZB^{+}22]. SAfeDJ [HDZ^{+}15]. Safety [BCNP24]. Salience [LZWL25, ZXZ^{+}24b]. Salience-Aware [LZWL25]. Saliency [CBSC18, HYLD20, LZD^{+}21, LYW^{+}24, NALM23, STSW23, YHG^{+}24, ZWL^{+}17, ZPY^{+}23, ZCHD23, ZXZ^{+}25, ZSZ^{+}23]. Saliency-based [HYLD20]. Salient [CCH^{+}24, HGGZ23, LDL24, LG24, TB05, WWS^{+}24, WLSX24, XZL^{+}22a, XWW^{+}21, ZHS20]. SAM [LQG25]. SAM-guided [LQG25]. SAMAF [BSSNF^{+}20]. Same [ZCX^{+}24]. Sample [DLD^{+}22, LJL^{+}24, LDL^{+}25, DCC^{+}13]. Sampler [WZWY23]. Samples [YZG^{+}20]. Sampling [FBSGR25, SLMJ24, ZGR21, LK07, WDCX07]. SAND [PHS^{+}20]. SARA [YJM^{+}19]. Sarcasm [OL25, TNZ^{+}25, ZYW^{+}24, ZZL25e]. Satellite [DLL^{+}24, FZW^{+}25b, HYT^{+}25]. Save [SLP15]. Saving [HSN^{+}14]. SCAE [LZZ^{+}24a]. SCAG [LYW^{+}25]. Scalable [GSM^{+}08, GFB^{+}14, PKLK23, RBBS25, SAF19, SZTL16, WAD^{+}18, WZTL19, XGZ^{+}24, ZZT24, ZZL^{+}24b, FKFB05, HH08a, HH08b, HH11c, JKKL08, RHS12, SAAH10, TCJ08, LCSX11]. Scale [BLW^{+}24, BQBLN18, FYW^{+}24, GJAF18, HDW^{+}18, LSK^{+}15, LGC^{+}24, REP^{+}19, WDC^{+}25, WWQ^{+}25a, YSY^{+}22, ZDZ^{+}23, ZSZ^{+}24b, ZWR^{+}20, ZWYS20, BBB^{+}23, CE10, CML^{+}13, HZC22, HJ23, LZWC23, ML11, SX11, SW18, TYY^{+}22, VGNL10, WMH^{+}22, WQTX25, Whi13, WLZ08, XWZ24a, XLS^{+}25, YW23, YP20, ZJZC20, ZTQX23, ZWS^{+}20, ZJSJ20, ZLLT13]. Scale-Adaptive [BQBLN18]. Scale-Semantic [ZSZ^{+}24b]. ScaleFFS [JKKL08]. Scales [WWX^{+}21]. Scaling [CHHH25, WCK05]. Scan [YZY^{+}25]. Scan-Order [YZY^{+}25]. Scanning [XLN^{+}21]. Scanning-Robust [XLN^{+}21]. Scattering [YLCC18]. Scenario [LYX23, LCL22, MOL^{+}22, NSM^{+}24]. Scenario-Aware [MOL^{+}22]. Scenario-relevant [LYX23]. Scene [CWXW25, CSW^{+}25b, GF17, GSDT21, HDH^{+}24, JCSL19, LWL24, LZDL25, LOJZ18, LGLZ20, LZL20b, LZL21b, LWZ^{+}25, SHT^{+}24, TWY24, UJLS22, WCLC18, WYMX23, WLH^{+}21, XFZ^{+}19, XWL^{+}25a, YZYX24, ZPY^{+}23, ZZTL24, ZCYL24, ZFFX24, PJ13, SG07]. Scene-Centric [CWXW25]. Scenes [CW25, LZH^{+}24a, LOJZ18, ZZB^{+}21, ZJL^{+}21]. scenic [ZYZ^{+}13]. Scent [MLQM14, LCSX11]. Scents [HC22]. SCeVE [VMP20]. Scheduling [SLP15, ZMH^{+}20, RWW05, WLQL12, ETF06]. Scheme [ARA^{+}23, CPSH14, CQA^{+}25, CXC^{+}17, GZH17, HXZ^{+}20, HSL^{+}20, LXB^{+}22, LZW^{+}21, RLXW24, SR22, TFY^{+}25, WFZ25, XJG^{+}22, XLZ^{+}23, YWF^{+}24, Zha19, ZWZL21, GZHD12, ILL08, LLHS12, LC12, PBS12, ZC12, WWGT09]. Schemes [DLO^{+}20, LZD^{+}22, MAGT23, WFC^{+}25b, HM10]. Sclerosis [WZ20]. Score [RTR21, HCS12]. Screen [BFAS15, CZY^{+}21, yHcCzH^{+}21, MFL^{+}16, WXZ^{+}23, ZWZL21, CBR14].

Screen-Sharing [yHcCzH^{+}21]. Screening [LWLC24]. Screens [NSJB17, KS09]. Scribble [LWZY22]. Scribble-Supervised [LWZY22]. Scribbles [YQC^{+}21]. Sculptural [YLHL22]. SD [ZLT^{+}25b]. SD-Meta [ZLT^{+}25b]. SDCN2 [SG22b]. SDN [BBZ18, KCC17, KJJ^{+}21, OE19, SP21]. SDN-Assisted [OE19, SP21]. SDN-Enabled [BBZ18]. Seam [Hon19, ZHG^{+}21]. Seamful [NHP^{+}16]. Search [AR15, CLP17, EFF^{+}25, GCZ^{+}24, HMOS17, HLX^{+}14, HJZ^{+}23, JLW^{+}18, KVB^{+}24, LGX^{+}24, LKM^{+}19, LVM^{+}21, MKS20, NN21, RNR^{+}22, TCS^{+}25, TJN14, TS22, WHF^{+}18, WLC^{+}20, WMW^{+}22, XZL^{+}22a, ZWZL22, ZXX22, ZWYS20, HTT^{+}11, JYZC12, LZP07, NW08, RW12, TTR12, YZL^{+}14, ZYM^{+}10, ZR13, ZLLT13]. search-from [ZR13]. Searchable [XJG^{+}22]. Searching [SWM^{+}24, HCW^{+}07]. Searching-Based [SWM^{+}24]. Second [ZGM^{+}20]. Second-Order [ZGM^{+}20]. Secret [Lin15, LZW^{+}21, WY23, XHYX23, YLLL21, YLS^{+}23]. Section [APRS^{+}19, HCM^{+}22, YNC18, YZRW22, BSH08, CDGJ09, GPHOH12, KZHC13, NLS13, SHOG12]. Secure [BWA13, DKJ^{+}21, HZL^{+}16, HXL24, LWZ^{+}21a, RTR21, SRAA17, SSSA23, YWF^{+}24, ZC12]. Securing [NCMM21, YMA17]. Security [ACC^{+}21, BCNP24, CQA^{+}24, GHS^{+}24, KJJ^{+}21, LQS21, SMK^{+}24, SZEST21, SLK21, SAL^{+}21a, WJS^{+}21, WWZ^{+}25, WHZ^{+}24, WDS21, ZPR^{+}25, Kan12]. See [SkFM18]. Seed [XXG^{+}21]. Seeing [LLL^{+}22]. seen [YZL^{+}14]. Segment [Ala21, TCJ08, YJM^{+}19]. Segment-aware [YJM^{+}19]. segment-event-object-based [TCJ08]. Segmentation [AM24, CHLW19, CH21, CYZ23, CLZ^{+}23a, CYZ^{+}24, FBG22, HYT^{+}25, HCY^{+}23, JXL^{+}25, LZP07, LXC^{+}24, LLZ^{+}20, LYL^{+}21b, LWZY22, LLL^{+}24a, LLL^{+}24b, LLWX24, LGW^{+}25, LDZ^{+}20, MYX^{+}23b, PA20, RTR21, RLAM25, SMN^{+}22, TWY24, TY23, TRRB20, WLFL23, WWL^{+}23, XYJ^{+}20, XXG^{+}21, XSL^{+}23, YXYB21, YTY^{+}24, YFLF19, ZCXL23, ZCD15, ZZW^{+}24, ZXZ^{+}24b, ZHG^{+}25, ML11]. Segments [TKK^{+}17]. Seizure [HAAM19, JGJ^{+}20]. Selected [PSS05]. Selecting [CCG20, VV11]. Selection [CNG22, Cla18, HZC^{+}16, HSZ^{+}18, LZZ23b, LLO^{+}20, MEA^{+}21, NTT20, TAPP^{+}15, WWL^{+}24, WHW18, XYL^{+}23, YBW^{+}25, YM24, YBO14, YHZ19, ZZLL17, AKO07, CE10]. Selective [DHT^{+}19, PZJL22, TYY^{+}22, WHY18]. Selectively [CQA^{+}24]. Self [AQL^{+}20, AMG23, CYZ^{+}24, DPL^{+}23, GHQ24, LMXJ21, LYZ^{+}24, LH22, LWHC24, LZF^{+}22, PSL^{+}24, QXG^{+}24, SLG^{+}24, SYX23, SZZ^{+}23, WZTL19, WZD^{+}20, WGL^{+}23, WFZ25, YML^{+}22, YWH^{+}17, YWW^{+}24, ZYW^{+}24, ZHY^{+}25, ZZY^{+}23, DCM13, GTP^{+}22]. Self-Adaptive [SLG^{+}24, ZYW^{+}24]. Self-attention [AQL^{+}20, YML^{+}22, YWW^{+}24, ZZY^{+}23]. Self-calibration [WGL^{+}23]. Self-contained [AMG23]. Self-Defense [WFZ25]. Self-Organizing [WZTL19]. Self-Paced [WZD^{+}20]. Self-refined [LMXJ21]. self-similarity [DCM13]. Self-Supervised [DPL^{+}23, LYZ^{+}24, LH22, LWHC24, LZF^{+}22, PSL^{+}24, QXG^{+}24, ZHY^{+}25, CYZ^{+}24, GHQ24, SYX23, SZZ^{+}23, GTP^{+}22]. Self-Trained [YWH^{+}17]. Selfies [HHLY19]. Seller [CSSZ19, CSSZ23]. Semantic [AAPLV25, AGCC25, CH21, DLZ^{+}17, DNL^{+}23, GSDT21, HH19, JLK^{+}24, LWY22, LZZ^{+}25, LX21, LNT^{+}21, LLYL22, LGW^{+}25, LYW^{+}25, LQX^{+}24, MHW^{+}19, MYX^{+}23b, NDC^{+}23, PA20, RBLJ25, RXY^{+}25, SSZ^{+}23, SLL^{+}21, TJN14, TGC23, TNZ^{+}25, WZNM14, WLCG21, WTL24, WXH^{+}25, WDC^{+}25, WLH^{+}21, WHW18, WWL^{+}23, WZW24, XWY21, XQG^{+}24, XGX^{+}25, XXG^{+}21, XDL^{+}21, XSL^{+}23, YZX16, YWG^{+}20, YLK^{+}20, YLL^{+}23, YZZZ25, YFLF19, ZZY^{+}14, ZLN^{+}16, ZML^{+}24, ZLJ^{+}25, ZCY^{+}19, ZZW^{+}19, ZSZ^{+}24b, ZLL^{+}22, ZXZ^{+}24b, ZHG^{+}25, FHP23, JP11, MTTH13, Yan10, ZI13, FLM^{+}06]. Semantic-aware [MYX^{+}23b, RXY^{+}25]. Semantic-Based [GSDT21, XQG^{+}24]. Semantic-driven [RBLJ25]. Semantic-rich [LQX^{+}24]. Semantically [YLL^{+}24, ZWZL22]. Semantics [GCTG24, GCC^{+}25, GLC05, LDZ^{+}24, TWL19, ZCX^{+}24, ZXD^{+}23a, SWH06]. Semantics-Augmented [LDZ^{+}24]. Semi [AdSGMS25, ASI^{+}21, FLG^{+}21, hHLC10, LLZ^{+}21, LLZ^{+}22, LZH^{+}24b, LWLC24, LJG^{+}24, MPTD22, NHZ^{+}24, SZB^{+}22, TSZ^{+}23, TSH^{+}25, WLFL23, ZZL^{+}17, ZLL^{+}24a, ZZW^{+}24, ZHY^{+}25, ZLZ22b, ZHLY11]. Semi-Automatic [AdSGMS25]. Semi-fragile [NHZ^{+}24]. Semi-Siamese [TSZ^{+}23]. Semi-Supervised [ASI^{+}21, LZH^{+}24b, LWLC24, MPTD22, ZZL^{+}17, ZHY^{+}25, ZLZ22b, FLG^{+}21, hHLC10, LLZ^{+}21, LLZ^{+}22, LJG^{+}24, SZB^{+}22, TSH^{+}25, WLFL23, ZLL^{+}24a, ZZW^{+}24, ZHLY11]. Semidiscrete [GLL^{+}24]. send [CBJ^{+}09]. Sensations [RD17]. sensed [CLN^{+}21]. Sensing [APV08, CCF^{+}24, EC16, GCF^{+}21, GHY^{+}24, HJMY15, HXL24, LL15, WDPX23, WGM25, ZDE16, ZSZ^{+}24b, DE12, JYZC12, MVW08]. Sensitive [AGC^{+}18, ACGH18]. Sensitivity [HZSC20, YDJ24]. Sensor [CWTG22, EC16, KEYY22, MC19, VCO15, YHQH17, CW10, FKFB05, GLWK19, Hae10, LCS09]. Sensor-based [CWTG22]. Sensor-enhanced [LCS09]. Sensor-Enriched [VCO15]. Sensor-Metadata [MC19]. Sensorial [GTLG14a, Ano13]. Sensors [CJZ^{+}24, Chu15, NRUT20, SVF12]. Sensory [RT14, RHE10]. Sentence [LYW^{+}23a, LZW^{+}19, LQD^{+}24, NCL^{+}23, XWH^{+}21]. Sentiment [CHHH25, CYM^{+}25, FFCM24, HWWL20, LLC^{+}24a, PN16, SSY20, SHSR24, WMM^{+}23, WZZ^{+}23, YV23, ZTR^{+}22, ZYW^{+}24, ZLJ^{+}25]. Sentiment-Oriented [FFCM24]. Separated [HYG^{+}25]. Sequence [BSSNF^{+}20, BR22, DLL^{+}24, MXH^{+}23, ZZW^{+}22]. Sequence-to-sequence [BSSNF^{+}20]. Sequences [AGCC25, LDZ^{+}20, HH08a, HH08b, LK07]. Sequential [HQF^{+}20, LZW^{+}23, LLL^{+}24a, LKM^{+}19, YP20]. serendipitous [CY11]. Series [LPC^{+}18, DŞB^{+}22, MON21]. Server [AS20, CE10, LDT^{+}18, SLP15, ZZL21, AH12, HB08, LJP08]. Server-side [AS20]. Servers [XHYX23]. Service [FAA18, GJAF18, LWZ^{+}21a, SSP21, REV^{+}12, ZCL^{+}12]. Services [CXW^{+}19, CZC15, PVWD18, SKVHC18a, SKVHC18b, XHZ^{+}21, XZL^{+}22b, ZHLJ22]. session [ZCL^{+}24]. Set [JXLD25, LDZ^{+}24, WDCX07]. Sets [PRGA18]. SETTI [GTP^{+}22]. SEVA} [LCS09]. Seventeen [HGKS24]. Sexism [MNPOF22]. Shadow [ZLZX23]. Shallow [SG22b, TSZ^{+}23]. Shape [AA23, HDH^{+}24, LKL25a, LCZ^{+}24a, NWNL21, SHF^{+}24, SWK^{+}22, YZS^{+}23, YCFX23]. Shape-driven [HDH^{+}24]. Shape-Texture [LCZ^{+}24a]. Shaped [FXF^{+}23, HWW25]. Shapes [ATMM24]. Shared [CS17, CLS17, CSSZ23, JCSL19, KFP^{+}25, LCS17, PHS^{+}20, XYJ^{+}20, YLK^{+}20, YCZ^{+}24]. Shares [SA16, YLLL21]. Sharing [FYZ^{+}21, yHcCzH^{+}21, LCL14, Lin15, LZW^{+}21, MFL^{+}16, SZZT18, WQTX25, WY23, XHYX23, YLLL21, YLS^{+}23, CBJ^{+}09, CY11, CBR14, CSJ^{+}08, CL12, SX13]. Sharpening [FDZ^{+}25]. Shift [LHC^{+}25]. Shifting [SHSR24]. Shoe [HLX^{+}14]. shooter [AH12]. Shopping [AWG^{+}15, JWF18]. Short [YZZZ25]. Shot [AM24, FJF^{+}25, GJY^{+}25, GGSV25, HHL^{+}22, LLH^{+}24, RYZ^{+}23, SS24, TY23, WDC^{+}25, XJB25, YZ23, ZLL^{+}24d, ZFC^{+}24, JMLL20, LWL24, LJW^{+}25, LNT^{+}21, PYBY25, SYY^{+}24, WLB^{+}25, WXJL25, XTL^{+}21, YBW^{+}25]. Shots [TWY24]. ShotVis [ZZC^{+}15]. Shoulder [XLH^{+}23]. Showdown [LKM^{+}19, LVM^{+}21]. Shuffle [SZL^{+}22]. Shuffle-invariant [SZL^{+}22]. Shuffled [MKS20]. shunted [FQJ^{+}25]. SHVC [SAF19]. Siamese [AY21, SED25, TSZ^{+}23, TCL^{+}23, ZWX24]. Side [XLF^{+}25, AS20]. Side-Channel [XLF^{+}25]. Sides [Hon19]. Sieve [DZW^{+}21]. SIFT [ZLLT13]. SigFormer [LLL^{+}24b]. Sightseeing [FMIS17]. SIGMM [RJ05]. Sign [GZLW18, HZPL21, TXW^{+}25]. Signal [DWS^{+}24, HCJ^{+}24, LLL^{+}24b, YP15]. Signal-based [DWS^{+}24]. Signal-guided [LLL^{+}24b]. Signals [LGT^{+}24, YHH^{+}22, YSY^{+}22, ZGD^{+}19, ZWF^{+}20, And13]. Signals-based [ZWF^{+}20]. Signatures [NY24, ASVE13]. Signed [CXX^{+}24]. Significant [XZY^{+}24a]. Similar [ZLL^{+}24d]. Similarities [CE23, RVTC25, SZW^{+}24]. Similarity [AGCC25, AR15, CZY^{+}21, FZZ^{+}25, HMOS17, KPK^{+}24, LKL25a, LCW16, LWLC24, LWWZ20, NSM^{+}24, PRGA18, SJC^{+}19, SHZ^{+}20, WDLW23, YZRW22, ZWZL22, DCM13, HOSS13, KS13, LZP07, NW08, VV11, WJQ^{+}22a]. Similarity-aware [NSM^{+}24]. Similarity-Preserving [SHZ^{+}20]. Similarly [ROST20]. Simple [DSL^{+}22]. Simplification [BLMP18, YK07]. Simplistic [GS19]. Simulate [RD17]. Simulating [SKW^{+}15]. Simulation [CGPCR18, CJP^{+}21, SHIE15, ZDZ^{+}23]. simulcast [HH11c]. Simultaneous [FMG20, LCC^{+}14b, CCG^{+}08, DCC^{+}13]. Single [CZL^{+}23a, CH15, CRL20, HJ23, HL24, HSR18, LSYM19, LCL^{+}21, LLZ^{+}20, LNT^{+}21, LLYL22, LKW^{+}22, LLP^{+}23, LNC^{+}25, QDX^{+}24, SHF^{+}24, SZB^{+}22, YW23, YXK^{+}25, ZSS20, ZWL^{+}24].

Single-domain [QDX^{+}24]. Single-Image [LNC^{+}25, YW23]. Single-shot [LNT^{+}21]. Single-stage [LLZ^{+}20]. singular [BWA13]. Site [SRDJ18]. Situation [HDZ^{+}15]. Situation-Aware [HDZ^{+}15]. Size [LHS^{+}21, PFC^{+}24]. Skeleton [AA23, CZZ^{+}24, CZW^{+}25, HLM^{+}22, NXL^{+}25, TLY^{+}22, WCG^{+}25, XW25, XLZ^{+}21, YJTN18, ZLP^{+}14, ZZW^{+}22]. Skeleton-Aware [CZW^{+}25]. Skeleton-Based [WCG^{+}25, XW25, XLZ^{+}21, YJTN18, TLY^{+}22]. Skeleton-Boundary-Guided [NXL^{+}25]. SKEPRID [YJTN18]. Sketch [HVC^{+}20, LLH^{+}24, LMF^{+}14, YHZ23]. Sketch-Based [LMF^{+}14]. Sketch-guided [HVC^{+}20]. Sketch-to-Photo [LLH^{+}24]. skew [SOC^{+}13]. Skews [AMMG16]. Skills [MON21]. skimming [LGX^{+}08]. Skin [RTR21]. Skip [ZLK^{+}19]. Skipped [PLZW18]. SLAM [SSZ^{+}23, SZZZ25]. slice [ZCXL23]. Slicing [AAT^{+}22, KN21b, YLLZ25]. Slicing-assigned [AAT^{+}22]. Slicing-based [YLLZ25]. Slide [SSP^{+}24]. Small [LLL23, QSY^{+}23, KS09]. Small-group [LLL23]. Smart [BCNP24, FYZ^{+}21, GZ20, GCL^{+}25, HJMY15, HCM^{+}22, JGJ^{+}20, LYW^{+}22, LYXA22, LCL22, NSJB17, RK15, XYJ^{+}20, YQC^{+}21, MVW08, SCFL14, PCB^{+}21]. Smartening [BPT^{+}15]. Smartphone [BFAS15, LL15, SCXC15, VCO15, ZZC^{+}15]. Smartphone-Based [SCXC15, ZZC^{+}15]. Smartphone-Controlled [BFAS15]. smell [GA12a]. SMIL [BB11]. Smooth [YSG^{+}06]. Smoothly [ACVM20]. Smoothness [FBSGR25, ZZTL24]. sMRI [KN21b]. SNIPPET [YJM^{+}24]. Social [APRS^{+}19, CSJC17, CSSZ23, CLN^{+}21, LCL14, LSK^{+}15, LWY22, LZW^{+}24, LQX^{+}24, MY15, NSK^{+}21, OL25, PRPO23, QZXH14, RUD23, SMTR22, SX13, SZZT18, TAS16, TVK18, TVZ^{+}19, WZZL25, XZH^{+}21, XWL^{+}25b, Yan17, YYY^{+}24, ZGL^{+}18, APV08, BRA^{+}09, BGP11, BLJX10, BJLX11, CBJ^{+}09, CYMW07, LCSX11, LSDK12, MHT^{+}08, RSB11, SX11, TBC^{+}11, WWL13, WZC^{+}13]. social-aware [WZC^{+}13]. Social-sensed [CLN^{+}21]. socialized [CL12]. Socializing [XZH^{+}21]. Socially [BCG13]. Socially-aware [BCG13]. Socio [SAL^{+}22]. Socio-historical [SAL^{+}22]. SOD [ZCHD23]. Soft [CZL^{+}23a, CFG^{+}24, WXJL25]. Soft-biometrics [CFG^{+}24]. Soft-label [WXJL25]. Software [ZLT^{+}25b, IB10, MPSR05]. Software-Defined [ZLT^{+}25b]. Solfège [SDJ17]. Solution [BQBLN18, HXH^{+}24, SRPH16]. Solutions [BLH^{+}25]. Songs [CYM^{+}25]. Sorting [HZW22]. Soul [HYSL20]. Soundscape [SRDJ18]. Source [GIL^{+}10, JGJ^{+}20, LZLJ22, MYX23a, WCL23, XJFS25, LSS^{+}24b]. Source-Channel [LZLJ22]. sourced [MC19]. sources [XC06]. Space [HLD18, JCB25, SZW^{+}24, XYJ^{+}20, XJFS25, YLK^{+}20, YDCZ23, YHH24, ZYCD24, CKRB23, ML11, RBLJ25]. Space-Time [ZYCD24]. Spaces [SGY^{+}24]. Spark [GJAF18]. Sparse [CJZ^{+}24, CZW^{+}25, FZW^{+}25a, GZL^{+}20, LLL^{+}24b, MBXZ24, TTR12, WCL^{+}22, YKQ^{+}21, YF22, ZZL^{+}17, ZDZ^{+}22, CML^{+}13]. Sparsity [DXY^{+}21, XG24, JGZ^{+}11, LYC^{+}12]. Sparsity-guided [XG24]. Spatial [FXMZ24, FQJ^{+}25, KMK^{+}21, LYZY21, LZJ^{+}20, LLC^{+}25, LKW^{+}22, RHE10, SGY^{+}23, TLY^{+}22, WLSN24, XJW^{+}22, YH14, YSZZ14, YYBX24, YFLF19, YCLH22]. Spatial-geometric [RHE10]. Spatial-Temporal [LLC^{+}25, WLSN24, XJW^{+}22, YSZZ14, YYBX24, YCLH22, FXMZ24, FQJ^{+}25, LYZY21]. Spatially [GHAM25]. Spatio [BR22, DVA21, HZS^{+}24, HYLD20, HLS^{+}22, LLJ^{+}20, LLZ^{+}24, LDZ^{+}20, DMSRL18, WLMZ25, WZL^{+}24, ZWLZ24, ZHG^{+}25, CWC10]. Spatio-Temporal [BR22, HLS^{+}22, LLJ^{+}20, DMSRL18, WLMZ25, ZHG^{+}25, DVA21, HZS^{+}24, HYLD20, LLZ^{+}24, LDZ^{+}20, WZL^{+}24, ZWLZ24, CWC10]. Spatiotemporal [HSG23, KFA^{+}24, LHL^{+}24, ZZL^{+}25a]. Speaker [HKYW14, LZX^{+}21, FYH10]. Speaker-Following [HKYW14]. Special [ACGH18, APRS^{+}19, BSH08, CCO15, CLS^{+}21, FSEW25, GZ20, GS11a, GTLG14a, GWZ25, GMWW25, HCM^{+}22, HT15, IPMD25, LPD^{+}24, MSG22, NLKB24, ODMD17b, PCH^{+}20, SWS21, SZEST21, SCXC15, SLK21, SKC24, SKVHC18a, TKPL20, VNDB24, WJXN24, WFZ^{+}21b, YNC18, YZRW22, ZZMS14, ZYZE19b, ZZX^{+}20, ZGD21, ZLZ^{+}22a, ZHLJ22, ZPR^{+}25, ZJSJ20, BLJX10, BJLX11, CCK06, CDGJ09, CBS08, GPHOH12, GDGC07, Hae10, Kan12, KZHC13, NLS13, SHOG12]. Specific [ZH18, ADCB07]. Spectator [Abd18]. Spectators [JSEI16]. Spectral [HZSC20, YGNT19]. Spectrum [DMF^{+}20, WBRZ17, ZLT^{+}25b]. Spectrum-Based [WBRZ17]. Speech [APM21, BRG24, BPFA24, DYL^{+}25, JD23, KP15, LCL^{+}24a, LLC^{+}24b, UFJ21, WWH17, YMA17, CL07]. Speech-based [UFJ21]. Speech-Related [LLC^{+}24b]. speech-to-text [CL07]. Speeds [YKW^{+}22]. Sperm [NDX^{+}21]. SPGAN [LLJC21]. Sphere [ZZZ^{+}22]. Spherical [LHZ^{+}23, LLA^{+}21]. Spider [KH13]. Spiking [LXZ^{+}25]. Splicing [WWX^{+}21]. SplitFed [SKA24]. Spontaneous [QRTM24, WHJ20]. Spoofing [LLJC21, LCL^{+}24a]. Sporting [AE22]. Sports [CED^{+}16, TCJ08, XC06, XXD^{+}08]. Spot [VVSV17]. Spott [VVSV17]. Spotting [LWZ^{+}25, WTL^{+}25]. Spread [ZLT^{+}25b]. Sprite [CDA12]. Square [SZFR23]. SSAT [XWZ^{+}24b]. ST [LLC^{+}25]. ST-360 [LLC^{+}25]. Stabilization [WLL23]. Stacked [ZHS20]. Stage [YXJ^{+}24, ZYCD24, LLZ^{+}20, ZMX^{+}23]. Stall [ZQRS18]. Stall-Aware [ZQRS18]. Standard [SK23]. State [CDA^{+}24, GTLG14b, LMS^{+}24, WDE25, WF25, LSDJ06]. State-of-the-Art [CDA^{+}24, WDE25, WF25]. States [LGT^{+}24, ZQKH19]. Statistic [ZYW^{+}25]. Statistical [GHP^{+}06, HH11b, ZLK^{+}19]. Statistically [ZDD^{+}24]. Statistics [LCZ^{+}24a, LGLZ20, TWB^{+}23, WCZ^{+}21]. Status [CXL^{+}22, MYGX21, PVWD18]. Stealthy [ZYW^{+}25]. Steganalysis [PJL23, YI14, LSQ11]. Steganalysis-resistant [PJL23]. Steganalytic [Ber18]. Steganographer [ZWR^{+}20]. Steganography [SLZ24, WYGC24, ZZL^{+}25d]. Stent [HLZ^{+}20]. step [ZLS^{+}25]. Stereo [Hon19, YF22, YP15, ZHG^{+}21, ZT22, ZPL^{+}23, DÇ07]. Stereoscopic [CED^{+}16, Chu15, LL23, ZHG^{+}21]. Still [LYLL23, ZLL20]. stills [TB05]. Stimulation [RD17]. Stimuli [LGF^{+}14]. Stitch [LXL^{+}23]. Stitching [ZZCZ21]. Stochastic [CXW^{+}19, WZZ^{+}24, GL12]. Stop [SGYX22]. Storage [EG17, WCL23]. stories [RSB11]. story [JSL07, JWL06]. story-oriented [JSL07]. Storytelling [WPRC18]. Strategies [CDZ^{+}17, SSSK18]. Strategy [HZC^{+}16, WHW18, WHY19, ZSZ^{+}25, FLZ^{+}12]. Stream [GKSB17, WHW^{+}21, XLZ^{+}21, XJW^{+}22, XZL^{+}25, ZYH^{+}24, ZHL^{+}25, KMP05, LQG25, LCL^{+}24b, QWH^{+}21, ZCAP08]. Streamed [CFN^{+}23, DCM13, HH11b]. Streaming [AH20, AS22a, AS20, BPM15, BBZ18, BYOZ20, BTBZ20, BLH^{+}25, BRZS18, BQBLN18, CED^{+}16, CKRB23, CNG22, CDZ^{+}17, DLC^{+}22, FZW^{+}25b, TCP^{+}20, GGB14, HSN^{+}14, HJWW19, HXZ^{+}20, HSL^{+}22, HSS^{+}25, JSEI16, KPL^{+}22, KFP^{+}25, KABB20, LYJ^{+}15, LCS17, LZLJ22, LTHT24, MAGT23, MN16, MATW17, NTT20, PB14, PFC^{+}16, PVWD18, PHS^{+}20, RXC14, SSP^{+}24, SS17, SWZ25, SYS17, TAT^{+}25, TAPP^{+}15, TNH^{+}21, VTP^{+}20, WCO15, WDE25, WCY^{+}18, WLW^{+}23, BLS^{+}19, YNY25, YM24, YJM^{+}19, Zha19, ZMH^{+}20, ZLP^{+}14, ZLOL18, Bag11, COM^{+}11, GS11b, GSM^{+}08, GKW08, HH08a, HH11a, ILL08, JC10, LLP06, LCK09, LLKL11, LWLZ13, LLSC12, LCT^{+}12, LH12, MCM^{+}09, REV^{+}12, SKSZ13, SAAH10, SZ12, Swa13, TSHP05, TNEcC08, WKST08, WWGT09, WCK05, WLZ08, WLQL12, WLZ12, ZSO13, ZC12]. Streams [BXMH15, GHAM25, HLM^{+}22, RSR25, WGM25, AKO07, HM10, JP11, LZP07, LSDK12, QS10, San11]. Stress [CXL^{+}22]. Stretching [SB23]. Stroke [AM24, RLAM25, CXS^{+}08]. Strongly [SSY20]. Strongly- [SSY20]. Structural [LZZ^{+}24a, WJQ^{+}22a]. Structural-similarity-based [WJQ^{+}22a]. Structure [CAJ19, CZY^{+}21, FH20, FNH22, HPLL23, LCL23a, LPY^{+}19, LPW^{+}22, LDQ^{+}25, LW24, LZW^{+}21, MGY22, WCLC18, WLCG21, YZJ^{+}25, YZY^{+}25, YFLF19, ZLF^{+}23, ZYC^{+}24, ZLL^{+}24d, ZHL^{+}25, LSDK12, SG07, XMST07]. Structure-Aware [CAJ19, WCLC18, ZHL^{+}25, LCL23a, LPW^{+}22, MGY22]. Structure-Guided [LDQ^{+}25]. Structure-Learning [FNH22]. Structure-Preserving [LPY^{+}19]. Structure-Texture [YZY^{+}25]. Structured [BH05, LYZX21, XLZ^{+}21, ZLP^{+}14, ZJL^{+}21, ZZCZ21, JKKL08]. Structures [Tas20, WFC^{+}25b]. Student [BPT^{+}15, XWW^{+}25]. Study [AS22b, HHH22, LSL^{+}20, MHF24, NN21, WWY^{+}21, ZDZ^{+}23, ZFC^{+}24, DY09, TSHP05, WKST08]. Studying [RD17]. Style [ADMM25, CSW^{+}25a, LCL23a, LCZ^{+}24c, MZH^{+}24, SZM^{+}19, SHF^{+}24, WLZF22, WWY^{+}25, WQTX25, QDX^{+}24, SWH06, ADMM25]. Style-based [ADMM25]. Style-FG [ADMM25]. Style-Mixed [SHF^{+}24]. StyleInject [ZBYZ25]. Styles [ACVM20, SWS^{+}22]. Stylistic [YMY^{+}21, ZGY^{+}24a]. Stylization [LTD^{+}21]. Stylized [ZFFX24]. Sub [JGY^{+}24]. Sub-Images [JGY^{+}24]. Subject [TSP^{+}24]. Subjective [APRS^{+}19, YJM^{+}24, ZSZ^{+}24a, NH10]. subset [AKO07]. Subspace [FBG22, XWDC23, YWH^{+}17, ZRZ^{+}21]. Subspaces [ZWH^{+}23]. substitutable [WLH^{+}25]. Substituted [HD19]. Subtitle [YHZ19]. Subtitles [HKYW14]. Subtle [HPH^{+}20, XWL^{+}25b]. Subtractive [KMK^{+}21]. success [GA12a]. successes [Row13]. Suggesting [JYZC12]. suggestion [ZYM^{+}10]. summaries [VM12, MA10]. Summarization [AE22, GWM^{+}14, HP17, LPW^{+}22, SSBT20, TVK18, ZZZ14, ZLL^{+}22, HTT^{+}11, LGX^{+}08, MTTH13, SXM^{+}06]. Summary [ZYW^{+}25]. sung [ML11]. Super [FYW^{+}24, GHR^{+}22, GCZ^{+}24, HJ23, HPLL23, JXTC21, LSYM19, LCL^{+}21, LZWL25, LZW^{+}23, MGY22, NNK25, SYM^{+}25, SGY^{+}24, WXZ^{+}23, WLTL23, WGL^{+}23, YW23, YLLL23, YWL^{+}24, ZSS20, ZYCD24, ZLF^{+}24]. Super-Resolution [FYW^{+}24, GHR^{+}22, GCZ^{+}24, HJ23, HPLL23, LSYM19, LZWL25, LZW^{+}23, SYM^{+}25, SGY^{+}24, WLTL23, YLLL23, YWL^{+}24, ZSS20, ZYCD24, JXTC21, LCL^{+}21, MGY22, WGL^{+}23, YW23, ZLF^{+}24]. Superpixel [Li23, LGC^{+}24]. Supervised [ASI^{+}21, CLZ^{+}23a, DPL^{+}23, DZWH23, DWL^{+}23, FZZ^{+}25, HLW^{+}24, JLZ^{+}21, LZZ^{+}23a, LZH^{+}24b, LYZ^{+}24, LWLC24, LZZ^{+}25, LH22, LWZY22, LGW^{+}25, LWHC24, LZF^{+}22, MPTD22, PSL^{+}24, QZXH14, QXG^{+}24, SSY20, TCS^{+}25, TYY^{+}18, TS22, WZTL19, XXG^{+}21, XZZL23, YZW^{+}24, ZWX24, ZZL^{+}17, ZJZC20, ZHY^{+}25, ZLZ22b, CYZ^{+}24, FLG^{+}21, GHQ24, hHLC10, LLZ^{+}21, LLZ^{+}22, LJG^{+}24, SYX23, SZZ^{+}23, SZB^{+}22, TSH^{+}25, WLFL23, ZZL^{+}24a, ZLL^{+}24a, ZZW^{+}24, ZHLY11, GTP^{+}22]. Supervision [HXH^{+}24]. Supplement [Ano11, Ano12, Ano14, Ano20, Sha21, TOM12]. supplemental [Dao17]. Support [CHHH18, CJP^{+}21, PB14, TGSF21, TB17, ZCS^{+}20, ARE13, ONH07]. Supporting [AWG^{+}15, FDKB11, MBK^{+}23]. Suppression [CZW^{+}24]. Surgery [CGPCR18, WJS^{+}21]. Surround [SZZZ25, ZDZ^{+}22]. Surround-view [ZDZ^{+}22]. Surveillance [ATS19, CF22, GZ20, GZH17, KWK25, KD18, LMC^{+}22, LLZ^{+}24, NAK15, SKR09, SVF12]. Survey [AGC^{+}18, AC19, CCD24, CLWW24, DOD23, DDL^{+}25, GF17, HLM^{+}24, KKGE18, LYW^{+}23a, LZG^{+}24, MAX^{+}24, NAK15, NSJB17, ODMD17a, PWX^{+}24, SAL^{+}21a, SS22, SKVHC18b, Wan21, WHX^{+}25, WF25, ZDE16, ZZL^{+}25c, ZWS^{+}20, ZLX^{+}25, ZGL^{+}18, HMVI13, KS13, SCFL14]. SVC [FSK^{+}15]. sweet [GA12a]. Switch [DMF17]. Switch-Based [DMF17]. Switching [APM21, LTHT24, Zha19, HH11c]. Symbiosis [MHF24]. Symmetrical [LSYM19]. Synchronization [Abd18, MLQM14, XW17, HNS13]. Synchronous [ZLS^{+}25]. Synchronously [GCC^{+}25]. Syncretic [YHH24]. Synergy [PSSZ24, XSL^{+}23]. Synergy-Enhanced [PSSZ24]. Syntactic [LYL^{+}25]. Syntactic-Conditional [LYL^{+}25]. Synthesis [ADMM25, CRL20, CSW^{+}25b, GYY^{+}25, GZT21, GHW^{+}24, HYW^{+}25, JXL^{+}25, JLL24, KBI^{+}23, LJL^{+}24, LLH^{+}24, LCH^{+}25, LLP^{+}23, LZL^{+}24c, TM25, WWL20, XCM^{+}23, MSL10, YK07]. Synthesising [AAA^{+}21]. Synthesized [IKTD25, LBD08]. Synthesizing [SSK20]. Synthetic [CSBJ24, DYL^{+}25, DBPM25, DŞB^{+}22, IPMD25, LXJ^{+}24, VF25, WF25, XQG^{+}23, ZDZ^{+}23]. System [ATS19, BBRS23, CHHH18, CJHH21, CS17, DMF17, GAH25, HDW^{+}18, HZW22, JGJ^{+}20, KT21, LWY22, LCC^{+}14a, LLJW15, LCC^{+}14b, LLO^{+}20, LGF^{+}14, MY15, MFL^{+}16, MDAE19, PCB^{+}21, RPE^{+}17, ROST20, SZZZ25, SAF19, SWM^{+}24, SWZ25, SRAA17, SSSA23, TB17, WJXN24, WLS^{+}24a, WCY^{+}18, WZZL25, YBO14, ZDZ^{+}22, ZWLZ24, ZCS^{+}20, ARE13, BBT^{+}05, CEE09, JWL06, JKKL08, KH13, ONH07, RWW05, SNC12, TCJ08, VGNL10, WT10, YH13, ZRCH08]. Systematic [HC22, TV07]. Systems [AAT^{+}22, ACC^{+}21, CHHH18, CSZ^{+}25, DKJ^{+}21, HC22, KWK25, KD18, LCY^{+}23, MN16, MSG22, NLKB24, NAK15, SS17, SCXC15, TFY^{+}25, TB17, WDS21, ZYL^{+}24, ZZC^{+}25b, ZTB20, CCK06, CPP^{+}13, CE10, CW10, DY09, GDGC07, HAE11, ILL08, JKKL08, She13, SKR09, Sun13, TSHP05, YH13, ZCL^{+}12, ZSO13, CPP^{+}14].

T [WLSX24, WWWL23, ZZL^{+}24c]. T-GAN [WWWL23]. T2C [YZD^{+}25]. Table [Ano11, Ano12, Ano14, Ano20, BLW^{+}24, Dao17, Sha21, TOM12]. Tactile [KTK^{+}17, MBK^{+}23]. Tag [YSZZ14, CBJ^{+}09, MPE^{+}11]. tagging [TCW^{+}13, YYS13]. Tags [PDD16, Yan17, LYC11]. Tail [TSZ^{+}23, RUD23]. Tailed [WWY^{+}24, LZZ^{+}24b]. Taking [FKW22, HHLY19]. Talking [LCH^{+}25, LW23, LLC^{+}24b, SLL^{+}24]. tamper [NC13]. Tampering [SRAA17]. Tape [CFP15]. Target [Cla18, GCC^{+}25, JXLD25]. Target-Centric [GCC^{+}25]. targeted [RW12]. Targets [CCG20]. Task [BBUD24, GGSV25, HMUC21, LLC^{+}21, MON21, PKLK23, WYM18, WWC^{+}24, WWZ24, WBR25, YXZQ22, ZCHD23, ZZC^{+}25b, AG13, CWTG22, GCF^{+}21, JXTC21, ZSLW23, ZZW^{+}19, ZJT^{+}23]. Task-Aware [GGSV25]. Task-centric [PKLK23]. Task-independent [MON21]. Task-oriented [BBUD24]. Tasks [Che24, CSŞ^{+}25, SAL21b]. Taste [RD17]. TCP [GKW08, TNEcC08, WKST08, WWGT09, WCK05]. TCP-friendly [WCK05]. TDMA [BOZ17]. Teacher [XWW^{+}25]. Teacher-Student-Based [XWW^{+}25]. Teaching [Hua23]. team [XC06]. TEC [WFC^{+}25a]. TEC-CNN [WFC^{+}25a]. Technique [MSS^{+}24, NPG^{+}22, SG22a]. Techniques [DZW^{+}21, IKTD25, LZL20a, TRK^{+}20, VVSV17, WF25, DY09]. Technologies [yHcCzH^{+}21, LLL^{+}25, SCXC15, FKFB05]. Technology [WDS21]. Teenagers [MNPOF22]. Tele [WJS^{+}21, WAK^{+}12, YWN^{+}10a, YWN^{+}10b]. tele-immersive [WAK^{+}12, YWN^{+}10a, YWN^{+}10b]. Tele-Surgery [WJS^{+}21]. Telehaptic [GNC17]. Teleoperation [CXC^{+}17]. Telephone [Cho13]. Teleportation [Cho13]. Television [VVSV17, CVV06, CBS08, MHT^{+}08, OBBW12, Cho13]. Tell [ZXX22]. Tells [CGW^{+}24]. Temporal [APC^{+}25, BR22, CFGW05, DZWH23, FBSGR25, HSR^{+}23, HLS^{+}22, ICZ^{+}22, LYW^{+}23a, LLJ^{+}20, LSXZ23, LCH^{+}25, LLZ^{+}25, LLC^{+}25, LZZ^{+}25, LPC^{+}18, LQD^{+}24, PWHP24, DMSRL18, SHZ^{+}20, TLY^{+}22, WLZ^{+}24a, WLSN24, WLMZ25, WGHX24, XZZL23, XLW23a, XJW^{+}22, YYX^{+}24, YHQH17, YSZZ14, YYBX24, YCLH22, ZLT^{+}25a, ZHG^{+}25, ZXS^{+}25, BZ05, CWC10, DVA21, FXMZ24, FQJ^{+}25, HZS^{+}24, HYLD20, HNS13, LYZY21, LK07, LLZ^{+}24, LDZ^{+}20, MHT^{+}13, QHR^{+}08, WCK05, WZL^{+}24, ZWLZ24]. Temporal-Consistent [WLZ^{+}24a]. ten [JPS05]. Tennis [BLW^{+}24]. Tensor [DLL^{+}20, FKCW20, FMG20, GYF^{+}21, LLS^{+}21, LYW^{+}22, WFC^{+}25a, ZYW^{+}25, ZYL^{+}24]. Tensor-based [FKCW20]. Tensor-Empowered [ZYL^{+}24]. Term [CLZ^{+}23b, SLBS20, ZZC^{+}25b, ZTX^{+}23]. terminals [CVV06, LC12]. Termination [ZLK^{+}19]. terminology [ENHN09]. testbed [HH12b]. Testing [REP^{+}19]. TEVL [MSC^{+}23]. TexIm [AGCC25]. Text [AGCC25, CLZ^{+}23a, DHP23, EFF^{+}25, FSEW25, FXF^{+}23, FZZ^{+}25, HCZ^{+}22, HZS^{+}24, HSR^{+}23, HDH^{+}24, HWWL20, HPLL23, JCB25, KBI^{+}23, LLS^{+}22, LZW^{+}19, LZL20b, LZL21b, LGW^{+}25, LWZ^{+}25, MKY^{+}24, MSFR25, PYZ25, RYZ^{+}23, SXY^{+}23, SCZJ24, TM25, THR^{+}22, WGHX24, XFZ^{+}19, YMY^{+}21, YLL^{+}23, YHP25, YZW^{+}24, YZD^{+}25, YBTX23, ZXX22, ZZG^{+}20, ZSZ^{+}24b, ZGY^{+}24b, ZBYZ25, ZWZ^{+}23b, ZFFX24, CL07, CZQ^{+}22, FHP23, JWL06, YLL^{+}24]. Text-Based [EFF^{+}25, FZZ^{+}25, CLZ^{+}23a, ZGY^{+}24b]. Text-Driven [SCZJ24]. Text-Guided [TM25, YZD^{+}25, ZWZ^{+}23b]. Text-Image [JCB25]. Text-Multimedia [FSEW25]. Text-Supervised [LGW^{+}25]. Text-to-Face [MKY^{+}24]. Text-to-Image [AGCC25, DHP23, ZBYZ25, YHP25]. Text-to-Motion [MSFR25]. Text-to-Video [HSR^{+}23, YZW^{+}24]. Text-Video [HCZ^{+}22, HZS^{+}24]. TextBlockV2 [LWZ^{+}25]. Texts [XZY^{+}25]. Textual [EL25, HZL^{+}25, LZL^{+}24c, MAE^{+}21, QSZ^{+}21, SSBT20, YMX^{+}16]. Texture [LCZ^{+}24a, LDQ^{+}25, YZS^{+}23, YZY^{+}25, ZCQ^{+}23, ZLF^{+}23]. them [Dao17]. theoretic [PBS12]. Theory [BYOZ20, Zho16, ZWM12a]. Thermal [KBB21, SZC24, TXWK24]. Things [AAA^{+}21, GYF^{+}21, LWZ^{+}21a, LQS21, SP21, SMTR22, SZEST21, SKC24, YWF^{+}24]. Things-based [SMTR22]. Thinking [WPRC18]. Thought [LLC^{+}24a]. thousands [ZR13]. Three [LYL23a, NWL^{+}20, NWNL21, ZCXL23].

Three-dimensional

[NWL^{+}20, NWNL21, ZCXL23].

Three-phase [LYL23a]. Three-View [ZCXL23]. Threshold [LWLC24, LZW^{+}21, LZD^{+}22]. Throughput [MATW17]. Thumbstick [Cla18]. Tibetan [Hua23]. Tier [MSKYJ21]. Tightly [SSZ^{+}23]. Tile [NTT20, VTP^{+}20, YM24]. Tile-based [VTP^{+}20]. Tiling [YM24]. Time [BPM15, BZDX^{+}18, CXL^{+}22, DG17a, DYSX14, DŞB^{+}22, DWC^{+}21, FHG^{+}17, GGA^{+}20, yHcCzH^{+}21, HRL^{+}24, LPC^{+}18, MON21, MDAE19, SAZ^{+}15, TWM^{+}25, WHL^{+}23, YZL^{+}24, ZYW^{+}25, ZYCD24, EGEM06, GLZW23, HCKL13, IB10, KK08, KMP05, KW11, MPM^{+}25, RSR25, SAAH10, SNC12, YSG^{+}06]. Time-Aware [DYSX14]. Time-aware-based [CXL^{+}22]. Time-series [DŞB^{+}22, MON21]. Timing [CSJC17]. TinyPredNet [DLL^{+}24]. TMIV [HHH22]. Today [LXL^{+}14]. Token [FXMZ24]. Tokens [GCTG24]. Tolerant [LZY^{+}24, LYL23a]. TOMCCAP [Geo05]. TOMM [Dao17]. Tomography [RLAM25]. Tongue [RD17]. Tool [HHLY19, SDJ17, BB11]. Tools [RABC24]. Top [GWM^{+}14, PLYM23, LMXJ21]. Top-Down [GWM^{+}14, PLYM23]. Topic [BXMH15, WXQC20, SX11]. Topic-aware [WXQC20]. Topics [YGNT19]. topographic [WYM07]. topography [VPSS^{+}13]. Topological [LCW16]. topologies [WLZ08]. Topology [GWWC25]. Touchable [CEE09]. Tourist [DNPG^{+}17]. Tours [VMP20]. TP [LYL23a]. TP-FER [LYL23a]. track [WYM07]. Tracking [AY21, CZW15, CJZ^{+}24, CLZ^{+}23b, GCC^{+}25, GZH17, GZL^{+}20, GQG^{+}24, HPZ25, LYZY21, LZL^{+}23, LYZ^{+}24, LW24, LL15, MHCG19, SZF^{+}25, TXWK24, WJS^{+}21, WHL^{+}23, WJQ^{+}22a, WZL^{+}24, XWZ^{+}24b, YYBX24, YCLH22, ZX14, ZWL^{+}24, ZYZ^{+}25a, ZZL^{+}24c, ADCB07, CCD07, FPH^{+}08, GDGC07, KO11, WT10]. Tracklet [YYBX24]. Tradeoff [LZZ^{+}24c]. Traffic [GGA^{+}20, KW11, XFQ^{+}21, GIL^{+}10]. Train [GYF^{+}21, LYW^{+}22, ZZC^{+}25a]. Trained [LZZ23b, YWH^{+}17, LWZ^{+}25]. Training [CGPCR18, FNH22, IKTD25, LXRL23, LJW^{+}25, PDD16, PKL25, SS24, TSZ^{+}23, ZSZ^{+}22, LFP^{+}22, LGX^{+}08, WDCX07]. Trans [BV25]. Trans-Convo-Former [BV25]. Transactions [BLJX10, Geo05, HGKS24]. Transcoding [HJWW19, RSR25, SLP15, LS05, LLSC12, LC12, LCT^{+}12]. Transductive [WXJL25]. Transfer [BTWZ22, DJWZ25, GSY^{+}24, HLZ^{+}20, JGJ^{+}20, JJML24, KN21b, LCL23a, LWP22, LWZW21, LZG^{+}24, MZH^{+}24, SZL^{+}24, SZM^{+}19, TLZL24, WLZF22, WXH^{+}25, WWY^{+}25, WQTX25, WHY19, WWWL23, WZW24, XYJ^{+}20, XZC^{+}25, YZXY15, ZCY^{+}19, TTR12, YYS13]. Transferability [WLMZ25]. Transferable [DYL^{+}25, WLCH22]. Transferrable [CSW^{+}25a]. Transform [FCL^{+}22, GGML22, LLA^{+}21, LQD^{+}24, WSW^{+}24, YWNW15]. Transform-Equivariant [LQD^{+}24]. Transformation [FCL^{+}22, LZW^{+}23, XJFS25, YZY^{+}25, GZHD12]. Transformation-guided [FCL^{+}22]. Transformations [GHQ24]. Transformer [AM24, CZ24, CSW^{+}25b, FPA24, FFCM24, GHZW25, HWW25, HCZ^{+}25, HPZ25, KFA^{+}24, LCL^{+}23b, LLG^{+}23, LML^{+}24, LZL^{+}25, LYDY24, LW24, LLL^{+}24a, LLL^{+}24b, LCL^{+}24b, LWW^{+}25, LHC^{+}25, MOL^{+}22, MAE^{+}21, PLB^{+}24, QRTM24, RTM^{+}24, TWFW24, TCL^{+}23, WCCN23, WLH^{+}25, XLW^{+}23b, XYC^{+}23, YLD^{+}25, YLLZ25, YXK^{+}25, YCZ^{+}23, YTL^{+}24b, ZTX^{+}23, ZHL^{+}25]. Transformer-Based [FFCM24, LLG^{+}23, TWFW24, ZHL^{+}25, CSW^{+}25b, LML^{+}24, QRTM24]. Transformers [FXMZ24, LYZ^{+}24, TJZ^{+}25].

Transition [HLW^{+}25]. Translation [Hua23, LLH^{+}24, LYH24, ZY21]. Transmission [BOZ17, EWSZ15, HSL^{+}22, ZZL^{+}24b, ZWZL21, CLC05, KW11, LC12, MC11, PBS12]. Transparent [LLWX24]. Transport [GLL^{+}24, OMP07, YSG^{+}06]. transport-level [OMP07]. Transportation [SWM^{+}24, WJXN24, ZWLZ24, ZYL^{+}24]. Treatment [ZCS^{+}20]. tree [LS05]. Trees [LQZH14]. Trends [Sin21, SLK21, VNDB24]. Tri [YZZZ25]. Tri-Semantic [YZZZ25]. Trident [LHL^{+}24]. Triggers [HYW^{+}25]. Trilinear [MSC^{+}23]. Triple [WCG^{+}25, YZS^{+}22, ZGY^{+}24a]. Triple-Network [YZS^{+}22]. Triplet [BHD24, JWF18, LWL24, TH22, YPSC22, ZYO20, ZGR21]. Triplet-Consistency [YPSC22]. Triplets [WZZ^{+}22]. TripRes [XFQ^{+}21]. Tropical [SYM^{+}25]. true [LWH^{+}22]. Trust [GHS^{+}24, LS21, MY15]. Trustworthy [NY24]. Try [FCL^{+}22, GZW^{+}24b, LZZ^{+}24c, RTM^{+}24, SLG^{+}24, WC23, WLL^{+}24]. Try-On [GZW^{+}24b, RTM^{+}24, WC23, WLL^{+}24, FCL^{+}22, LZZ^{+}24c, SLG^{+}24]. Tryon [PKL25]. TSFormer [LW24]. TSK [JGJ^{+}20]. TSVD [LYW^{+}22]. TT-TSVD [LYW^{+}22]. TTV [SG22a]. Tucker [ZYL^{+}17]. Tumor [JXL^{+}25, KMK^{+}21]. Tuning [JJML24, SHWC19, ZBYZ25, FZYY18, TJZ^{+}25]. Turning [WXCW24]. TV [HH12b, HH11c, KS09, UTK^{+}08, WWL13, YH13]. twenty [Eff13, She13, Tur13]. Twice [ZSYW23]. Twin [TGSF21]. Twins [AE22, LCL22, NY24, WCL23]. Twitter [DYSX14]. Two [HLM^{+}22, HSL^{+}20, LWLZ13, LQG25, LCL^{+}24b, NWL^{+}20, QWH^{+}21, SJC^{+}25, WY23, XWL^{+}25b, YXJ^{+}24, YZS^{+}22, ZMX^{+}23, ZYH^{+}24]. Two-Branch [SJC^{+}25]. Two-dimensional [NWL^{+}20]. Two-in-One [WY23]. Two-Layer [HSL^{+}20]. Two-Level [YZS^{+}22]. Two-person [LCL^{+}24b]. Two-Stage [YXJ^{+}24, ZMX^{+}23]. Two-Stream [ZYH^{+}24, LQG25, LCL^{+}24b, QWH^{+}21].

U [FHF^{+}24, HWW25, PA20, WWL20, YXYB21]. U-Net [FHF^{+}24, PA20, WWL20, YXYB21]. U-shaped [HWW25]. UAV [PSN^{+}24, YCLH22, ZZC^{+}25b]. UAV-Enabled [ZZC^{+}25b]. Ubiquitous [BPT^{+}15, SSTK07, ARE13, CTGP08]. UDA [JCB25]. UEFPN [QSY^{+}23]. UGC [TFHM24]. UHD [GGML22]. UID2021 [HLY^{+}23]. Unaligned [MXH^{+}23]. Unaware [YWC^{+}25]. Unbiased [JLK^{+}24, PLZT22, SZC24, XLW^{+}23b, YLYB24]. Uncertainty [ASI^{+}21, HMUC21, KN21a]. Uncertainty-Aware [ASI^{+}21]. Unconditional [AAPLV25]. unconscious [And13]. Uncoupled [WDLW23]. Under-Exposure [HSR18]. Underlying [SZW^{+}24]. Understanding [APRS^{+}19, BBT^{+}05, CS22, GF17, GMWW25, HLTH24, LCL14, RBLJ25, TVZ^{+}19, UJLS22, VGNL10, YNC18, YZX16, ZMZ^{+}23]. Underwater [DBPM25, HLY^{+}23, HZZ^{+}24, LXJ^{+}24, SBS23, WCZ^{+}21, YSSF24]. unequal [CLC05]. Unexpanded [SA16]. Uni [LFP^{+}22]. Uni-EDEN [LFP^{+}22]. Unified [FKCW20, FMG20, JCB25, MKY^{+}24, QSY^{+}23, TZZ25, XGX^{+}25, YSXH16, ZZY^{+}14, ZZT24, ZSX^{+}25, ZFC^{+}24, YGHH12]. Uniform [FBSGR25, GHZW25, LQZH14]. Uniforms [XWH^{+}25]. Unifying [TCL^{+}23, YTY^{+}24]. Union [LXL^{+}23]. UniQRNet [YTY^{+}24]. United [HL24]. Units [LLC^{+}24b]. Universal [DYN^{+}25, LFP^{+}22, TYL^{+}25, WLZF22, ZYZ^{+}25a, ZLL^{+}24b]. Unlabeled [ASI^{+}21]. Unleashing [EAK^{+}25]. Unpleasant [KYVE14]. Unposed [LLP^{+}23]. Unrelated [XLH^{+}23]. Unsupervised [BR22, DWS^{+}24, DFXY20, FZYY18, FDBP23, KLS^{+}18, LSL^{+}24, LLP^{+}23, LLN^{+}24, LZL^{+}24b, LWHC24, PZJL22, PRGA18, SO22, TDW^{+}23, WTD^{+}21, WWQ^{+}25b, WLSZ22, XHL^{+}24a, ZLY^{+}24, ZXT^{+}25, ZLL20, ZZZ^{+}22, ZLL^{+}22]. Untrimmed [LZF^{+}22]. Unveil [RVTC25]. Unwanted [MY15]. Up-conversion [HYLD20]. Up-Fusion [WRK14]. Update [YWLZ23, LW07]. Updates [ZWZL21]. Updating [YF22]. Upgrading [LLO^{+}20]. Uplink [EWSZ15]. Upsampling [LSS^{+}24a, YQZP24]. Urban [Ala21, GCF^{+}21, LB15, QDX^{+}24]. Usage [REP^{+}19, MMW10]. Use [AdSGMS25, HLY21, YSZ15, CCK06, SSTK07]. Usefulness [AdSGMS25, Tas24]. User [BC15, CE23, CYM^{+}25, CSZ^{+}25, CS17, CSSZ23, EG17, FHH22, FHG^{+}17, GA12b, HDW^{+}18, KPK^{+}24, LCY^{+}23, LLJW15, LCC^{+}14b, MZL^{+}18, ROST20, RVTC25, SLZ^{+}21, TYY^{+}18, XLF^{+}25, YSXH16, YBO14, YCGM14, ZZZ14, ZGR21, And13, CKRB23, GG06, ZYM^{+}10, ZCY^{+}13, ZSO13, XWZ^{+}24b]. User-Click-Data-Based [TYY^{+}18]. user-generated [ZCY^{+}13]. User-Preference-Aware [LLJW15]. Users [ROST20, CY11, HWY^{+}11]. Using [ARA^{+}23, ASLA18, ASI^{+}21, AM24, AS20, ATMM24, AY21, BZDX^{+}18, CWTG22, CI23, CGGC20, CJZ^{+}24, CSJC17, CNG22, Chu15, FYW^{+}24, FNH22, FSK^{+}15, FHG^{+}17, GGSV25, GJAF18, GZL^{+}20, GWWC25, GSDT21, HZC22, HLM^{+}22, HD19, HMOS17, HH11c, HHH22, HSL^{+}20, IKTD25, JLL^{+}21, KFA^{+}24, KP15, KEYY22, LYLL23, LQZH14, LCS17, LLJC21, Li23, LZL^{+}24a, LCC^{+}14a, LYW^{+}23b, LHJ^{+}24, LLWX24, MAZ22, MON21, MZZ^{+}20, MHCG19, MAE^{+}21, MATW17, NCMM21, NRUT20, PHS^{+}20, PS17, QRTM24, QJ23, RYZ^{+}23, RSE16, RHS^{+}20, SR22, SSK20, SL22, SLBS20, SO22, SK23, SVA^{+}21, SRAA17, SZFR23, TBC^{+}11, TRRB20, VVSV17, WYM07, WHJ20, WSLM18, XWY21, XJG^{+}22, XWZ24a, XZLZ23, Yan17, YLZ^{+}21, YWNW15, YW23, YLD^{+}25, YM24, YZY^{+}25, YP15, YZ23, ZZC^{+}23, ZCL^{+}24, ZCY^{+}19, ZZL21, ZGL^{+}18, APV08, ASVE13, AH12, BBUD24, BAK13, CDA12, CLS17, CSSZ19, CSSZ23, DL14, ETF06, GFB^{+}14]. using [GHP^{+}06, HCW^{+}07, HOSS13, HCS12, JC08, JJML24, LCL^{+}23b, LHL^{+}24, MSL10, MVW08, NR24, NT08, REV^{+}12, RW12, SMK^{+}24, TWB^{+}23, TS20, TNEcC08, YQH12]. Utility [RT14]. Utilization [SSP^{+}24]. Utilizing [LLC^{+}24b]. UV [XJFS25]. UV-Space [XJFS25]. UVC [TZZ25].

V [HP17, LSS^{+}24a, RSR25]. V-JAUNE [HP17]. V-PCC [RSR25]. V-PCC-Coded [LSS^{+}24a]. Validation [ACC^{+}21, DMSRL18, SEK12]. Value [AR15, AMT25, SZHY19, BWA13]. VANET [Ala21, AS22a, CNG22]. vantage [VV11]. Variable [KMSW18, LCZ^{+}24c, HH11b]. variable-bit-rate [HH11b]. Variable-to-Variable [KMSW18]. Variables [MGS18]. Variate [LPC^{+}18]. Variation [FDBP23]. Variational [FFCM24, LLL^{+}24a, XLZ^{+}24, ZYT^{+}23]. Vault [DKJ^{+}21]. VBR [KW11, SWZ25]. VBR-encoded [SWZ25]. Vector [CZW15, HYLD20, LH20, TGSF21]. Vectors [Ala21]. Vehicle [CWW^{+}25, LLL^{+}22]. Vehicular [BOZ17, LB15]. Vein [CJHH21]. Verb [PRH14]. Verification [GDM^{+}23, LLL^{+}22, Lin15, LLC^{+}21, ZWYS20, ZLC^{+}23, ASVE13, ZLLT13]. Vernacular [SAL^{+}22]. Versatile [CI23, ZZL^{+}23]. Version [BCP14]. versus [MGP19]. Vertical [ZHG^{+}21]. Via [HWHL18, WKS^{+}23, XLT^{+}23, ZZW^{+}24, AWG^{+}15, CDL^{+}20, CZX^{+}24, CLZ^{+}23b, DCC^{+}13, FBSGR25, FXMZ24, GLL^{+}24, GHQ24, GLL^{+}22, GLC05, GCF^{+}21, HPH^{+}20, HDH^{+}24, HH25, HLW^{+}25, JLL^{+}25, JMLL20, KTK^{+}17, LT14, LCW16, LLSX20, LML^{+}24, LLH^{+}24, LWL24, LJW^{+}25, LLL^{+}22, LYDY24, LTL^{+}24, LL15, LWZH19, LLP^{+}23, LW23, LLL^{+}24a, LZYY24, LLZ^{+}24, LDL^{+}25, LXZ^{+}25, LYD^{+}21, LWZW21, LWW^{+}25, MXH^{+}23, MYX^{+}23b, NLN^{+}13, NSM^{+}24, OL25, PZJL22, PSL^{+}24, PWY^{+}25, QSZ^{+}21, QDX^{+}24, QZXH14, QLGL24, SCZ^{+}25, SZM^{+}21, SGY^{+}23, SLMJ24, SYM^{+}25, SHWC19, TBC^{+}11, TYY^{+}18, TYY^{+}22, TXWK24, TLZL24, WKST08, WWGT09, WCL^{+}22, WLTL23, WYGC24, WWZ24, WWY^{+}25, WXJL25, WC12, WHW18, WLL^{+}19, WXQC20, WHW^{+}21, WZZ^{+}22, WSX^{+}23, WDLW23, XYJ^{+}20, XQG^{+}24, XSL^{+}23, XLZ^{+}24, XZC^{+}25, YNC18, YTOH22, YWG^{+}20, YPSC22, YSSF24, YMS^{+}24, YP20, YDJ24, YHP25, YNLZ22, ZHG^{+}21, ZCL^{+}22, ZX14, ZZTL24, ZZL^{+}24a, ZHY^{+}25, ZLC^{+}25, ZYLN21, ZSS^{+}23, ZWR^{+}20, ZLF^{+}23, ZZL^{+}24c, ZZL^{+}25d, ZWF^{+}20, ZFFX24, ZZL25e]. Vibrotactile [CJP^{+}21, PRH14]. ViCoFace [LXT^{+}24]. VIdeo [MA10, AGC^{+}18, ACGH18, ABR17, AH20, AS22a, AS20, dCAGF^{+}25, ATS19, BDV08, BPM15, BFAS15, BWP^{+}24, BRA^{+}09, BYOZ20, BLH^{+}25, BRZS18, BTWZ22, BZDX^{+}18, CPSH14, CPP^{+}14, CE23, CISF^{+}25, CHHH18, CI23, CZC15, CHLW19, CYZ23, CPL^{+}23, CLZ^{+}23a, CZZ^{+}24, CYZ^{+}24, CCH^{+}24, CWL24, CQA^{+}24, CWXW25, CNG22, VTD22, DLZ^{+}17, DYSX14, DNL^{+}23, DOD23, DG17b, EWSZ15, FBSGR25, FXMZ24, TCP^{+}20, FFCM24, GKSB17, GZ20, GLL^{+}22, GHAM25, GGSV25, GF17, GWM^{+}14, HCZ^{+}22, HZS^{+}24, HSR^{+}23, HZH24, HCM^{+}25, HWY^{+}11, HLTH24, HWLC19, HHH22, HKYW14, HJWW19, HZL^{+}21, HYLG23, HCWM14, HDW^{+}18, HSL^{+}22, HSG23, HP17, IVS^{+}20, JW21, JLL24, JC08, KWK25, KPL^{+}22, KKGE18, KFP^{+}25, KLS^{+}18, KD18, KO11, KZGH15, LGT^{+}24, LCL23a, LYJ^{+}15, LCL14, LB15, LLSX20, LLJ^{+}20, LLS^{+}21, LWH^{+}22, LPW^{+}22, LLS^{+}22, LSXZ23, LHZ^{+}23, LLL^{+}25, LZWL25, LLC^{+}25, LZL^{+}25, LZLJ22, LJZ^{+}22, LMF^{+}14, LLA^{+}21, LWZ21b, LMC^{+}22]. Video [MBXZ24, MOL^{+}22, MSC^{+}23, MGS18, MN16, MKS20, NTT20, NN21, NNK25, NDC^{+}23, NCL^{+}23, PYZ^{+}20, PSSZ24, PFC^{+}16, PVWD18, PWHP24, RSAH25, REV^{+}12, RBBS25, RLY^{+}21, SG22a, SGS21, SS17, SAF19, SHZ^{+}20, SXY^{+}23, SWZ25, SkFM18, SK23, SSSK18, SYS17, SLP15, SGYX22, SCZJ24, TWL19, TDW^{+}23, TZ24, TZZ25, TFHM24, TB17, TV07, TESU16, VVSV17, VTP^{+}20, WZNM14, WCO15, WAD^{+}18, WDPX23, WZWY23, WLL23, WLSN24, WWL^{+}24, WBR25, WCY^{+}18, WLW^{+}23, WGHX24, WGM25, BLS^{+}19, YSXH16, YZX16, YYX^{+}24, YZJ^{+}25, YLD^{+}25, YNY25, YZZZ25, YM24, YJM^{+}19, YH14, YWL^{+}24, YSZZ14, YSZ15, YZW^{+}24, YZL^{+}24, YGMY24, ZQRS18, ZCL^{+}22, ZCD15, Zha19, ZHD^{+}23, ZYCD24, ZZW^{+}24, ZSX^{+}25, ZYJZ25, ZZL^{+}25a, ZLC^{+}25, ZYH^{+}24, ZLL^{+}22, Zho16, ZLOL18, ZGL^{+}18, ZZCZ21, ZZL^{+}23, ZXY^{+}20, ADCB07, CPP^{+}13, CE10, CTGP08, CBJ^{+}09, CEE09, CL12, CWC10, EGEM06, FKFB05, GL08, HH08a, HB08, HH08b]. video [HH11a, HH11c, Hua13, IB10, JCC^{+}10, JSL07, KH13, KS09, LLHS12, LWL^{+}12, LWLZ13, LK07, LCS09, LC12, LCT^{+}12, LH12, LGX^{+}08, MDMK06, MPSR05, MTTH13, MC11, NH10, PBS12, QHR^{+}08, QS10, RHS12, RW12, SAAH10, SOC^{+}13, SVF12, SWH06, Swa13, TC08, TTR12, TCJ08, VM12, WC12, WAK^{+}12, XC06, XXD^{+}08, YSG^{+}06, YYS13, YZL^{+}14, LKM^{+}19, LVM^{+}21]. Video-Based [CWL24, WDPX23]. Video-elicited [LGT^{+}24]. Video-language [MSC^{+}23]. video-on-demand [CE10, TC08]. Video-Text [LLS^{+}22, SXY^{+}23]. Videoconferencing [DMF17, BBT^{+}05]. videography [HWG07]. Videos [APC^{+}25, AMG23, BBRS23, BUS^{+}21, BLW^{+}24, CED^{+}16, DLC^{+}22, FHH22, GZH17, GGML22, LYW^{+}23a, LCS17, LHL^{+}24, LZC^{+}19, LNT^{+}21, LZF^{+}22, LQX^{+}24, PDD22, QJ23, SZL^{+}22, SWZ25, SLBS20, TJN14, TFHM24, WLZ^{+}24a, WKN25, WKE16, XZH^{+}21, YYY^{+}24, ZZZ14, ZSZ^{+}23, ZMZ^{+}23, DCO10, HH12a, HTT^{+}11, HH11b, HZW^{+}11, KW11, SXM^{+}06, YZL^{+}14]. View [DYN^{+}25, GLW20, GZT21, HLY21, LZZ^{+}24a, LKM17, LKW^{+}22, LLP^{+}23, RS16, RE24, SZZZ25, SLNL20, TVZ^{+}19, ZCXL23, ZXS^{+}24, ZFC^{+}24, AE22, ATM06, GHQ24, JYZC12, JMLL20, LWLZ13, WQTX25, YCFX23, ZDZ^{+}22, ZPL^{+}23]. View-Based [GLW20]. View-Invariant [RS16, RE24]. ViewCast [YWN^{+}10b, YWN^{+}10a]. Viewer [ZMH^{+}20]. Viewing [CE23, YHZ19, NH10, SGW08]. Viewpoint [LSL^{+}24, HNS13]. Viewport [VTD22, LHZ^{+}23, LZL^{+}25, NTT20, QJ23, YWC^{+}25]. Viewport-Adaptive [NTT20]. Viewport-based [QJ23]. Viewport-Unaware [YWC^{+}25]. Views [CZC15, LYZ^{+}18]. Viola [TWB^{+}23]. violation [SEK12]. Violin [TWB^{+}23]. Virality [CSJC17, WWW^{+}22]. Virtual [BLMP18, CGPCR18, EG17, FCL^{+}22, GEL^{+}15, GCL^{+}25, GZW^{+}24b, HEA14, HWG07, HLD18, HLZ^{+}20, LZZ^{+}24c, MSKYJ21, PKL25, PB14, RTM^{+}24, SCX^{+}25, SHIE15, SLG^{+}24, VTP^{+}20, WC23, WPRC18, WLL^{+}24, XWZ24a, YLHL22, DL14, MCM^{+}09, MRS^{+}07, WLHL13, ZWM12b, ZCT^{+}07, HHLY19]. Virtualization [AAT^{+}22]. VirtualLoc [XWZ24a]. VIS [YHH24]. VisActive [KIT^{+}24]. VISCOUNTH [BBB^{+}23]. Visibility [CH15]. Visible [CZL23b, DSL^{+}25, HZH24, HCZ^{+}25, LLS^{+}24, LYW^{+}24, NR24, WWQ^{+}25a, YLYB24, ZGZ^{+}22]. Visible-Infrared [CZL23b, DSL^{+}25, HZH24, HCZ^{+}25, WWQ^{+}25a, YLYB24, ZGZ^{+}22, LLS^{+}24]. Vision [CLP^{+}23, CDA^{+}24, GCC^{+}25, GGSV25, HLW^{+}21, HCZ^{+}25, HZL^{+}25, KUH^{+}22, LFP^{+}22, LL15, NNK25, PSN^{+}24, QCW^{+}25, TLZL24, WDR^{+}24, ZHS^{+}25, CCD07]. Vision-aided [PSN^{+}24]. Vision-and-Language [CLP^{+}23, HZL^{+}25]. Vision-Language [GCC^{+}25, GGSV25, LFP^{+}22, NNK25, QCW^{+}25, WDR^{+}24]. Visual [BYM^{+}18, BBB^{+}23, Chu15, DNL^{+}23, DMF^{+}20, EC16, EL25, EFF^{+}25, FZYW20, FHP23, GCC^{+}25, GUH^{+}20, GZL^{+}20, GSDT21, HWZL24, HH19, HLTH24, HLZ^{+}21, HLW^{+}25, HZL^{+}25, KRKK14, KIT^{+}24, LGX^{+}24, LPY^{+}19, LZH^{+}20, LXB^{+}22, LLG^{+}23, LJL^{+}24, LML^{+}24, LCC^{+}14a, LXT^{+}24, LZW^{+}21, LGY^{+}22, LZD^{+}22, LLZ^{+}24, LDL^{+}25, LYW^{+}25, MBK^{+}23, MAE^{+}21, RXY^{+}25, SZZZ25, SMN^{+}22, SA16, SZF^{+}25, SZLL17, SZFR23, TWM^{+}25, TWFW24, TWL19, TGC23, TS22, WZNM14, WFZ^{+}21b, WTD^{+}21, WLS^{+}24b, WTL24, WSLC24, WSLZ25, WLH^{+}25, WSLM18, WHZ^{+}24, WJ15b, WLL^{+}19, WFC^{+}25b, XGX^{+}25, XCC^{+}23, XWZ24a, YMX^{+}16, YWC24, YJM^{+}24, YDJ24, YSZ15, YZYX24, YZG^{+}20, YHZ19, ZYO20, ZYM^{+}10, ZHL19, ZWZL22, ZLZ^{+}22a, ZPY^{+}23, ZYT^{+}23, ZMX^{+}23, ZLX24, ZGY^{+}24a, ZXZ^{+}25, ZZW^{+}19, ZZC^{+}25a, ZSZ^{+}25, ZWYS20, ZHG^{+}25, ZMZ^{+}23, BPFA24, FYH10, JYZC12, JCC^{+}10, JC08, LML^{+}13, LXZ^{+}25, NW08, WYK12, ZWH^{+}23, ZG08, ZSZ^{+}23]. Visual-concept-based [KIT^{+}24]. Visual-Consistent [LXT^{+}24]. Visual-Inertial-Surround-View [SZZZ25]. Visual-Language [RXY^{+}25, ZMX^{+}23]. Visual-linguistic-stylistic [ZGY^{+}24a]. Visual-Semantic [HH19, XGX^{+}25, FHP23]. visual-similarity-based [NW08]. Visual-Textual [HZL^{+}25, YMX^{+}16].

Visualization [HAAM19, HLZ^{+}20,

SMK^{+}24, ZZC^{+}15, DÇ07, KTM^{+}06]. VL [ZMX^{+}23]. VL-NMS [ZMX^{+}23]. Vlogs [ICZ^{+}22]. VlogSense [BGP11]. Vocabulary [LGW^{+}25]. vocal [MSL10]. Vocoder [LCL^{+}24a]. VoD [CSJ^{+}08, HDW^{+}18, LCC^{+}14b, WLQL12, ZCL^{+}12, ZSO13]. Voice [CZX^{+}24, CGW^{+}24, JTZ^{+}16, KDC08]. Voice-Based [CZX^{+}24]. Voice-Face [CGW^{+}24]. VoiceStyle [CZX^{+}24]. Volume [Ano11, Ano12, Ano14, Ano20, HLZ^{+}20, Sha21, ZHL19, ZLC^{+}25, TOM12]. Volumetric [BLH^{+}25, LZL^{+}25, LYH^{+}25]. VoteNet [WXW^{+}22]. Voxel [LZL^{+}25]. VPFormer [LZL^{+}25]. VQA [PLZT22, XJB25, ZGY^{+}24b]. VR [CKRB23, CXX^{+}24, TCP^{+}20, LSCH25, LZH^{+}24a, LLA^{+}21, ROST20, XLF^{+}25, YM24]. VR-EXP [TCP^{+}20]. VRVul [SCX^{+}25]. VRVul-Discovery [SCX^{+}25]. vs [NN21, YSZ15]. VTON [WLL^{+}24]. Vulnerability [SCX^{+}25, TWM^{+}25]. VVC [LJ25].

Waiting [SAAH10]. Waiting-time [SAAH10]. Wakeup [ZXT^{+}25]. Wakeup-Darkness [ZXT^{+}25]. Walking [YKW^{+}22]. warehouse [ATM06]. Warp [FCL^{+}22]. Warping [LYLL23]. Watch [ATS19, CTGP08]. Watch-and-comment [CTGP08]. Watching [FHH22]. Watermarking [BA20, DG17b, GKSB17, HD19, HLS^{+}22, LLA^{+}21, LYH^{+}25, NR24, NPG^{+}22, RS16, WWZ^{+}25, WYY^{+}25, YWNW15, vRPPP23, AP10, BWA13, GZHD12, LLHS12, MB08, NC13, XMST07]. Watermarks [NHZ^{+}24, MB08]. Wave [CKRB23, LLYH14, LZH^{+}24a]. Wavelength [SBS23]. Wavelength-based [SBS23]. Wavelet [BA20, LLA^{+}21]. Wavelet-domain [BA20]. Weak [WKE16]. Weakly [CLZ^{+}23a, DZWH23, DWL^{+}23, FZZ^{+}25, JLZ^{+}21, LZZ^{+}23a, LZZ^{+}25, SSY20, TCS^{+}25, TYY^{+}18, XXG^{+}21, XZZL23, YZW^{+}24, ZZL^{+}24a]. Weakly-Supervised [LZZ^{+}25]. Wearable [VLE25]. web [HZW^{+}11, SX11, YGHH12, ZLLT13, AP13, BYM^{+}18, CL07, FLM^{+}06, GHS^{+}24, GJAF18, HTT^{+}11, LWL^{+}12, LZY^{+}24, LOJZ18, MGCH13, QSZ^{+}21]. Web-accessible [AP13]. Web-based [CL07, MGCH13]. web-scale [SX11, GJAF18]. Web3 [WAC^{+}25, CDA^{+}24]. webcams [AP13]. WebRTC [APM21, BPFA24, CPCM21]. WebRTC-based [BPFA24]. websites [SX13, LXL^{+}18]. Weight [LS21, SKA24, ZZL^{+}24c]. Weight-Based [SKA24]. Weighted [LLZ^{+}21, LZH^{+}24b, LLS^{+}24, MAZ22, WLSX24, XZY^{+}24a, YWG^{+}20]. Weights [AAS^{+}20, PDD16]. Well [HHLY19]. Well-Posed [HHLY19]. we’ve [Cha13]. Whale [TRRB20]. Where [ZZB^{+}21]. Whole [GWWC25]. Whole-Body [GWWC25]. Wide [LZL^{+}24a, CY11, WLZ12]. Wide-Receptive-Field [LZL^{+}24a]. Wikipedia [ENHN09]. Wild [HSL^{+}22, ZDZ^{+}23, ZSZ^{+}24a]. Will [RABC24, WWW^{+}22]. Window [ZZL^{+}24b]. Wire [YLHL22]. Wireless [AAT^{+}22, BPM15, KEYY22, KD18, WCO15, BDV08, CY11, GS11b, HNL08, HH11a, LCK09, YSG^{+}06]. Wirelessly [ZZC^{+}25b]. wise [DVA21, GQSG24, YSF^{+}21, ZLL^{+}24a]. within [MRS^{+}07]. without [LW07]. WLANs [KW11, PBS12]. WMD [ATS19]. Wolf [AAS^{+}20]. Word [MSL10]. Words [WCX^{+}14, WZNM14, XJG^{+}22]. Words-and-Pictures [WCX^{+}14]. Work [MHF24]. Workshop [CHHH18, Hua23, TB17]. Workshops [ZTB20]. World [XHL^{+}24b, ZCYL24, LXJ^{+}24, LWH^{+}23, WLHL13]. worldwide [SSS13]. Worth [HZH24]. Wow [LXL^{+}14]. writing [LT14]. WTRPNet [XHL^{+}21].

WWW [HCW^{+}07].

xCos [LLC^{+}21]. XMT [SG07]. XOR

[CQA^{+}25].

Y-Net [CH21]. Years

[HGKS24, Cha13, Eff13, JPS05, Row13, She13, Tur13, Whi13, ZR13]. YOLO [GHY^{+}24]. YouTube [BGP11, DYSX14, GGA^{+}20, KLS^{+}18, KZGH15].

Zero [DLZ^{+}17, GJY^{+}25, GGSV25, LWL24, LJW^{+}25, LYH^{+}25, RYZ^{+}23, SS24, WDC^{+}25, WYY^{+}25, XTL^{+}21, XJB25]. Zero-Shot [GJY^{+}25, GGSV25, RYZ^{+}23, SS24, WDC^{+}25, XJB25, LWL24, LJW^{+}25, XTL^{+}21]. Zero-Watermarking [LYH^{+}25, WYY^{+}25]. Zigzag [ZZL^{+}24b]. Zoomable [WCO15].

References

Alwaely:2023:GGB

[AA23] Basheer Alwaely and Charith Abhayaratne. GHOSM: Graphbased hybrid outline and skeleton modelling for shape recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):86:1–86:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                                    URL https://dl.acm.org/
                                                    doi/10.1145/3554922.

Alkhariji:2021:SPD

[AAA^{+}21] Lamya Alkhariji, Nada Alhirabi, Mansour Naser Alraja, Mahmoud Barhamgi, Omer Rana, and Charith Perera. Synthesising privacy by design knowledge toward explainable Internet of Things application designing in healthcare. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):62:1– 62:29, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3434186.

Airale:2025:AGS

[AAPLV25] Louis Airale, Xavier Alameda-Pineda, Stéphane Lathuilière, and Dominique Vaufreydaz. Autoregressive GAN for semantic unconditional head motion generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):14:1–14:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635154.

Agrawal:2020:HWB

[AAS^{+}20] Utkarsh Agrawal, Jatin Arora, Rahul Singh, Deepak Gupta, Ashish Khanna, and Aditya Khamparia. Hybrid wolf– bat algorithm for optimization of connection weights in multi-layer perceptron. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):37:1– 37:20, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3350532.

Alharbi:2022:NSA

[AAT^{+}22] Abdullah Alharbi, Mohammed Aljebreen, Amr Tolba, Konstantinos A. Lizos, Saied Abd El-Atty, and Farid Shawki. A normalized slicing-assigned virtualization method for 6g-based wireless communication systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):134:1–134:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3546077.

Abdallah:2018:AHD

[Abd18] Maha Abdallah. Aesthetic highlight detection in movies based on synchronization of spectators’ reactions. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):68:1– 68:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Akhtar:2017:COV

[ABR17] Shahid Akhtar, Andre Beck, and Ivica Rimac. Caching online video: Analysis and proposed algorithm. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):48:1–48:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ahmad:2019:HDF

[AC19] Kashif Ahmad and Nicola Conci. How deep features have improved event recognition in multimedia: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):39:1–39:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3306240.

Amato:2021:SPV

[ACC^{+}21] Flora Amato, Valentina Casola, Giovanni Cozzolino, Alessandra De Benedictis, Nicola Mazzocca, and Francesco Moscato. A security and privacy validation methodology for e-health systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):67:1–67:22, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3412373.

Abdallah:2018:ISI

[ACGH18] Maha Abdallah, Kuan-Ta Chen, Carsten Griwodz, and Cheng-Hsin Hsu. Introduction to the special issue on delay-sensitive video computing in the cloud. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):53:1–53:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

AlvesdeAlmeida:2020:RPS

[ACVM20] Marcos Alves de Almeida, Carolina Coimbra Vieira, Pedro Olmo Stancioli Vaz De Melo, and Renato Martins Assunção. Random playlists smoothly commuting between styles. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–20, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3361742.

Agrafiotis:2007:TEC

[ADCB07] D. Agrafiotis, S. J. C. Davies, N. Canagarajah, and D. R. Bull. Towards efficient contextspecific video coding based on gaze-tracking analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4):4:1–4:15, December 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

AMEUR:2025:SFS

[ADMM25] Zoubida AMEUR, Claire-Hélène Demarty, Olivier Le Meur, and Daniel Ménard. Style-FG: a style-based framework for film grain analysis and synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):252:1–252:24, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Abreu:2025:AUE

[AdSGMS25] Raphael Abreu, Joel dos Santos, Gheorghita Ghinea, and Débora C. Muchaluat-Saade. Assessing usefulness, ease of use, and recognition performance of semi-automatic mulsemedia authoring. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):253:1– 253:19, September 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Aloufi:2022:MDT

[AE22] Samah Aloufi and Abdulmotaleb El Saddik. MMSUM digital twins: a multi-view multi-modality summarization framework for sporting events. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):5:1–5:25, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3462777.

Ademoye:2013:IRT

[AG13] Oluwakemi A. Ademoye and Gheorghita Ghinea. Information recall task impact in olfaction-enhanced multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3): 17:1–17:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Abdallah:2018:DSV

[AGC^{+}18] Maha Abdallah, Carsten Griwodz, Kuan-Ta Chen, Gwendal Simon, Pin-Chun Wang, and Cheng-Hsin Hsu. Delaysensitive video computing in the cloud: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):54:1– 54:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ansar:2025:TFT

[AGCC25] Wazib Ansar, Saptarsi Goswami, Amlan Chakrabarti, and Basabi Chakraborty. TexIm FAST: Text-to-image encoding for semantic similarity evaluation of disproportionate sequences. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):168:1–168:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Armitage:2012:ROF

[AH12] Grenville Armitage and Amiel Heyde. REED: Optimizing first person shooter game server discovery using network coordinates. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2):20:1–20:??, May 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Akpinar:2020:PPP

[AH20] Kutalmis Akpinar and Kien A. Hua. PPNet: Privacy protected CDN–ISP collaboration for QoS-aware multi-CDN adaptive video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2):49:1–49:23, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3379983.

Atrey:2007:GOO

[AKO07] Pradeep K. Atrey, Mohan S.

Kankanhalli, and John B. Oommen. Goal-oriented optimal subset selection of correlated multimedia streams. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(1): ??, February 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Alaya:2021:PBD

[Ala21] Bechir Alaya. Payoff-based dynamic segment replication and graph classification method with attribute vectors adapted to urban VANET. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):85:1–85:22, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3440018.

Alshehri:2024:ISS

[AM24] Fatima Alshehri and Ghulam Muhammad. Ischemic stroke segmentation by transformer and convolutional neural network using few-shot learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (12):394:1–394:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3699513.

Ahmad:2018:EDM

[AMC^{+}18] Kashif Ahmad, Mohamed Lamine Mekhalfi, Nicola Conci, Farid Melgani, and Francesco De Natale. Ensemble of deep models for event recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2):51:1– 51:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Amoroso:2025:PCD

[AMC^{+}25] Roberto Amoroso, Davide Morelli, Marcella Cornia, Lorenzo Baraldi, Alberto Del Bimbo, and Rita Cucchiara. Parents and children: Distinguishing multimodal deepfakes from natural images. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):11:1– 11:??, January 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3665497.

Ayoughi:2023:SCE

[AMG23] Melika Ayoughi, Pascal Mettes, and Paul Groth. Self-contained entity discovery from captioned videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):177:1–177:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3583138.

Ademoye:2016:AME

[AMMG16] Oluwakemi A. Ademoye, Niall Murray, Gabriel-Miro Muntean, and Gheorghita Ghinea. Audio masking effect on intercomponent skews in olfactionenhanced multimedia presentations. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):51:1–51:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Aoun:2025:MBL

[AMT25] Alain Aoun, Mahmoud Masadeh, and Sofiène Tahar. ML-based load value approximator for efficient multimedia processing. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (7):210:1–210:18, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Andre:2013:EUU

[And13] Elisabeth Andre. Exploiting unconscious user signals in multimodal human-computer interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):48:1–48:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Anonymous:2011:TCO

[Ano11] Anonymous. Table of contents: Online supplement volume 7S, number 1. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(4):34:1–34:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Anonymous:2012:TCO

[Ano12] Anonymous. Table of contents: Online supplement volume 8S, number 1. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2):16:1–16:??, May 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Anonymous:2013:CPM

[Ano13] Anonymous. Call for papers: Multiple sensorial (MulSeMedia) multi-modal media: Advances and applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3):15:1– 15:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Anonymous:2014:TCO

[Ano14] Anonymous. Table of contents: Online supplement volume 10, number 1s. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(3):22:1–22:??, April 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Anonymous:2020:TCO

[Ano20] Anonymous. Table of contents: Online supplement volume 16, number 1s. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):74:1– 74:5, September 2020. ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3409367.

Agarwal:2010:BRW

[AP10] Parag Agarwal and Balakrishnan Prabhakaran. Blind robust watermarking of 3D motion data. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(1):2:1–2:??, February 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Abrams:2013:WAG

[AP13] Austin Abrams and Robert Pless. Web-accessible geographic integration and calibration of webcams. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1):8:1–8:??, February 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ahmad:2025:CLC

[APC^{+}25] Wasim Ahmad, Yan-Tsung Peng, Yuan-Hao Chang, Gaddisa Olani Ganfure, and Sarwar Khan. CapST: Leveraging capsule networks and temporal attention for accurate model attribution in deep-fake videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (4):123:1–123:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Alahmadi:2021:ABS

[APM21] Mohannad Alahmadi, Peter Pocta, and Hugh Melvin. An adaptive bitrate switching algorithm for speech applications in context of WebRTC. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):133:1– 133:21, November 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3458751.

Amirpour:2022:ELF

[APP^{+}22] Hadi Amirpour, Antonio Pinheiro, Manuela Pereira, Fernando J. P. Lopes, and Mohammad Ghanbari. Efficient light field image compression with enhanced random access. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):44:1– 44:18, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3471905.

Alameda-Pineda:2019:SSM

[APRS^{+}19] Xavier Alameda-Pineda, Miriam Redi, Mohammad Soleymani, Nicu Sebe, Shih-Fu Chang, and Samuel Gosling. Special section on multimodal understanding of social, affective, and subjective attributes. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1s):11:1–11:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3292061.

Adams:2008:SUS

[APV08] Brett Adams, Dinh Phung, and Svetha Venkatesh. Sensing and using social context. ACM Transactions on Multimedia Computing, Communications, and Applications, 5 (2):11:1–11:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ainam:2020:EAF

[AQL^{+}20] Jean-Paul Ainam, Ke Qin, Guisong Liu, Guangchun Luo, and Brighter Agyemang. Enforcing affinity feature learning through self-attention for person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1):16:1– 16:22, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377352.

Antaris:2015:SSC

[AR15] Stefanos Antaris and Dimitrios Rafailidis. Similarity search over the cloud based on image descriptors’ dimensions value cardinalities. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):51:1–51:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ahmed:2023:DBC

[ARA^{+}23] Fawad Ahmed, Muneeb Ur Rehman, Jawad Ahmad, Muhammad Shahbaz Khan, Wadii Boulila, Gautam Srivastava, Jerry Chun-Wei Lin, and William J. Buchanan. A DNA based colour image encryption scheme using a convolutional autoencoder. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):128:1–128:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3570165.

Alam:2013:MHB

[ARE13] Kazi Masudul Alam, Abu Saleh Md Mahfujur Rahman, and Abdulmotaleb El Saddik. Mobile haptic e-book system to support 3D immersive reading in ubiquitous environments.

ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4): 27:1–27:??, August 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Altamimi:2020:QFD

[AS20] Sa’di Altamimi and Shervin Shirmohammadi. QoE-fair DASH video streaming using server-side reinforcement learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):68:1–68:21, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3397227.

Alaya:2022:MVE

[AS22a] Bechir Alaya and Lamaa Sellami. Multilayer video encoding for QoS managing of video streaming in VANET environment. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):82:1–82:19, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491433.

Anand:2022:CSD

[AS22b] Ashima Anand and Amit Kumar Singh. A comprehensive study of deep learningbased covert communication. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2s):118:1–118:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3508365.

Alizadehsani:2021:UAS

[ASI^{+}21] Roohallah Alizadehsani, Danial Sharifrazi, Navid Hoseini Izadi, Javad Hassannataj Joloudari, Afshin Shoeibi, Juan M. Gorriz, Sadiq Hussain, Juan E. Arco, Zahra Alizadeh Sani, Fahime Khozeimeh, Abbas Khosravi, Saeid Nahavandi, Sheikh Mohammed Shariful Islam, and U. Rajendra Acharya. Uncertainty-aware semi-supervised method using large unlabeled and limited labeled COVID-19 data. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):104:1– 104:24, October 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3462635.

Akputu:2018:ERU

[ASLA18] Oryina Kingsley Akputu, Kah Phooi Seng, Yunli Lee, and Li-Minn Ang. Emotion recognition using multiple kernel learning toward e-learning applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):1:1–1:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Alsulaiman:2013:IVB

[ASVE13] Fawaz A. Alsulaiman, Nizar Sakr, Julio J. Valdés, and Abdulmotaleb El Saddik. Identity verification based on handwritten signatures with haptic information using genetic programming. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(2):11:1–11:??, May 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Arigon:2006:HMP

[ATM06] Anne-Muriel Arigon, Anne Tchounikine, and Maryvonne Miquel. Handling multiple points of view in a multimedia data warehouse. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(3):199–218, August 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Alvarez:2024:GES

[ATMM24] Luis Alvarez, Agustín Trujillo, Nelson Monzón, and Jean-Michel Morel. Generation and editing of 2D shapes using a branched representation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):369:1– 369:??, December 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3691635.

Atrey:2019:WMD

[ATS19] Pradeep K. Atrey, Bakul Trehan, and Mukesh K. Saini. Watch me from distance (WMD): a privacy-preserving long-distance video surveillance system. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):37:1–37:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3312574.

Adams:2005:IIM

[AVJ05] Brett Adams, Svetha Venkatesh, and Ramesh Jain. IMCE: Integrated media creation environment. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(3):211–247, August 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ahn:2015:SHG

[AWG^{+}15] Junho Ahn, James Williamson, Mike Gartrell, Richard Han, Qin Lv, and Shivakant Mishra. Supporting healthy grocery shopping via mobile augmented reality. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):16:1–16:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

An:2021:MTU

[AY21] Na An and Wei Qi Yan. Multitarget tracking using Siamese neural networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):75:1–75:16, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3441656.

Abdussalam:2023:NNC

[AYH^{+}23] Amr Abdussalam, Zhongfu Ye, Ammar Hawbani, Majjed Al-Qatf, and Rashid Khan. NumCap: a number-controlled multi-caption image captioning network. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):148:1–148:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3576927.

Bhowmik:2020:EDA

[BA20] Deepayan Bhowmik and Charith Abhayaratne. Embedding distortion analysis in waveletdomain watermarking. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–24, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3357333.

Bagchi:2011:FAD

[Bag11] Susmit Bagchi. A fuzzy algorithm for dynamically adaptive multimedia streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(2):11:1– 11:??, February 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bhatt:2013:RPB

[BAK13] Chidansh A. Bhatt, Pradeep K. Atrey, and Mohan S. Kankanhalli. A reward-and-punishment-based approach for concept detection using adaptive ontology rules. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(2):10:1–10:??, May 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bouyakoub:2011:SBI

[BB11] Samia Bouyakoub and Abdelkader Belkhir. SMIL builder: an incremental authoring tool for SMIL Documents. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(1):2:1–2:??, January 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Becattini:2023:VLS

[BBB^{+}23] Federico Becattini, Pietro Bongini, Luana Bulla, Alberto Del Bimbo, Ludovica Marinucci, Misael Mongiovì, and Valentina Presutti. VIS-COUNTH: a large-scale multilingual visual question answering dataset for cultural heritage. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):193:1–193:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3590773.

Becattini:2024:HPE

[BBL^{+}24] Federico Becattini, Carmen Bisogni, Vincenzo Loia, Chiara Pero, and Fei Hao. Head pose estimation patterns as deepfake detectors. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):342:1–342:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3612928.

Bai:2023:DDI

[BBRS23] Chongyang Bai, Maksim Bolonkin, Viney Regunath, and V. S. Subrahmanian. DIPS: a dyadic impression prediction system for group interaction videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1s):43:1–43:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3532865.

Baker:2005:UPC

[BBT^{+}05] H. Harlyn Baker, Nina Bhatti, Donald Tanguay, Irwin Sobel, Dan Gelb, Michael E. Goss, W. Bruce Culbertson, and Thomas Malzbender. Understanding performance in Coliseum, an immersive videoconferencing system. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(2):190–210, May 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Baldrati:2024:CIR

[BBUD24] Alberto Baldrati, Marco Bertini, Tiberio Uricchio, and Alberto Del Bimbo. Composed image retrieval using contrastive learning and taskoriented CLIP-based features. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (3):62:1–62:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3617597.

Bentaleb:2018:OSO

[BBZ18] Abdelhak Bentaleb, Ali C. Begen, and Roger Zimmermann. ORL–SDN: Online reinforcement learning for SDN-enabled HTTP adaptive streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (3):71:1–71:??, August 2018.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bianco:2015:UPM

[BC15] Simone Bianco and Gianluigi Ciocca. User preferences modeling and learning for pleasing photo collage generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (1):6:1–6:??, August 2015. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bulterman:2013:SAM

[BCG13] Dick C. A. Bulterman, Pablo Cesar, and Rodrigo Laiola Guimarães. Socially-aware multimedia authoring: Past, present, and future. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):35:1– 35:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bisogni:2024:IEB

[BCNP24] Carmen Bisogni, Lucia Cascone, Michele Nappi, and Chiara Pero. IoT-enabled biometric security: Enhancing smart car safety with depthbased head pose estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (6):155:1–155:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3639367.

Botta:2014:PCI

[BCP14] Marco Botta, Davide Cavagnino, and Victor Pomponiu. Protecting the content integrity of digital imagery with fidelity preservation: An improved version. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(3):29:1–29:??, April 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Babich:2008:VQE

[BDV08] Fulvio Babich, Marco D’orlando, and Francesca Vatta. Video quality estimation in wireless IP networks: Algorithms and applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1):3:1–3:??, January 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Berretti:2018:IAS

[Ber18] Stefano Berretti. Improved audio steganalytic feature and its applications in audio forensics. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (2):43:1–43:??, May 2018. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Baldauf:2015:ISG

[BFAS15] Matthias Baldauf, Peter Fröhlich, Florence Adegeye, and Stefan Suette. Investigating onscreen gamepad designs for Smartphone-controlled video games. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):22:1–22:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Biel:2011:VCB

[BGP11] Joan-Isaac Biel and Daniel Gatica-Perez. VlogSense: Conversational behavior and social attention in YouTube. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):33:1– 33:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bulterman:2005:SMA

[BH05] Dick C. A. Bulterman and Lynda Hardman. Structured multimedia authoring. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1):89–109, February 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Beuve:2024:HLD

[BHD24] Nicolas Beuve, Wassim Hamidouche, and Olivier Déforges. Hierarchical learning and dummy triplet loss for efficient deepfake detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):89:1–89:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3626101.

Boll:2011:ISI

[BJLX11] Susanne Boll, Ramesh Jain, Jiebo Luo, and Dong Xu. Introduction to special issue on social media. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):25:1–25:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bentaleb:2025:SCO

[BLH^{+}25] Abdelhak Bentaleb, May Lim, Sarra Hammoudi, Saad Harous, and Roger Zimmermann. Solutions, challenges, and opportunities in volumetric video streaming: an architectural perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):187:1–187:35, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Boll:2010:CPA

[BLJX10] Susanne Boll, Jiebo Luo, Ramesh Jain, and Dong Xu. Call for papers: ACM Transactions on Multimedia Computing, Communications and Applications special issue on social media. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):22:1–22:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bahirat:2018:DEM

[BLMP18] Kanchan Bahirat, Chengyuan Lai, Ryan P. Mcmahan, and Balakrishnan Prabhakaran. Designing and evaluating a mesh simplification algorithm for virtual reality. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):63:1–63:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yahia:2019:HBF

[BLS^{+}19] Mariem Ben Yahia, Yannick Le Louedec, Gwendal Simon, Loutfi Nuaymi, and Xavier Corbillon. HTTP/2based frame discarding for lowlatency adaptive video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):18:1–18:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3280854.

Bian:2024:PAL

[BLW^{+}24] Jiang Bian, Xuhong Li, Tao Wang, Qingzhong Wang, Jun Huang, Chen Liu, Jun Zhao, Feixiang Lu, Dejing Dou, and Haoyi Xiong. P^{2}ANet: a largescale benchmark for dense action detection from table tennis match broadcasting videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):118:1–118:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3633516.

Bharati:2017:ETC

[BOZ17] Sailesh Bharati, Hassan Aboubakr Omar, and Weihua Zhuang. Enhancing transmission collision detection for distributed TDMA in vehicular networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (3s):37:1–37:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Biondi:2022:CRC

[BPB^{+}22] Niccoló Biondi, Federico Pernici, Matteo Bruni, Daniele Mugnai, and Alberto Del Bimbo. CL^{2}R: Compatible lifelong learning representations. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2s):132:1–132:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3564786.

Bingol:2024:QEW

[BPFA24] Gülnaziye Bingöl, Simone Porcu, Alessandro Floris, and Luigi Atzori. QoE estimation of WebRTC-based audio-visual conversations from facial and speech features. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):130:1–130:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638251.

Baik:2015:EMR

[BPM15] Eilwoo Baik, Amit Pande, and Prasant Mohapatra. Efficient MAC for real-time video streaming over wireless LAN. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (4):50:1–50:??, April 2015. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bental:2015:SSL

[BPT^{+}15] Diana S. Bental, Eliza Papadopoulou, Nicholas K. Taylor, M. Howard Williams, Fraser R. Blackmun, Idris S. Ibrahim, Mei Yii Lim, Ioannis Mimtsoudis, Stuart W. Whyte, and Edel Jennings. Smartening up the student learning experience with ubiquitous media. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (1s):23:1–23:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bruneau-Queyreix:2018:PNS

[BQBLN18] Joachim Bruneau-Queyreix, Jordi Mongay Batalla, Mathias Lacaud, and Daniel Negru. PMS: a novel scale-adaptive and quality-adaptive hybrid P2P/ multisource solution for live streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):35:1–35:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Buckchash:2022:GLG

[BR22] Himanshu Buckchash and Balasubramanian Raman. GraSP: Local Grassmannian spatiotemporal patterns for unsupervised pose sequence recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):79:1–79:23, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3491227.

Benevenuto:2009:VIO

[BRA^{+}09] Fabrício Benevenuto, Tiago Rodrigues, Virgilio Almeida, Jussara Almeida, and Keith Ross. Video interactions in online video social networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(4): 30:1–30:??, October 2009. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bhattacharjee:2024:ESM

[BRG24] Mrinmoy Bhattacharjee, Prasanna Mahadeva S. R., and Prithwijit Guha. Exploration of speech and music information for movie genre classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (8):241:1–241:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3664197.

Bhat:2018:SNA

[BRZS18] Divyashri Bhat, Amr Rizk, Michael Zink, and Ralf Steinmetz. SABR: Network-assisted content distribution for QoEdriven ABR video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (2s):32:1–32:??, May 2018. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bailey:2008:SSA

[BSH08] Brian P. Bailey, Nicu Sebe, and Alan Hanjalic. Special section from the ACM Multimedia Conference 2007. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1):1:1–1:??, October 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bao:2024:CLP

[BSQ^{+}24] Yongtang Bao, Chunjian Su, Yutong Qi, Yanbing Geng, and Haojie Li. Category-level pose estimation and iterative refinement for monocular RGB-D image. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):374:1–374:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3695877.

Bhattacharya:2011:HAA

[BSS11] Subhabrata Bhattacharya, Rahul Sukthankar, and Mubarak Shah. A holistic approach to aesthetic enhancement of photographs. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):21:1–21:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Baez-Suarez:2020:SSS

[BSSNF^{+}20] Abraham Báez-Suárez, Nolan Shah, Juan Arturo Nolazco-Flores, Shou-Hsuan S. Huang, Omprakash Gnawali, and Weidong Shi. SAMAF: Sequenceto-sequence autoencoder model for audio fingerprinting. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):43:1– 43:23, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380828.

Bentaleb:2020:PAA

[BTBZ20] Abdelhak Bentaleb, Christian Timmerer, Ali C. Begen, and Roger Zimmermann. Performance analysis of ACTE: a bandwidth prediction method for lowlatency chunked streaming.

ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2s):69:1–69:24, July 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3387921.

Bi:2022:DTE

[BTWZ22] An-Qi Bi, Xiao-Yang Tian, Shui-Hua Wang, and Yu-Dong Zhang. Dynamic transfer exemplar based facial emotion recognition model toward online video. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):121:1–121:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3538385.

Becattini:2021:DPA

[BUS^{+}21] Federico Becattini, Tiberio Uricchio, Lorenzo Seidenari, Lamberto Ballan, and Alberto Del Bimbo. Am I done? Predicting action progress in videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):119:1–119:24, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3402447.

Bhayana:2025:TCF

[BV25] Divya Arora Bhayana and Om Prakash Verma. Transconvo-former net for hierarchical prediction of household images. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):291:1–291:21, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bhatnagar:2013:SRI

[BWA13] Gaurav Bhatnagar, Q. M. Jonathan Wu, and Pradeep K. Atrey. Secure randomized image watermarking based on singular value decomposition. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(1):4:1– 4:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bandung:2024:IVD

[BWP^{+}24] Yoanes Bandung, Mokhamad Arfan Wicaksono, Sean Pribadi, Armein Z. R. Langi, and Dion Tanjung. IoT video delivery optimization through machine learning-based frame resolution adjustment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):277:1–277:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3665929.

Bao:2015:CPE

[BXMH15] Bing-Kun Bao, Changsheng Xu, Weiqing Min, and Mohammod Shamim Hossain. Crossplatform emerging topic detection and elaboration from multimedia streams. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):54:1–54:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bai:2018:ADA

[BYM^{+}18] Yalong Bai, Kuiyuan Yang, Tao Mei, Wei-Ying Ma, and Tiejun Zhao. Automatic data augmentation from massive Web images for deep visual recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):69:1–69:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Bentaleb:2020:DDQ

[BYOZ20] Abdelhak Bentaleb, Praveen Kumar Yadav, Wei Tsang Ooi, and Roger Zimmermann. DQ-DASH: a queuing theory approach to distributed adaptive video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1):4:1–4:24, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3371040.

Buchanan:2005:ATL

[BZ05] M. Cecelia Buchanan and Polle T. Zellweger. Automatic temporal layout mechanisms revisited. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1):60–88, February 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Burger:2018:GAV

[BZDX^{+}18] Valentin Burger, Thomas Zinner, Lam Dinh-Xuan, Florian Wamser, and Phuoc Tran-Gia. A generic approach to video buffer modeling using discretetime analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):33:1–33:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2019:SAD

[CAJ19] Zhineng Chen, Shanshan Ai, and Caiyan Jia. Structureaware deep learning for product image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):4:1– 4:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3231742.

Cesar:2009:FTE

[CBJ^{+}09] Pablo Cesar, Dick C. A. Bulterman, Jack Jansen, David Geerts, Hendrik Knoche, and William Seager. Fragment, tag, enrich, and send: Enhancing social sharing of video. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3): 19:1–19:??, August 2009. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chandra:2014:HPM

[CBR14] Surendar Chandra, John Boreczky, and Lawrence A. Rowe. High performance many-to-many intranet screen sharing with DisplayCast. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):19:1–19:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cesar:2008:ISI

[CBS08] Pablo Cesar, Dick C. A. Bulterman, and Luiz Fernando Gomes Soares. Introduction to special issue: Human-centered television — directions in interactive digital television research. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(4): 24:1–24:??, October 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cornia:2018:PMA

[CBSC18] Marcella Cornia, Lorenzo Baraldi, Giuseppe Serra, and Rita Cucchiara. Paying more attention to saliency: Image captioning with saliency and context attention. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2):48:1–48:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chu:2017:POI

[CC17] Wei-Ta Chu and Chih-Hao Chiu. Predicting occupation from images by combining face and body context information. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):7:1–7:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Colombo:2007:RTR

[CCD07] Carlo Colombo, Dario Comanducci, and Alberto Del Bimbo. Robust tracking and remapping of eye appearance with passive computer vision. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4):2:1–2:20, December 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Capasso:2024:CSM

[CCD24] Paola Capasso, Giuseppe Cattaneo, and Maria De Marsico. A comprehensive survey on methods for image integrity. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):347:1–347:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3633203.

Chen:2024:ISC

[CCF^{+}24] Ning Chen, Zhipeng Cheng, Xuwei Fan, Zhang Liu, Bangzhen Huang, Yifeng Zhao, Lianfen Huang, Xiaojiang Du, and Mohsen Guizani. Integrated sensing, communication, and computing for costeffective multimodal federated perception. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):237:1–237:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3661313.

Chen:2008:ASD

[CCG^{+}08] Songqing Chen, Shiping Chen, Huiping Guo, Bo Shen, and Sushil Jajodia. Achieving simultaneous distribution control and privacy protection for Internet media delivery. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2): 9:1–9:??, May 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Claypool:2020:IMD

[CCG20] Mark Claypool, Andy Cockburn, and Carl Gutwin. The impact of motion and delay on selecting game targets with a mouse. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):73:1–73:24, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3390464.

Chen:2024:DMP

[CCH^{+}24] Baian Chen, Zhilei Chen, Xiaowei Hu, Jun Xu, Haoran Xie, Jing Qin, and Mingqiang Wei. Dynamic message propagation network for RGB-D and video salient object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (1):18:1–18:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3597612.

Candan:2006:ISI

[CCK06] K. Selçuk Candan, Augusto Celentano, and Wolfgang Klas. Introduction to special issue on the use of context in multimedia information systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(3): 173–176, August 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2015:ISI

[CCO15] Kuan-Ta Chen, Songqing Chen, and Wei Tsang Ooi. Introduction to the special issue on MM-Sys 2014 and NOSSDAV 2014. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (2s):41:1–41:??, February 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2012:SGU

[CDA12] Yi Chen, Abhidnya A. Deshpande, and Ramazan S. Aygüun. Sprite generation using sprite fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2):22:1–22:??, May 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2024:WMS

[CDA^{+}24] Hongzhou Chen, Haihan Duan, Maha Abdallah, Yufeng Zhu, Yonggang Wen, Abdulmotaleb El Saddik, and Wei Cai. Web3 Metaverse: State-of-the-art and vision. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):101:1–101:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3630258.

Candan:2009:ISS

[CDGJ09] K. Selçuk Candan, Alberto Del Bimbo, Carsten Griwodz, and Alejandro Jaimes. Introduction to the special section for the best papers of ACM Multimedia 2008. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3):18:1–18:??, August 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2020:AAC

[CDL^{+}20] Hui Chen, Guiguang Ding, Zijia Lin, Sicheng Zhao, Xiaopeng Gu, Wenyuan Xu, and Jungong Han. ACMNet: Adaptive confidence matching network for human behavior analysis via cross-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):27:1–27:21, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3362065.

Cofano:2017:DPE

[CDZ^{+}17] Giuseppe Cofano, Luca De Cicco, Thomas Zinner, Anh Nguyen-Ngoc, Phuoc Tran-Gia, and Saverio Mascolo. Design and performance evaluation of network-assisted control strategies for HTTP adaptive streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):42:1–42:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Carlsson:2010:SSL

[CE10] Niklas Carlsson and Derek L. Eager. Server selection in largescale video-on-demand systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(1): 1:1–1:??, February 2010. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Carlsson:2023:CUS

[CE23] Niklas Carlsson and Derek Eager. Cross-user similarities in viewing behavior for 360^{∘} video and caching implications. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):152:1– 152:??, September 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3507917.

Calagari:2016:DPS

[CED^{+}16] Kiana Calagari, Tarek Elgamal, Khaled Diab, Krzysztof Templin, Piotr Didyk, Wojciech Matusik, and Mohamed Hefeeda. Depth personalization and streaming of stereoscopic sports videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(3):41:1–41:??, June 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cha:2009:TVS

[CEE09] Jongeun Cha, Mohamad Eid, and Abdulmotaleb El Saddik. Touchable 3D video system. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(4): 29:1–29:??, October 2009. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cucchiara:2022:FGH

[CF22] Rita Cucchiara and Matteo Fabbri. Fine-grained human analysis under occlusions and perspective constraints in multimedia surveillance. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):32:1– 32:23, February 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3476839.

Carletti:2024:FSB

[CFG^{+}24] Vincenzo Carletti, Pasquale Foggia, Antonio Greco, Alessia Saggese, and Mario Vento. Facial soft-biometrics obfuscation through adversarial attacks. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):331:1–331:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3656474.

Cooper:2005:TEC

[CFGW05] Matthew Cooper, Jonathan Foote, Andreas Girgensohn, and Lynn Wilcox. Temporal event clustering for digital photo collections. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(3):269–288, August 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

CardiaNeto:2023:LSA

[CFN^{+}23] João Baptista Cardia Neto, Claudio Ferrari, Aparecido Nilceu Marana, Stefano Berretti, and Alberto Del Bimbo. Learning streamed attention network from descriptor images for cross-resolution 3D face recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):30:1–30:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3527158.

Canazza:2015:ATM

[CFP15] Sergio Canazza, Carlo Fantozzi, and Niccol‘o Pretto. Accessing tape music documents on mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):20:1–20:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chaudhary:2020:IRC

[CGGC20] Chandramani Chaudhary, Poonam Goyal, Navneet Goyal, and Yi-Ping Phoebe Chen. Image retrieval for complex queries using knowledge embedding. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):13:1–13:23, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3375786.

Chen:2015:TAT

[CGNG15] Shannon Chen, Zhenhuan Gao, Klara Nahrstedt, and Indranil Gupta. 3DTI amphitheater: Towards 3DTI broadcasting. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (2s):47:1–47:??, February 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cecil:2018:NBV

[CGPCR18] J. Cecil, Avinash Gupta, M. Pirela-Cruz, and Parmesh Ramanathan. A network-based virtual reality simulation training approach for orthopedic surgery. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):77:1–77:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cheng:2024:VFH

[CGW^{+}24] Harry Cheng, Yangyang Guo, Tianyi Wang, Qi Li, Xiaojun Chang, and Liqiang Nie. Voiceface homogeneity tells deepfake. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):76:1–76:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3625231.

Chen:2015:AVR

[CH15] Bo-Hao Chen and Shih-Chia Huang. An advanced visibility restoration algorithm for single hazy images. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):53:1–53:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2021:NDB

[CH21] Yizhen Chen and Haifeng Hu.

Y-Net: Dual-branch joint network for semantic segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):137:1–137:22, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460940.

Chang:2013:HFW

[Cha13] Shih-Fu Chang. How far we’ve come: Impact of 20 years of multimedia information retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):42:1–42:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cheung:2024:LPF

[Che24] Ming Cheung. Learning from the past: Fast NAS for tasks and datasets. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):70:1–70:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3618000.

Cesar:2018:BPA

[CHHH18] Pablo Cesar, Cheng-Hsin Hsu, Chun-Ying Huang, and Pan Hui. Best papers of the ACM Multimedia Systems (MM-Sys) Conference 2017 and the ACM Workshop on Network and Operating System Support for Digital Audio and Video (NOSSDAV) 2017. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):60:1– 60:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2025:ACG

[CHHH25] Jili Chen, Qionghao Huang, Changqin Huang, and Xiaodi Huang. Actual cause-guided adaptive gradient scaling for balanced multimodal sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):179:1–179:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2019:ACV

[CHLW19] Yadang Chen, Chuanyan Hao, Alex X. Liu, and Enhua Wu. Appearance-consistent video object segmentation based on a multinomial event model. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):40:1–40:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3321507.

Chou:2013:AIC

[Cho13] Philip A. Chou. Advances in immersive communication: (1) Telephone, (2) Television, (3) Teleportation. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):41:1–41:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chan:2023:LDF

[CHS^{+}23] Patrick P. K. Chan, Xiaoman Hu, Haorui Song, Peng Peng, and Keke Chen. Learning disentangled features for person re-identification under clothes changing. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):180:1–180:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3584359.

Chu:2015:VCS

[Chu15] Chung-Hua Chu. Visual comfort for stereoscopic 3D by using motion sensors on 3D mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):14:1–14:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chan:2023:UFH

[CI23] Ka-Hou Chan and Sio-Kei Im. Using four hypothesis probability estimators for CABAC in versatile video coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):40:1– 40:??, February 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3531015.

Ceballos:2018:IEC

[CIPE18] Rodrigo Ceballos, Beatrice Ionascu, Wanjoo Park, and Mohamad Eid. Implicit emotion communication: EEG classification and haptic feedback. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):3:1–3:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Celard:2025:NMD

[CISF^{+}25] Pedro Celard, Eva Lorenzo Iglesias, Jose Manuel Sorribes-Fernández, Lourdes Borrajo, and Adrián Seara Vieira. New metrics and dataset for biological development video generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):12:1–12:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3653456.

Chen:2021:EAM

[CJHH21] Yung-Yao Chen, Sin-Ye Jhong, Chih-Hsien Hsia, and Kai-Lung Hua. Explainable AI: a multispectral palm-vein identification system with new augmentation features. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):111:1–111:21, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468873.

Chehabeddine:2021:BMH

[CJP^{+}21] Said Chehabeddine, Muhammad Hassan Jamil, Wanjoo Park, Dianne L. Sefo, Peter M. Loomer, and Mohamad Eid. Bimanual haptic-based periodontal simulation with finger support and vibrotactile feedback. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1):28:1–28:17, April 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3421765.

Chen:2024:FBH

[CJZ^{+}24] Xiaowei Chen, Xiao Jiang, Lishuang Zhan, Shihui Guo, Qunsheng Ruan, Guoliang Luo, Minghong Liao, and Yipeng Qin. Full-body human motion reconstruction with sparse joint tracking using flexible sensors. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):44:1–44:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3564700.

Chakareski:2023:MWF

[CKRB23] Jacob Chakareski, Mahmudur Khan, Tanguy Ropitault, and Steve Blandino. Millimeter wave and free-space-optics for future dual-connectivity 6DOF mobile multi-user VR streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):57:1–57:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3544494.

Chen:2007:EMO

[CL07] Herng-Yow Chen and Sheng-Wei Li. Exploring many-to-one speech-to-text correlation for Web-based language learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(3):13:1–13:??, August 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cheng:2012:EIC

[CL12] Xu Cheng and Jiangchuan Liu. Exploring interest correlation for peer-to-peer socialized video sharing. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1):5:1–5:??, January 2012.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Claypool:2018:GID

[Cla18] Mark Claypool. Game input with delay-moving target selection with a game controller thumbstick. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):57:1–57:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cai:2005:LUL

[CLC05] Jianfei Cai, Xiangjun Li, and Chang Wen Chen. Layered unequal loss protection with preinterleaving for fast progressive image transmission over packet-loss channels. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(4):338–353, November 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cui:2021:SSI

[CLN^{+}21] Chaoran Cui, Peiguang Lin, Xiushan Nie, Muwei Jian, and Yilong Yin. Social-sensed image aesthetics assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):103:1–103:19, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3414843.

Chiu:2017:AAS

[CLP17] Chih-Yi Chiu, Yu-Cyuan Liou, and Amorntip Prayoonwong. Approximate asymmetric search for binary embedding codes. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):3:1–3:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2023:BVL

[CLP^{+}23] Jingwen Chen, Jianjie Luo, Yingwei Pan, Yehao Li, Ting Yao, Hongyang Chao, and Tao Mei. Boosting vision-and-language navigation with direction guiding and backtracing. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):9:1–9:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3526024.

Chen:2025:PNM

[CLQQ25] Jiaxin Chen, Xin Liao, Zhenxing Qian, and Zheng Qin. PRest-Net: Multi-domain probability estimation network for robust image forgery detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):89:1–89:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3711930.

Cheung:2017:ECF

[CLS17] Ming Cheung, Xiaopeng Li, and James She. An efficient computation framework for connection discovery using shared images. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):58:1–58:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cheng:2021:ISI

[CLS^{+}21] Wen-Huang Cheng, Jiaying Liu, Nicu Sebe, Junsong Yuan, and Hong-Han Shuai. Introduction to the special issue on explainable AI on multimedia computing. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):108:1–108:2, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3489522.

Chen:2024:GLB

[CLWW24] Honghua Chen, Zhiqi Li, Mingqing Wei, and Jun Wang. Geometric and learning-based mesh denoising: a comprehensive survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):85:1–85:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3625098.

Chen:2021:DAC

[CLZ^{+}21] Bingzhi Chen, Yishu Liu, Zheng Zhang, Yingjian Li, Zhao Zhang, Guangming Lu, and Hongbing Yu. Deep active context estimation for automated COVID-19 diagnosis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):101:1–101:22, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3457124.

Chen:2023:WST

[CLZ^{+}23a] Weidong Chen, Guorong Li, Xinfeng Zhang, Shuhui Wang, Liang Li, and Qingming Huang. Weakly supervised text-based actor-action video segmentation by clip-level multi-instance learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):12:1–12:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3514250.

Chu:2023:RLT

[CLZ^{+}23b] Binfei Chu, Yiting Lin, Bineng Zhong, Zhenjun Tang, Xianxian Li, and Jing Wang. Robust long-term tracking via localizing occluders. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):89:1–89:??, April 2023. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3557896.

Chen:2013:LSM

[CML^{+}13] Xiangyu Chen, Yadong Mu, Hairong Liu, Shuicheng Yan, Yong Rui, and Tat-Seng Chua. Large-scale multilabel propagation based on efficient sparse graph construction. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(1):6:1– 6:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chowdhury:2022:DGS

[CNG22] Debanjan Roy Chowdhury, Sukumar Nandi, and Diganta Goswami. Distributed gateway selection for video streaming in VANET using IP multicast. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):81:1–81:24, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3491388.

Cheng:2011:MPM

[COM^{+}11] Wei Cheng, Wei Tsang Ooi, Sebastien Mondet, Romulus Grigoras, and Géraldine Morin. Modeling progressive mesh streaming: Does data dependency matter? ACM Transactions on Multimedia Computing, Communications, and Applications, 7(2):10:1–10:??, February 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cinar:2021:IJB

[CPCM21] Yusuf Cinar, Peter Pocta, Desmond Chambers, and Hugh Melvin. Improved jitter buffer management for WebRTC. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1):30:1–30:20, April 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3410449.

Chen:2023:RAC

[CPL^{+}23] Jingwen Chen, Yingwei Pan, Yehao Li, Ting Yao, Hongyang Chao, and Tao Mei. Retrieval augmented convolutional encoder-decoder networks for video captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1s):48:1–48:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3539225.

Carbunar:2013:FNA

[CPP^{+}13] Bogdan Carbunar, Rahul Potharaju, Michael Pearce, Venugopal Vasudevan, and Michael Needham. A framework for network aware caching for video on demand systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4): 30:1–30:??, August 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). See errata [CPP^{+}14].

Carbunar:2014:EFN

[CPP^{+}14] Bogdan Carbunar, Rahul Potharaju, Michael Pearce, Venugopal Vasudevan, and Michael Needham. Errata for: A Framework for Network Aware Caching for Video on Demand Systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):23:1– 23:??, September 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). See [CPP^{+}13].

Calagari:2014:AAL

[CPSH14] Kiana Calagari, Mohammad Reza Pakravan, Shervin Shirmohammadi, and Mohamed Hefeeda. ALP: Adaptive loss protection scheme with constant overhead for interactive video applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2):25:1–25:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2024:SSE

[CQA^{+}24] Chen Chen, Lingfeng Qu, Hadi Amirpour, Xingjun Wang, Christian Timmerer, and Zhihong Tian. On the security of selectively encrypted HEVC video bitstreams. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):288:1–288:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672568.

Chen:2025:CAR

[CQA^{+}25] Fan Chen, Lingfeng Qu, Hadi Amirpour, Christian Timmerer, and Hongjie He. Counterfeiting attacks on an RDH-EI scheme based on blockpermutation and Co-XOR. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (4):127:1–127:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2020:LLF

[CRL20] Bin Chen, Lingyan Ruan, and Miu-Ling Lam. LF-GAN: 4D light field synthesis from a single RGB image. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):2:1–2:20, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3366371.

Cheung:2017:ASU

[CS17] Ming Cheung and James She. An analytic system for user gender identification through user shared images. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):30:1– 30:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cetinic:2022:UCA

[CS22] Eva Cetinic and James She.

Understanding and creating art with AI: Review and outlook. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2):66:1–66:22, May 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3475799.

Clement:2024:SDH

[CSBJ24] Nathan Clement, Alan Schoen, Arnold Boedihardjo, and Andrew Jenkins. Synthetic data and hierarchical object detection in overhead imagery. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):117:1–117:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635309.

Cheng:2008:GIP

[CSJ^{+}08] Bin Cheng, Lex Stein, Hai Jin, Xiaofei Liao, and Zheng Zhang. GridCast: Improving peer sharing for P2P VoD. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(4):26:1–26:??, October 2008. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic).

Cheung:2017:PVT

[CSJC17] Ming Cheung, James She, Alvin Junus, and Lei Cao. Prediction of virality timing using cascades in social media. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):2:1–2:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Constantin:2025:EGA

[CSŞ^{+}25] Mihai Gabriel Constantin, Dan-Cristian Stanciu, Liviu-Daniel Ştefan, Mihai Dogariu, Dan Mihăilescu, George Ciobanu, Matt Bergeron, Winston Liu, Konstantin Belov, Octavian Radu, and Bogdan Ionescu. Exploring generative adversarial networks for augmenting network intrusion detection tasks. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):16:1–16:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3689636.

Cheung:2019:DOC

[CSSZ19] Ming Cheung, James She, Weiwei Sun, and Jiantao Zhou. Detecting online counterfeit-goods seller using connection discovery. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):35:1–35:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3311785.

Cheung:2023:SNA

[CSSZ23] Ming Cheung, Weiwei Sun, James She, and Jiantao Zhou. Social network analytic-based online counterfeit seller detection using user shared images. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1):23:1–23:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524135.

Cui:2025:CTR

[CSW^{+}25a] Jia Cui, Jinchen Shen, Jialin Wei, Shiyu Liu, Zhaojia Ye, Shijian Luo, and Zhen Qin. Community transferrable representation learning for image style classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):172:1– 172:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cui:2025:LLE

[CSW^{+}25b] Xiao Cui, Qi Sun, Min Wang, Li Li, Wengang Zhou, and Houqiang Li. LayoutEnc: Leveraging enhanced layout representations for transformerbased complex scene synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (4):120:1–120:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2025:DAM

[CSZ^{+}25] Xiaolin Chen, Xuemeng Song, Jianhui Zuo, Yinwei Wei, Liqiang Nie, and Tat-Seng Chua. Domain-aware multimodal dialog systems with distribution-based user characteristic modeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):66:1–66:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3704811.

Cornia:2022:MFA

[CTBC22] Marcella Cornia, Matteo Tomei, Lorenzo Baraldi, and Rita Cucchiara. Matching faces and attributes between the artistic and the real domain: the PersonArt approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):77:1–77:23, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490033.

Cattelan:2008:WCP

[CTGP08] Renan G. Cattelan, Cesar Teixeira, Rudinei Goularte, and Maria Da Graça C. Pimentel. Watch-and-comment as a paradigm toward ubiquitous interactive video editing. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(4): 28:1–28:??, October 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cesar:2006:GAH

[CVV06] Pablo Cesar, Petri Vuorimaa, and Juha Vierinen. A graphics architecture for high-end interactive television terminals. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(4): 343–357, November 2006. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chetty:2010:MSF

[CW10] Girija Chetty and Matthew White. Multimedia sensor fusion for retrieving identity in biometric access control systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(4):26:1–26:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cao:2025:DBC

[CW25] Yuan Cao and Dong Wang.

Dual-branch cross-layer information flow network for camouflaged object detection in complex scenes. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):294:1–294:19, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chiu:2010:FMH

[CWC10] Chih-Yi Chiu, Hsin-Min Wang, and Chu-Song Chen. Fast min-hashing indexing and robust spatio-temporal matching for detecting video copies. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(2): 10:1–10:??, March 2010. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2024:OMG

[CWL24] Peilin Chen, Shiqi Wang, and Zhu Li. Occupancy map guided attributes artifacts removal for video-based point cloud compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):387:1–387:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3697351.

Cao:2022:SBH

[CWTG22] Jie Cao, Youquan Wang, Haicheng Tao, and Xiang Guo. Sensor-based human activity recognition using graph LSTM and multi-task classification model. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):139:1–139:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561387.

Cao:2025:DRR

[CWW^{+}25] Anqi Cao, Zhijing Wan, Xiao Wang, Wei Liu, Wei Wang, Zheng Wang, and Xin Xu. Diversity-representativeness replay and knowledge alignment for lifelong vehicle reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):63:1–63:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3702998.

Chen:2025:TSC

[CWXW25] Jiayue Chen, Xiaomeng Wang, Tong Xu, and Shiwei Wu. Towards scene-centric multi-level interest mining for video recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):92:1–92:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3712600.

Cizmeci:2017:MSM

[CXC^{+}17] Burak Cizmeci, Xiao Xu, Rahul Chaudhari, Christoph Bachhuber, Nicolas Alt, and Eckehard Steinbach. A multiplexing scheme for multimodal teleoperation. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(2):21:1–21:??, May 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2022:MFT

[CXL^{+}22] Min Chen, Wenjing Xiao, Miao Li, Yixue Hao, Long Hu, and Guangming Tao. A multifeature and time-aware-based stress evaluation mechanism for mental status adjustment. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1s):39:1–39:18, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3462763.

Chen:2008:DDN

[CXS^{+}08] Yinpeng Chen, Weiwei Xu, Hari Sundaram, Thanassis Rikakis, and Sheng-Min Liu. A dynamic decision network framework for online media adaptation in stroke rehabilitation. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1):4:1–4:??, October 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cao:2019:SOG

[CXW^{+}19] Tengfei Cao, Changqiao Xu, Mu Wang, Zhongbai Jiang, Xingyan Chen, Lujie Zhong, and Luigi Alfredo Grieco. Stochastic optimization for green multimedia services in dense 5G networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):79:1– 79:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3328996.

Chen:2024:QQC

[CXX^{+}24] Shuangmin Chen, Rui Xu, Jian Xu, Shiqing Xin, Changhe Tu, Chenglei Yang, and Lin Lu. QuickCSGModeling: Quick CSG operations based on fusing signed distance fields for VR modeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(7):189:1–189:??, July 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3599729.

Cao:2025:GGA

[CXY^{+}25] Zongsheng Cao, Qianqian Xu, Zhiyong Yang, Yuan He, Xiaochun Cao, and Qingming Huang. GAHE: Geometryaware embedding for hyperrelational knowledge graph representation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):205:1–205:26, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chandra:2011:EAS

[CY11] Surendar Chandra and Xuwen Yu. An empirical analysis of serendipitous media sharing among campus-wide wireless users. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(1):6:1–6:??, January 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2025:BSA

[CYM^{+}25] Yipei Chen, Hua Yuan, Baojun Ma, Limin Wang, and Yu Qian. Beyond songs: Analyzing user sentiment through music playlists and multimodal data. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):82:1–82:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3708346.

Chen:2007:DSI

[CYMW07] Datong Chen, Jie Yang, Robert Malkin, and Howard D. Wactlar. Detecting social interactions of the elderly in a nursing home environment. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(1): ??, February 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cai:2024:TAC

[CYW24] Chen Cai, Kim-Hui Yap, and Suchen Wang. Toward attribute-controlled fashion image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):280:1–280:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3671000.

Chen:2023:CCF

[CYZ23] Zhen Chen, Ming Yang, and Shiliang Zhang. Complementary coarse-to-fine matching for video object segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):203:1–203:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3596496.

Chen:2024:BBA

[CYZ^{+}24] Ying Chen, Rui Yao, Yong Zhou, Jiaqi Zhao, Bing Liu, and Abdulmotaleb El Saddik. Black-box attack against selfsupervised video object segmentation models with contrastive loss. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):57:1–57:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617502.

Chen:2024:MCT

[CZ24] Chengxin Chen and Pengyuan Zhang. Modality-collaborative transformer with hybrid feature reconstruction for robust emotion recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):146:1–146:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3640343.

Chen:2015:ADF

[CZC15] Liang Chen, Yipeng Zhou, and Dah Ming Chiu. Analysis and detection of fake views in online video services. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2s):44:1–44:??, February 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Cao:2023:DKP

[CZL^{+}23a] Gaofeng Cao, Fei Zhou, Kanglin Liu, Anjie Wang, and Leidong Fan. A decoupled kernel prediction network guided by soft mask for single image HDR reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):79:1–79:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3550277.

Chen:2023:IFD

[CZL23b] Xiumei Chen, Xiangtao Zheng, and Xiaoqiang Lu. Identity feature disentanglement for visible-infrared person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):201:1–201:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3595183.

Cheng:2022:CMG

[CZQ^{+}22] Yuhao Cheng, Xiaoguang Zhu, Jiuchao Qian, Fei Wen, and Peilin Liu. Cross-modal graph matching network for imagetext retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):95:1–95:23, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3499027.

Cao:2025:DDD

[CZS^{+}25] Xuanqing Cao, Wengang Zhou, Qi Sun, Weilun Wang, Li Li, and Houqiang Li. DISA: Disentangled dual-branch framework for affordance-aware human insertion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):98:1–98:??, March 2025. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3715140.

Chen:2015:PMV

[CZW15] Ke Chen, Zhong Zhou, and Wei Wu. Progressive motion vector clustering for motion estimation and auxiliary tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (3):33:1–33:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chan:2024:AFF

[CZW^{+}24] Sixian Chan, Xianpeng Zeng, Xinhua Wang, Jie Hu, and Cong Bai. Auxiliary feature fusion and noise suppression for HOI detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):299:1–299:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3674980.

Chen:2025:SAG

[CZW^{+}25] Kaixin Chen, Lin Zhang, Zhong Wang, Shengjie Zhao, and Yicong Zhou. Skeleton-aware graph-based adversarial networks for human pose estimation from sparse IMUs. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (4):106:1–106:??, April 2025.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Chen:2024:VVB

[CZX^{+}24] Wuyang Chen, Boqing Zhu, Kele Xu, Yong Dou, and Dawei Feng. VoiceStyle: Voice-based face generation via cross-modal prototype contrastive learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):279:1–279:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3671002.

Chen:2021:FRS

[CZY^{+}21] Chenglizhao Chen, Hongmeng Zhao, Huan Yang, Teng Yu, Chong Peng, and Hong Qin. Full-reference screen content image quality assessment by fusing multilevel structure similarity. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):94:1–94:21, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3447393.

Czekierda:2021:AOO

[CZZ21] Lukasz Czekierda, Krzysztof Zieliński, and Sławomir Zieliński. Automated orchestration of online educational collaboration in cloud-based environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1):31:1–31:26, April 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3412381.

Chen:2023:PHD

[CZZ^{+}23] Haozhe Chen, Hang Zhou, Jie Zhang, Dongdong Chen, Weiming Zhang, Kejiang Chen, Gang Hua, and Nenghai Yu. Perceptual hashing of deep convolutional neural networks for model copy detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):123:1– 123:??, May 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3572777.

Chen:2024:AAL

[CZZ^{+}24] Shuqin Chen, Xian Zhong, Yi Zhang, Lei Zhu, Ping Li, Xiaokang Yang, and Bin Sheng. Action-aware linguistic skeleton optimization network for non-autoregressive video captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):326:1–326:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3679203.

Dao:2017:TCM

[Dao17] Minh Son Dao. This is the table of contents for the most recent online-only supplemental issue TOMM 13(3s). Please find this supplemental issue in the ACM Digital Library and enjoy reading them! ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):47:1–47:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

delBimbo:2016:PEC

[dB16] Alberto del Bimbo. From the past Editor-In-Chief. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(3):37:1– 37:??, June 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

DeDivitiis:2023:DFF

[DBBD23] Lavinia De Divitiis, Federico Becattini, Claudio Baecchi, and Alberto Del Bimbo. Disentangling features for fashion recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):39:1–39:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531017.

Desai:2025:RRS

[DBPM25] Chaitra Desai, Sujay Benur, Ujwala Patil, and Uma Mudenagudi. RSUIGM: Realistic synthetic underwater image generation with image formation model. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):10:1–10:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3656473.

Djenouri:2024:EAG

[DBSL24] Youcef Djenouri, Asma Belhadi, Gautam Srivastava, and Jerry Chun-Wei Lin. An efficient and accurate GPU-based deep learning model for multimedia recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):40:1– 40:??, February 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524022.

Duchowski:2007:FGC

[DÇ07] Andrew T. Duchowski and Arzu Çöltekin. Foveated gazecontingent displays for peripheral LOD management, 3D visualization, and stereo imaging. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4): 6:1–6:18, December 2007. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Araujo:2025:DVP

[dCAGF^{+}25] Gabriel de Castro Araújo, Henrique Domingues Garcia, Mylène C. Q. Farias, Ravi Prakash, and Marcelo M. Carvalho. A 360-degree video player for dynamic video editing applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):249:1–249:23, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Dong:2013:RIA

[DCC^{+}13] Jian Dong, Bin Cheng, Xiangyu Chen, Tat-Seng Chua, Shuicheng Yan, and Xi Zhou. Robust image annotation via simultaneous feature and sample outlier pursuit. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4):24:1–24:??, August 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Doherty:2013:SSA

[DCM13] Jonathan Doherty, Kevin Curran, and Paul Mckevitt. A selfsimilarity approach to repairing large dropouts of streamed music. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3): 20:1–20:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

DeOliveira:2010:LND

[DCO10] Rodrigo De Oliveira, Mauro Cherubini, and Nuria Oliver. Looking at near-duplicate videos from a human-centric perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):15:1–15:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Deb:2024:AIC

[DDK24] Subhrajyoti Deb, Abhilash Das, and Nirmalya Kar. An applied image cryptosystem on Moore’s automaton operating on δ(q_{k})/F_{2}. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):52:1–52:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3614433.

Du:2025:SCI

[DDL^{+}25] Longye Du, Shuaiyu Deng, Ying Li, Jun Li, and Qi Tian. A survey on composed image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):202:1–202:27, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Dornaika:2012:IRF

[DE12] Fadi Dornaika and James H. Elder. Image registration for foveated panoramic sensing. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2): 17:1–17:??, May 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ding:2020:AEU

[DFXY20] Yuhang Ding, Hehe Fan, Mingliang Xu, and Yi Yang. Adaptive exploration for unsupervised person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):3:1–3:19, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3369393.

Demirbilek:2017:MLB

[DG17a] Edip Demirbilek and Jean-Charles Grégoire. Machine learning-based parametric audiovisual quality prediction models for real-time communications. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(2):16:1–16:??, May 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Dutta:2017:EFC

[DG17b] Tanima Dutta and Hari Prabhat Gupta. An efficient framework for compressed domain watermarking in P frames of high-efficiency video coding (HEVC)-encoded video. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):12:1– 12:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Deng:2023:LGL

[DHP23] Zijun Deng, Xiangteng He, and Yuxin Peng. LFR-GAN: Local feature refinement based generative adversarial network for text-to-image generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):207:1– 207:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3589002.

Do:2019:SDC

[DHT^{+}19] Thanh-Toan Do, Tuan Hoang, Dang-Khoa Le Tan, Huu Le, Tam V. Nguyen, and Ngai-Man Cheung. From selective deep convolutional features to compact binary representations for image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):43:1–43:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3314051.

Dongye:2025:DLH

[DJWZ25] Xiaonuo Dongye, Haiyan Jiang, Dongdong Weng, and Zhenliang Zhang. Demonstrative learning for human-agent knowledge transfer. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):64:1– 64:??, February 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3703838.

Dong:2021:SCL

[DKJ^{+}21] Xingbo Dong, Soohyong Kim, Zhe Jin, Jung Yeon Hwang, Sangrae Cho, and Andrew Beng Jin Teoh. Secure chaffless fuzzy vault for face identification systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):79:1–79:22, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3442198.

Deng:2014:DLB

[DL14] Yunhua Deng and Rynson W. H. Lau. Dynamic load balancing in distributed virtual environments using heat diffusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):16:1–16:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Duanmu:2022:BQE

[DLC^{+}22] Zhengfang Duanmu, Wentao Liu, Diqi Chen, Zhuoran Li, Zhou Wang, Yizhou Wang, and Wen Gao. A Bayesian qualityof-experience model for adaptive streaming videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):141:1– 141:??, October 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3491432.

Duan:2022:NMS

[DLD^{+}22] Mingxing Duan, Kenli Li, Jiayan Deng, Bin Xiao, and Qi Tian. A novel multi-sample generation method for adversarial attacks. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):112:1–112:21, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3506852.

Dong:2018:LMK

[DLL^{+}18] Husheng Dong, Ping Lu, Chunping Liu, Yi Ji, and Shengrong Gong. Learning multiple kernel metrics for iterative person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):78:1–78:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Duan:2020:FEM

[DLL^{+}20] Mingxing Duan, Kenli Li, Xiangke Liao, Keqin Li, and Qi Tian. Features-enhanced multi-attribute estimation with convolutional tensor correlation fusion network. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1–23, January 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3355542.

Dai:2024:TLF

[DLL^{+}24] Kuai Dai, Xutao Li, Huiwei Lin, Yin Jiang, Xunlai Chen, Yunming Ye, and Di Xian. TinyPredNet: a lightweight framework for satellite image sequence prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (5):142:1–142:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3638773.

Duan:2020:EFE

[DLO^{+}20] Mingxing Duan, Kenli Li, Aijia Ouyang, Khin Nandar Win, Keqin Li, and Qi Tian. EGroupNet: a feature-enhanced network for age estimation with novel age group schemes. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):42:1–42:23, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3379449.

DeBoer:2017:SRZ

[DLZ^{+}17] Maaike H. T. De Boer, Yi-Jie Lu, Hao Zhang, Klamer Schutte, Chong-Wah Ngo, and Wessel Kraaij. Semantic reasoning in zero example video event retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):60:1–60:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Daronco:2017:DRA

[DMF17] Stefano D’aronco, Sergio Mena, and Pascal Frossard. Distributed rate allocation in switch-based multiparty videoconferencing system. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):41:1– 41:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Duan:2020:VAA

[DMF^{+}20] Huiyu Duan, Xiongkuo Min, Yi Fang, Lei Fan, Xiaokang Yang, and Guangtao Zhai. Visual attention analysis and prediction on human faces for children with autism spectrum disorder. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1–23, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3337066.

Gomes:2024:AAG

[dMGRT24] Pedro de Medeiros Gomes, Silvia Rossi, and Laura Toni. AGAR — attention graph-RNN for adaptative motion prediction of point clouds of deformable objects. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):245:1–245:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3662183.

Santos:2018:HAS

[DMSRL18] Joel A. F. Dos Santos, Débora C. Muchaluat-Saade, Cécile Roisin, and Nabil Layaïda. A hybrid approach for spatio-temporal validation of declarative multimedia documents. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(4):86:1–86:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Dong:2023:SEG

[DNL^{+}23] Shanshan Dong, Tianzi Niu, Xin Luo, Wu Liu, and Xinshun Xu. Semantic embedding guided attention with explicit visual feature fusion for video captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):68:1–68:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3550276.

Dang-Nguyen:2017:MRD

[DNPG^{+}17] Duc-Tien Dang-Nguyen, Luca Piras, Giorgio Giacinto, Giulia Boato, and Francesco G. B. De Natale. Multimodal retrieval with diversification and relevance feedback for tourist attraction images. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):49:1–49:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Dong:2023:VFI

[DOD23] Jiong Dong, Kaoru Ota, and Mianxiong Dong. Video frame interpolation: a comprehensive survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):78:1–78:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3556544.

DelBimbo:2006:CBR

[DP06] Alberto Del Bimbo and Pietro Pala. Content-based retrieval of 3D models. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(1):20–43, February 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ding:2023:BRD

[DPL^{+}23] Xuewei Ding, Yingwei Pan, Yehao Li, Ting Yao, Dan Zeng, and Tao Mei. Boosting relationship detection in images with multi-granular self-supervised learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):88:1–88:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3556978.

Dogariu:2022:GRS

[DŞB^{+}22] Mihai Dogariu, Liviu-Daniel Ştefan, Bogdan Andrei Boteanu, Claudiu Lamba, Bomi Kim, and Bogdan Ionescu. Generation of realistic synthetic financial time-series. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):96:1–96:27, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3501305.

Dai:2022:FFS

[DSL^{+}22] Hanbin Dai, Hailin Shi, Wu Liu, Linfang Wang, Yinglu Liu, and Tao Mei. FasterPose: a faster simple baseline for human pose estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):103:1–103:16, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3503464.

Ding:2025:VIP

[DSL^{+}25] Hao Ding, Jing Sun, Rui Long, Xiaoping Jiang, Hongling Shi, Yuting Qin, Zongze Li, and Jian-Jin Li. Visibleinfrared person re-identification based on feature decoupling and refinement. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):265:1–265:16, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Dhiman:2021:PWS

[DVA21] Chhavi Dhiman, Dinesh Kumar Vishwakarma, and Paras Agarwal. Part-wise spatio-temporal attention driven CNN-based 3D human action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (3):86:1–86:24, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3441628.

Du:2021:RTE

[DWC^{+}21] Gaoming Du, Jiting Wu, Hongfang Cao, Kun Xing, Zhenmin Li, Duoli Zhang, and Xiaolei Wang. A real-time effective fusion-based image defogging architecture on FPGA. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (3):93:1–93:21, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3446241.

Du:2023:WSH

[DWL^{+}23] Yongchao Du, Min Wang, Zhenbo Lu, Wengang Zhou, and Houqiang Li. Weakly supervised hashing with reconstructive cross-modal attention. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):208:1–208:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589185.

Dai:2024:UDA

[DWS^{+}24] Qingfeng Dai, Yongkang Wong, Guofei Sun, Yanwei Wang, Zhou Zhou, Mohan S. Kankanhalli, Xiangdong Li, and Weidong Geng. Unsupervised domain adaptation by causal learning for biometric signalbased HCI. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):49:1–49:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3583885.

Du:2021:IGS

[DXY^{+}21] Yong Du, Yangyang Xu, Taizhong Ye, Qiang Wen, Chufeng Xiao, Junyu Dong, Guoqiang Han, and Shengfeng He. Invertible grayscale with sparsity enforcing priors. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):97:1– 97:17, August 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3451993.

Deng:2009:SCA

[DY09] Robert H. Deng and Yanjiang Yang. A study of content authentication in proxy-enabled multimedia delivery systems: Model, techniques, and applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(4):28:1–28:??, October 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Deng:2025:ITA

[DYL^{+}25] Jiacheng Deng, Dengpan Ye, Jizhi Li, Ziyi Liu, Long Tang, and Yunming Zhang. The interpretable and transferable adversarial attack against synthetic speech detectors. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):135:1– 135:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Du:2025:EEI

[DYN^{+}25] Qianqian Du, Hui Yin, Lang Nie, Yanting Liu, and Jin Wan. EnIter: Enhancing iterative multi-view depth estimation with universal contextual hints. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):169:1–169:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Deng:2014:TFP

[DYSX14] Zhengyu Deng, Ming Yan, Jitao Sang, and Changsheng Xu. Twitter is faster: Personalized time-aware video recommendation from Twitter to YouTube. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (2):31:1–31:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Duan:2023:MGL

[DYZ^{+}23] Wei Duan, Yi Yu, Xulong Zhang, Suhua Tang, Wei Li, and Keizo Oyama. Melody generation from lyrics with local interpretability. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):124:1–124:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572031.

Ding:2021:MSA

[DZW^{+}21] Yaoling Ding, Liehuang Zhu, An Wang, Yuan Li, Yongjuan Wang, Siu Ming Yiu, and Keke Gai. A multiple sieve approach based on artificial intelligent techniques and correlation power analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):71:1– 71:21, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3433165.

Dou:2023:MTP

[DZWH23] Peng Dou, Ying Zeng, Zhuoqun Wang, and Haifeng Hu. Multiple temporal pooling mechanisms for weakly supervised temporal action localization. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):108:1–108:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3567828.

ElSaddik:2025:UCM

[EAK^{+}25] Abdulmotaleb El Saddik, Jamil Ahmad, Mustaqeem Khan, Saad Abouzahir, and Wail Gueaieb. Unleashing creativity in the metaverse: Generative AI and multimodal content. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):186:1–186:43, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ebrahim:2016:MIB

[EC16] Mansoor Ebrahim and Wai Chong Chia. Multiview image block compressive sensing with joint multiphase decoding for visual sensor network. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):30:1–30:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Effelsberg:2013:PLB

[Eff13] Wolfgang Effelsberg. A personal look back at twenty years of research in multimedia content analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):43:1–43:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ergasti:2025:MPM

[EFF^{+}25] Alex Ergasti, Tomaso Fontanini, Claudio Ferrari, Massimo Bertozzi, and Andrea Prati. MARS: Paying more attention to visual attributes for textbased person search. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):281:1– 281:22, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Engelbrecht:2017:PDS

[EG17] Herman A. Engelbrecht and John S. Gilmore. Pithos: Distributed storage for massive multi-user virtual environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):31:1–31:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Eide:2006:RTV

[EGEM06] Viktor S. Wold Eide, Ole-Christoffer Granmo, Frank Eliassen, and Jørgen Andreas Michaelsen. Real-time video content analysis: QoS-aware application composition and parallel processing. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(2):149–172, May 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ehsani:2025:ETQ

[EL25] Sina Ehsani and Jian Liu. Elevating textual question answering with on-demand visual augmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):287:1–287:25, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Erdmann:2009:IEB

[ENHN09] Maike Erdmann, Kotaro Nakayama, Takahiro Hara, and Shojiro Nishio. Improving the extraction of bilingual terminology from Wikipedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(4):31:1–31:??, October 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Etsion:2006:PPU

[ETF06] Yoav Etsion, Dan Tsafrir, and Dror G. Feitelson. Process prioritization using output production: Scheduling for multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(4):318–342, November 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

ElEssaili:2015:QBC

[EWSZ15] Ali El Essaili, Zibin Wang, Eckehard Steinbach, and Liang Zhou. QoE-based cross-layer optimization for uplink video transmission. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1):2:1–2:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Floris:2018:QAO

[FAA18] Alessandro Floris, Arslan Ahmad, and Luigi Atzori. QoEaware OTT-ISP collaboration in service management: Architecture and approaches. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):36:1– 36:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Francis:2022:GRS

[FBG22] Jobin Francis, M. Baburaj, and Sudhish N. George. An l_{1/2} and graph regularized subspace clustering method for robust image segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):53:1– 53:24, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3476514.

Ferrari:2023:CRR

[FBGD23] Claudio Ferrari, Federico Becattini, Leonardo Galteri, and Alberto Del Bimbo. (Compress and Restore) N: a robust defense against adversarial attacks on image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1s):26:1–26:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524619.

Fani:2025:TSV

[FBSGR25] Davoud Fani, Aliasghar Beheshti-Shirazi, Mohammad Ghanbari, and Esmatollah Rezaei. On temporal smoothness of video reconstruction quality in the DCVS via non-uniform sampling. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):156:1–156:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fincato:2022:TWD

[FCL^{+}22] Matteo Fincato, Marcella Cornia, Federico Landi, Fabio Cesari, and Rita Cucchiara. Transform, warp, and dress: a new transformation-guided model for virtual try-on. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):62:1– 62:24, May 2022. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491226.

Fontanini:2023:UDM

[FDBP23] Tomaso Fontanini, Luca Donati, Massimo Bertozzi, and Andrea Prati. Unsupervised discovery and manipulation of continuous disentangled factors of variation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):189:1–189:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3591358.

Feng:2011:SRI

[FDKB11] Wu-Chi Feng, Thanh Dang, John Kassebaum, and Tim Bauman. Supporting region-of-interest cropping through constrained compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):17:1– 17:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fan:2025:GHQ

[FDZ^{+}25] Bing Fan, Feng Ding, Guopu Zhu, Jiwu Huang, Sam Kwong, Pradeep K. Atrey, and Siwei Lyu. Generating higher-quality anti-forensics DeepFakes with adversarial sharpening mask. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):180:1–180:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Franti:2023:DPC

[FF23] Pasi Fränti and Nancy Fazal. Design principles for content creation in location-based games. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):165:1–165:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3583689.

Fu:2024:SOT

[FFCM24] Fengyi Fu, Shancheng Fang, Weidong Chen, and Zhendong Mao. Sentiment-oriented transformer-based variational autoencoder network for live video commenting. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):104:1– 104:??, April 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3633334.

Feng:2020:LJS

[FH20] Shenming Feng and Haifeng Hu. Learning joint structure for human pose estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):85:1– 85:17, September 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3392302.

Feng:2024:ADD

[FHF^{+}24] Yuan Feng, Yaojun Hu, Pengfei Fang, Sheng Liu, Yanhong Yang, and Shengyong Chen. Asymmetric dual-decoder U-Net for joint rain and haze removal. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):87:1–87:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3628451.

FujiiPontello:2017:MUR

[FHG^{+}17] Luciana Fujii Pontello, Pedro H. F. Holanda, Bruno Guilherme, João Paulo V. Cardoso, Olga Goussevskaia, and Ana Paula Couto Da Silva. Mixtape: Using real-time user feedback to navigate large media collections. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):50:1–50:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fan:2022:MUE

[FHH22] Ching-Ling Fan, Tse-Hou Hung, and Cheng-Hsin Hsu. Modeling the user experience of watching 360^{∘} videos with head-mounted displays. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):3:1– 3:23, January 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3463825.

Feihong:2023:THQ

[FHK^{+}23] Lu Feihong, Chen Hang, Li Kang, Deng Qiliang, Zhao Jian, Zhang Kaipeng, and Han Hong. Toward high-quality face-mask occluded restoration. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1):24:1–24:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524137.

Feng:2023:MMK

[FHP23] Duoduo Feng, Xiangteng He, and Yuxin Peng. MKVSE: Multimodal knowledge enhanced visual-semantic embedding for image-text retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):162:1– 162:??, September 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3580501.

Fang:2025:CLD

[FJF^{+}25] Jiajie Fang, Mengjuan Jiang, Jiaqing Fan, Bangjun Wang, and Fanzhang Li. Complementarily learning decoupled category-region-aware prototype for few-shot classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):206:1–206:22, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Feng:2020:UTB

[FKCW20] Zhen-Hua Feng, Josef Kittler, Bill Christmas, and Xiao-Jun Wu. A unified tensorbased active appearance model. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–22, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3338841.

Feng:2005:PSL

[FKFB05] Wu-Chi Feng, Ed Kaiser, Wu Chang Feng, and Mikael Le Baillif. Panoptes: scalable lowpower video sensor networking technologies. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(2):151–167, May 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Farhat:2022:CCC

[FKW22] Farshid Farhat, Mohammad Mahdi Kamani, and James Z. Wang. CAPTAIN: Comprehensive composition assistance for photo taking. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):14:1–14:24, January 2022. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462762.

Fu:2021:DGL

[FLG^{+}21] Sichao Fu, Weifeng Liu, Weili Guan, Yicong Zhou, Dapeng Tao, and Changsheng Xu. Dynamic graph learning convolutional networks for semisupervised classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):4:1– 4:13, March 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3412846.

Ferrara:2006:SWO

[FLM^{+}06] Alfio Ferrara, Luca A. Ludovico, Stefano Montanelli, Silvana Castano, and Goffredo Haus. A Semantic Web ontology for context-based classification and retrieval of music resources. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(3):177–198, August 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Feng:2012:CAO

[FLZ^{+}12] Hui Feng, Hefei Ling, Fuhao Zou, Weiqi Yan, and Zhengding Lu. A collusion attack optimization strategy for digital fingerprinting. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2S):36:1–36:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Francis:2020:UTF

[FMG20] Jobin Francis, Baburaj M., and Sudhish N. George. A unified tensor framework for clustering and simultaneous reconstruction of incomplete imaging data. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):92:1–92:24, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3399806.

Franti:2017:MMO

[FMIS17] Pasi Fränti, Radu Mariescu-Istodor, and Lahari Sengupta. O-Mopsi: Mobile orienteering game for sightseeing, exercising, and education. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):56:1– 56:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Feng:2022:CSL

[FNH22] Shenming Feng, Xingzhong Nong, and Haifeng Hu. Cascaded structure-learning network with using adversarial training for robust facial landmark detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):49:1–49:20, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474595.

Floris:2024:CMP

[FPA24] Alessandro Floris, Simone Porcu, and Luigi Atzori. Controlling media player with hands: a transformer approach and a quality of experience assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):132:1–132:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638560.

Franke:2008:TAC

[FPH^{+}08] Ingmar S. Franke, Sebastian Pannasch, Jens R. Helmert, Robert Rieger, Rainer Groh, and Boris M. Velichkovsky. Towards attention-centered interfaces: an aesthetic evaluation of perspective with eye tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3): 18:1–18:??, August 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Feng:2025:ASS

[FQJ^{+}25] Yuchao Feng, Mengjie Qin, Jiawei Jiang, Jintao Lai, and Jianwei Zheng. Axial-shunted spatial-temporal conversation for change detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):141:1– 141:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Falcon:2025:ISI

[FSEW25] Alex Falcon, Giuseppe Serra, Sergio Escalera, and Michael Wray. Introduction to the special issue on text-multimedia retrieval: Retrieving multimedia data by means of natural language. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):274:1–274:4, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fu:2015:QBS

[FSK^{+}15] Bo Fu, Dirk Staehle, Gerald Kunzmann, Eckehard Steinbach, and Wolfgang Kellerer. QoE-based SVC layer dropping in LTE networks using content-aware layer priorities. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (1):7:1–7:??, August 2015. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fang:2014:DGI

[FSX14] Quan Fang, Jitao Sang, and Changsheng Xu. Discovering geo-informative attributes for location recognition and exploration. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):19:1–19:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Feng:2015:CAC

[FWLA15] Fangxiang Feng, Xiaojie Wang, Ruifan Li, and Ibrar Ahmad. Correspondence autoencoders for cross-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):26:1– 26:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fu:2023:LPA

[FXF^{+}23] Zilong Fu, Hongtao Xie, Shancheng Fang, Yuxin Wang, Mengting Xing, and Yongdong Zhang. Learning pixel affinity pyramid for arbitrary-shaped text detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):29:1–29:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3524617.

Feng:2024:EVT

[FXMZ24] Zhanzhou Feng, Jiaming Xu, Lei Ma, and Shiliang Zhang. Efficient video transformers via spatial-temporal token merging for action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):120:1– 120:??, April 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3633781.

Friedland:2010:DAS

[FYH10] Gerald Friedland, Chuohao Yeo, and Hayley Hung. Dialocalization: Acoustic speaker diarization and visual localization as joint optimization problem. ACM Transactions on Multimedia Computing, Communications, and Applications, 6 (4):27:1–27:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fang:2024:IAS

[FYW^{+}24] Jing Fang, Yinbo Yu, Zhongyuan Wang, Xin Ding, and Ruimin Hu. An image arbitrary-scale super-resolution network using frequency-domain information. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):81:1–81:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3616376.

Fu:2021:FAA

[FYY^{+}21] Yunfei Fu, Hongchuan Yu, Chih-Kuo Yeh, Tong-Yee Lee, and Jian J. Zhang. Fast accurate and automatic brushstroke extraction. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):44:1–44:24, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429742.

Fang:2021:PPM

[FYZ^{+}21] Liming Fang, Changchun Yin, Juncen Zhu, Chunpeng Ge, M. Tanveer, Alireza Jolfaei, and Zehong Cao. Privacy protection for medical data sharing in smart healthcare. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):100:1– 100:18, January 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3408322.

Fan:2025:SRR

[FZW^{+}25a] Linlin Fan, Mingliang Zhou, Xuekai Wei, Yong Feng, Tao Xiang, Bin Fang, Zhaowei Shang, Fan Jia, Xu Zhuang, Huayan Pu, and Jun Luo. Sparse reduced-rank fully connected layers with its applications in detection and classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):155:1–155:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fang:2025:SML

[FZW^{+}25b] Hao Fang, Haoyuan Zhao, Feng Wang, Yi Ching Chou, Long Chen, Jianxin Shi, and Jiangchuan Liu. Streaming media over LEO satellite networking: a measurementbased analysis and optimization. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):251:1–251:24, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fan:2020:RAN

[FZYW20] Hehe Fan, Linchao Zhu, Yi Yang, and Fei Wu. Recurrent attention network with reinforced generator for visual dialog. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):78:1–78:16, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3390891.

Fan:2018:UPR

[FZYY18] Hehe Fan, Liang Zheng, Chenggang Yan, and Yi Yang. Unsupervised person re-identification: Clustering and fine-tuning. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (4):83:1–83:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Fu:2025:SRC

[FZZ^{+}25] Ao Fu, Jiaqi Zhao, Yong Zhou, Wenliang Du, Rui Yao, and Abdulmotaleb El Saddik. Similarity regulation and calibration alignment for weakly supervised text-based person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):96:1–96:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3711861.

Ghinea:2012:SSS

[GA12a] Georghita Ghinea and Oluwakemi Ademoye. The sweet smell of success: Enhancing multimedia applications with olfaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1): 2:1–2:??, January 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ghinea:2012:UPM

[GA12b] Gheorghita Ghinea and Oluwakemi Ademoye. User perception of media content association in olfaction-enhanced multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 8 (4):52:1–52:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guerrini:2017:IFR

[GAB^{+}17] Fabrizio Guerrini, Nicola Adami, Sergio Benini, Alberto Piacenza, Julie Porteous, Marc Cavazza, and Riccardo Leonardi. Interactive film recombination. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (4):52:1–52:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ghorbanpour:2025:RRD

[GAH25] Ali Ghorbanpour, Mohammad Amin Arab, and Mohamed Hefeeda. RDIAS: Robust and decentralized image authentication system. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):303:1–303:28, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ge:2025:BVC

[GCC^{+}25] Jiawei Ge, Jiuxin Cao, Xiangmei Chen, Xuelin Zhu, Weijia Liu, Chang Liu, Kun Wang, and Bo Liu. Beyond visual cues: Synchronously exploring target-centric semantics for vision-language tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):136:1–136:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guan:2021:UPS

[GCF^{+}21] Weili Guan, Zhaozheng Chen, Fuli Feng, Weifeng Liu, and Liqiang Nie. Urban perception: Sensing cities via a deep interactive multi-task learning framework. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):13:1–13:20, March 2021. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3424115.

Gramoli:2025:GED

[GCL^{+}25] Lysa Gramoli, Julien Cumin, Jérémy Lacoche, Anthony Foulonneau, Bruno Arnaldi, and Valérie Gouranton. Generating and evaluating data of daily activities with an autonomous agent in a virtual smart home. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):13:1–13:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665331.

Garcia:2024:SNF

[GCTG24] Roberto García, Ana Cediel, Mercè Teixidó, and Rosa Gil. Semantics and non-fungible tokens for copyright management on the metaverse and beyond. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (7):186:1–186:??, July 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3585387.

Guo:2024:DLS

[GCZ^{+}24] Kehua Guo, Liang Chen, Xiangyuan Zhu, Xiaoyan Kui, Jian Zhang, and Heyuan Shi. Double-layer search and adaptive pooling fusion for reference-based image superresolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):15:1–15:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604937.

Ghinea:2007:ISI

[GDGC07] Gheorghita Ghinea, Chabane Djeraba, Stephen Gulliver, and Kara Pernice Coyne. Introduction to special issue on eyetracking applications in multimedia systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4):1:1–1:4, December 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gao:2023:NGA

[GDM^{+}23] Honghao Gao, Baobin Dai, Huaikou Miao, Xiaoxian Yang, Ramon J. Duran Barroso, and Hussain Walayat. A novel GAPG approach to automatic property generation for formal verification: The GAN perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):16:1–16:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3517154.

Gaddam:2015:COM

[GEL^{+}15] Vamsidhar Reddy Gaddam, Ragnhild Eg, Ragnar Langseth, Carsten Griwodz, and Pål Halvorsen. The cameraman operating my virtual camera is artificial: Can the machine be as good as a human? ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):56:1–56:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Georganas:2005:EBA

[Geo05] Nicolas D. Georganas. Editorial: The birth of the ACM Transactions on Multimedia Computing, Communications and Applications (TOMC-CAP). ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1):1–2, February 2005. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Grant:2017:CSU

[GF17] Jason M. Grant and Patrick J. Flynn. Crowd scene understanding from video: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (2):19:1–19:??, May 2017. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gonina:2014:SMC

[GFB^{+}14] Ekaterina Gonina, Gerald Friedland, Eric Battenberg, Penporn Koanantakool, Michael Driscoll, Evangelos Georganas, and Kurt Keutzer. Scalable multimedia content analysis on parallel platforms using Python. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):18:1–18:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gulliver:2006:DUP

[GG06] Stephen R. Gulliver and Gheorghita Ghinea. Defining user perception of distributed multimedia quality. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(4):241–257, November 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gutterman:2020:RRT

[GGA^{+}20] Craig Gutterman, Katherine Guo, Sarthak Arora, Trey Gilliland, Xiaoyang Wang, Les Wu, Ethan Katz-Bassett, and Gil Zussman. Requet: Realtime QoE metric detection for encrypted YouTube traffic. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2s):71:1–71:28, July 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3394498.

Gaeta:2014:DDI

[GGB14] Rossano Gaeta, Marco Grangetto, and Lorenzo Bovio. DIP: Distributed Identification of Polluters in P2P live streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (3):24:1–24:??, April 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guarnera:2024:MDD

[GGB24] Luca Guarnera, Oliver Giudice, and Sebastiano Battiato. Mastering deepfake detection: a cutting-edge approach to distinguish GAN and diffusionmodel images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):343:1–343:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3652027.

Guo:2022:ATA

[GGML22] Yang Guo, Wei Gao, Siwei Ma, and Ge Li. Accelerating transform algorithm implementation for efficient intra coding of 8K UHD videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):113:1–113:20, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3507970.

Gragnaniello:2025:VFR

[GGSV25] Diego Gragnaniello, Antonio Greco, Carlo Sansone, and Bruno Vento. Video fire recognition using zero-shot visionlanguage models guided by a task-aware object detector. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):288:1–288:24, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gokarn:2025:RMR

[GHAM25] Ila Gokarn, Yigong Hu, Tarek Abdelzaher, and Archan Misra. RA-MOSAIC: Resource adaptive edge AI optimization over spatially multiplexed video streams. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):247:1–247:25, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gopalan:2006:SAC

[GHP^{+}06] Kartik Gopalan, Lan Huang, Gang Peng, Tzi-Cker Chiueh, and Yow-Jian Lin. Statistical admission control using delay distribution measurements. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(4): 258–281, November 2006. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gao:2024:SSM

[GHQ24] Xiang Gao, Wei Hu, and Guo-Jun Qi. Self-supervised multiview learning via auto-encoding 3D transformations. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):19:1– 19:??, January 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3597613.

Guo:2022:DIE

[GHR^{+}22] Kehua Guo, Min Hu, Sheng Ren, Fangfang Li, Jian Zhang, Haifu Guo, and Xiaoyan Kui. Deep illumination-enhanced face super-resolution network for low-light images. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):87:1– 87:19, August 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3495258.

Gebre:2024:ETS

[GHS^{+}24] Daniel Gebre, Siem Hadish, Aron Sbhatu, Moayad Aloqaily, and Mohsen Guizani. Establishing trust and security in decentralized metaverse: a Web 3.0 approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):389:1–389:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3696454.

Gu:2024:PPR

[GHW^{+}24] Xiaoling Gu, Jie Huang, Yongkang Wong, Jun Yu, Jianping Fan, Pai Peng, and Mohan S. Kankanhalli. PAINT: Photo-realistic fashion design synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):48:1–48:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3545610.

Guo:2024:DYC

[GHY^{+}24] Zhenjiang Guo, Xiaohai He, Yu Yang, Linbo Qing, and Honggang Chen. DAG-YOLO: a context-feature adaptive fusion rotating detection network in remote sensing images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):304:1–304:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674978.

Guo:2025:ULT

[GHZW25] Xiang Guo, Ruimin Hu, Dong Liang Zhu, and Mei Wang. Uniform light transformer for person re-identification under complex illumination. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):272:1–272:18, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gomes:2010:STA

[GIL^{+}10] João V. P. Gomes, Pedro R. M. Inácio, Branka Lakic, Mário M. Freire, Henrique J. A. Da Silva, and Paulo P. Monteiro. Source traffic analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):21:1–21:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gudmundsson:2018:PWS

[GJAF18] Gylfi Tór Gudmundsson, Björn Tór Jónsson, Laurent Amsaleg, and Michael J. Franklin. Prototyping a Web-scale multimedia retrieval service using Spark. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (3s):65:1–65:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ge:2025:FGA

[GJY^{+}25] Shiping Ge, Zhiwei Jiang, Yafeng Yin, Cong Wang, Zifeng Cheng, and Qing Gu. Finegrained alignment network for zero-shot cross-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):275:1–275:24, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gaj:2017:DCR

[GKSB17] Sibaji Gaj, Aditya Kanetkar, Arijit Sur, and Prabin Kumar Bora. Drift-compensated robust watermarking algorithm for H.265/HEVC video stream. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (1):11:1–11:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Goel:2008:LLA

[GKW08] Ashvin Goel, Charles Krasic, and Jonathan Walpole. Lowlatency adaptive streaming over TCP. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3):20:1–20:??, August 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gleicher:2008:RCI

[GL08] Michael L. Gleicher and Feng Liu. Re-cinematography: Improving the camerawork of casual video. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1):2:1–2:??, October 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gopinathan:2011:OLM

[GL11] Ajay Gopinathan and Zongpeng Li. Optimal layered multicast. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(2):7:1–7:??, February 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gopinathan:2012:ASO

[GL12] Ajay Gopinathan and Zongpeng Li. Algorithms for stochastic optimization of multicast content delivery with network coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(4):56:1–56:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Goh:2005:SFD

[GLC05] Kingshy Goh, Beitao Li, and Edward Y. Chang. Semantics and feature discovery via confidence-based ensemble. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(2): 168–189, May 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gao:2025:MIM

[GLGG25] Lei Gao, Kai Liu, Zheng Guo, and Ling Guan. Mathematicsinspired models: a green and interpretable learning paradigm for multimedia computing. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):197:1–197:22, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ge:2022:DVD

[GLL^{+}22] Shiming Ge, Fanzhao Lin, Chenyu Li, Daichi Zhang, Weiping Wang, and Dan Zeng.

Deepfake video detection via predictive representation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):115:1–115:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3536426.

Gao:2024:IMC

[GLL^{+}24] Jing Gao, Peng Li, Asif Ali Laghari, Gautam Srivastava, Thippa Reddy Gadekallu, Sidra Abbas, and Jianing Zhang. Incomplete multiview clustering via semidiscrete optimal transport for multimedia data mining in IoT. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(6):158:1–158:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3625548.

Grigorev:2020:DDD

[GLT^{+}20] Aleksei Grigorev, Shaohui Liu, Zhihong Tian, Jianxin Xiong, Seungmin Rho, and Jiang Feng. Delving deeper in drone-based person re-id by employing deep decision forest and attributes fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):25:1–25:15, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3360050.

Gao:2020:EDL

[GLW20] Zan Gao, Yinming Li, and Shaohua Wan. Exploring deep learning for view-based 3D model retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1):18:1– 18:21, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377876.

Gan:2019:MSF

[GLWK19] Tian Gan, Junnan Li, Yongkang Wong, and Mohan S. Kankanhalli. A multi-sensor framework for personal presentation analytics. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):30:1–30:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3300941.

Gao:2023:RTI

[GLZW23] Qiqi Gao, Jie Li, Tiejun Zhao, and Yadong Wang. Real-time image enhancement with attention aggregation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):98:1–98:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3564607.

Guo:2025:ISI

[GMWW25] Dan Guo, Troy McDaniel, Shuhui Wang, and Meng Wang. Introduction to the special issue on deep learning for robust human body language understanding. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):103:1–103:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gokhale:2017:CCN

[GNC17] Vineet Gokhale, Jayakrishnan Nair, and Subhasis Chaudhuri. Congestion control for networkaware telehaptic communication. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(2):17:1–17:??, May 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guo:2023:MBR

[GNC^{+}23] Yangyang Guo, Liqiang Nie, Harry Cheng, Zhiyong Cheng, Mohan Kankanhalli, and Alberto Del Bimbo. On modality bias recognition and reduction. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):103:1–103:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3565266.

Gatica-Perez:2012:ISS

[GPHOH12] Daniel Gatica-Perez, Gang Hua, Wei Tsang Ooi, and Pål Halvorsen. Introduction to the special section of best papers of ACM Multimedia 2011. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):38:1– 38:??, September 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guo:2024:FDN

[GQG^{+}24] Wen Guo, Wuzhou Quan, Junyu Gao, Tianzhu Zhang, and Changsheng Xu. Feature disentanglement network: Multi-object tracking needs more differentiated features. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (3):83:1–83:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3626825.

Guo:2024:PDP

[GQSG24] Jintao Guo, Lei Qi, Yinghuan Shi, and Yang Gao. PLACE Dropout: a progressive layerwise and channel-wise dropout for domain generalization. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (3):65:1–65:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3624015.

Ghandeharizadeh:2011:CPS

[GS11a] Shahram Ghandeharizadeh and Shahin Shayandeh. Call for papers: Special issue on 3D mobile multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(4):41:1–41:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ghandeharizadeh:2011:DCC

[GS11b] Shahram Ghandeharizadeh and Shahin Shayandeh. Domical cooperative caching for streaming media in wireless home networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(4):40:1–40:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gupta:2018:AGM

[GS18] Abhinav Gupta and Divya Singhal. Analytical global median filtering forensics based on moment histograms. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2):44:1– 44:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gupta:2019:SGM

[GS19] Abhinav Gupta and Divya Singhal. A simplistic global median filtering forensics based on frequency domain analysis of image residuals. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):71:1– 71:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3321508.

Galteri:2022:LLB

[GSB^{+}22] Leonardo Galteri, Lorenzo Seidenari, Pietro Bongini, Marco Bertini, and Alberto Del Bimbo. LANBIQUE: LANguage-based Blind Image QUality Evaluation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):116:1– 116:??, June 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3538649.

Gupta:2021:VSB

[GSDT21] Shikha Gupta, Krishan Sharma, Dileep Aroor Dinesh, and Veena Thenkanidiyoor. Visual semantic-based representation learning using deep CNNs for scene recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):53:1–53:24, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436494.

Gill:2008:SDM

[GSM^{+}08] Phillipa Gill, Liqi Shi, Anirban Mahanti, Zongpeng Li, and Derek L. Eager. Scalable on-demand media streaming for heterogeneous clients. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1): 8:1–8:??, October 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gou:2024:HMA

[GSY^{+}24] Jianping Gou, Liyuan Sun, Baosheng Yu, Shaohua Wan, and Dacheng Tao. Hierarchical multi-attention transfer for knowledge distillation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):51:1–51:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3568679.

Ghinea:2014:ISI

[GTLG14a] Gheorghita Ghinea, Christian Timmerer, Weisi Lin, and Stephen Gulliver. Introduction to special issue on multiple sensorial (MulSeMedia) multimodal media: Advances and applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):9:1–9:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ghinea:2014:MSA

[GTLG14b] Gheorghita Ghinea, Christian Timmerer, Weisi Lin, and Stephen R. Gulliver. Mulsemedia: State of the art, perspectives, and challenges. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):17:1– 17:??, September 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Golmaryami:2022:SSS

[GTP^{+}22] Marjan Golmaryami, Rahim Taheri, Zahra Pooranian, Mohammad Shojafar, and Pei Xiao. SETTI: a Self-supervised AdvErsarial Malware DeTection ArchiTecture in an IoT environment. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):122:1–122:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3536425.

Gelli:2020:LVE

[GUH^{+}20] Francesco Gelli, Tiberio Uricchio, Xiangnan He, Alberto Del Bimbo, and Tat-Seng Chua. Learning visual elements of images for discovery of brand posts. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):56:1–56:21, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3385413.

Guan:2014:TAV

[GWM^{+}14] Genliang Guan, Zhiyong Wang, Shaohui Mei, Max Ott, Mingyi He, and David Dagan Feng. A top-down approach for video summarization. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1):4:1–4:??, August 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guo:2025:AHM

[GWWC25] Fubin Guo, Qi Wang, Qingshan Wang, and Sheng Chen. Accurate hand modeling in whole-body mesh reconstruction using joint-level features and kinematic-aware topology. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (7):211:1–211:23, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Griwodz:2025:ISI

[GWZ25] Carsten Griwodz, Mea Wang, and Roger Zimmermann. Introduction to the special issue on MMSys 2023 and NOSSDAV 2023. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):244:1–244:4, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gati:2021:DPT

[GYF^{+}21] Nicholaus J. Gati, Laurence T.

Yang, Jun Feng, Yijun Mo, and Mamoun Alazab. Differentially private tensor train deep computation for Internet of Multimedia Things. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):95:1–95:20, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3421276.

Guan:2012:EMM

[GYN12] Wei Guan, Suya You, and Ulrich Newmann. Efficient matchings and mobile augmented reality. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):47:1–47:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gan:2025:GAN

[GYY^{+}25] Yan Gan, Chenxue Yang, Mao Ye, Renjie Huang, and Deqiang Ouyang. Generative adversarial networks with learnable auxiliary module for image synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):111:1–111:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gao:2020:ISI

[GZ20] Honghao Gao and Yudong Zhang. Introduction to the special issue on smart communications and networking for future video surveillance. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2s):58:1–58:2, July 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3398382.

Guo:2017:EMD

[GZH17] Jianting Guo, Peijia Zheng, and Jiwu Huang. An efficient motion detection and tracking scheme for encrypted surveillance videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):61:1–61:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gao:2012:RHC

[GZHD12] Xifeng Gao, Caiming Zhang, Yan Huang, and Zhigang Deng. A robust high-capacity affine-transformation-invariant scheme for watermarking 3D geometric models. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2S):34:1–34:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guo:2020:RVT

[GZL^{+}20] Changyong Guo, Zhaoxin Zhang, Jinjiang Li, Xuesong Jiang, Jun Zhang, and Lei Zhang. Robust visual tracking using kernel sparse coding on multiple covariance descriptors. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):20:1–20:22, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3360308.

Guo:2018:OEL

[GZLW18] Dan Guo, Wengang Zhou, Houqiang Li, and Meng Wang. Online early-late fusion based on adaptive HMM for sign language recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):8:1–8:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gao:2021:FVS

[GZT21] Wei Gao, Linjie Zhou, and Lvfang Tao. A fast view synthesis implementation method for light field applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):134:1– 134:20, November 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3459098.

Gong:2024:MMM

[GZW^{+}24a] Wenjuan Gong, Yue Zhang, Wei Wang, Peng Cheng, and Jordi Gonzàlez. Meta-MMFNet: Meta-learning-based multi-model fusion network for micro-expression recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):39:1–39:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3539576.

Gu:2024:RAF

[GZW^{+}24b] Xiaoling Gu, Junkai Zhu, Yongkang Wong, Zizhao Wu, Jun Yu, Jianping Fan, and Mohan Kankanhalli. Recurrent appearance flow for occlusion-free virtual try-on. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):239:1–239:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3659581.

Guo:2025:MMM

[GZW^{+}25] Xu Guo, Tong Zhang, Fuyun Wang, Xudong Wang, Xiaoya Zhang, Xin Liu, and Zhen Cui. MMHCL: Multi-modal hypergraph contrastive learning for recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):298:1–298:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Guo:2024:IDB

[GZZ^{+}24] Qiang Guo, Zhi Zhang, Mingliang Zhou, Hong Yue, Huayan Pu, and Jun Luo. Image defogging based on regional gradient constrained prior. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):64:1–64:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617834.

Hossain:2019:ADL

[HAAM19] M. Shamim Hossain, Syed Umar Amin, Mansour Alsulaiman, and Ghulam Muhammad. Applying deep learning for epilepsy seizure detection and brain mapping visualization. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1s):10:1–10:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3241056.

Haenselmann:2010:FSI

[Hae10] Thomas Haenselmann. Foreword to the special issue on multimedia sensor fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(4):24:1– 24:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hossain:2011:MAQ

[HAE11] M. Anwar Hossain, Pradeep K. Atrey, and Abdulmotaleb El Saddik. Modeling and assessing quality of information in multisensor multimedia monitoring systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(1):3:1–3:??, January 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hanjalic:2013:MRM

[Han13] Alan Hanjalic. Multimedia retrieval that matters. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):44:1– 44:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hlavacs:2008:HVP

[HB08] Helmut Hlavacs and Shelley Buchinger. Hierarchical video patching with optimal server bandwidth. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1):8:1–8:??, January 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Holloman:2022:DSS

[HC22] Amanda K. Holloman and Chris S. Crawford. Defining scents: a systematic literature review of olfactory-based computing systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):15:1–15:22, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3470975.

Hsu:2025:MAF

[HC25] Wei-Yen Hsu and Yu-Chieh Chen. Multi-attribute featureaware network for facial expression recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):154:1–154:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2024:GCL

[HCJ^{+}24] Wuliang Huang, Yiqiang Chen, Xinlong Jiang, Teng Zhang, and Qian Chen. GJFusion: a channel-level correlation construction method for multimodal physiological signal fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):60:1–60:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617503.

Ho:2013:IPC

[HCKL13] Edmond S. L. Ho, Jacky C. P. Chan, Taku Komura, and Howard Leung. Interactive partner control in close interactions for real-time applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3): 21:1–21:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hossain:2022:SSA

[HCM^{+}22] M. Shamim Hossain, Rita Cucchiara, Ghulam Muhammad, Diana P. Tobón, and Abdulmotaleb El Saddik. Special section on AI-empowered multimedia data analytics for smart healthcare. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):38:1–38:2, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505281.

Hodzic:2025:DFC

[HCM^{+}25] Kerim Hodžić, Mirsad Cosovic, Sasa Mrdovic, Jason J. Quinlan, and Darijo Raca. DashRe-Streamer: Framework for creation of impaired video clips under realistic network conditions. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):15:1–15:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3640016.

Huang:2012:TAM

[HCS12] Jiun-Long Huang, Shih-Chuan Chiu, and Man-Kwan Shan. Towards an automatic music arrangement framework using score reduction. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1):8:1–8:??, January 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

He:2007:CSW

[HCW^{+}07] Xiaofei He, Deng Cai, Ji-Rong Wen, Wei-Ying Ma, and Hong-Jiang Zhang. Clustering and searching WWW images using link and page layout analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(2): 10:1–10:??, May 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2014:PVR

[HCWM14] Qinghua Huang, Bisheng Chen, Jingdong Wang, and Tao Mei. Personalized video recommendation through graph propagation. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(4):32:1–32:??, June 2014. ISSN 1551-6857 (print), 1551-6865 (electronic).

Hung:2023:FCN

[HCY^{+}23] Heyu Huang, Runmin Cong, Lianhe Yang, Ling Du, Cong Wang, and Sam Kwong. Feedback chain network for hippocampus segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):133:1– 133:??, June 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3571744.

Han:2022:AMG

[HCZ^{+}22] Ning Han, Jingjing Chen, Hao Zhang, Huanwen Wang, and Hao Chen. Adversarial multigrained embedding network for cross-modal text-video retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):63:1–63:23, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3483381.

Hua:2025:LAR

[HCZ^{+}25] Xuecheng Hua, Ke Cheng, Gege Zhu, Hu Lu, Yuanquan Wang, and Shitong Wang. Local-aware residual attention vision transformer for visible-infrared person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):146:1–146:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hosny:2019:RCI

[HD19] Khalid M. Hosny and Mohamed M. Darwish. Resilient color image watermarking using accurate quaternion radial substituted Chebyshev moments. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):46:1–46:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3325193.

Hu:2024:STR

[HDH^{+}24] Yijie Hu, Bin Dong, Kaizhu Huang, Lei Ding, Wei Wang, Xiaowei Huang, and Qiu-Feng Wang. Scene text recognition via dual-path network with shape-driven attention alignment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):107:1–107:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3633517.

Huang:2018:UBA

[HDW^{+}18] Lei Huang, Bowen Ding, Aining Wang, Yuedong Xu, Yipeng Zhou, and Xiang Li. User behavior analysis and video popularity prediction on a largescale VoD system. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):67:1–67:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hu:2015:SCC

[HDZ^{+}15] Xiping Hu, Junqi Deng, Jidi Zhao, Wenyan Hu, Edith C.- H. Ngai, Renfei Wang, Johnny Shen, Min Liang, Xitong Li, Victor C. M. Leung, and Yu-Kwong Kwok. SAfeDJ: a crowd-cloud codesign approach to situation-aware music delivery for drivers. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):21:1–21:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hamam:2014:QEM

[HEA14] Abdelwahab Hamam, Abdulmotaleb El Saddik, and Jihad Alja’am. A quality of experience model for haptic virtual environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(3):28:1–28:??, April 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hui:2023:MGC

[HGGZ23] Shuaixiong Hui, Qiang Guo, Xiaoyu Geng, and Caiming Zhang. Multi-guidance CNNs for salient object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):117:1–117:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3570507.

Hussain:2024:SYA

[HGKS24] Walayat Hussain, Honghao Gao, Rafiul Karim, and Abdulmotaleb El Saddik. Seventeen years of the ACM Transactions on Multimedia Computing, Communications and Applications: a bibliometric overview. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):297:1–297:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3660347.

Hefeeda:2008:RDO

[HH08a] Mohamed Hefeeda and Cheng-Hsin Hsu. Rate-distortion optimized streaming of fine-grained scalable video sequences. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1):2:1–2:??, January 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hsu:2008:ACR

[HH08b] Cheng-Hsin Hsu and Mohamed Hefeeda. On the accuracy and complexity of rate-distortion models for fine-grained scalable video sequences. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2):15:1–15:??, May 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hsu:2011:FCL

[HH11a] Cheng-Hsin Hsu and Mohamed Hefeeda. A framework for cross-layer optimization of video streaming in wireless networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(1):5:1–5:??, January 2011.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hsu:2011:SMV

[HH11b] Cheng-Hsin Hsu and Mohamed Hefeeda. Statistical multiplexing of variable-bit-rate videos streamed to mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(2): 12:1–12:??, February 2011. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hsu:2011:USS

[HH11c] Cheng-Hsin Hsu and Mohamed Hefeeda. Using simulcast and scalable video coding to efficiently control channel switching delay in mobile TV broadcast networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(2):8:1–8:??, February 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hamza:2012:EEM

[HH12a] Ahmed Hamza and Mohamed Hefeeda. Energy-efficient multicasting of multiview 3D videos to mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):45:1–45:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hefeeda:2012:DET

[HH12b] Mohamed Hefeeda and Cheng-Hsin Hsu. Design and evaluation of a testbed for mobile TV networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1):3:1–3:??, January 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

He:2019:ICV

[HH19] Chen He and Haifeng Hu. Image captioning with visualsemantic double attention. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1):26:1–26:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3292058.

Hu:2025:DPC

[HH25] Qianjiang Hu and Wei Hu. Dynamic point cloud denoising via gradient fields. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):132:1–132:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hao:2022:DLL

[HHGW22] Shijie Hao, Xu Han, Yanrong Guo, and Meng Wang. Decoupled low-light image enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):92:1–92:19, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498341.

Hsu:2022:OIV

[HHH22] Chih-Fan Hsu, Tse-Hou Hung, and Cheng-Hsin Hsu. Optimizing immersive video coding configurations using deep learning: a case study on TMIV. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1):19:1–19:25, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3471191.

He:2022:RLD

[HHL^{+}22] Jun He, Richang Hong, Xueliang Liu, Mingliang Xu, and Qianru Sun. Revisiting local descriptor for improved few-shot classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):127:1–127:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3511917.

Hoi:2010:SSD

[hHLC10] Steven C. h. Hoi, Wei Liu, and Shih-Fu Chang. Semisupervised distance metric learning for collaborative image retrieval and clustering. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3): 18:1–18:??, August 2010. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hu:2019:VPI

[HHLY19] Chuan-Shen Hu, Yi-Tsung Hsieh, Hsiao-Wei Lin, and Mei-Chen Yeh. Virtual Portraitist: an intelligent tool for taking well-posed selfies. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):12:1– 12:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3288760.

Hsu:2023:RMS

[HJ23] Wei-Yen Hsu and Pei-Wen Jian. Recurrent multi-scale approximation-guided network for single image super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):194:1–194:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3592613.

Hao:2015:LEP

[HJMY15] Fei Hao, Mingjie Jiao, Geyong Min, and Laurence T. Yang. Launching an efficient participatory sensing campaign: a smart mobile device-based approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):18:1–18:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hu:2019:OCT

[HJWW19] Han Hu, Yichao Jin, Yonggang Wen, and Cedric Westphal. Orchestrating caching, transcoding and request routing for adaptive video streaming over ICN. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):24:1–24:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3289184.

Huang:2023:BPL

[HJZ^{+}23] Wenxin Huang, Xuemei Jia, Xian Zhong, Xiao Wang, Kui Jiang, and Zheng Wang. Beyond the parts: Learning coarse-to-fine adaptive alignment representation for person search. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):105:1–105:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3565886.

Hu:2014:SFV

[HKYW14] Yongtao Hu, Jan Kautz, Yizhou Yu, and Wenping Wang. Speaker-following video subtitles. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2):32:1–32:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hsu:2024:CDA

[HL24] Wei-Yen Hsu and Hsien-Wen Lin. Context-detail-aware united network for single image deraining. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):144:1–144:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3639407.

Huang:2025:DRG

[HLC^{+}25] Xin Huang, Dong Liang, Hongrui Cai, Yunfeng Bai, Juyong Zhang, Feng Tian, and Jinyuan Jia. Double reference guided interactive 2D and 3D caricature generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):9:1–9:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3655624.

Hou:2018:NHC

[HLD18] Xueshi Hou, Yao Lu, and Sujit Dey. Novel hybrid-cast approach to reduce bandwidth and latency for cloud-based virtual space. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):58:1–58:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Han:2022:STS

[HLM^{+}22] Qing Han, Huiting Liu, Weidong Min, Tiemei Huang, Deyu Lin, and Qi Wang. 3D skeleton and two streams approach to person re-identification using optimized region matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2s):129:1–129:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3538490.

Hou:2024:DLL

[HLM^{+}24] Sujuan Hou, Jiacheng Li, Weiqing Min, Qiang Hou, Yanna Zhao, Yuanjie Zheng, and Shuqiang Jiang. Deep learning for logo detection: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):72:1–72:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3611309.

Hui:2022:STC

[HLS^{+}22] Chen Hui, Shaohui Liu, Wuzhen Shi, Feng Jiang, and Debin Zhao. Spatio-temporal context based adaptive camcorder recording watermarking. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):144:1–144:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3503160.

Hou:2024:TLF

[HLTH24] Wenxuan Hou, Guangyao Li, Yapeng Tian, and Di Hu. Toward long form audio-visual video understanding. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):284:1– 284:??, September 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3672079.

He:2021:ACO

[HLW^{+}21] Zhaoliang He, Hongshan Li, Zhi Wang, Shutao Xia, and Wenwu Zhu. Adaptive compression for online computer vision: an edge reinforcement learning approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):118:1–118:23, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447878.

Han:2024:SHO

[HLW^{+}24] Kai Han, Yu Liu, Rukai Wei, Ke Zhou, Jinhui Xu, and Kun Long. Supervised hierarchical online hashing for crossmodal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):103:1–103:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632527.

Hu:2025:LRR

[HLW^{+}25] Xiaobo Hu, Youfang Lin, Jinwen Wang, Yue Liu, Shuo Wang, Hehe Fan, and Kai Lv. Learning robust representations via bidirectional transition for visual reinforcement learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):292:1–292:24, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2014:CSA

[HLX^{+}14] Junshi Huang, Si Liu, Junliang Xing, Tao Mei, and Shuicheng Yan. Circle & search: Attribute-aware shoe retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (1):3:1–3:??, August 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

He:2021:MFU

[HLY21] Xin He, Qiong Liu, and You Yang. Make full use of priors: Cross-view optimized filter for multi-view depth enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):127:1–127:19, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3408293.

Hou:2023:UUI

[HLY^{+}23] Guojia Hou, Yuxuan Li, Huan Yang, Kunqian Li, and Zhenkuan Pan. UID2021: an underwater image dataset for evaluation of no-reference quality assessment metrics. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (4):151:1–151:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3578584.

Huang:2020:NTF

[HLZ^{+}20] Chenxi Huang, Yisha Lan, Guokai Zhang, Gaowei Xu, Landu Jiang, Nianyin Zeng, Jenhong Tan, E. Y. K. Ng, Yongqiang Cheng, Ningzhi Han, Rongrong Ji, and Yonghong Peng. A new transfer function for volume visualization of aortic stent and its application to virtual endoscopy. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2s):65:1–65:14, July 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3373358.

Hu:2021:AEN

[HLZ^{+}21] Yutao Hu, Xuhui Liu, Baochang Zhang, Jungong Han, and Xianbin Cao. Alignment enhancement network for finegrained visual categorization. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1s):12:1–12:20, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3446208.

Hefeeda:2010:ASM

[HM10] Mohamed Hefeeda and Kianoosh Mokhtarian. Authentication schemes for multimedia streams: Quantitative analysis and comparison. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(1):6:1–6:??, February 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hellmann:2025:GGB

[HMB^{+}25] Fabio Hellmann, Silvan Mertes, Mohamed Benouis, Alexander Hustinx, Tzung-Chien Hsieh, Cristina Conati, Peter Krawitz, and Elisabeth André. GANonymization: a GAN-based face anonymization framework for preserving emotional expressions. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):6:1–6:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3641107.

Houle:2017:QEC

[HMOS17] Michael E. Houle, Xiguo Ma, Vincent Oria, and Jichao Sun. Query expansion for contentbased similarity search using local and global features. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):25:1– 25:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hama:2021:EUM

[HMUC21] Kenta Hama, Takashi Matsubara, Kuniaki Uehara, and Jianfei Cai. Exploring uncertainty measures for image-caption embedding-and-retrieval task. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (2):46:1–46:19, June 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3425663.

Hendrikx:2013:PCG

[HMVI13] Mark Hendrikx, Sebastiaan Meijer, Joeri Van Der Velden, and Alexandru Iosup. Procedural content generation for games: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1):1:1–1:??, February 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

He:2008:EED

[HNL08] Wenbo He, Klara Nahrstedt, and Xue Liu. End-to-end delay control of multimedia applications over multihop wireless links. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(2):16:1–16:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2013:ETM

[HNS13] Zixia Huang, Klara Nahrstedt, and Ralf Steinmetz. Evolution of temporal multimedia synchronization principles: a historical viewpoint. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):34:1–34:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hong:2019:ASS

[Hon19]  Richang Hong.  Advanced

stereo seam carving by considering occlusions on both sides. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):69:1–69:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3321513.

Houle:2013:API

[HOSS13] Michael E. Houle, Vincent Oria, Shin’ichi Satoh, and Jichao Sun. Annotation propagation in image databases using similarity graphs. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(1):7:1–7:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hussein:2017:VJF

[HP17] Fairouz Hussein and Massimo Piccardi. V-JAUNE: a framework for joint action recognition and video summarization. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (2):20:1–20:??, May 2017. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hong:2020:CSF

[HPH^{+}20] Xiaopeng Hong, Wei Peng, Mehrtash Harandi, Ziheng Zhou, Matti Pietikäinen, and Guoying Zhao. Characterizing subtle facial movements via Riemannian manifold. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1– 24, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3342227.

Huang:2023:TIS

[HPLL23] Cong Huang, Xiulian Peng, Dong Liu, and Yan Lu. Text image super-resolution guided by text structure and embedding priors. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):199:1–199:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3595924.

Huang:2020:RHR

[HPW20] Xin Huang, Yuxin Peng, and Zhang Wen. RCE-HIL: Recognizing cross-media entailment with heterogeneous interactive learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1):5:1–5:21, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3365003.

Huang:2025:DTD

[HPZ25] Yuanzhou Huang, Songwei Pei, and Rui Zeng. DQFormer: Transformer with decoupled query augmentations for endto-end multi-object tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):174:1–174:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hu:2019:CAA

[HQF^{+}19] Jun Hu, Shengsheng Qian, Quan Fang, Xueliang Liu, and Changsheng Xu. A^{2} CMHNE: Attention-aware collaborative multimodal heterogeneous network embedding. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):45:1–45:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3321506.

Huang:2020:MPA

[HQF^{+}20] Xiaowen Huang, Shengsheng Qian, Quan Fang, Jitao Sang, and Changsheng Xu. Metapath augmented sequential recommendation with contextual co-attention network. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):52:1– 52:24, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3382180.

Huang:2025:NFR

[HQF^{+}25] Lin Huang, Chuan Qin, Guorui Feng, Xiangyang Luo, and Xinpeng Zhang. New framework of robust image encryption. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):91:1–91:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3712601.

Huang:2025:AEM

[HQLW25] Jiannan Huang, Mengxue Qu, Longfei Li, and Yunchao Wei. AdGPT: Explore meaningful advertising with ChatGPT. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (4):133:1–133:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2024:RTA

[HRL^{+}24] Junjian Huang, Hao Ren, Shulin Liu, Yong Liu, Chuanlu Lv, Jiawen Lu, Changyong Xie, and Hong Lu. Realtime attentive dilated U-net for extremely dark image enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):231:1–231:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3654668.

Huang:2023:FTF

[HSG23] Zhijie Huang, Jun Sun, and Xiaopeng Guo. FastCNN: Towards fast and accurate spatiotemporal network for HEVC compressed video enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):111:1–111:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3569583.

Hu:2020:ATL

[HSL^{+}20] Xinjue Hu, Jingming Shan, Yu Liu, Lin Zhang, and Shervin Shirmohammadi. An adaptive two-layer light field compression scheme using GNN-based reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):72:1–72:23, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3395620.

Huang:2022:OTV

[HSL^{+}22] Jiawei Huang, Qichen Su, Weihe Li, Zhuoran Liu, Tao Zhang, Sen Liu, Ping Zhong, Wanchun Jiang, and Jianxin Wang. Opportunistic transmission for video streaming over wild Internet. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):140:1–140:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3488722.

Hoque:2014:SEM

[HSN^{+}14] Mohammad Asharful Hoque, Matti Siekkinen, Jukka K. Nurminen, Sasu Tarkoma, and Mika Aalto. Saving energy in mobile devices for on-demand multimedia streaming — a cross-layer approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(3):25:1– 25:??, April 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hu:2018:UER

[HSR18] Wei Hu, Mozhdeh Seifi, and Erik Reinhard. Over- and under-exposure reconstruction of a single plenoptic capture. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (2):52:1–52:??, May 2018. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hao:2023:FGT

[HSR^{+}23] Jiachang Hao, Haifeng Sun, Pengfei Ren, Yiming Zhong, Jingyu Wang, Qi Qi, and Jianxin Liao. Fine-grained text-to-video temporal grounding from coarse boundary. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):157:1–157:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579825.

Huang:2025:CEC

[HSS^{+}25] I-Chun Huang, Yuang Shi, Yuan-Chun Sun, Wei Tsang Ooi, Chun-Ying Huang, and Cheng-Hsin Hsu. Composing error concealment pipelines for dynamic 3D point cloud streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):157:1–157:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2022:IBP

[HSX22] Xiaowen Huang, Jitao Sang, and Changsheng Xu. Imagebased personality questionnaire design. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):91:1–91:??, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3503489.

Huang:2018:MSH

[HSZ^{+}18] Min Huang, Song-Zhi Su, Hong-Bo Zhang, Guo-Rong Cai, Dongying Gong, Donglin Cao, and Shao-Zi Li. Multifeature selection for 3D human action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2):45:1–45:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hung:2015:ISI

[HT15] Hayley Hung and George Toderici. Introduction to: Special issue on extended best papers from ACM Multimedia 2014. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):24:1–24:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hong:2011:BSE

[HTT^{+}11] Richang Hong, Jinhui Tang, Hung-Khoon Tan, Chong-Wah Ngo, Shuicheng Yan, and Tat-Seng Chua. Beyond search: Event-driven summarization for Web videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(4):35:1–35:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hua:2013:OVD

[Hua13] Kien A. Hua. Online video delivery: Past, present, and future. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):39:1–39:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2023:TMM

[Hua23] Xin Huang. On teaching mode of MTI translation workshop based on IPT corpus for Tibetan areas of China. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):31:1– 31:??, February 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3527173.

Ho:2020:SGD

[HVC^{+}20] Trang-Thi Ho, John Jethro Virtusio, Yung-Yao Chen, Chih-Ming Hsu, and Kai-Lung Hua. Sketch-guided deep portrait generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):88:1–88:18, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3396237.

Heck:2007:VV

[HWG07] Rachel Heck, Michael Wallick, and Michael Gleicher. Virtual videography. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(1):??, February 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2018:EHD

[HWHL18] Shao Huang, Weiqiang Wang, Shengfeng He, and Rynson W. H. Lau. Egocentric hand detection via dynamic region growing. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):10:1–10:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hsu:2019:LMC

[HWLC19] Chih-Fan Hsu, Yu-Shuen Wang, Chin-Laung Lei, and Kuan-Ta Chen. Look at me! Correcting eye gaze in live video communication. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):38:1–38:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3311784.

Huang:2023:LEF

[HWW^{+}23] Baojin Huang, Zhongyuan Wang, Guangcheng Wang, Zhen Han, and Kui Jiang. Local eyebrow feature attention network for masked face recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):114:1–114:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3569943.

Hu:2025:DNI

[HWW25] Cong Hu, Xiao-Zhong Wei, and Xiao-Jun Wu. DIRformer: a novel image restoration approach based on U-shaped transformer and diffusion models. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):57:1–57:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3703632.

Huang:2020:ABM

[HWWL20] Feiran Huang, Kaimin Wei, Jian Weng, and Zhoujun Li. Attention-based modalitygated networks for image-text sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):79:1– 79:19, September 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3388861.

Hong:2011:VAE

[HWY^{+}11] Richang Hong, Meng Wang, Xiao-Tong Yuan, Mengdi Xu, Jianguo Jiang, Shuicheng Yan, and Tat-Seng Chua. Video accessibility enhancement for hearing-impaired users. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):24:1– 24:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hao:2024:HHC

[HWZL24] Dongze Hao, Qunbo Wang, Xinxin Zhu, and Jing Liu. HCCL: Hierarchical counterfactual contrastive learning for robust visual question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):301:1–301:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673902.

Hu:2024:OBS

[HXH^{+}24] Hengtong Hu, Lingxi Xie, Xinyue Huo, Richang Hong, and Qi Tian. One-bit supervision for image classification: Problem, solution, and beyond. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):113:1–113:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3633779.

Huang:2024:SLC

[HXL24] Hui Huang, Di Xiao, and Jia Liang. Secure low-complexity compressive sensing with preconditioning prior regularization reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):116:1– 116:??, April 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635308.

Hu:2020:ACA

[HXZ^{+}20] Shenghong Hu, Min Xu, Haimin Zhang, Chunxia Xiao, and Chao Gui. Affective content-aware adaptation scheme on QoE optimization of adaptive streaming over HTTP. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–18, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3328997.

Huang:2021:KDE

[HYG^{+}21] Yi Huang, Xiaoshan Yang, Junyu Gao, Jitao Sang, and Changsheng Xu. Knowledgedriven egocentric multimodal activity recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):133:1– 133:133, January 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3409332.

He:2025:DSB

[HYG^{+}25] Peng He, Jun Yu, Chengjie Ge, Ye Yu, Wei Xu, Lei Wang, Tianyu Liu, and Zhen Kan. Domain-separated bottleneck attention fusion framework for multimodal emotion recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):110:1–110:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

He:2020:STS

[HYLD20] Jiale He, Gaobo Yang, Xin Liu, and Xiangling Ding. Spatiotemporal saliency-based motion vector refinement for frame rate up-conversion. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):55:1–55:18, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3382506.

Hu:2023:DIP

[HYLG23] Yuzhang Hu, Wenhan Yang, Jiaying Liu, and Zongming Guo. Deep inter prediction with error-corrected auto-regressive network for video coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):33:1– 33:??, February 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3528173.

Hou:2020:SDE

[HYSL20] Yuxin Hou, Hongxun Yao, Xiaoshuai Sun, and Haoran Li. Soul Dancer: Emotionbased human action generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–19, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3340463.

Hu:2025:GAR

[HYT^{+}25] Wenmiao Hu, Yifang Yin, Ying Kiat Tan, An Tran, Hannes Kruppa, and Roger Zimmermann. GAN-assisted road segmentation from satellite imagery. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):5:1–5:??, January 2025. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635153.

Han:2025:EBF

[HYW^{+}25] Xiaoxuan Han, Songlin Yang, Wei Wang, Ziwen He, and Jing Dong. Exploiting backdoors of face synthesis detection with natural triggers. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):51:1–51:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3677380.

Hu:2016:ADA

[HZC^{+}16] Yao Hu, Chen Zhao, Deng Cai, Xiaofei He, and Xuelong Li. Atom decomposition with adaptive basis selection strategy for matrix completion. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (3):43:1–43:??, June 2016. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Han:2022:HIR

[HZC22] Xian-Hua Han, Yinqiang Zheng, and Yen-Wei Chen. Hyperspectral image reconstruction using multi-scale fusion learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1):16:1–16:21, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3477396.

He:2024:FMW

[HZH24] Qiaolin He, Zhijie Zheng, and Haifeng Hu. A feature map is worth a video frame: Rethinking convolutional features for visible-infrared person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):59:1–59:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617375.

Hu:2016:SND

[HZL^{+}16] Xianjun Hu, Weiming Zhang, Ke Li, Honggang Hu, and Nenghai Yu. Secure nonlocal denoising in outsourced images. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (3):40:1–40:??, June 2016. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hu:2021:MLM

[HZL^{+}21] Hezhen Hu, Wengang Zhou, Xingze Li, Ning Yan, and Houqiang Li. MV2Flow: Learning motion representation for fast compressed video action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):102:1–102:19, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3422360.

Huang:2025:AAV

[HZL^{+}25] Bowen Huang, Yanwei Zheng, Chuanlin Lan, Dongchen Sui, Xinpeng Zhao, Xiao Zhang, Mengbai Xiao, and Dongxiao Yu. Action-aware visualtextual alignment for longinstruction vision-and-language navigation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):270:1–270:22, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Hu:2021:GLE

[HZPL21] Hezhen Hu, Wengang Zhou, Junfu Pu, and Houqiang Li. Global-local enhancement network for NMF-aware sign language recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):80:1–80:19, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436754.

Han:2024:BNL

[HZS^{+}24] Ning Han, Yawen Zeng, Chuhao Shi, Guangyi Xiao, Hao Chen, and Jingjing Chen. BiC-Net: Learning efficient spatio-temporal relation for text-video retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):86:1– 86:??, March 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3627103.

Han:2020:HRR

[HZSC20] Xian-Hua Han, Yinqiang Zheng, Jiande Sun, and Yen-Wei Chen. Hyperspectral reconstruction with redundant camera spectral sensitivity functions. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):57:1–57:15, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386313.

Hu:2011:RAI

[HZW^{+}11] Weiming Hu, Haiqiang Zuo, Ou Wu, Yunfei Chen, Zhongfei Zhang, and David Suter. Recognition of adult images, videos, and web page bags. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S (1):28:1–28:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2022:SFM

[HZW22] Wei Huang, Yuze Zhang, and Shaohua Wan. A sorting fuzzy min-max model in an embedded system for atrial fibrillation detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):126:1–126:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3554737.

Huang:2024:UOI

[HZZ^{+}24] Jiafeng Huang, Tianjun Zhang, Shengjie Zhao, Lin Zhang, and Yicong Zhou. An underwater organism image dataset and a lightweight module designed for object detection networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):147:1–147:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3640465.

Iqbal:2024:DAB

[IAJ^{+}24] Farkhund Iqbal, Ahmed Abbasi, Abdul Rehman Javed, Ahmad Almadhor, Zunera Jalil, Sajid Anwar, and Imad Rida. Data augmentationbased novel deep learning method for deepfaked images detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):339:1–339:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3592615.

Ivanov:2010:RTH

[IB10] Yuri V. Ivanov and C. J. Bleakley. Real-time H.264 video encoding in software with fast mode decision and dynamic complexity control. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(1):5:1–5:??, February 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ignat:2022:WDI

[ICZ^{+}22] Oana Ignat, Santiago Castro, Yuhang Zhou, Jiajun Bao, Dandan Shan, and Rada Mihalcea. When did it happen? Duration-informed temporal localization of narrated actions in vlogs. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):142:1–142:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3495211.

Idris:2025:SIT

[IKTD25] Azeez Idris, Mohammed Khaleel, Wallapak Tavanapong, and Piet C. De Groen. Synthesized image training techniques: On improving model performance using confusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):4:1–4:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3641856.

Ip:2008:RRS

[ILL08] Alan T. S. Ip, John C. S. Lui, and Jiangchuan Liu. A revenuerewarding scheme of providing incentive for cooperative proxy caching for media streaming systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1):5:1–5:??, January 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ionescu:2025:ISI

[IPMD25] Bogdan Ionescu, Ioannis Patras, Henning Müller, and Alberto Del Bimbo. Introduction to the special issue on realistic synthetic data: Generation, learning, evaluation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):1:1–1:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3703593.

Illahi:2020:CGF

[IVS^{+}20] Gazi Karam Illahi, Thomas Van Gemert, Matti Siekkinen, Enrico Masala, Antti Oulasvirta, and Antti Ylä-Jääski. Cloud gaming with foveated video encoding. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1):7:1– 7:24, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3369110.

Jie:2008:VGD

[JC08] Li Jie and James J. Clark.

Video game design using an eye-movement-dependent model of visual attention. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3):22:1– 22:??, August 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jin:2010:DMN

[JC10] Xing Jin and S.-H. Gary Chan.

Detecting malicious nodes in peer-to-peer streaming by peerbased monitoring. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(2):9:1–9:??, March 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jiao:2025:UTI

[JCB25] Yifan Jiao, Chenglong Cai, and Bing-Kun Bao. Unified text-image space alignment with cross-modal prompting in CLIP for UDA. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):86:1–86:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3715699.

Jiang:2010:AVA

[JCC^{+}10] Wei Jiang, Courtenay Cotton, Shih-Fu Chang, Dan Ellis, and Alexander C. Loui. Audiovisual atoms for generic video concept classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):14:1– 14:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jiao:2022:GGL

[JCF^{+}22] Yingying Jiao, Haipeng Chen, Runyang Feng, Haoming Chen, Sifan Wu, Yifang Yin, and Zhenguang Liu. GLPose: Global-local representation learning for human pose estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2s):128:1–128:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3519305.

Jiang:2019:DPR

[JCSL19] Shuqiang Jiang, Gongwei Chen, Xinhang Song, and Linhu Liu. Deep patch representations with shared codebook for scene classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):5:1– 5:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3231738.

Jaiswal:2023:CNN

[JD23] Rahul Kumar Jaiswal and Rajesh Kumar Dubey. CAQoE: a novel no-reference contextaware speech quality prediction metric. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):35:1–35:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529394.

Jiang:2020:SDM

[JGJ^{+}20] Yizhang Jiang, Xiaoqing Gu, Dingcheng Ji, Pengjiang Qian, Jing Xue, Yuanpeng Zhang, Jiaqi Zhu, Kaijian Xia, and Shitong Wang. Smart diagnosis: a multiple-source transfer TSK fuzzy system for EEG seizure identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2s):59:1–59:21, July 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3340240.

Jia:2024:HBS

[JGY^{+}24] Jun Jia, Zhongpai Gao, Yiwei Yang, Wei Sun, Dandan Zhu, Xiaohong Liu, Xiongkuo Min, and Guangtao Zhai. Hidden barcode in sub-images with invisible locating marker. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):302:1– 302:??, October 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3674976.

Ji:2011:MFL

[JGZ^{+}11] Rongrong Ji, Yue Gao, Bineng Zhong, Hongxun Yao, and Qi Tian. Mining flickr landmarks by modeling reconstruction sparsity. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):31:1–31:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jonker:2024:DPE

[JJML24] Simon Jonker, Malthe Jelstrup, Weizhi Meng, and Brooke Lampe. Detecting post editing of multimedia images using transfer learning and fine tuning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(6):154:1–154:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3633284.

John:2025:MCF

[JK25] Vijay John and Yasutomo Kawanishi. Multimodal cascaded framework with multimodal latent loss functions robust to missing modalities. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):255:1–255:21, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jung:2008:SSL

[JKKL08] Dawoon Jung, Jaegeuk Kim, Jin-Soo Kim, and Joonwon Lee. ScaleFFS: a scalable log-structured flash file system for mobile multimedia systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1): 9:1–9:??, October 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jabeen:2023:RMA

[JLA^{+}23] Summaira Jabeen, Xi Li, Muhammad Shoib Amin, Omar Bourahla, Songyuan Li, and Abdul Jabbar. A review on methods and applications in multimodal deep learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):76:1–76:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3545572.

Jin:2024:USR

[JLK^{+}24] Xuanyu Jin, Ni Li, Wanzeng Kong, Jiajia Tang, and Bing Yang. Unbiased semantic representation learning based on causal disentanglement for domain generalization. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):249:1– 249:??, August 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3659953.

Jia:2021:RGL

[JLL^{+}21] Wei Jia, Li Li, Zhu Li, Xiang Zhang, and Shan Liu. Residual-guided in-loop filter using convolution neural network. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):139:1–139:19, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460820.

Jin:2022:AAA

[JLL^{+}22] Xin Jin, Xinning Li, Hao Lou, Chenyu Fan, Qiang Deng, Chaoen Xiao, Shuai Cui, and Amit Kumar Singh. Aesthetic attribute assessment of images numerically on mixed multi-attribute datasets. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):138:1– 138:??, October 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3547144.

Jia:2024:ENC

[JLL24] Zhaoyang Jia, Yan Lu, and Houqiang Li. Exploring neighbor correspondence matching for multiple-hypotheses video frame synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):111:1–111:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3633780.

Ji:2025:BFM

[JLL^{+}25] Wei Ji, Li Li, Zheqi Lv, Wenqiao Zhang, Mengze Li, Zhen Wan, Wenqiang Lei, and Roger Zimmermann. Backpropagationfree multi-modal on-device model adaptation via clouddevice collaboration. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):69:1– 69:??, February 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3706422.

Jin:2025:NIC

[JLLL25] Yili Jin, Jiahao Li, Bin Li, and Yan Lu. Neural image compression with regional decoding. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):73:1–73:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3708347.

Jiang:2018:DMP

[JLW^{+}18] Yu-Gang Jiang, Minjun Li, Xi Wang, Wei Liu, and Xian-Sheng Hua. DeepProduct: Mobile product search with portable deep features. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2):50:1– 50:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ji:2021:MPG

[JLZ^{+}21] Ruyi Ji, Zeyu Liu, Libo Zhang, Jianwei Liu, Xin Zuo, Yanjun Wu, Chen Zhao, Haofeng Wang, and Lin Yang. Multi-peak graph-based multiinstance learning for weakly supervised object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):70:1– 70:21, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3432861.

Jiang:2020:FSF

[JMLL20] Shuqiang Jiang, Weiqing Min, Yongqiang Lyu, and Linhu Liu. Few-shot food recognition via multi-view representation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):87:1–87:20, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3391624.

Jonna:2023:DDK

[JMS23] Sankaraganesh Jonna, Moushumi Medhi, and Rajiv Ranjan Sahay. Distill-DBDGAN: Knowledge distillation and adversarial learning framework for defocus blur detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):87:1–87:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3557897.

Jin:2011:KDH

[JP11] Yohan Jin and Balakrishnan Prabhakaran. Knowledge discovery from 3D human motion streams through semantic dimensional reduction. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(2):9:1–9:??, February 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jain:2005:GEI

[JPS05] Ramesh Jain, Thomas Plagemann, and Ralf Steinmetz. Guest editorial: The International ACM Multimedia Conference 1993 — ten years after. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1): 14–15, February 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jia:2016:WGB

[JSEI16] Adele Lu Jia, Siqi Shen, Dick H. J. Epema, and Alexandru Iosup. When game becomes life: The creators and spectators of online game replays and live streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):47:1–47:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jung:2007:NBA

[JSL07] Byunghee Jung, Junehwa Song, and Yoonjoon Lee. A narrativebased abstraction framework for story-oriented video. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(2):11:1– 11:??, May 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jiang:2016:CVI

[JTZ^{+}16] Yijing Jiang, Shanyu Tang, Liping Zhang, Muzhou Xiong, and Yau Jim Yip. Covert voice over Internet protocol communications with packet loss based on fractal interpolation. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (4):54:1–54:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ji:2021:MIM

[JW21] Wanting Ji and Ruili Wang. A multi-instance multi-label dual learning approach for video captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):72:1–72:18, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446792.

Jiang:2018:DBC

[JWF18] Shuhui Jiang, Yue Wu, and Yun Fu. Deep bidirectional crosstriplet embedding for online clothing shopping. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):5:1–5:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jiang:2021:BDC

[JWH21] Weitao Jiang, Weixuan Wang, and Haifeng Hu. Bi-directional co-attention network for image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):125:1–125:20, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460474.

Joshi:2006:SPE

[JWL06] Dhiraj Joshi, James Z. Wang, and Jia Li. The Story Picturing Engine—a system for automatic text illustration. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(1):68–89, February 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jin:2024:CBN

[JWW^{+}24] Yiting Jin, Jie Wu, Wanliang Wang, Yidong Yan, Jiawei Jiang, and Jianwei Zheng. Cascading blend network for image inpainting. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):14:1–14:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3608952.

Ji:2025:ITS

[JXL^{+}25] Hefeng Ji, Jing Xiao, Jiefan Lin, Jimin Liu, and Haoyong Yu. Intelligent tumor synthesis based on medical image knowledge for liver tumor segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):264:1–264:23, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jing:2025:INT

[JXLD25] Taotao Jing, Haifeng Xia, Hongfu Liu, and Zhengming Ding. Interpretable novel target discovery through open-set domain adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):147:1–147:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Jin:2021:MTL

[JXTC21] Xin Jin, Jianfeng Xu, Kazuyuki Tasaka, and Zhibo Chen. Multi-task learning-based allin-one collaboration framework for degraded image superresolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):21:1–21:21, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3417333.

Jiang:2025:MLC

[JYZ^{+}25] Wei Jiang, Jiayu Yang, Yongqi Zhai, Feng Gao, and Ronggang Wang. MLIC++: Linear complexity multi-reference entropy modeling for learned image compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):142:1–142:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ji:2012:AQS

[JYZC12] Rongrong Ji, Felix X. Yu, Tongtao Zhang, and Shih-Fu Chang. Active query sensing: Suggesting the best query view for mobile visual search. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):40:1–40:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kua:2020:ACA

[KABB20] Jonathan Kua, Grenville Armitage, Philip Branch, and Jason But. Adaptive chunklets and AQM for higherperformance content streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (4):1–24, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3344381.

Kankanhalli:2012:ISI

[Kan12] Mohan S. Kankanhalli. Introduction to special issue on multimedia security. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2S):31:1–31:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kieu:2021:BLD

[KBB21] My Kieu, Andrew D. Bagdanov, and Marco Bertini. Bottom-up and layerwise domain adaptation for pedestrian detection in thermal images. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1):32:1–32:19, April 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3418213.

Kumar:2023:AFS

[KBI^{+}23] Puneet Kumar, Gaurav Bhatt, Omkar Ingle, Daksh Goyal, and Balasubramanian Raman. Affective feedback synthesis towards multimodal text and image data. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):190:1–190:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589186.

Kleinrouweler:2017:SAP

[KCC17] Jan Willem Kleinrouweler, Sergio Cabrero, and Pablo Cesar. An SDN architecture for privacy-friendly networkassisted DASH. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):44:1–44:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kong:2018:EVE

[KD18] Lingchao Kong and Rui Dai.

Efficient video encoding for automatic video analysis in distributed wireless surveillance systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):72:1–72:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kolan:2008:NLV

[KDC08] Prakash Kolan, Ram Dantu, and João W. Cangussu. Nuisance level of a voice call.

ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1): 6:1–6:??, October 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kizilkaya:2022:EFF

[KEYY22] Burak Kizilkaya, Enver Ever, Hakan Yekta Yatbaz, and Adnan Yazici. An effective forest fire detection framework using heterogeneous wireless multimedia sensor networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):47:1– 47:21, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473037.

Kaddar:2024:DDU

[KFA^{+}24] Bachir Kaddar, Sid Ahmed Fezza, Zahid Akhtar, Wassim Hamidouche, Abdenour Hadid, and Joan Serra-Sagristá. Deepfake detection using spatiotemporal transformer. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):345:1–345:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3643030.

Kimura:2025:EAV

[KFP^{+}25] Bruno Kimura, Simone Ferlin, Thomas Paiva, Toktam Mahmoodi, Anna Brunstrom, and Ozgu Alay. Evaluating adaptive video streaming over multipath QUIC with shared bottleneck detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):246:1–246:25, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Khodabakhshi:2013:SSF

[KH13] Naghmeh Khodabakhshi and Mohamed Hefeeda. Spider: a system for finding 3D video copies. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1):7:1–7:??, February 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Khaleel:2024:VVC

[KIT^{+}24] Mohammed Khaleel, Azeez Idris, Wallapak Tavanapong, Jacob R. Pratt, Junghwan Oh, and Piet C. de Groen. Vis-Active: Visual-concept-based active learning for image classification under class imbalance. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):84:1–84:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3617999.

Krishnan:2021:SEQ

[KJJ^{+}21] Prabhakar Krishnan, Kurunandan Jain, Pramod George Jose, Krishnashree Achuthan, and Rajkumar Buyya. SDN enabled QoE and security framework for multimedia applications in 5G networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):39:1– 39:29, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3377390.

Komogortsev:2008:PRT

[KK08] Oleg V. Komogortsev and Javed I. Khan. Predictive real-time perceptual compression based on eye-gaze-position analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3):23:1–23:??, August 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kiess:2018:SCA

[KKGE18] Johannes Kiess, Stephan Kopf, Benjamin Guthier, and Wolfgang Effelsberg. A survey on content-aware image and video retargeting. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):76:1–76:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Koch:2018:CYU

[KLS^{+}18] Christian Koch, Moritz Lode, Denny Stohr, Amr Rizk, and Ralf Steinmetz. Collaborations on YouTube: From unsupervised detection to the impact on video and channel popularity. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (4):89:1–89:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kang:2024:OCM

[KLX^{+}24] Xiao Kang, Xingbo Liu, Wen Xue, Xiushan Nie, and Yilong Yin. Online cross-modal hashing with dynamic prototype. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (8):252:1–252:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3665249.

Kumar:2021:DDE

[KMK^{+}21] Ambeshwar Kumar, Ramachandran Manikandan, Utku Kose, Deepak Gupta, and Suresh C. Satapathy. Doctor’s dilemma: Evaluating an explainable subtractive spatial lightweight convolutional neural network for brain tumor diagnosis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):105:1– 105:26, October 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3457187.

Kum:2005:RTM

[KMP05] Sang-Uok Kum and Ketan Mayer-Patel. Real-time multidepth stream compression.

ACM Transactions on Multimedia Computing, Communications, and Applications, 1(2): 128–150, May 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kirchhoffer:2018:PDV

[KMSW18] Heiner Kirchhoffer, Detlev Marpe, Heiko Schwarz, and Thomas Wiegand. Properties and design of variable-to-variable length codes. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):75:1– 75:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

K:2021:AML

[KN21a] Mythili K. and Manish Narwaria. Assessment of machine learning-based audiovisual quality predictors: Why uncertainty matters. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):45:1– 45:22, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3430376.

Kumar:2021:ESE

[KN21b] S. Sambath Kumar and M. Nandhini. Entropy slicing extraction and transfer learning classification for early diagnosis of Alzheimer diseases with sMRI. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (2):40:1–40:22, June 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3383749.

Korshunov:2011:VQF

[KO11] Pavel Korshunov and Wei Tsang Ooi. Video quality for face detection, recognition, and tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):14:1–14:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kotharu:2008:PQR

[KP08] Phani S. Kotharu and B. Prabhakaran. Partial query resolution for animation authoring. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1): 4:1–4:??, January 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kim:2015:ERD

[KP15] Yelin Kim and Emily Mower Provost. Emotion recognition during speech using dynamics of multiple regions of the face. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (1s):25:1–25:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kang:2024:DGN

[KPK^{+}24] Yan Kang, Bin Pu, Yongqi Kou, Yun Yang, Jianguo Chen, Khan Muhammad, Po Yang, Lida Xu, and Mohammad Hijji. A deep graph network with multiple similarity for user clustering in human-computer interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):46:1–46:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549954.

Karagkioules:2022:OLA

[KPL^{+}22] Theodoros Karagkioules, Georgios S. Paschos, Nikolaos Liakopoulos, Attilio Fiandrotti, Dimitrios Tsilimantos, and Marco Cagnazzo. Online learning for adaptive video streaming in mobile networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):2:1– 2:22, January 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3460819.

Katti:2014:OEE

[KRKK14] Harish Katti, Anoop Kolar Rajagopal, Mohan Kankanhalli, and Ramakrishnan Kalpathi. Online estimation of evolving human visual interest. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1):8:1– 8:??, August 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Knoche:2009:BPS

[KS09] H. Knoche and M. A. Sasse.

The big picture on small screens delivering acceptable video quality in mobile TV. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3): 20:1–20:??, August 2009. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Knees:2013:SMS

[KS13] Peter Knees and Markus Schedl. A survey of music similarity and recommendation from music context data. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (1):2:1–2:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kasyap:2021:PPD

[KT21] Harsh Kasyap and Somanath Tripathy. Privacy-preserving decentralized learning framework for healthcare system. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):68:1–68:24, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3426474.

Karafotias:2017:IER

[KTK^{+}17] Georgios Karafotias, Akiko Teranishi, Georgios Korres, Friederike Eyssel, Scandar Copti, and Mohamad Eid. Intensifying emotional reactions via tactile gestures in immersive films. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):29:1–29:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kahol:2006:MCH

[KTM^{+}06] Kanav Kahol, Priyamvada Tripathi, Troy Mcdaniel, Laura Bratton, and Sethuraman Panchanathan. Modeling context in haptic perception, rendering, and visualization. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(3):219–240, August 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Khan:2022:PPM

[KUH^{+}22] Abbas Khan, Ijaz Ul Haq, Tanveer Hussain, Khan Muhammad, Mohammad Hijji, Muhammad Sajjad, Victor Hugo C. De Albuquerque, and Sung Wook Baik. PMAL: a proxy model active learning approach for vision based industrial applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):123:1–123:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3534932.

Kawai:2024:RBH

[KVB^{+}24] Vinicius Sato Kawai, Lucas Pascotti Valem, Alexandro Baldassin, Edson Borin, Daniel Carlos Guimarães Pedronette, and Longin Jan Latecki. Rank-based hashing for effective and efficient nearest neighbor search for image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):329:1–329:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3659580.

Kuo:2011:TPQ

[KW11] Wen-Kuang Kuo and Kuo-Wei Wu. Traffic prediction and QoS transmission of real-time live VBR videos in WLANs. ACM Transactions on Multimedia Computing, Communications, and Applications, 7 (4):36:1–36:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kansal:2025:IPR

[KWK25] Kajal Kansal, Yongkang Wong, and Mohan Kankanhalli. Implications of privacy regulations on video surveillance systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):184:1–184:27, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kou:2025:PFF

[KWL^{+}25] Feifei Kou, Bingwei Wang, Haisheng Li, Chuangying Zhu, Lei Shi, Jiwei Zhang, and Limei Qi. Potential features fusion network for multimodal fake news detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):87:1–87:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3711866.

Kroupi:2014:ECP

[KYVE14] Eleni Kroupi, Ashkan Yazdani, Jean-Marc Vesin, and Touradj Ebrahimi. EEG correlates of pleasant and unpleasant odor perception. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):13:1–13:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Krishnappa:2015:CCV

[KZGH15] Dilip Kumar Krishnappa, Michael Zink, Carsten Griwodz, and Pål Halvorsen. Cache-centric video recommendation: an approach to improve the efficiency of YouTube caches. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):48:1–48:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Kompatsiaris:2013:ISS

[KZHC13] Ioannis (Yiannis) Kompatsiaris, Wenjun (Kevin) Zeng, Gang Hua, and Liangliang Cao. Introduction to the special section of best papers of ACM multimedia 2012. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):50:1–50:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lathey:2015:IEE

[LA15] Ankita Lathey and Pradeep K.

Atrey. Image enhancement in encrypted domain over cloud. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (3):38:1–38:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2015:QQG

[LB15] Yang Li and Azzedine Boukerche. QuGu: a quality guaranteed video dissemination protocol over urban vehicular ad hoc networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):55:1–55:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2008:FSE

[LBD08] Rui Li, Bir Bhanu, and Anlei Dong. Feature synthesized EM algorithm for image retrieval.

ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2): 10:1–10:??, May 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2012:NVT

[LC12] Shujie Liu and Chang Wen Chen. A novel 3D video transcoding scheme for adaptive 3D video transmission to heterogeneous terminals. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):43:1– 43:??, September 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2014:LAM

[LCC^{+}14a] Chih-Wei Lin, Kuan-Wen Chen, Shen-Chi Chen, Cheng-Wu Chen, and Yi-Ping Hung. Large-area, multilayered, and high-resolution visual monitoring using a dual-camera system. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (2):30:1–30:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2014:DUB

[LCC^{+}14b] Ning Liu, Huajie Cui, S.- H. Gary Chan, Zhipeng Chen, and Yirong Zhuang. Dissecting user behaviors for a simultaneous live and VoD IPTV system. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (3):23:1–23:??, April 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2025:ELD

[LCH^{+}25] Haojie Li, Hao Chen, Yining Huang, Tianshui Chen, and Shuangping Huang. Enhancing lip dynamic authenticity: Learning 3D temporal representations for talking head synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):295:1–295:21, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2009:PBR

[LCK09] Mingzhe Li, Mark Claypool, and Robert Kinicki. Playout buffer and rate optimization for streaming over IEEE 802.11 wireless networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3):26:1–26:??, August 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2014:UVS

[LCL14] Haitao Li, Xu Cheng, and Jiangchuan Liu. Understanding video sharing propagation in social networks: Measurement and analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(4):33:1–33:??, June 2014. ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2021:LSI

[LCL^{+}21] Yanchun Li, Jianglian Cao, Zhetao Li, Sangyoon Oh, and Nobuyoshi Komuro. Lightweight single image super-resolution with dense connection distillation network. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):9:1–9:17, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3414838.

Lv:2022:SCC

[LCL22] Zhihan Lv, Dongliang Chen, and Haibin Lv. Smart city construction and management by digital twins and BIM big data in COVID-19 scenario. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2s):117:1–117:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3529395.

Le:2023:SAV

[LCL23a] Thi-Ngoc-Hanh Le, Ya-Hsuan Chen, and Tong-Yee Lee. Structure-aware video style transfer with map art. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):131:1– 131:??, June 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3572030.

Li:2023:DLH

[LCL^{+}23b] Weixin Li, Tiantian Cao, Chang Liu, Xue Tian, Ya Li, Xiaojie Wang, and Xuan Dong. Dual-lens HDR using guided 3D exposure CNN and guided denoising transformer. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):158:1– 158:??, September 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3579167.

Li:2024:VDS

[LCL^{+}24a] Fan Li, Yanxiang Chen, Haiyang Liu, Zuxing Zhao, Yuanzhi Yao, and Xin Liao. Vocoder detection of spoofing speech based on GAN fingerprints and domain generalization. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(6):157:1–157:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3630751.

Liu:2024:TSM

[LCL^{+}24b] Yao Liu, Gangfeng Cui, Jiahui Luo, Xiaojun Chang, and Lina Yao. Two-stream multi-level dynamic point transformer for two-person interaction recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):145:1–145:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3639470.

Liu:2009:SSE

[LCS09] Xiaotao Liu, Mark Corner, and Prashant Shenoy. SEVA: Sensor-enhanced video annotation. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3):24:1–24:??, August 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2017:DSF

[LCS17] Xiaopeng Li, Ming Cheung, and James She. A distributed streaming framework for connection discovery using shared videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):59:1–59:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2011:SSC

[LCSX11] Yu-Ru Lin, K. Selçcuk Candan, Hari Sundaram, and Lexing Xie. SCENT: Scalable compressed monitoring of evolving multirelational social networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S (1):29:1–29:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2012:QOV

[LCT^{+}12] Yanwei Liu, Song Ci, Hui Tang, Yun Ye, and Jinxia Liu. QoEoriented 3D video transcoding for mobile streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):42:1– 42:??, September 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2016:MCR

[LCW16] Xuelong Li, Mulin Chen, and Qi Wang. Measuring collectiveness via refined topological similarity. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):34:1–34:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lei:2023:LUD

[LCY^{+}23] Fei Lei, Zhongqi Cao, Yuning Yang, Yibo Ding, and Cong Zhang. Learning the user’s deeper preferences for multimodal recommendation systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):138:1–138:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3573010.

Li:2024:DST

[LCZ^{+}24a] Yixuan Li, Peilin Chen, Hanwei Zhu, Keyan Ding, Leida Li, and Shiqi Wang. Deep shapetexture statistics for completely blind image quality evaluation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):382:1–382:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3694977.

Liu:2024:DDE

[LCZ^{+}24b] Caixia Liu, Yali Chen, Minhong Zhu, Chenhui Hao, Haisheng Li, and Xiaochuan Wang. DE-GAN: Detail-enhanced generative adversarial network for monocular depth-based 3D reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):365:1–365:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3690826.

Lv:2024:SVI

[LCZ^{+}24c] Kai Lv, Haobo Chen, Chuyang Zhao, Kai Tu, Junru Chen, Yadong Li, Boxun Li, and Youfang Lin. Style variable and irrelevant learning for generalizable person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):281:1–281:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3671003.

Li:2025:MMH

[LD25] Ying Li and Yuxiang Ding.

MoHGCN: Momentum hypergraph convolution network for cross-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):277:1– 277:21, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2024:GMM

[LDL24] Yefan Li, Fuqing Duan, and Ke Lu. Gated multi-modal edge refinement network for light field salient object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):300:1–300:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674836.

Liu:2025:BMC

[LDL^{+}25] Dongjun Liu, Weichen Dai, Honggang Liu, Hangjie Yi, and Wanzeng Kong. Brain-machine cross-modal alignment via sample relational learning for visual classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):182:1–182:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2025:TSG

[LDQ^{+}25] Runing Li, Jiangyan Dai, Qibing Qin, Chengduan Wang, Huihui Zhang, and Yugen Yi. Texture and structure-guided dual-attention mechanism for image inpainting. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):116:1–116:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2018:CES

[LDT^{+}18] Yusen Li, Yunhua Deng, Xueyan Tang, Wentong Cai, Xiaoguang Liu, and Gang Wang. Cost-efficient server provisioning for cloud gaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (3s):55:1–55:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Luo:2020:STS

[LDZ^{+}20] Guoliang Luo, Zhigang Deng, Xin Zhao, Xiaogang Jin, Wei Zeng, Wenqiang Xie, and Hyewon Seo. Spatio-temporal segmentation based adaptive compression of dynamic mesh sequences. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1):14:1–14:24, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3377475.

Liu:2024:MMR

[LDZ^{+}24] Huaijin Liu, Jixiang Du, Yong Zhang, Hongbo Zhang, and Jiandian Zeng. MSSA: Multirepresentation semantics-augmented set abstraction for 3D object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):328:1–328:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3686157.

Li:2022:UEU

[LFP^{+}22] Yehao Li, Jiahao Fan, Yingwei Pan, Ting Yao, Weiyao Lin, and Tao Mei. Uni-EDEN: Universal encoder-decoder network by multi-granular visionlanguage pre-training. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):48:1– 48:16, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473140.

Liao:2024:RFM

[LG24] Guibiao Liao and Wei Gao.

Rethinking feature mining for light field salient object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):307:1–307:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3676967.

Lu:2024:MDE

[LGC^{+}24] Junxin Lu, Yongbin Gao, Jieyu Chen, Jeng-Neng Hwang, Hamido Fujita, and Zhijun Fang. Monocular depth and ego-motion estimation with scale based on superpixel and normal constraints. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):303:1– 303:??, October 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3674977.

Luque:2014:IMS

[LGF^{+}14] Francisco Pedro Luque, Iris Galloso, Claudio Feijoo, Carlos Alberto Martín, and Guillermo Cisneros. Integration of multisensorial stimuli and multimodal interaction in a hybrid 3DTV system. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):16:1– 16:??, September 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2020:BIQ

[LGLZ20] Yutao Liu, Ke Gu, Xiu Li, and Yongbing Zhang. Blind image quality assessment by natural scene statistics and perceptual characteristics. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):91:1–91:91, September 2020. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3414837.

Li:2024:DPP

[LGM^{+}24] Zhenyu Li, Shanshan Gao, Deqian Mao, Shouwen Song, Lei Li, and Yuanfeng Zhou. Deep plug-and-play non-iterative cluster for 3D global feature extraction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):323:1–323:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3679204.

Latifzadeh:2024:EDA

[LGT^{+}24] Kayhan Latifzadeh, Nima Gozalpour, V. Javier Traver, Tuukka Ruotsalo, Aleksandra Kawala-Sterniuk, and Luis A. Leiva. Efficient decoding of affective states from videoelicited EEG signals: an empirical investigation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):317:1–317:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663669.

Liu:2025:MGC

[LGW^{+}25] Yajie Liu, Pu Ge, Guodong Wang, Qingjie Liu, and Di Huang. Multi-grained contrastive learning for textsupervised open-vocabulary semantic segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):81:1– 81:??, March 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3711868.

Luo:2008:IFH

[LGX^{+}08] Hangzai Luo, Yuli Gao, Xiangyang Xue, Jinye Peng, and Jianping Fan. Incorporating feature hierarchy and boosting to achieve more effective classifier training and conceptoriented video summarization and skimming. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1):1:1–1:??, January 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lei:2024:BDV

[LGX^{+}24] Si-Chao Lei, Yue-Jiao Gong, Xiao-Lin Xiao, Yi cong Zhou, and Jun Zhang. Boosting diversity in visual search with Pareto non-dominated reranking. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):79:1–79:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3625296.

Liu:2022:AQR

[LGY^{+}22] Yibing Liu, Yangyang Guo, Jianhua Yin, Xuemeng Song, Weifeng Liu, Liqiang Nie, and Min Zhang. Answer questions with right image regions: a visual attention regularization approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):93:1–93:18, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498340.

Lou:2012:QDD

[LH12] Xiaosong Lou and Kai Hwang.

Quality of data delivery in peer-to-peer video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 8s (1):12:1–12:??, February 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2020:EIH

[LH20] Shiguang Liu and Ziqing Huang. Efficient image hashing with geometric invariant vector distance for copy detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–22, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3355394.

Liu:2022:DSS

[LH22] Zhe Liu and Xian-Hua Han. Deep self-supervised hyperspectral image reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):149:1–149:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3510373.

Lu:2025:MAC

[LHC^{+}25] Xiusheng Lu, Yanbin Hao, Lechao Cheng, Sicheng Zhao, Yutao Liu, and Mingli Song. Mixed attention and channel shift transformer for efficient action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):93:1– 93:??, March 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3712594.

Lv:2014:MHF

[LHF^{+}14] Zhihan Lv, Alaa Halawani, Shengzhong Feng, Haibo Li, and Shafiq Ur Réhman. Multimodal hand and foot gesture interaction for handheld devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):10:1–10:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2024:PIP

[LHJ^{+}24] Qingzhi Liu, Yuchen Huang, Chenglu Jin, Xiaohan Zhou, Ying Mao, Cagatay Catal, and Long Cheng. Privacy and integrity protection for IoT multimodal data using machine learning and blockchain. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (6):153:1–153:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3638769.

Lin:2024:DDV

[LHL^{+}24] Kaihan Lin, Weihong Han, Shudong Li, Zhaoquan Gu, Huimin Zhao, and Yangyang Mei. Detecting deepfake videos using spatiotemporal Trident network. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):340:1–340:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3623639.

Liu:2021:MIC

[LHS^{+}21] Xiangbin Liu, Jiesheng He, Liping Song, Shuai Liu, and Gautam Srivastava. Medical image classification based on an adaptive size deep learning model. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):102:1–102:18, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3465220.

Liu:2024:BDB

[LHY^{+}24] Xingyu Liu, Zhongyun Hua, Shuang Yi, Yushu Zhang, and Yicong Zhou. Bi-directional block encoding for reversible data hiding over encrypted images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):149:1–149:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3638771.

Li:2023:SCE

[LHZ^{+}23] Jie Li, Ling Han, Chong Zhang, Qiyue Li, and Zhi Liu. Spherical convolution empowered viewport prediction in 360 video multicast with limited FoV feedback. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):3:1–3:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3511603.

Li:2023:DMO

[Li23] Yang Li. Detection of moving object using superpixel fusion network. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):160:1–160:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579998.

Lin:2015:DVS

[Lin15] Pei-Yu Lin. Double verification secret sharing mechanism based on adaptive pixel pair matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(3):36:1–36:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2025:MMR

[LJ25] Pengyu Li and Cheolkon Jung.

MRFGNet: Multiscale reference frame generation network for VVC inter-coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):261:1– 261:20, September 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Luo:2024:TES

[LJG^{+}24] Xiao Luo, Wei Ju, Yiyang Gu, Yifang Qin, Siyu Yi, Daqing Wu, Luchen Liu, and Ming Zhang. Toward effective semisupervised node classification with hybrid curriculum pseudolabeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):82:1–82:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3626528.

Li:2024:ASS

[LJL^{+}24] Chuanhao Li, Chenchen Jing, Zhen Li, Yuwei Wu, and Yunde Jia. Adversarial sample synthesis for visual question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):378:1–378:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3688848.

Lim:2008:DPP

[LJP08] Seung-Ho Lim, Yo-Won Jeong, and Kyu Ho Park. Data placement and prefetching with accurate bit rate control for interactive media server. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3):21:1– 21:??, August 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2025:CIP

[LJW^{+}25] Suyi Li, Chenyi Jiang, Shidong Wang, Yang Long, Zheng Zhang, and Haofeng Zhang. Contextual interaction via primitive-based adversarial training for compositional zeroshot learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):276:1–276:24, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2022:NCN

[LJZ^{+}22] Kai Lin, Chuanmin Jia, Xinfeng Zhang, Shanshe Wang, Siwei Ma, and Wen Gao. NR-CNN: Nested-residual guided CNN in-loop filtering for video coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):102:1–102:22, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502723.

Liu:2007:CAT

[LK07] Tiecheng Liu and John R.

Kender. Computational approaches to temporal sampling of video sequences. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(2):7:1–7:??, May 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2018:REN

[LK18] Xiaodan Lin and Xiangui Kang. Robust electric network frequency estimation with rank reduction and linear prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (4):84:1–84:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lee:2025:FSS

[LKL25a] Seongmin Lee, Jiwoo Kang, and Sanghoon Lee. 3D facial shape similarity with deep perceptual representations. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):183:1– 183:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2025:MAO

[LKL^{+}25b] Qianxing Li, Dehui Kong, Jinghua Li, Dongpan Chen, and Baocai Yin. Multianchor offset representation based coarse-to-fine diffusion model for human pose estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):122:1–122:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lisanti:2017:MKC

[LKM17] Giuseppe Lisanti, Svebor Karaman, and Iacopo Masi. Multichannelkernel canonical correlation analysis for cross-view person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(2):13:1–13:??, May 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lokoc:2019:ISS

[LKM^{+}19] Jakub Lokoc, Gregor Kovalcík, Bernd Münzer, Klaus Schöffmann, Werner Bailer, Ralph Gasser, Stefanos Vrochidis, Phuong Anh Nguyen, Sitapa Rujikietgumjorn, and Kai Uwe Barthel. Interactive search or sequential browsing? A detailed analysis of the Video Browser Showdown 2018. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):29:1– 29:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3295663.

Liu:2022:SRP

[LKW^{+}22] Caixia Liu, Dehui Kong, Shaofan Wang, Jinghua Li, and Baocai Yin. A spatial relationship preserving adversarial network for 3D reconstruction from a single depth view. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):110:1–110:22, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3506733.

Liu:2015:ECA

[LL15] Kaikai Liu and Xiaolin Li.

Enabling context-aware indoor augmented reality via Smartphone sensing and vision tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):15:1–15:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2023:MGF

[LL23] Jiazhi Liu and Feng Liu.

Modified 2D-ghost-free stereoscopic display with depth-of-field effects. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):47:1–47:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3534964.

Liu:2021:DVV

[LLA^{+}21] Yanwei Liu, Jinxia Liu, Antonios Argyriou, Siwei Ma, Liming Wang, and Zhen Xu. 360degree VR video watermarking based on spherical wavelet transform. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):38:1–38:23, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3425605.

Lin:2011:PCI

[LLC11] Pei-Yu Lin, Jung-San Lee, and Chin-Chen Chang. Protecting the content integrity of digital imagery with fidelity preservation. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):15:1–15:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2021:XEC

[LLC^{+}21] Yu-Sheng Lin, Zhe-Yu Liu, Yu-An Chen, Yu-Siang Wang, Ya-Liang Chang, and Winston H.

Hsu. xCos: an explainable cosine metric for face verification task. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):112:1–112:16, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469288.

Li:2024:MPC

[LLC^{+}24a] Yan Li, Xiangyuan Lan, Haifeng Chen, Ke Lu, and Dongmei Jiang. Multimodal PEAR chain-of-thought reasoning for multimodal sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):286:1–286:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672398.

Liu:2024:MFT

[LLC^{+}24b] Zhilei Liu, Xiaoxing Liu, Sen Chen, Jiaxing Liu, Longbiao Wang, and Chongke Bi. Multimodal fusion for talking face generation utilizing speechrelated facial action units. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):292:1–292:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672565.

Li:2025:SST

[LLC^{+}25] Jiaxi Li, Jingwei Liao, Bo Chen, Anh Nguyen, Aditi Tiwari, Qian Zhou, Zhisheng Yan, and Klara Nahrstedt. ST-360: Spatial-temporal filteringbased low-latency 360-degree video analytics framework. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):248:1–248:25, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2017:RHA

[LLCH17] Shih-Yao Lin, Yen-Yu Lin, Chu-Song Chen, and Yi-Ping Hung. Recognizing human actions with outlier frames by observation filtering and completion. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):28:1–28:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2023:TBV

[LLG^{+}23] Kun Li, Jiaxiu Li, Dan Guo, Xun Yang, and Meng Wang. Transformer-based visual grounding with crossmodality interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):183:1– 183:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3587251.

Li:2024:FSF

[LLH^{+}24] Yongkang Li, Qifan Liang, Zhen Han, Wenjun Mai, and Zhongyuan Wang. Few-shot face sketch-to-photo synthesis via global-local asymmetric image-to-image translation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):325:1–325:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672400.

Li:2012:RIB

[LLHS12] Jian Li, Hongmei Liu, Jiwu Huang, and Yun Q. Shi. Reference index-based H.264 video watermarking scheme. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2S):33:1– 33:??, September 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2020:STD

[LLJ^{+}20] Yonggang Li, Chunping Liu, Yi Ji, Shengrong Gong, and Haibao Xu. Spatio-temporal deep residual network with hierarchical attentions for video event recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):62:1–62:21, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3378026.

Li:2021:SFF

[LLJC21] Yidong Li, Wenhua Liu, Yi Jin, and Yuanzhouhan Cao. SP-GAN: Face forgery using spoofing generative adversarial networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):19:1–19:20, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3432817.

Lin:2015:AMD

[LLJW15] Yin-Tzu Lin, I-Ting Liu, Jyh-Shing Roger Jang, and Ja-Ling Wu. Audio musical dice game: a user-preference-aware medley generating system. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):52:1–52:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2011:GDO

[LLKL11] Frederick W. B. Li, Rynson W. H. Lau, Danny Kilis, and Lewis W. F. Li. Game-on-demand:: an online game engine based on geometry streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):19:1–19:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liang:2021:FDI

[LLL^{+}21] Wei Liang, Jing Long, Kuan-Ching Li, Jianbo Xu, Nanjun Ma, and Xia Lei. A fast defogging image recognition algorithm based on bilateral hybrid filtering. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):42:1–42:16, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3391297.

Liang:2022:SCP

[LLL^{+}22] Liqian Liang, Congyan Lang, Zun Li, Jian Zhao, Tao Wang, and Songhe Feng. Seeing crucial parts: Vehicle model verification via a discriminative representation model. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):28:1– 28:22, February 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3474596.

Lin:2023:IPG

[LLL23] Yun-Shao Lin, Yi-Ching Liu, and Chi-Chun Lee. An interaction-process-guided framework for small-group performance prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):58:1–58:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3558768.

Liu:2024:DIC

[LLL^{+}24a] Bing Liu, Jinfu Lu, Mingming Liu, Hao Liu, Yong Zhou, and Dongping Yang. Diverse image captioning via panoptic segmentation and sequential conditional variational transformer. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):376:1–376:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3695878.

Liu:2024:SSS

[LLL^{+}24b] Qi Liu, Xinchen Liu, Kun Liu, Xiaoyan Gu, and Wu Liu. Sig-Former: Sparse signal-guided transformer for multi-modal action segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):234:1– 234:??, August 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3657296.

Li:2025:ANN

[LLL^{+}25] Yue Li, Junru Li, Chaoyi Lin, Kai Zhang, Li Zhang, Franck Galpin, Thierry Dumas, Hongtao Wang, Muhammed Coban, Jacob Ström, Du Liu, and Kenneth Andersson. Advanced neural network-based video coding technologies for intra prediction and in-loop filtering. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (7):213:1–213:23, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2024:FUC

[LLN^{+}24] Xingbo Liu, Jiamin Li, Xiushan Nie, Xuening Zhang, and Yilong Yin. Fast unsupervised cross-modal hashing with robust factorization and dual projection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):370:1–370:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3694684.

Liu:2020:UNA

[LLO^{+}20] Fangyu Liu, Rémi Lebret, Didier Orel, Philippe Sordet, and Karl Aberer. Upgrading the newsroom: an automated image selection system for news articles. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):81:1–81:28, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3396520.

Li:2006:MSP

[LLP06] H. Li, M. Li, and B. Prabhakaran. Middleware for streaming 3D progressive meshes over lossy networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(4):282–317, November 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2023:NVS

[LLP^{+}23] Bingzheng Liu, Jianjun Lei, Bo Peng, Chuanbo Yu, Wanqing Li, and Nam Ling. Novel view synthesis from a single unposed image via unsupervised learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):186:1–186:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3587467.

Li:2021:TAE

[LLS^{+}21] Yang Li, Guangcan Liu, Yubao Sun, Qingshan Liu, and Shengyong Chen. 3D tensor auto-encoder with application to video compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):48:1– 48:18, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3431768.

Li:2022:GGA

[LLS^{+}22] Kunpeng Li, Chang Liu, Mike Stopa, Jun Amano, and Yun Fu. Guided graph attention learning for video-text matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):131:1–131:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3538533.

Li:2024:DWG

[LLS^{+}24] Chenghua Li, Zongze Li, Jing Sun, Yun Zhang, Xiaoping Jiang, and Fan Zhang. Dynamic weighted gradient reversal network for visible-infrared person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):12:1– 12:??, January 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3607535.

Liu:2012:BET

[LLSC12] Dongyu Liu, Fei Li, Bo Shen, and Songqing Chen. Building an efficient transcoding overlay for P2P streaming to heterogeneous devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 8s(1):10:1–10:??, February 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2020:IVR

[LLSX20] Mading Li, Jiaying Liu, Xiaoyan Sun, and Zhiwei Xiong. Image/video restoration via multiplanar autoregressive model and low-rank optimization. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–23, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341728.

Li:2013:ENO

[LLW^{+}13] Zechao Li, Jing Liu, Meng Wang, Changsheng Xu, and Hanqing Lu. Enhancing news organization for convenient retrieval and browsing. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(1):1:1– 1:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2024:TDC

[LLWX24] Boqian Liu, Haojie Li, Zhihui Wang, and Tianfan Xue. Transparent depth completion using segmentation features. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):373:1– 373:??, December 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3694978.

Luo:2014:ICH

[LLYH14] Da Luo, Weiqi Luo, Rui Yang, and Jiwu Huang. Identifying compression history of wave audio and its applications.

ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (3):30:1–30:??, April 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2022:SGS

[LLYL22] Yunfei Liu, Yu Li, Shaodi You, and Feng Lu. Semantic guided single image reflection removal. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):151:1–151:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3510821.

Lin:2020:SSI

[LLZ^{+}20] Feng Lin, Bin Li, Wengang Zhou, Houqiang Li, and Yan Lu. Single-stage instance segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):86:1–86:19, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3387926.

Li:2021:SSL

[LLZ^{+}21] Zhixin Li, Lan Lin, Canlong Zhang, Huifang Ma, Weizhong Zhao, and Zhiping Shi. A semisupervised learning approach based on adaptive weighted fusion for automatic image annotation. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):37:1–37:23, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3426974.

Li:2022:DIG

[LLZ^{+}22] Jinfeng Li, Weifeng Liu, Yicong Zhou, Jun Yu, Dapeng Tao, and Changsheng Xu. Domaininvariant graph for adaptive semi-supervised domain adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):72:1–72:18, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3487194.

Liu:2023:DDD

[LLZ^{+}23] Hui Liu, Shanshan Li, Jicheng Zhu, Kai Deng, Meng Liu, and Liqiang Nie. DDIFN: a dual-discriminator multi-modal medical image fusion network. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (4):145:1–145:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3574136.

Liu:2024:PAR

[LLZ^{+}24] Zhenyu Liu, Da Li, Xinyu Zhang, Zhang Zhang, Peng Zhang, Caifeng Shan, and Jungong Han. Pedestrian attribute recognition via spatio-temporal relationship learning for visual surveillance. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(6):159:1–159:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632624.

Li:2025:MFA

[LLZ^{+}25] Qiang Li, Di Liu, Guang Zu, Sen Li, Hui Sun, and Jianzhong Wang. Multigranularity feature aggregation and cross-level boundary modeling for temporal action detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):90:1– 90:??, March 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3712598.

Liu:2022:PPM

[LMC^{+}22] Changming Liu, Xiaojing Ma, Sixing Cao, Jiayun Fu, and Bin B. Zhu. Privacy-preserving motion detection for HEVCcompressed surveillance video. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1):23:1–23:27, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3472669.

Liu:2014:SBA

[LMF^{+}14] Yong-Jin Liu, Cui-Xia Ma, Qiufang Fu, Xiaolan Fu, Sheng-Feng Qin, and Lexing Xie. A sketch-based approach for interactive organization of video clips. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1):2:1–2:??, August 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2013:RAM

[LML^{+}13] Heng Liu, Tao Mei, Houqiang Li, Jiebo Luo, and Shipeng Li. Robust and accurate mobile visual localization and its applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):51:1–51:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2024:EVR

[LML^{+}24] Jingyu Li, Zhendong Mao, Hao Li, Weidong Chen, and Yongdong Zhang. Exploring visual relationships via transformerbased graphs for enhanced image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):133:1–133:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638558.

Liu:2014:MIK

[LMLC14] Xianglong Liu, Yadong Mu, Bo Lang, and Shih-Fu Chang. Mixed image-keyword query adaptive hashing over multilabel images. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):22:1–22:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Loveleen:2024:EDH

[LMS^{+}24] Gaur Loveleen, Bhandari Mohan, Bhadwal Singh Shikhar, Jhanjhi Nz, Mohammad Shorfuzzaman, and Mehedi Masud. Explanation-driven HCI model to examine the mini-mental state for Alzheimer’s disease. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):41:1–41:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3527174.

Liang:2024:RFO

[LMX24] Shuang Liang, Wentao Ma, and Chi Xie. Relation with free objects for action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):58:1–58:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3617596.

Li:2021:DBS

[LMXJ21] Honglin Li, Xiaoyang Mao, Mengdi Xu, and Xiaogang Jin. Deep-based self-refined facetop coordination. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):95:1–95:23, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446970.

Liu:2025:EIN

[LNC^{+}25] Yue Liu, Zhangkai Ni, Peilin Chen, Shiqi Wang, Xinfeng Zhang, Hanli Wang, and Sam Kwong. EIN: Exposureinduced network for singleimage HDR reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):302:1–302:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2023:MCM

[LNH23] Zhiming Liu, Kai Niu, and Zhiqiang He. ML-CookGAN: Multi-label generative adversarial network for food image generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):85:1–85:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3554738.

Liu:2021:SSS

[LNT^{+}21] Xinfang Liu, Xiushan Nie, Junya Teng, Li Lian, and Yilong Yin. Single-shot semantic matching network for moment localization in videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (3):84:1–84:14, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3441577.

Liu:2018:CBC

[LOJZ18] Chang Liu, Wei Tsang Ooi, Jinyuan Jia, and Lei Zhao. Cloud Baking: Collaborative scene illumination for dynamic Web3D scenes. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):59:1–59:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2018:RMV

[LPC^{+}18] Sicong Liu, Silvestro Roberto Poccia, K. Selçuk Candan, Maria Luisa Sapino, and Xiaolan Wang. Robust multivariate temporal features of multi-variate time series. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):7:1– 7:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lv:2024:SID

[LPD^{+}24] Zhihan Lv, Fabio Poiesi, Qi Dong, Jaime Lloret, and Houbing Song. Special issue on deep learning for intelligent human computer interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):38:1–38:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3605151.

Langroodi:2015:DCA

[LPS15] Mohsen Jamali Langroodi, Joseph Peters, and Shervin Shirmohammadi. Decodercomplexity-aware encoding of motion compensation for multiple heterogeneous receivers. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (2s):46:1–46:??, February 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2022:CFH

[LPW^{+}22] Wenxu Li, Gang Pan, Chen Wang, Zhen Xing, and Zhenjun Han. From coarse to fine: Hierarchical structure-aware video summarization. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):37:1–37:16, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485472.

Li:2019:LCB

[LPY^{+}19] Yehao Li, Yingwei Pan, Ting Yao, Hongyang Chao, Yong Rui, and Tao Mei. Learning click-based deep structurepreserving embeddings with visual attention. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):78:1–78:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3328994.

Liu:2024:TEC

[LQD^{+}24] Daizong Liu, Xiaoye Qu, Jianfeng Dong, Pan Zhou, Zichuan Xu, Haozhao Wang, Xing Di, Weining Lu, and Yu Cheng. Transform-equivariant consistency learning for temporal sentence grounding. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):106:1–106:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3634749.

Li:2025:SGT

[LQG25] Chenghao Li, Lei Qi, and Xin Geng. A SAM-guided two-stream lightweight model for anomaly detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):67:1– 67:??, February 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3706574.

Li:2018:LLP

[LQH18] Kai Li, Guo-Jun Qi, and Kien A. Hua. Learning label preserving binary codes for multimedia retrieval: a general approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):2:1–2:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lv:2021:ASI

[LQS21] Zhihan Lv, Liang Qiao, and Houbing Song. Analysis of the security of Internet of Multimedia Things. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):97:1–97:16, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3398201.

Lyu:2024:ISI

[LQX^{+}24] Yuanjie Lyu, Penggang Qin, Tong Xu, Chen Zhu, and Enhong Chen. InteractNet: Social interaction recognition for semantic-rich videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):240:1– 240:??, August 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3663668.

Lei:2014:FND

[LQZH14] Yanqiang Lei, Guoping Qiu, Ligang Zheng, and Jiwu Huang. Fast near-duplicate image detection using uniform randomized trees. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(4):35:1–35:??, June 2014. ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2005:CEM

[LS05] Keqiu Li and Hong Shen. Coordinated enroute multimedia object caching in transcoding proxies for tree networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(3): 289–314, August 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lv:2021:TMF

[LS21] Zhihan Lv and Houbing Song.

Trust mechanism of feedback trust weight in multimedia network. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):140:1–140:26, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3391296.

Lee:2025:ACV

[LSCH25] Kuan-Yu Lee, Ashutosh Singla, Pablo Cesar, and Cheng-Hsin Hsu. Adaptive cloud VR gaming optimized by gamer QoE models. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):256:1–256:24, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lew:2006:CBM

[LSDJ06] Michael S. Lew, Nicu Sebe, Chabane Djeraba, and Ramesh Jain. Content-based multimedia information retrieval: State of the art and challenges. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(1): 1–19, February 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2012:DMS

[LSDK12] Yu-Ru Lin, Hari Sundaram, Munmun De Choudhury, and Aisling Kelliher. Discovering multirelational structure in social media streams. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1):4:1–4:??, January 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2015:CSN

[LSK^{+}15] Li-Jia Li, David A. Shamma, Xiangnan Kong, Sina Jafarpour, Roelof Van Zwol, and Xuanhui Wang. CelebrityNet: a social network constructed from large-scale online celebrity images. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1):3:1–3:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2020:BDC

[LSL^{+}20] Jiaying Liu, Sijie Song, Chunhui Liu, Yanghao Li, and Yueyu Hu. A benchmark dataset and comparison study for multimodal human action analytics. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2):41:1–41:24, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3365212.

Li:2024:VDG

[LSL^{+}24] Zongyi Li, Yuxuan Shi, Hefei Ling, Jiazhong Chen, Boyuan Liu, Runsheng Wang, and Chengxin Zhao. Viewpoint disentangling and generation for unsupervised object Re-ID. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):141:1–141:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3632959.

Liu:2020:EEA

[LSN^{+}20] Jinhuan Liu, Xuemeng Song, Liqiang Nie, Tian Gan, and Jun Ma. An end-to-end attention-based neural model for complementary clothing matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–16, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3368071.

Liu:2011:DBA

[LSQ11] Qingzhong Liu, Andrew H. Sung, and Mengyu Qiao. Derivative-based audio steganalysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):18:1–18:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2024:UAV

[LSS^{+}24a] Ting-Lan Lin, Bing-Wei Su, Po-Cheng Shen, Ding-Yuan Chen, Chi-Fu Liang, Yan-Cheng Chen, Yangming Wen, and Mohammad Shahid. Upsampling algorithm for V-PCC-Coded 3D point clouds. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):367:1– 367:??, December 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3690641.

Liu:2024:PPM

[LSS^{+}24b] Jing Liu, Litao Shang, Yuting Su, Weizhi Nie, Xin Wen, and Anan Liu. Privacy-preserving multi-source cross-domain recommendation based on knowledge graph. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):148:1–148:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639706.

Liu:2025:LLP

[LSX^{+}25] Xin Liu, Qiya Song, Lin Xiao, Chun Wang, and Xieping Gao. LPIC: Learnable prompts and ID-guided contrastive learning for multimodal recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):273:1–273:16, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2023:LBR

[LSXZ23] Chen Li, Li Song, Rong Xie, and Wenjun Zhang. Local bidirection recurrent network for efficient video deblurring with the fused temporal merge module. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):170:1–170:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587468.

Li:2019:SRC

[LSYM19] Xianguo Li, Yemei Sun, Yanli Yang, and Changyun Miao. Symmetrical residual connections for single image superresolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):19:1–19:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3282445.

Lee:2014:NDH

[LT14] Ya-Lin Lee and Wen-Hsiang Tsai. A new data hiding method via revision history records on collaborative writing platforms. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):20:1–20:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2021:DAM

[LTD^{+}21] Minxuan Lin, Fan Tang, Weiming Dong, Xiao Li, Changsheng Xu, and Chongyang Ma. Distribution aligned multimodal and multi-domain image stylization. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):96:1–96:17, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3450525.

Lorenzi:2024:MDC

[LTHT24] Daniele Lorenzi, Farzad Tashtarian, Hermann Hellwagner, and Christian Timmerer. MEDUSA: a dynamic codec switching approach in HTTP adaptive streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):319:1–319:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3656175.

Liang:2024:RHG

[LTL^{+}24] Xiaoping Liang, Zhenjun Tang, Zhixin Li, Mengzhu Yu, Hanyun Zhang, and Xianquan Zhang. Robust hashing via global and local invariant features for image copy detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):2:1–2:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3600234.

Lin:2023:FAC

[LTX^{+}23] Jiaying Lin, Xin Tan, Ke Xu, Lizhuang Ma, and Rynson W. H. Lau. Frequency-aware camouflaged object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (2):61:1–61:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3545609.

Lokoc:2021:RIS

[LVM^{+}21] Jakub Lokoc, Patrik Veselý, Frantisek Mejzlík, Gregor Kovalcík, Tomás Soucek, Luca Rossetto, Klaus Schoeffmann, Werner Bailer, Cathal Gurrin, Loris Sauter, Jaeyub Song, Stefanos Vrochidis, Jiaxin Wu, and Björn óR Jónsson. Is the reign of interactive search eternal? Findings from the Video Browser Showdown 2020. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (3):91:1–91:26, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3445031.

Loschky:2007:HLC

[LW07] Lester C. Loschky and Gary S. Wolverton. How late can you update gaze-contingent multiresolutional displays without detection? ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4):7:1–7:10, December 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2023:TFG

[LW23] Shiguang Liu and Huixin Wang. Talking face generation via facial anatomy. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):125:1–125:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3571746.

Lin:2024:TTS

[LW24] Jiayu Lin and Yuan-Gen Wang.

TSFormer: Tracking structure transformer for image inpainting. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):381:1–381:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3696452.

Luo:2020:FFI

[LWH20] Xiaofan Luo, Fukoeng Wong, and Haifeng Hu. FIN: Feature integrated network for object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):48:1–48:18, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3381086.

Li:2022:CAP

[LWH^{+}22] Ran Li, Wei Wei, Peinan Hao, Jian Su, and Fengyuan Sun. Context-aware pseudotrue video interpolation at 6G edge. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):133:1–133:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3555313.

Liu:2023:GML

[LWH^{+}23] Deyin Liu, Lin (Yuanbo) Wu, Richang Hong, Zongyuan Ge, Jialie Shen, Farid Boussaid, and Mohammed Bennamoun. Generative metric learning for adversarially robust open-world person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):20:1– 20:??, January 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3522714.

Lou:2024:SSC

[LWHC24] Xulei Lou, Tinghui Wu, Haifeng Hu, and Dihu Chen. Self-supervised consistency based on joint learning for unsupervised person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (1):27:1–27:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3612926.

Lin:2008:NNB

[LWL08] Tsungnan Lin, Chiapin Wang, and Po-Chiang Lin. A neuralnetwork-based context-aware handoff algorithm for multimedia computing. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3):17:1–17:??, August 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2012:VPA

[LWL^{+}12] Guangda Li, Meng Wang, Zheng Lu, Richang Hong, and Tat-Seng Chua. In-video product annotation with Web information mining. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(4):55:1–55:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2024:ZSS

[LWL24] Jiankai Li, Yunhong Wang, and Weixin Li. Zero-shot scene graph generation via triplet calibration and reduction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):5:1–5:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3604284.

Liang:2024:BSS

[LWLC24] Zechen Liang, Yuan-Gen Wang, Wei Lu, and Xiaochun Cao. Boosting semi-supervised learning with dual-threshold screening and similarity learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):287:1–287:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3672563.

Li:2013:TDI

[LWLZ13] Baochun Li, Zhi Wang, Jiangchuan Liu, and Wenwu Zhu. Two decades of Internet video streaming: a retrospective view. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):33:1–33:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2022:FEA

[LWP22] Shiguang Liu, Huixin Wang, and Min Pei. Facial-expression-aware emotional color transfer based on convolutional neural network. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):8:1–8:19, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3464382.

Lu:2025:KDP

[LWW^{+}25] Chenyang Lu, Zhikai Wei, Huapeng Wu, Le Sun, and Tianming Zhan. KANformer: Dual-priors-guided low-light enhancement via KAN and transformer. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):271:1–271:20, September 2025. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic).

Lv:2020:FSM

[LWWZ20] Chenlei Lv, Zhongke Wu, Xingce Wang, and Mingquan Zhou. 3D facial similarity measurement and its application in facial organization. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):82:1– 82:20, September 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3397765.

Li:2022:DES

[LWY22] Yongrui Li, Zengfu Wang, and Jun Yu. Densely enhanced semantic network for conversation system in social media. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):101:1–101:24, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3501799.

Liang:2021:FBS

[LWZ^{+}21a] Haoran Liang, Jun Wu, Xi Zheng, Mengshi Zhang, Jianhua Li, and Alireza Jolfaei. Fog-based secure service discovery for Internet of Multimedia Things: a cross-blockchain approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (3s):96:1–96:23, January 2021.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3415151.

Liu:2021:VDB

[LWZ21b] Shiguang Liu, Huixin Wang, and Xiaoli Zhang. Video decolorization based on the CNN and LSTM neural network. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (3):88:1–88:18, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3446619.

Lyu:2025:TTP

[LWZ^{+}25] Jiahao Lyu, Jin Wei, Gangyan Zeng, Zeng Li, Enze Xie, Wei Wang, Can Ma, and Yu Zhou. TextBlockV2: Towards precisedetection-free scene text spotting with pre-trained language model. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):175:1–175:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2019:CMF

[LWZH19] Xueliang Liu, Meng Wang, Zheng-Jun Zha, and Richang Hong. Cross-modality feature learning via convolutional autoencoder. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):7:1– 7:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3231740.

Lu:2021:EFD

[LWZW21] Siyuan Lu, Di Wu, Zheng Zhang, and Shui-Hua Wang. An explainable framework for diagnosis of COVID-19 pneumonia via transfer learning and discriminant correlation analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):103:1–103:16, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3449785.

Liu:2022:SSM

[LWZY22] Xiaoming Liu, Shuo Wang, Ying Zhang, and Quan Yuan. Scribble-supervised meibomian glands segmentation in infrared images. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):88:1–88:23, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3497747.

Liu:2021:AAB

[LX21] Xiaoxiao Liu and Qingyang Xu.

Adaptive attention-based highlevel semantic introduction for image caption. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):128:1–128:22, January 2021. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3409388.

Li:2022:IKB

[LXB^{+}22] Qun Li, Fu Xiao, Bir Bhanu, Biyun Sheng, and Richang Hong. Inner knowledge-based Img2Doc scheme for visual question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):76:1– 76:21, August 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3489142.

Li:2024:MGH

[LXC^{+}24] Jiachen Li, Qing Xie, Xiaojun Chang, Jinyu Xu, and Yongjian Liu. Mutually-guided hierarchical multi-modal feature learning for referring image segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):393:1–393:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3698771.

Liu:2025:MMP

[LXC^{+}25] Wei Liu, Xin Xu, Hua Chang, Xin Yuan, and Zheng Wang. Mix-modality person reidentification: a new and practical paradigm. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):112:1–112:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2024:UIQ

[LXJ^{+}24] Xinyue Li, Haiyong Xu, Gangyi Jiang, Mei Yu, Ting Luo, Xuebo Zhang, and Hongwei Ying. Underwater image quality assessment from synthetic to real-world: Dataset and objective method. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):71:1–71:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624983.

Liu:2014:WYB

[LXL^{+}14] Luoqi Liu, Junliang Xing, Si Liu, Hui Xu, Xi Zhou, and Shuicheng Yan. “Wow! You Are So Beautiful Today!”. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):20:1–20:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2018:TPA

[LXL^{+}18] Zhenguang Liu, Yingjie Xia, Qi Liu, Qinming He, Chao Zhang, and Roger Zimmermann. Toward personalized activity level prediction in community question answering Websites. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):41:1–41:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2023:CSL

[LXL^{+}23] Yajing Liu, Zhiwei Xiong, Ya Li, Yuning Lu, Xinmei Tian, and Zheng-Jun Zha. Categorystitch learning for union domain generalization. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):25:1– 25:??, January 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3524136.

Li:2023:EEH

[LXRL23] Ziqiang Li, Pengfei Xia, Xue Rui, and Bin Li. Exploring the effect of high-frequency components in GANs training. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):153:1–153:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3578585.

Ling:2024:VLD

[LXT^{+}24] Jun Ling, Han Xue, Anni Tang, Rong Xie, and Li Song. Vi-CoFace: Learning disentangled latent motion representations for visual-consistent face reenactment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):392:1–392:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3698769.

Liu:2025:TEE

[LXZ^{+}25] Xu Liu, Na Xia, Jinxing Zhou, Zhangbin Li, and Dan Guo. Towards energy-efficient audiovisual classification via multimodal interactive spiking neural network. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):144:1–144:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2011:EOM

[LYC11] Yu-Ching Lin, Yi-Hsuan Yang, and Homer H. Chen. Exploiting online music tags for music emotion classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S (1):26:1–26:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2012:LRC

[LYC^{+}12] Xiaobai Liu, Shuicheng Yan, Bin Cheng, Jinhui Tang, Tat-Sheng Chua, and Hai Jin. Label-to-region with continuity-biased bi-layer sparsity priors. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(4):50:1–50:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2025:CAG

[LYC^{+}25] Yuanman Li, Lanhao Ye, Haokun Cao, Wei Wang, and Zhongyun Hua. Cascaded adaptive graph representation learning for image copy–move forgery detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):56:1–56:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3669905.

Liu:2012:ILC

[LYCJ12] Xiaobai Liu, Shuicheng Yan, Tat-Seng Chua, and Hai Jin. Image label completion by pursuing contextual decomposability. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2): 21:1–21:??, May 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lu:2021:CIC

[LYD^{+}21] Huimin Lu, Rui Yang, Zhenrong Deng, Yonglin Zhang, Guangwei Gao, and Rushi Lan. Chinese image captioning via fuzzy attention-based DenseNet-BiLSTM. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):14:1– 14:18, March 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3422668.

Liang:2024:CCM

[LYDY24] Xiao Liang, Erkun Yang, Cheng Deng, and Yanhua Yang. CrossFormer: Crossmodal representation learning via heterogeneous graph transformer. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):380:1–380:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3688801.

Lyu:2024:DDB

[LYH24] Jun Lyu, Shouang Yan, and M. Shamim Hossain. DBGAN: Dual branch generative adversarial network for multi-modal MRI translation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):235:1–235:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3657298.

Liu:2025:ADC

[LYH^{+}25] Xiyao Liu, Cundian Yang, Jianbiao He, Hui Fang, Gerald Schaefer, Jian Zhang, Yuesheng Zhu, and Shichao Zhang.

Attack-defending contrastive learning for volumetric medical image zero-watermarking. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):62:1–62:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3702230.

Liu:2013:IRQ

[LYJ^{+}13] Dong Liu, Shuicheng Yan, Rong-Rong Ji, Xian-Sheng Hua, and Hong-Jiang Zhang. Image retrieval with queryadaptive hashing. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1):2:1–2:??, February 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lee:2015:LAR

[LYJ^{+}15] Suk Kyu Lee, Seungho Yoo, Jongtack Jung, Hwangnam Kim, and Jihoon Ryoo. Linkaware reconfigurable point-to-point video streaming for mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1):9:1–9:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2021:NNB

[LYL^{+}21a] Yue Li, Yan Yi, Dong Liu, Li Li, Zhu Li, and Houqiang Li. Neural-network-based crosschannel intra prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):77:1– 77:23, August 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3434250.

Liu:2021:ADA

[LYL^{+}21b] Yiding Liu, Siyu Yang, Bin Li, Wengang Zhou, Jizheng Xu, Houqiang Li, and Yan Lu. Affinity derivation for accurate instance segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):34:1– 34:20, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3407090.

Li:2023:TFE

[LYL23a] Junjie Li, Jin Yuan, and Zhiyong Li. TP-FER: an effective three-phase noise-tolerant recognizer for facial expression recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):113:1–113:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570329.

Liu:2023:DML

[LYL^{+}23b] Hao Liu, Zhaoyu Yan, Bing Liu, Jiaqi Zhao, Yong Zhou, and Abdulmotaleb El Saddik. Distilled meta-learning for multi-class incremental learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):149:1–149:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3576045.

Liu:2025:SCD

[LYL^{+}25] Bing Liu, Wenjie Yang, Mingming Liu, Hao Liu, Yong Zhou, and Peng Liu. Syntacticconditional diffusion networks for controllable image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):259:1–259:25, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Le:2023:ASN

[LYLL23] Thi-Ngoc-Hanh Le, Chih-Kuo Yeh, Ying-Chi Lin, and Tong-Yee Lee. Animating still natural images using warping. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):4:1–4:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3511894.

Liu:2022:TTM

[LYW^{+}22] Debin Liu, Laurence T. Yang, Puming Wang, Ruonan Zhao, and Qingchen Zhang. TT-TSVD: a multi-modal tensor train decomposition with its application in convolutional neural networks for smart healthcare. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):41:1–41:17, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491223.

Lan:2023:STS

[LYW^{+}23a] Xiaohan Lan, Yitian Yuan, Xin Wang, Zhi Wang, and Wenwu Zhu. A survey on temporal sentence grounding in videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (2):51:1–51:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3532626.

Liu:2023:TNR

[LYW^{+}23b] Yun Liu, Xiaohua Yin, Zuliang Wan, Guanghui Yue, and Zhi Zheng. Toward a noreference omnidirectional image quality evaluation by using multi-perceptual features. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (2):72:1–72:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3549544.

Liu:2024:AAS

[LYW^{+}24] Ziyi Liu, You Yang, Kejun Wu, Qiong Liu, Xinghua Xu, Xiaoxuan Ma, and Jiang Tang. ASI-Fusion: an adaptive saliency injection-based infrared and visible image fusion network. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):275:1–275:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3665893.

Liu:2025:SSC

[LYW^{+}25] Zheng Liu, Kunyu Yang, Yu Weng, Zheng He, Xuan Liu, and Honghao Gao. SCAG: Semantic co-occurring attention guided alignment for knowledge-based visual question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):207:1–207:20, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2023:CSR

[LYX23] Hao Liu, Xiaoshan Yang, and Changsheng Xu. Counterfactual scenario-relevant knowledge-enriched multi-modal emotion reasoning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):176:1– 176:??, October 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3583690.

Lv:2022:DLB

[LYXA22] Zhihan Lv, Zengchen Yu, Shuxuan Xie, and Atif Alamri. Deep learning-based smart predictive evaluation for interactive multimedia-enabled smart healthcare. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):43:1–43:20, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468506.

Liu:2023:PCQ

[LYXY23] Yipeng Liu, Qi Yang, Yiling Xu, and Le Yang. Point cloud quality assessment: Dataset construction and learningbased no-reference metric. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):80:1–80:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3550274.

Li:2018:PMB

[LYZ^{+}18] Yue Li, Gaobo Yang, Yapei Zhu, Xiangling Ding, and Rongrong Gong. Probability model-based early merge mode decision for dependent views coding in 3D-HEVC. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(4):85:1– 85:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2024:MAS

[LYZ^{+}24] Shenglan Li, Rui Yao, Yong Zhou, Hancheng Zhu, Jiaqi Zhao, Zhiwen Shao, and Abdulmotaleb El Saddik. Motionaware self-supervised RGBT tracking with multi-modality hierarchical transformers. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):391:1– 391:??, December 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3698399.

Li:2021:PBS

[LYZX21] Yaoyu Li, Hantao Yao, Tianzhu Zhang, and Changsheng Xu. Part-based structured representation learning for person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):134:1–134:22, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3412384.

Lan:2021:STR

[LYZY21] Xiangyuan Lan, Zifei Yang, Wei Zhang, and Pong C. Yuen. Spatial-temporal regularized multi-modality correlation filters for tracking with re-detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):57:1–57:16, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3430257.

Liu:2024:DCF

[LYZZ24] Chunpu Liu, Guanglei Yang, Wangmeng Zuo, and Tianyi Zang. DPDFormer: a coarseto-fine model for monocular depth estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):139:1–139:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3638559.

Liu:2019:DCN

[LZC^{+}19] Jiawei Liu, Zheng-Jun Zha, Xuejin Chen, Zilei Wang, and Yongdong Zhang. Dense 3Dconvolutional neural network for person re-identification in videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):8:1–8:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3231741.

Lin:2021:RRN

[LZD^{+}21] Feng Lin, Wengang Zhou, Jiajun Deng, Bin Li, Yan Lu, and Houqiang Li. Residual refinement network with attribute guidance for precise saliency detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):81:1–81:19, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3440694.

Liu:2022:CEC

[LZD^{+}22] Zuquan Liu, Guopu Zhu, Feng Ding, Xiangyang Luo, Sam Kwong, and Peng Li. Contrastenhanced color visual cryptography for (k, n) threshold schemes. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):148:1–148:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3508394.

Li:2025:GHH

[LZDL25] Nianzi Li, Guijuan Zhang, Ping Du, and Dianjie Lu. GP-HSI: Human-scene interaction with geometric and physical constraints. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):113:1–113:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Luo:2022:ERU

[LZF^{+}22] Dezhao Luo, Yu Zhou, Bo Fang, Yucan Zhou, Dayan Wu, and Weiping Wang. Exploring relations in untrimmed videos for self-supervised learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):35:1– 35:21, February 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3473342.

Lv:2024:CTI

[LZG^{+}24] Chenlei Lv, Dan Zhang, Shengling Geng, Zhongke Wu, and Hui Huang. Color transfer for images: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):258:1– 258:??, August 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635152.

Liu:2025:MFI

[LZG^{+}25] Junle Liu, Yun Zhang, Zixi Guo, Xiaoxia Huang, and Gangyi Jiang. Multiscale feature importance-based bit allocation for end-to-end feature coding for machines. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):263:1– 263:19, September 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2020:HCR

[LZH^{+}20] Liang Li, Xinge Zhu, Yiming Hao, Shuhui Wang, Xingyu Gao, and Qingming Huang. A hierarchical CNN-RNN approach for visual emotion classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1–17, January 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3359753.

Li:2024:AAD

[LZH^{+}24a] Mi Li, Wei Zhang, Bin Hu, Jiaming Kang, Yuqi Wang, and Shengfu Lu. Automatic assessment of depression and anxiety through encoding pupilwave from HCI in VR scenes. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):42:1–42:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3513263.

Li:2024:DWA

[LZH^{+}24b] Mingyu Li, Tao Zhou, Zhuo Huang, Jian Yang, Jie Yang, and Chen Gong. Dynamic weighted adversarial learning for semi-supervised classification under intersectional class mismatch. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):115:1–115:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3635310.

Li:2020:SPG

[LZJ^{+}20] Zhaoju Li, Zongwei Zhou, Nan Jiang, Zhenjun Han, Junliang Xing, and Jianbin Jiao. Spatial preserved graph convolution networks for person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):26:1–26:14, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3362988.

Li:2020:CRT

[LZL20a] Miaopeng Li, Zimeng Zhou, and Xinguo Liu. Cross refinement techniques for markerless human motion capture. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):6:1–6:18, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372207.

Liu:2020:LAB

[LZL20b] Zhandong Liu, Wengang Zhou, and Houqiang Li. AB-LSTM: Attention-based bidirectional LSTM model for scene text detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–23, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3356728.

Li:2021:MHP

[LZL^{+}21a] Jianshu Li, Jian Zhao, Congyan Lang, Yidong Li, Yunchao Wei, Guodong Guo, Terence Sim, Shuicheng Yan, and Jiashi Feng. Multi-human parsing with a graph-based generative adversarial model. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):29:1–29:21, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3418217.

Liu:2021:MML

[LZL21b] Zhandong Liu, Wengang Zhou, and Houqiang Li. MFECN: Multi-level feature enhanced cumulative network for scene text detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):78:1–78:22, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3440087.

Li:2023:PEE

[LZL^{+}23] Rui Li, Baopeng Zhang, Wei Liu, Zhu Teng, and Jianping Fan. PANet: an end-to-end network based on relative motion for online multiobject tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):197:1–197:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3595379.

Li:2024:LCA

[LZL^{+}24a] Fengyong Li, Huajun Zhai, Teng Liu, Xinpeng Zhang, and Chuan Qin. Learning compressed artifact for JPEG manipulation localization using wide-receptive-field network. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):315:1–315:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3678883.

Liu:2024:PFM

[LZL^{+}24b] Jiabei Liu, Weiming Zhuang, Yuanyuan Liu, Yonggang Wen, Jun Huang, and Wei Lin. Personalized federated mutual learning for unsupervised camera-aware person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):385:1–385:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3696453.

Liu:2024:CCS

[LZL^{+}24c] Linlin Liu, Haijun Zhang, Qun Li, Jianghong Ma, and Zhao Zhang. Collocated clothing synthesis with GANs aided by textual information: a multimodal framework. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):26:1–26:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3614097.

Li:2025:VLT

[LZL^{+}25] Jie Li, Zhixia Zhao, Qiyue Li, Zhixin Li, Pengyuan Zhou, Zhi Liu, Hao Zhou, and Zhu Li. VPFormer: Leveraging transformer with voxel integration for viewport prediction in volumetric video. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):158:1–158:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Lin:2022:JSC

[LZLJ22] Jinzhi Lin, Yun Zhang, Na Li, and Hongling Jiang. Joint source-channel decoding of polar codes for HEVC-based video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):100:1–100:23, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502208.

Li:2007:SRM

[LZP07] Chuanjun Li, S. Q. Zheng, and B. Prabhakaran. Segmentation and recognition of motion streams by similarity search. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(3): 16:1–16:??, August 2007. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liu:2020:EFA

[LZT^{+}20] Zhiwei Liu, Xiangyu Zhu, Ming Tang, Zhen Lei, and Jinqiao Wang. Efficient face alignment with fast normalization and contour fitting loss. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–16, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3338842.

Liu:2019:MII

[LZW^{+}19] Ruoyu Liu, Yao Zhao, Shikui Wei, Liang Zheng, and Yi Yang. Modality-invariant image-text embedding for image-sentence matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):27:1–27:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3300939.

Liu:2021:NKT

[LZW^{+}21] Zuquan Liu, Guopu Zhu, Yuan-Gen Wang, Jianquan Yang, and Sam Kwong. A novel (t, s, k, n)- threshold visual secret sharing scheme based on access structure partition. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):118:1–118:21, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418212.

Liu:2023:SHL

[LZW^{+}23] Yuqing Liu, Xinfeng Zhang, Shanshe Wang, Siwei Ma, and Wen Gao. Sequential hierarchical learning with distribution transformation for image super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):44:1–44:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532864.

Liu:2024:GRA

[LZW^{+}24] Jun Liu, Jiantao Zhou, Haiwei Wu, Weiwei Sun, and Jinyu Tian. Generating robust adversarial examples against online social networks (OSNs). ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):98:1– 98:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632528.

Li:2023:MSE

[LZWC23] Lei Li, Zhiyuan Zhou, Suping Wu, and Yongrong Cao. Multi-scale edge-guided learning for 3D reconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):109:1–109:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3568678.

Li:2025:EEB

[LZWL25] Na Li, Zichen Zhu, Sheng Wei, and Yao Liu. EVASR: Edgebased salience-aware superresolution for enhanced video quality and power efficiency. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):245:1–245:24, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2021:LFS

[LZX^{+}21] Jiguo Li, Xinfeng Zhang, Jizheng Xu, Siwei Ma, and Wen Gao. Learning to fool the speaker recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):109:1– 109:21, October 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3468673.

Liu:2025:MPP

[LZXP25] Zhiyuan Liu, Qi Zou, Xixia Xu, and Yanting Pei. Multiperson pose estimation with feature enhancement and decoupling based on contrastive learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):181:1–181:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Li:2020:DLA

[LZXY20] Mengyan Li, Zhaoyu Zhang, Guochen Xie, and Jun Yu. A deep learning approach for face hallucination guided by facial boundary responses. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):17:1–17:23, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377874.

Liang:2024:NTH

[LZY^{+}24] Chao Liang, Linchao Zhu, Zongxin Yang, Wei Chen, and Yi Yang. Noise-tolerant hybrid prototypical learning with noisy Web data. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):308:1–308:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672396.

Liu:2024:HFM

[LZYY24] Jinliang Liu, Zhedong Zheng, Zongxin Yang, and Yi Yang. High fidelity makeup via 2D and 3D identity preservation net. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):230:1–230:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3656475.

Li:2023:HAW

[LZZ^{+}23a] Bo Li, Yong Zhang, Chengyang Zhang, Xinglin Piao, and Baocai Yin. Hypergraph association weakly supervised crowd counting. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):195:1–195:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3594670.

Li:2023:IIT

[LZZ23b] Yue Li, Li Zhang, and Kai Zhang. iDAM: Iteratively trained deep in-loop filter with adaptive model selection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):34:1– 34:??, February 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3529107.

Li:2024:SSC

[LZZ^{+}24a] Mengran Li, Ronghui Zhang, Yong Zhang, Xinglin Piao, Shiyu Zhao, and Baocai Yin. SCAE: Structural contrastive auto-encoder for incomplete multi-view representation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):282:1–282:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3672078.

Liang:2024:NHN

[LZZ^{+}24b] Rongjiao Liang, Shichao Zhang, Wenzhen Zhang, Guixian Zhang, and Jinyun Tang. Nonlocal hybrid network for long-tailed image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):108:1–108:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3630256.

Liu:2024:AVT

[LZZ^{+}24c] Yu Liu, Mingbo Zhao, Zhao Zhang, Yuping Liu, and Shuicheng Yan. Arbitrary virtual try-on network: Characteristics preservation and tradeoff between body and clothing. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):123:1–123:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3636426.

Lin:2025:TSC

[LZZ^{+}25] Kang Lin, Wei Zhou, Zhijie Zheng, Dihu Chen, and Tao Su. Temporal and semantic correlation network for weakly-supervised temporal action localization. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):150:1–150:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Liang:2023:HNR

[LZZS23] Shuang Liang, Anjie Zhu, Jiasheng Zhang, and Jie Shao. Hyper-node relational graph attention network for multimodal knowledge graph completion. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):62:1–62:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3545573.

Money:2010:EEL

[MA10] Arthur G. Money and Harry Agius. ELVIS: Entertainment-Led VIdeo Summaries. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):17:1– 17:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Messina:2021:FGV

[MAE^{+}21] Nicola Messina, Giuseppe Amato, Andrea Esuli, Fabrizio Falchi, Claudio Gennaro, and Stéphane Marchand-Maillet. Fine-grained visual textual alignment for cross-modal retrieval using transformer encoders. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):128:1–128:23, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3451390.

Menon:2023:EEM

[MAGT23] Vignesh V. Menon, Hadi Amirpour, Mohammad Ghanbari, and Christian Timmerer. EMES: Efficient multiencoding schemes for HEVCbased adaptive bitrate streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):129:1–129:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3575659.

Miller:2017:QBL

[MATW17] Konstantin Miller, Abdel-Karim Al-Tamimi, and Adam Wolisz. QoE-based low-delay live streaming using throughput predictions. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):4:1–4:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Manzoor:2024:MRL

[MAX^{+}24] Muhammad Arslan Manzoor, Sarah Albarri, Ziting Xian, Zaiqiao Meng, Preslav Nakov, and Shangsong Liang. Multimodality representation learning: a survey on evolution, pretraining and its applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):74:1–74:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3617833.

Masud:2022:CNN

[MAZ22] Mehedi Masud, Mohammed F.

Alhamid, and Yin Zhang. A convolutional neural network model using weighted loss function to detect diabetic retinopathy. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):40:1–40:16, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3470976.

Mohanty:2008:IWB

[MB08] Saraju P. Mohanty and Bharat K.

Bhargava. Invisible watermarking based on creation and robust insertion-extraction of image adaptive watermarks. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(2):12:1– 12:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mackowski:2023:MPI

[MBK^{+}23] Michał Maćkowski, Piotr Brzoza, Mateusz Kawulok, Rafał Meisel, and Dominik Spinczyk. Multimodal presentation of interactive audio-tactile graphics supporting the perception of visual information by blind people. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):167:1–167:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3586076.

Ma:2024:ANR

[MBXZ24] Jingwei Ma, Kangkang Bian, Yang Xu, and Lei Zhu. ANAGL: a noise-resistant and anti-sparse graph learning for micro-video recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):278:1–278:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3670407.

Milani:2011:CAE

[MC11] Simone Milani and Giancarlo Calvagno. A cognitive approach for effective coding and transmission of 3D video. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):23:1– 23:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

More:2019:PLA

[MC19] Amit More and Subhasis Chaudhuri. A pseudo-likelihood approach for geo-localization of events from crowd-sourced sensor-metadata. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):75:1–75:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3321701.

Mondet:2009:CPP

[MCM^{+}09] Sebastien Mondet, Wei Cheng, Geraldine Morin, Romulus Grigoras, Frederic Boudon, and Wei Tsang Ooi. Compact and progressive plant models for streaming in networked virtual environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3):21:1–21:??, August 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Miao:2019:DLS

[MDAE19] Yu Miao, Haiwei Dong, Jihad Mohamad Al Jaam, and Abdulmotaleb El Saddik. A deep learning system for recognizing facial expression in realtime. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):33:1–33:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3311747.

Madhwacharyula:2006:MHV

[MDMK06] Chitra L. Madhwacharyula, Marc Davis, Philippe Mulhem, and Mohan S. Kankanhalli. Metadata handling: a video perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(4):358–388, November 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mohammed:2021:MAF

[MEA^{+}21] Mazin Abed Mohammed, Mohamed Elhoseny, Karrar Hameed Abdulkareem, Salama A. Mostafa, and Mashael S. Maashi. A multi-agent feature selection and hybrid classification model for Parkinson’s disease diagnosis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):74:1–74:22, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3433180.

Miao:2016:HFL

[MFL^{+}16] Dan Miao, Jingjing Fu, Yan Lu, Shipeng Li, and Chang Wen Chen. A high-fidelity and lowinteraction-delay screen sharing system. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(3):44:1–44:??, June 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mallik:2013:MOR

[MGCH13] Anupama Mallik, Hiranmay Ghosh, Santanu Chaudhury, and Gaurav Harit. MOWL: an ontology representation language for Web-based multimedia applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(1):8:1–8:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Motamedi:2017:PPF

[MGG17] Mohammad Motamedi, Philipp Gysel, and Soheil Ghiasi. PLACID: a platform for FPGA-based accelerator creation for DCNNs. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):62:1–62:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mou:2019:AVG

[MGP19] Wenxuan Mou, Hatice Gunes, and Ioannis Patras. Alone versus in-a-group: a multi-modal framework for automatic affect recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):47:1–47:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3321509.

Mazaheri:2018:LMC

[MGS18] Amir Mazaheri, Boqing Gong, and Mubarak Shah. Learning a multi-concept video retrieval model with multiple latent variables. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2):46:1–46:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ma:2022:SAM

[MGY22] Haoyu Ma, Bingchen Gong, and Yizhou Yu. Structureaware meta-fusion for image super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):60:1–60:25, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3477553.

Mesfin:2019:UET

[MHCG19] Gebremariam Mesfin, Nadia Hussain, Alexandra Covaci, and Gheorghita Ghinea. Using eye tracking and heart-rate activity to examine crossmodal correspondences QoE in Mulsemedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):34:1–34:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3303080.

Mahmud:2024:SHA

[MHF24] Bahar Mahmud, Guan Hong, and Bernard Fong. A study of human–AI symbiosis for creative work: Recent developments and future directions in deep learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):47:1–47:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542698.

Min:2025:MFL

[MHL^{+}25] Weiqing Min, Xingjian Hong, Yuxin Liu, Mingyu Huang, Ying Jin, Pengfei Zhou, Leyi Xu, Yilin Wang, Shuqiang Jiang, and Yong Rui. Multimodal food learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):196:1– 196:28, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Metcalf:2008:EPL

[MHT^{+}08] Crysta Metcalf, Gunnar Harboe, Joe Tullio, Noel Massey, Guy Romano, Elaine M. Huang, and Frank Bentley. Examining presence and lightweight messaging in a social television experience. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(4):27:1– 27:??, October 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mou:2013:CBC

[MHT^{+}13] Luntian Mou, Tiejun Huang, Yonghong Tian, Menglin Jiang, and Wen Gao. Content-based copy detection through multimodal feature representation and temporal pyramid matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(1):5:1–5:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ma:2019:PFC

[MHW^{+}19] Ruijun Ma, Haifeng Hu, Weixuan Wang, Jia Xu, and Zhengming Li. Photorealistic face completion with semantic parsing and face identitypreserving features. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):28:1– 28:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3300940.

Mishra:2021:DBR

[MKC21] Prerna Mishra, Santosh Kumar, and Mithilesh Kumar Chaube. Dissimilarity-based regularized learning of charts. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):131:1–131:23, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458884.

Mettes:2020:SIB

[MKS20] Pascal Mettes, Dennis C.

Koelma, and Cees G. M. Snoek. Shuffled ImageNet banks for video event detection and search. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):44:1–44:21, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3377875.

Messaoudi:2017:PAG

[MKSB17] Farouk Messaoudi, Adlen Ksentini, Gwendal Simon, and Philippe Bertin. Performance analysis of game engines on mobile and fixed devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):57:1– 57:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ma:2024:PBM

[MKY^{+}24] Yiyang Ma, Haowei Kuang, Huan Yang, Jianlong Fu, and Jiaying Liu. Prompt-based modality bridging for unified text-to-face generation and manipulation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):386:1–386:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3694974.

Maddage:2011:BSS

[ML11] Namunu C. Maddage and Haizhou Li. Beat space segmentation and octave scale cepstral feature for sung language recognition in pop music. ACM Transactions on Multimedia Computing, Communications, and Applications, 7 (4):37:1–37:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mei:2012:ITC

[MLHL12] Tao Mei, Lusong Li, Xian-Sheng Hua, and Shipeng Li. ImageSense: Towards contextual image advertising. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1):6:1–6:??, January 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mao:2022:MGD

[MLJ^{+}22] Aihua Mao, Yuan Liang, Jianbo Jiao, Yongtuo Liu, and Shengfeng He. Mask-guided deformation adaptive network for human parsing. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):11:1–11:20, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3467889.

Murray:2014:MSE

[MLQM14] Niall Murray, Brian Lee, Yuansong Qiao, and Gabriel-Miro Muntean. Multiple-scent enhanced multimedia synchronization. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):12:1–12:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Marshall:2010:OCM

[MMW10] Damien Marshall, Séamus Mcloone, and Tomás Ward. Optimizing consistency by maximizing bandwidth usage in distributed interactive applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(4):30:1–30:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Merani:2016:ASP

[MN16] Maria Luisa Merani and Laura Natali. Adaptive streaming in P2P live video systems: a distributed rate control approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (3):46:1–46:??, June 2016. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Morillo:2022:EIP

[MNPOF22] Pedro Morillo, José J. Navarro-Pérez, Juan M. Orduña, and Marcos Fernández. Evaluation of an intervention program based on mobile apps to learn sexism prevention in teenagers. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2):45:1–45:20, May 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3471139.

Man:2022:SAR

[MOL^{+}22] Xin Man, Deqiang Ouyang, Xiangpeng Li, Jingkuan Song, and Jie Shao. Scenarioaware recurrent transformer for goal-directed video captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):104:1–104:17, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3503927.

Mawalim:2021:TIR

[MON21] Candy Olivia Mawalim, Shogo Okada, and Yukiko I. Nakano. Task-independent recognition of communication skills in group interaction using timeseries modeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):122:1–122:27, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450283.

Mandel:2011:CTI

[MPE^{+}11] Michael I. Mandel, Razvan Pascanu, Douglas Eck, Yoshua Bengio, Luca M. Aiello, Rossano Schifanella, and Filippo Menczer. Contextual tag inference. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):32:1–32:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mameli:2025:RER

[MPM^{+}25] Marco Mameli, Marina Paolanti, Adriano Mancini, Primo Zingaretti, and Roberto Pierdicca. RenderGAN: Enhancing realtime rendering efficiency with deep learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):99:1–99:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3712263.

Mayer-Patel:2005:BSM

[MPSR05] Ketan Mayer-Patel, Brian C. Smith, and Lawrence A. Rowe. The Berkeley software MPEG-1 video decoder. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1):110–125, February 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mugnai:2022:FGA

[MPTD22] Daniele Mugnai, Federico Pernici, Francesco Turchini, and Alberto Del Bimbo. Fine-grained adversarial semisupervised learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):34:1– 34:19, February 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3485473.

Murray:2007:AEG

[MRS^{+}07] Norman Murray, Dave Roberts, Anthony Steed, Paul Sharkey, Paul Dickerson, and John Rae. An assessment of eye-gaze potential within immersive virtual environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4):8:1–8:17, December 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Meyer:2011:MRL

[MRS11] Marek Meyer, Christoph Rensing, and Ralf Steinmetz. Multigranularity reuse of learning resources. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(1):1:1–1:??, January 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Man:2023:TTE

[MSC^{+}23] Xin Man, Jie Shao, Feiyu Chen, Mingxing Zhang, and Heng Tao Shen. TEVL: Trilinear encoder for video-language representation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):168:1–168:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3585388.

Messina:2025:JDL

[MSFR25] Nicola Messina, Jan Sedmidubsky, Fabrizio Falchi, and Tomás Rebok. Joint-dataset learning and cross-consistent regularization for text-to-motion retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):289:1–289:24, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

MontenegroMarin:2022:ISI

[MSG22] Carlos Enrique Montenegro Marin, Dinesh Jackson Samuel, and Nallappan Gunasekaran. Introduction to the special issue on 6G enabled interactive multimedia communication systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):133:1–133:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3567835.

Mehrabi:2021:MTC

[MSKYJ21] Abbas Mehrabi, Matti Siekkinen, Teemu Kämäräinen, and Antti Ylä-Jääski. Multi-tier CloudVR: Leveraging edge computing in remote rendered virtual reality. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):49:1–49:24, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429441.

Maddage:2010:WLA

[MSL10] Namunu C. Maddage, Khe Chai Sim, and Haizhou Li. Word level automatic alignment of music and lyrics using vocal synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):19:1–19:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Mahto:2024:RCP

[MSS^{+}24] D. K. Mahto, A. K. Singh, K. N. Singh, O. P. Singh, and A. K. Agrawal. Robust copyright protection technique with high-embedding capacity for color images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):334:1–334:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580502.

Mei:2013:NLS

[MTTH13] Tao Mei, Lin-Xie Tang, Jinhui Tang, and Xian-Sheng Hua. Near-lossless semantic video summarization and its applications to video analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3): 16:1–16:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Moncrieff:2007:OAB

[MVW07] Simon Moncrieff, Svetha Venkatesh, and Geoff West. Online audio background determination for complex audio environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(2): 8:1–8:??, May 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Moncrieff:2008:DPA

[MVW08] Simon Moncrieff, Svetha Venkatesh, and Geoff West. Dynamic privacy assessment in a smart house environment using multimodal sensing. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(2):10:1–10:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ma:2024:HLD

[MWL^{+}24] Zeyu Ma, Siwei Wang, Xiao Luo, Zhonghui Gu, Chong Chen, Jinxing Li, Xian-Sheng Hua, and Guangming Lu. HARR: Learning discriminative and high-quality hash codes for image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):134:1– 134:??, May 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3627162.

Mai:2023:MGU

[MXH^{+}23] Sijie Mai, Songlong Xing, Jiaxuan He, Ying Zeng, and Haifeng Hu. Multimodal graph for unaligned multimodal sequence analysis via graph convolution and graph pooling. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (2):54:1–54:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3542927.

Mao:2024:IRB

[MXZ^{+}24] Yangjun Mao, Jun Xiao, Dong Zhang, Meng Cao, Jian Shao, Yueting Zhuang, and Long Chen. Improving referencebased distinctive image captioning with contrastive rewards. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):390:1–390:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3694683.

Ma:2015:PUC

[MY15] Sixuan Ma and Zheng Yan.

PSNController: an unwanted content control system in pervasive social networking based on trust management. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):17:1– 17:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ma:2021:HSP

[MYGX21] Xuan Ma, Xiaoshan Yang, Junyu Gao, and Changsheng Xu. Health status prediction with local-global heterogeneous behavior graph. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):129:1–129:21, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457893.

Ma:2023:MSK

[MYX23a] Xuan Ma, Xiaoshan Yang, and Changsheng Xu. Multi-source knowledge reasoning graph network for multi-modal commonsense inference. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):141:1–141:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3573201.

Mei:2023:MSS

[MYX^{+}23b] Haiyang Mei, Letian Yu, Ke Xu, Yang Wang, Xin Yang, Xiaopeng Wei, and Rynson W. H. Lau. Mirror segmentation via semanticaware contextual contrasted feature learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):100:1–100:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3566127.

Min:2017:FPT

[MZGY17] Xiongkuo Min, Guangtao Zhai, Ke Gu, and Xiaokang Yang. Fixation prediction through multimodal analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):6:1– 6:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ma:2024:RRA

[MZH^{+}24] Yingnan Ma, Chenqiu Zhao, Bingran Huang, Xudong Li, and Anup Basu. RAST: Restorable arbitrary style transfer. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):143:1–143:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3638770.

Ma:2018:CUB

[MZL^{+}18] Ming Ma, Lei Zhang, Jiangchuan Liu, Zhi Wang, Haitian Pang, Lifeng Sun, Weihua Li, Guangling Hou, and Kaiyan Chu. Characterizing user behaviors in mobile personal livecast: Towards an edge computingassisted paradigm. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):66:1– 66:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Meng:2020:ARU

[MZZ^{+}20] Quanling Meng, Heyan Zhu, Weigang Zhang, Xuefeng Piao, and Aijie Zhang. Action recognition using form and motion modalities. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):22:1–22:16, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3350840.

Natarajan:2015:MCC

[NAK15] Prabhu Natarajan, Pradeep K. Atrey, and Mohan Kankanhalli. Multi-camera coordination and control in surveillance systems: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(4):57:1–57:??, April 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Nousias:2023:DSM

[NALM23] Stavros Nousias, Gerasimos Arvanitis, Aris Lalos, and Konstantinos Moustakas. Deep saliency mapping for 3D meshes and applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):71:1–71:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3550073.

Naskar:2013:GTL

[NC13] Ruchira Naskar and Rajat Subhra Chakraborty. A generalized tamper localization approach for reversible watermarking algorithms. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3):19:1– 19:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Niu:2023:VCL

[NCL^{+}23] Tian-Zi Niu, Zhen-Duo Chen, Xin Luo, Peng-Fei Zhang, Zi Huang, and Xin-Shun Xu. Video captioning by learning from global sentence and looking ahead. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):171:1–171:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587252.

Namasudra:2021:SMU

[NCMM21] Suyel Namasudra, Rupak Chakraborty, Abhishek Majumder, and Nageswara Rao Moparthi. Securing multimedia by using DNA-based encryption in the cloud computing environment. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):99:1–99:19, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3392665.

Niu:2023:SEV

[NDC^{+}23] Tian-Zi Niu, Shan-Shan Dong, Zhen-Duo Chen, Xin Luo, Shanqing Guo, Zi Huang, and Xin-Shun Xu. Semantic enhanced video captioning with multi-feature fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):182:1– 182:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3588572.

Ni:2021:LCL

[NDX^{+}21] Tongguang Ni, Yan Ding, Jing Xue, Kaijian Xia, Xiaoqing Gu, and Yizhang Jiang. Local constraint and label embedding multi-layer dictionary learning for sperm head classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):100:1–100:16, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3458927.

Nystrom:2010:ECO

[NH10] Marcus Nyström and Kenneth Holmqvist. Effect of compressed offline foveated video on viewing behavior and subjective quality. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(1):4:1–4:??, February 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Nilsson:2016:ASD

[NHP^{+}16] Tommy Nilsson, Carl Hogsden, Charith Perera, Saeed Aghaee, David J. Scruton, Andreas Lund, and Alan F. Blackwell. Applying seamful design in location-based mobile museum applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):56:1–56:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Neekhara:2024:FSF

[NHZ^{+}24] Paarth Neekhara, Shehzeen Hussain, Xinqiao Zhang, Ke Huang, Julian McAuley, and Farinaz Koushanfar. FaceSigns: Semifragile watermarks for media authentication. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):337:1–337:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3640466.

Namasudra:2024:ISI

[NLKB24] Suyel Namasudra, Pascal Lorenz, Seifedine Kadry, and Syed Ahmad Chan Bukhari. Introduction to the special issue on DNA-centric modeling and practice for next-generation computing and communication systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):31:1–31:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578364.

Nguyen:2013:TDA

[NLN^{+}13] Tam V. Nguyen, Si Liu, Bingbing Ni, Jun Tan, Yong Rui, and Shuicheng Yan. Towards decrypting attractiveness via multi-modality cues. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4): 28:1–28:??, August 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Nahrstedt:2013:ISS

[NLS13] Klara Nahrstedt, Rainer Lienhart, and Malcolm Slaney. Introduction to the special section on the 20th anniversary of the ACM International Conference on Multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):32:1– 32:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Nie:2021:MMI

[NLW^{+}21] Weizhi Nie, Qi Liang, Yixin Wang, Xing Wei, and Yuting Su. MMFN: Multimodal information fusion networks for 3D model classification and retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):131:1–131:22, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3410439.

Nguyen:2021:ISV

[NN21] Phuong-Anh Nguyen and Chong-Wah Ngo. Interactive search vs. automatic search: an extensive study on video retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):47:1– 47:24, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429457.

Nishimura:2025:VLM

[NNK25] Taichi Nishimura, Shota Nakada, and Masayoshi Kondo. Visionlanguage models learn super images for efficient partially relevant video retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):282:1– 282:22, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Natgunanathan:2022:BBA

[NPG^{+}22] Iynkaran Natgunanathan, Purathani Praitheeshan, Longxiang Gao, Yong Xiang, and Lei Pan. Blockchain-based audio watermarking technique for multimedia copyright protection in distribution networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (3):86:1–86:23, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3492803.

Nadimpalli:2024:PDD

[NR24] Aakash Varma Nadimpalli and Ajita Rattani. ProActive Deep-Fake detection using GANbased visible watermarking. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (11):344:1–344:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3625547.

Noori:2020:HAR

[NRUT20] Farzan Majeed Noori, Michael Riegler, Md Zia Uddin, and Jim Torresen. Human activity recognition from multiple sensors data using multi-fusion representations and CNNs. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2):45:1–45:19, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3377882.

Ng:2017:WSD

[NSJB17] Pai Chet Ng, James She, Kang Eun Jeon, and Matthias Baldauf. When smart devices interact with pervasive screens: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):55:1–55:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Nandanwar:2021:NFB

[NSK^{+}21] Lokesh Nandanwar, Palaiahnakote Shivakumara, Divya Krishnani, Raghavendra Ramachandra, Tong Lu, Umapada Pal, and Mohan Kankanhalli. A new foregroundbackground based method for behavior-oriented social media image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):132:1– 132:25, November 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3458051.

Nie:2024:CSI

[NSM^{+}24] Xiushan Nie, Yang Shi, Ziyu Meng, Jin Huang, Weili Guan, and Yilong Yin. Complex scenario image retrieval via deep similarity-aware hashing. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):93:1– 93:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624016.

Nakayama:2008:ECR

[NT08] Minoru Nakayama and Yosiyuki Takahasi. Estimation of certainty for responses to multiplechoice questionnaires using eye movements. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(2):14:1–14:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Nguyen:2020:ETS

[NTT20] Duc V. Nguyen, Huyen T. T. Tran, and Truong Cong Thang. An evaluation of tile selection methods for viewport-adaptive streaming of 360-degree video. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):8:1–8:24, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3373359.

Nguyen:2008:OIV

[NW08] Giang Phuong Nguyen and Marcel Worring. Optimization of interactive visual-similarity-based search. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1):7:1–7:??, January 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Nie:2020:HHG

[NWL^{+}20] Weizhi Nie, Weijie Wang, Anan Liu, Yuting Su, and Jie Nie. HGAN: Holistic generative adversarial networks for two-dimensional imagebased three-dimensional object retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–24, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3344684.

Nie:2021:PPF

[NWNL21] Jie Nie, Zhi-Qiang Wei, Weizhi Nie, and An-An Liu. PGNet: Progressive feature guide learning network for three-dimensional shape recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):87:1–87:17, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3443708.

Niu:2025:SBG

[NXL^{+}25] Yuzhen Niu, Yeyuan Xu, Yuezhou Li, Jiabang Zhang, and Yuzhong Chen. Skeletonboundary-guided network for camouflaged object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (3):83:1–83:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3711869.

Nouma:2024:TED

[NY24] Saif E. Nouma and Attila A.

Yavuz. Trustworthy and efficient digital twins in postquantum era with hybrid hardware-assisted signatures. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (6):156:1–156:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3638250.

Olsen:2012:ITN

[OBBW12] Dan R. Olsen, Derek Bunn, Trent Boulter, and Robert Walz. Interactive television news. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2):19:1–19:??, May 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ota:2017:DLM

[ODMD17a] Kaoru Ota, Minh Son Dao, Vasileios Mezaris, and Francesco G. B. De Natale. Deep learning for mobile multimedia: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):34:1–34:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ota:2017:ISI

[ODMD17b] Kaoru Ota, Minh Son Dao, Vasileios Mezaris, and Francesco G. B. De Natale. Introduction to special issue on deep learning for mobile multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (3s):33:1–33:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ozcelik:2019:CDA

[OE19] Ihsan Mert Ozcelik and Cem Ersoy. Chunk duration-aware SDN-assisted DASH. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):82:1– 82:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3337681.

Ou:2025:MMS

[OL25] Lisong Ou and Zhixin Li.

Multi-modal sarcasm detection on social media via multigranularity information fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (3):95:1–95:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3715139.

Ott:2007:OAT

[OMP07] David E. Ott and Ketan Mayer-Patel. An open architecture for transport-level protocol coordination in distributed multimedia applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(3):17:1–17:??, August 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Okada:2019:MDG

[ONAGP19] Shogo Okada, Laurent Son Nguyen, Oya Aran, and Daniel Gatica-Perez. Modeling dyadic and group impressions with intermodal and interperson features. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):13:1– 13:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3265754.

Oshima:2007:PDS

[ONH07] Chika Oshima, Kazushi Nishimoto, and Norihiro Hagita. A piano duo support system for parents to lead children to practice musical performances. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(2): 9:1–9:??, May 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Punn:2020:IUN

[PA20] Narinder Singh Punn and Sonali Agarwal. Inception U-Net architecture for semantic segmentation to identify nuclei in microscopy cell images. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):12:1–12:15, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3376922.

Pazzi:2014:PPP

[PB14] Richard W. Pazzi and Azzedine Boukerche. PROPANE: a progressive panorama streaming protocol to support interactive 3D virtual environment exploration on graphics-constrained devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1):5:1–5:??, August 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pala:2019:RFM

[PB19] Pietro Pala and Stefano Berretti.

Reconstructing 3D face models by incremental aggregation and refinement of depth frames. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1):23:1–23:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3287309.

Patras:2012:CTS

[PBS12] Paul Patras, Albert Banchs, and Pablo Serrano. A control theoretic scheme for efficient video transmission over IEEE 802.11e EDCA WLANs. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3): 29:1–29:??, July 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pan:2021:SDE

[PCB^{+}21] Yingwei Pan, Yue Chen, Qian Bao, Ning Zhang, Ting Yao, Jingen Liu, and Tao Mei. Smart Director: an event-driven directing system for live broadcasting. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):119:1–119:18, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3448981.

Pala:2020:ISI

[PCH^{+}20] Pietro Pala, Liming Chen, Di Huang, Xiaoming Liu, and Stefanos Zafeiriou. Introduction to the special issue on face analysis applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–2, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3359624.

Papapanagiotou:2016:ICB

[PDD16] Vasileios Papapanagiotou, Christos Diou, and Anastasios Delopoulos. Improving conceptbased image retrieval with training weights computed from tags. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):32:1–32:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

PeresRebelo:2022:IAI

[PDD22] Ana Daniela Peres Rebelo, Guedes De Oliveira Inês, and D. E. Verboom Damion. The impact of artificial intelligence on the creativity of videos.

ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):9:1–9:27, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3462634.

Petrangeli:2016:QDR

[PFC^{+}16] Stefano Petrangeli, Jeroen Famaey, Maxim Claeys, Steven Latré, and Filip De Turck. QoE-driven rate adaptation heuristic for fair adaptive video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):28:1–28:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Peng:2024:JCJ

[PFC^{+}24] Yuxiang Peng, Chong Fu, Guixing Cao, Wei Song, Junxin Chen, and Chiu-Wing Sham. JPEG-compatible joint image compression and encryption algorithm with file size preservation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):105:1–105:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3633459.

Peng:2024:DDL

[PHH^{+}24] Yitao Peng, Lianghua He, Die Hu, Yihang Liu, Longzhen Yang, and Shaohua Shang. Decoupling deep learning for enhanced image recognition interpretability. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):309:1–309:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3674837.

Pham:2020:ESR

[PHS^{+}20] Stefan Pham, Patrick Heeren, Calvin Schmidt, Daniel Silhavy, and Stefan Arbanowski. Evaluation of shared resource allocation using SAND for ABR streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):70:1–70:18, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3388926.

Park:2013:ISL

[PJ13] Jong-Seung Park and Ramesh Jain. Identification of scene locations from geotagged images. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1): 5:1–5:??, February 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Peng:2023:LDS

[PJL23] Fei Peng, Wenyan Jiang, and Min Long. A low distortion and steganalysis-resistant reversible data hiding for 2D engineering graphics. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):53:1–53:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3539661.

Pathak:2025:GRT

[PKL25] Sanhita Pathak, Vinay Kaushik, and Brejesh Lall. Garment recycle training and conditional garment-person outline attention-guided virtual tryon. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):266:1–266:26, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Park:2023:SCQ

[PKLK23] Jae Hyun Park, Sanghoon Kim, Joo Chan Lee, and Jong Hwan Ko. Scalable color quantization for task-centric image compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):82:1–82:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3551389.

Pan:2024:IAL

[PLB^{+}24] Jinwang Pan, Xianming Liu, Yuanchao Bai, Deming Zhai, Junjun Jiang, and Debin Zhao. Illumination-aware low-light image enhancement with transformer and auto-knee curve. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (8):254:1–254:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3664653.

Pan:2023:BTO

[PLYM23] Yingwei Pan, Yehao Li, Ting Yao, and Tao Mei. Bottomup and top-down object inference networks for image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):161:1–161:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3580366.

Pan:2022:CIK

[PLZT22] Yonghua Pan, Zechao Li, Liyan Zhang, and Jinhui Tang. Causal inference with knowledge distilling and curriculum learning for unbiased VQA. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (3):67:1–67:23, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3487042.

Pan:2018:AFP

[PLZW18] Zhaoqing Pan, Jianjun Lei, Yajuan Zhang, and Fu Lee Wang.

Adaptive fractional-pixel motion estimation skipped algorithm for efficient HEVC motion estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):12:1–12:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pang:2016:OQA

[PN16] Lei Pang and Chong-Wah Ngo.

Opinion question answering by sentiment clip localization. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (2):31:1–31:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Peng:2019:CGC

[PQ19] Yuxin Peng and Jinwei Qi. CM-GANs: Cross-modal generative adversarial networks for common representation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1):22:1–22:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3284750.

PascottiValem:2018:USL

[PRGA18] Lucas Pascotti Valem, Carlos Renan De Oliveira, Daniel Carlos Guimarães Pedronette, and Jurandy Almeida. Unsupervised similarity learning through rank correlation and kNN sets. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(4):80:1–80:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Prasad:2014:DVC

[PRH14] Manoj Prasad, Murat Russell, and Tracy A. Hammond. Designing vibrotactile codes to communicate verb phrases. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):11:1–11:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Puig:2023:FFP

[PRPO23] Joan Manuel Marquès Puig, Helena Rifà-Pous, and Samia Oukemeni. From false-free to privacy-oriented communitarian microblogging social networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):83:1–83:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3555354.

Poellabauer:2005:FCD

[PS05] Christian Poellabauer and Karsten Schwan. Flexible cross-domain event delivery for quality-managed multimedia applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(3):248–268, August 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pouladzadeh:2017:MMF

[PS17] Parisa Pouladzadeh and Shervin Shirmohammadi. Mobile multifood recognition using deep learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):36:1–36:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Peng:2024:SSM

[PSL^{+}24] Bo Peng, Lin Sun, Jianjun Lei, Bingzheng Liu, Haifeng Shen, Wanqing Li, and Qingming Huang. Self-supervised monocular depth estimation via binocular geometric correlation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):250:1–250:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663570.

Padhy:2024:MVA

[PSN^{+}24] Ram Prasad Padhy, Pankaj Kumar Sa, Fabio Narducci, Carmen Bisogni, and Sambit Bakshi. Monocular vision-aided depth measurement from RGB images for autonomous UAV navigation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):37:1–37:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3550485.

Plagemann:2005:SPA

[PSS05] Thomas Plagemann, Prashant Shenoy, and John R. Smith. Selected papers from the ACM Multimedia Conference 2003. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(2): 127, May 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Peng:2024:HSE

[PSSZ24] Min Peng, Xiaohu Shao, Yu Shi, and Xiangdong Zhou. Hierarchical synergy-enhanced multimodal relational network for video question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):91:1–91:??, April 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3630101.

Petrangeli:2018:QEC

[PVWD18] Stefano Petrangeli, Jeroen Van Der Hooft, Tim Wauters, and Filip De Turck. Quality of experience-centric management of adaptive video streaming services: Status and challenges. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (2s):31:1–31:??, May 2018. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pi:2024:EVA

[PWHP24] Ruoyan Pi, Peng Wu, Xiangteng He, and Yuxin Peng. EOGT: Video anomaly detection with enhanced object information and global temporal dependency. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):320:1–320:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3662185.

Peng:2024:DLB

[PWX^{+}24] Yunjie Peng, Jinlin Wu, Boqiang Xu, Chunshui Cao, Xu Liu, Zhenan Sun, and Zhiqiang He. Deep learning based occluded person reidentification: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):73:1– 73:??, March 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3610534.

Peng:2025:API

[PWY^{+}25] Heqi Peng, Yunhong Wang, Ruijie Yang, Beichen Li, Rui Wang, and Yuanfang Guo. AED-PADA: Improving generalizability of adversarial example detection via principal adversarial domain adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):70:1–70:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3706061.

Pan:2025:GNG

[PYBY25] Xiao Pan, Zongxin Yang, Shuai Bai, and Yi Yang. GD-NeRF: Generative detail compensation for one-shot generalizable neural radiance fields. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):299:1–299:24, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pan:2020:FLB

[PYZ^{+}20] Zhaoqing Pan, Xiaokai Yi, Yun Zhang, Hui Yuan, Fu Lee Wang, and Sam Kwong. Framelevel bit allocation optimization based on video content characteristics for HEVC. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):15:1–15:20, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380827.

Pan:2025:RII

[PYZ25] Renjie Pan, Hua Yang, and Xiangyu Zhao. ReAL: Improving image-text retrieval with authentic negative repository learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):279:1–279:22, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pan:2008:CBM

[PZ08] Leon Pan and Chang N. Zhang.

A criterion-based multilayer access control approach for multimedia applications and the implementation considerations. ACM Transactions on Multimedia Computing, Communications, and Applications, 5 (2):17:1–17:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Pang:2022:FUP

[PZJL22] Bo Pang, Deming Zhai, Junjun Jiang, and Xianming Liu. Fully unsupervised person reidentification via selective contrastive learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):64:1–64:15, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485061.

Qu:2025:AHL

[QCW^{+}25] Xiaoye Qu, Qiyuan Chen, Wei Wei, Jiashuo Sun, Daizong Liu, and Jianfeng Dong. Alleviating hallucination in large visionlanguage models with active retrieval augmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):258:1– 258:22, September 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Qi:2024:DSD

[QDX^{+}24] Lei Qi, Peng Dong, Tan Xiong, Hui Xue, and Xin Geng. DoubleAUG: Singledomain generalized object detector in urban via color perturbation and dual-style memory. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):126:1–126:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3634683.

Qian:2021:KAM

[QHFX21] Shengsheng Qian, Jun Hu, Quan Fang, and Changsheng Xu. Knowledge-aware multimodal adaptive graph convolutional networks for fake news detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):98:1–98:23, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3451215.

Qi:2008:CMV

[QHR^{+}08] Guo-Jun Qi, Xian-Sheng Hua, Yong Rui, Jinhui Tang, Tao Mei, Meng Wang, and Hong-Jiang Zhang. Correlative multilabel video annotation with temporal kernels. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1):3:1–3:??, October 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Qin:2023:QEC

[QJ23] Qipu Qin and Cheolkon Jung.

Quality enhancement of compressed 360-degree videos using viewport-based deep neural networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):74:1–74:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3551641.

Qin:2024:MGP

[QLGL24] Tai Qin, Ge Li, Wei Gao, and Shan Liu. Multi-grained point cloud geometry compression via dual-model prediction with extended octree. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):289:1–289:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3671001.

Qi:2012:OBI

[QLSQ12] Heng Qi, Keqiu Li, Yanming Shen, and Wenyu Qu. Objectbased image retrieval with kernel on adjacency matrix and local combined features. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(4):54:1– 54:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Qiao:2023:HPI

[QLW^{+}23] Yu Qiao, Yuhao Liu, Ziqi Wei, Yuxin Wang, Qiang Cai, Guofeng Zhang, and Xin Yang. Hierarchical and progressive image matting. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):52:1–52:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3540201.

Qiu:2024:LOP

[QLW^{+}24] Heqian Qiu, Hongliang Li, Qingbo Wu, Hengcan Shi, Lanxiao Wang, Fanman Meng, and Linfeng Xu. Learning offset probability distribution for accurate object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (5):131:1–131:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3637214.

Qayyum:2024:SFB

[QRTM24] Abdul Qayyum, Imran Razzak, M. Tanveer, and Moona Mazher. Spontaneous facial behavior analysis using deep transformer-based framework for child-computer interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):43:1–43:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3539577.

Qudah:2010:EDD

[QS10] Bashar Qudah and Nabil J.

Sarhan. Efficient delivery of on-demand video streams to heterogeneous receivers. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):20:1– 20:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Qiao:2023:UUE

[QSY^{+}23] Ziteng Qiao, Dianxi Shi, Xiaodong Yi, Yanyan Shi, Yuhui Zhang, and Yangyang Liu. UEFPN: Unified and enhanced feature pyramid networks for small object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):95:1– 95:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561824.

Qi:2021:CAW

[QSZ^{+}21] Lianyong Qi, Houbing Song, Xuyun Zhang, Gautam Srivastava, Xiaolong Xu, and Shui Yu. Compatibility-aware Web API recommendation for mashup creation via textual description mining. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):20:1–20:19, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3417293.

Qi:2021:GNT

[QWH^{+}21] Lei Qi, Lei Wang, Jing Huo, Yinghuan Shi, and Yang Gao. GreyReID: a novel two-stream deep framework with RGBgrey information for person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):27:1–27:22, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3419439.

Qiao:2024:SSL

[QXG^{+}24] Shanbao Qiao, Neal N. Xiong, Yongbin Gao, Zhijun Fang, Wenjun Yu, Juan Zhang, and Xiaoyan Jiang. Self-supervised learning of depth and egomotion for 3D perception in human computer interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):45:1–45:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3588571.

Qian:2014:SEC

[QZXH14] Shengsheng Qian, Tianzhu Zhang, Changsheng Xu, and M. Shamim Hossain. Social event classification via boosted multimodal supervised latent Dirichlet allocation. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2):27:1– 27:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rime:2024:HWY

[RABC24] Jemily Rime, Alan Archer-Boyd, and Tom Collins. How will you pod? Implications of creators’ perspectives for designing innovative podcasting tools. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):69:1–69:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3625099.

Rota:2025:SRL

[RBBS25] Claudio Rota, Marco Buzzelli, Simone Bianco, and Raimondo Schettini. Scalable residual Laplacian network for HEVCcompressed video restoration. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):164:1–164:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ren:2025:SDC

[RBLJ25] Peng Ren, Yunfeng Bai, Xiaoheng Li, and Jinyuan Jia. Semantic-driven cross-space graph interaction network for fine-grained 3D point cloud understanding. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):163:1–163:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ranasinghe:2017:DLS

[RD17] Nimesha Ranasinghe and Ellen Yi-Luen Do. Digital lollipop: Studying electrical stimulation on the human tongue to simulate taste sensations. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):5:1– 5:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Roy:2024:CLV

[RE24] Shuvendu Roy and Ali Etemad. Contrastive learning of viewinvariant representations for facial expressions recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):97:1–97:??, April 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3632960.

Roberto:2019:DLS

[REP^{+}19] Pierdicca Roberto, Frontoni Emanuele, Zingaretti Primo, Mancini Adriano, Loncarski Jelena, and Paolanti Marina. Design, large-scale usage testing, and important metrics for augmented reality gaming applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):41:1–41:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3311748.

Riiser:2012:VSU

[REV^{+}12] Haakon Riiser, Tore Endestad, Paul Vigmostad, Carsten Griwodz, and Pâl Halvorsen. Video streaming using a locationbased bandwidth-lookup service for bitrate planning. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3):24:1– 24:??, July 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rahman:2021:MMP

[RHAG21] MD Abdur Rahman, M. Shamim Hossain, Nabil A. Alrajeh, and B. B. Gupta. A multimodal, multimedia point-of-care deep learning framework for COVID-19 diagnosis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):18:1–18:24, March 2021. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3421725.

Rahman:2010:SGA

[RHE10] Abu Saleh Md Mahfujur Rahman, M. Anwar Hossain, and Abdulmotaleb El Saddik. Spatial-geometric approach to physical mobile interaction based on accelerometer and IR sensory data fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 6 (4):28:1–28:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Roodaki:2012:NMD

[RHS12] Hoda Roodaki, Mahmoud Reza Hashemi, and Shervin Shirmohammadi. A new methodology to derive objective quality assessment metrics for scalable multiview 3D video coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):44:1–44:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rout:2020:ICA

[RHS^{+}20] Ranjeet Kumar Rout, Sk. Sarif Hassan, Sanchit Sindhwani, Hari Mohan Pandey, and Saiyed Umer. Intelligent classification and analysis of essential genes using quantitative methods. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):38:1–38:21, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3343856.

Rowe:2005:ASR

[RJ05] Lawrence A. Rowe and Ramesh Jain. ACM SIGMM Retreat report on future directions in multimedia research. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1): 3–13, February 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rawat:2015:CAP

[RK15] Yogesh Singh Rawat and Mohan S. Kankanhalli. Contextaware photography learning for smart mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):19:1– 19:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Raju:2025:ABF

[RLAM25] Chintha Sri Pothu Raju, Rabul Hussain Laskar, Zulfiqar Ali, and Ghulam Muhammad. Attention-based fusion for stroke lesion segmentation on computed tomography perfusion data. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):121:1–121:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ren:2024:HCC

[RLXW24] Yongjun Ren, Zhiying Lv, Neal N. Xiong, and Jin Wang. HCNCT: a cross-chain interaction scheme for the blockchainbased metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(7):188:1–188:??, July 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3594542.

Ruan:2021:CDI

[RLY^{+}21] Weijian Ruan, Chao Liang, Yi Yu, Zheng Wang, Wu Liu, Jun Chen, and Jiayi Ma. Correlation discrepancy insight network for video re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):120:1–120:21, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3402666.

Ren:2022:EES

[RNR^{+}22] Ruyong Ren, Shaozhang Niu, Hua Ren, Shubin Zhang, Tengyue Han, and Xiaohai Tong. ESRNet: Efficient search and recognition network for image manipulation detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):111:1–111:23, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3506853.

Rossi:2020:DUB

[ROST20] Silvia Rossi, Cagri Ozcinar, Aljosa Smolic, and Laura Toni. Do users behave similarly in VR? Investigation of the user influence on the system design. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2):46:1–46:26, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3381846.

Rowe:2013:LFY

[Row13] Lawrence A. Rowe. Looking forward 10 years to multimedia successes. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):37:1–37:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Riegler:2017:ACA

[RPE^{+}17] Michael Riegler, Konstantin Pogorelov, Sigrun Losada Eskeland, Peter Thelin Schmidt, Zeno Albisser, Dag Johansen, Carsten Griwodz, Pål Halvorsen, and Thomas De Lange. From annotation to computer-aided diagnosis: Detailed evaluation of a medical multimedia system. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):26:1–26:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rana:2016:DBV

[RS16] Shuvendu Rana and Arijit Sur. Depth-based viewinvariant blind 3D image watermarking. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):48:1–48:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rashed:2025:RPE

[RSAH25] Ammar Rashed, Shervin Shirmohammadi, Ihab Amer, and Mohamed Hefeeda. A review of player engagement estimation in video games: Challenges and opportunities. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):192:1–192:33, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rabbath:2011:ACP

[RSB11] Mohamad Rabbath, Philipp Sandhaus, and Susanne Boll. Automatic creation of photo books from stories in social media. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):27:1–27:??, 2011. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ravi:2016:FAL

[RSE16] Hareesh Ravi, A. V. Subramanyam, and Sabu Emmanuel. Forensic analysis of linear and nonlinear image filtering using quantization noise. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(3):39:1– 39:??, June 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rudolph:2025:TVP

[RSR25] Michael Rudolph, Stefan Schneegass, and Amr Rizk. Transcoding V-PCC point cloud streams in real-time. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):250:1–250:22, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rainer:2014:GUM

[RT14] Benjamin Rainer and Christian Timmerer. A generic utility model representing the quality of sensory experience. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):14:1– 14:??, September 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ren:2024:CIT

[RTM^{+}24] Bin Ren, Hao Tang, Fanyang Meng, Ding Runwei, Philip H. S. Torr, and Nicu Sebe. Cloth interactive transformer for virtual try-on. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):92:1–92:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617374.

Rajput:2021:SBS

[RTR21] Amitesh Singh Rajput, Vishesh Kumar Tanwar, and Balasubramanian Raman. -score-based secure biomedical model for effective skin lesion segmentation over eHealth cloud. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):65:1– 65:19, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3430806.

Ricci:2023:MLA

[RUD23] Simone Ricci, Tiberio Uricchio, and Alberto Del Bimbo. Metalearning advisor networks for long-tail and noisy labels in social image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):169:1– 169:??, October 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3584360.

Rossi:2025:CAU

[RVTC25] Silvia Rossi, Irene Viola, Laura Toni, and Pablo Cesar. A clustering approach to unveil user similarities in 6 df extended reality applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):254:1–254:27, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rooij:2012:ETS

[RW12] Ork De Rooij and Marcel Worring. Efficient targeted search using a focus and context video browser. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(4):51:1–51:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Rachovides:2007:CIM

[RWP07] Dorothy Rachovides, James Walkerdine, and Peter Phillips. The conductor interaction method. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4):9:1–9:23, December 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Reddy:2005:DSM

[RWW05] A. L. N. Reddy, Jim Wyllie, and K. B. R. Wijayaratne. Disk scheduling in a multimedia I/O system. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1):37–59, February 2005. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ren:2014:BGO

[RXC14] Dongni Ren, Yisheng Xu, and S.-H. Gary Chan. Beyond 1Mbps global overlay live streaming: The case of proxy helpers. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2):26:1–26:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ruan:2025:LSA

[RXY^{+}25] Haoxian Ruan, Zhihua Xu, Zhijing Yang, Yongyi Lu, Jinghui Qin, and Tianshui Chen. Learning semanticaware representation in visuallanguage models for multi-label recognition with partial labels. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):79:1–79:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3708991.

Rao:2023:DPZ

[RYZ^{+}23] Yunbo Rao, Ziqiang Yang, Shaoning Zeng, Qifeng Wang, and Jiansu Pu. Dual projective zero-shot learning using text descriptions. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):10:1–10:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3514247.

Shivani:2016:PVC

[SA16] Shivendra Shivani and Suneeta Agarwal. Progressive visual cryptography with unexpanded meaningful shares. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):50:1– 50:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sarhan:2010:WTP

[SAAH10] Nabil J. Sarhan, Mohammad A. Alsmirat, and Musab Al-Hadrusi. Waiting-time prediction in scalable on-demand video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(2):11:1–11:??, March 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shen:2019:LCS

[SAF19] Liquan Shen, Ping An, and Guorui Feng. Low-complexity scalable extension of the highefficiency video coding (SHVC) encoding system. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):44:1–44:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3313185.

Singh:2021:SHD

[SAL^{+}21a] A. K. Singh, A. Anand, Z. Lv, H. Ko, and A. Mohan. A survey on healthcare data: a security perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):59:1– 59:26, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3422816.

Sun:2021:ARO

[SAL21b] Lu Sun, Hussein Al Osman, and Jochen Lang. An augmented reality online assistance platform for repair tasks. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):50:1– 50:23, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3429285.

Stacchio:2022:THA

[SAL^{+}22] Lorenzo Stacchio, Alessia Angeli, Giuseppe Lisanti, Daniela Calanca, and Gustavo Marfia. Toward a holistic approach to the socio-historical analysis of vernacular photos. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):146:1– 146:??, October 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3507918.

Santini:2011:ECQ

[San11] Simone Santini. Efficient computation of queries on feature streams. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(4):38:1–38:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shen:2015:HFM

[SAZ^{+}15] Liquan Shen, Ping An, Zhaoyang Zhang, Qianqian Hu, and Zhengchuan Chen. A 3D– HEVC fast mode decision algorithm for real-time applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(3):34:1–34:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Srinivas:2023:CBN

[SB23] Kankanala Srinivas and Ashish Kumar Bhandari. Context-based novel histogram bin stretching algorithm for automatic contrast enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):204:1– 204:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3597303.

Sharma:2023:WBA

[SBS23] Prasen Sharma, Ira Bisht, and Arijit Sur. Wavelength-based attributed deep neural network for underwater image restoration. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):2:1–2:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3511021.

Seidenari:2017:DAD

[SBU^{+}17] Lorenzo Seidenari, Claudio Baecchi, Tiberio Uricchio, Andrea Ferracani, Marco Bertini, and Alberto Del Bimbo. Deep artwork detection and retrieval for automatic context-aware audio guides. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):35:1–35:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sarto:2024:TRA

[SCB^{+}24] Sara Sarto, Marcella Cornia, Lorenzo Baraldi, Alessandro Nicolosi, and Rita Cucchiara. Towards retrievalaugmented architectures for image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):242:1–242:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663667.

She:2014:CID

[SCFL14] James She, Jon Crowcroft, Hao Fu, and Flora Li. Convergence of interactive displays with smart mobile devices for effective advertising: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (2):17:1–17:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sha:2025:VDB

[SCX^{+}25] Letian Sha, Xiao Chen, Fu Xiao, Zhong Wang, Zhangbo Long, Qianyu Fan, and Jiankuo Dong. VRVul-discovery: BiLSTMbased vulnerability discovery for virtual reality devices in metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):53:1–53:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3677609.

She:2015:ISI

[SCXC15] James She, Alvin Chin, Feng Xia, and Jon Crowcroft. Introduction to: Special issue on Smartphone-based interactive technologies, systems, and applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):11:1–11:??, October 2015. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic).

Shao:2025:MER

[SCZ^{+}25] Zhiwen Shao, Yifan Cheng, Fan Zhang, Xuehuai Shi, Canlin Li, Lizhuang Ma, and Dit-Yan Yeung. Micro-expression recognition via fine-grained dynamic perception. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):301:1–301:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Song:2024:TDV

[SCZJ24] Xue Song, Jingjing Chen, Bin Zhu, and Yu-Gang Jiang. Textdriven video prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):296:1– 296:??, September 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3675171.

Schramm:2017:ATS

[SDJ17] Rodrigo Schramm, Helena De Souza Nunes, and Cláudio Rosito Jung. Audiovisual tool for Solfège assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):9:1– 9:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Singh:2021:EEB

[SDK^{+}21] Ashima Singh, Arwinder Dhillon, Neeraj Kumar, M. Shamim Hossain, Ghulam Muhammad, and Manoj Kumar. eDiaPredict: an ensemble-based framework for diabetes prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):66:1–66:26, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3415155.

Samrouth:2025:SNB

[SED25] Khouloud Samrouth, Pia El Housseini, and Olivier Deforges. Siamese network-based detection of deepfake impersonation attacks with a person of interest approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):77:1–77:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3708352.

Sachan:2012:ALV

[SEK12] Amit Sachan, Sabu Emmanuel, and Mohan S. Kankanhalli. Aggregate licenses validation for digital rights violation detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2S):37:1–37:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sakr:2007:RCB

[SG07] Ziad Sakr and Nicolas D. Georganas. Robust content-based MPEG-4 XMT scene structure authentication and multimedia content location. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(3):18:1–18:??, August 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

S:2022:TRL

[SG22a] Baiju P. S. and Sudhish N. George. TTV regularized LRTA technique for the estimation of haze model parameters in video dehazing. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):4:1–4:22, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3465454.

Singh:2022:SSD

[SG22b] Gurinder Singh and Puneet Goyal. SDCN2: a shallow densely connected CNN for multi-purpose image manipulation detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):150:1–150:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3510462.

Sharma:2021:HQF

[SGS21] Prasen Kumar Sharma, Sujoy Ghosh, and Arijit Sur. High-quality frame recurrent video de-raining with multicontextual adversarial network. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (2):56:1–56:24, June 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3444974.

Shipman:2008:AVG

[SGW08] Frank Shipman, Andreas Girgensohn, and Lynn Wilcox. Authoring, viewing, and generating hypervideo: an overview of Hyper-Hitchcock. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(2):15:1–15:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shu:2023:CAP

[SGY^{+}23] Zhenyu Shu, Ling Gao, Shun Yi, Fangyu Wu, Xin Ding, Ting Wan, and Shiqing Xin. Context-aware 3D points of interest detection via spatial attention mechanism. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):202:1– 202:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3597026.

Sun:2024:DDC

[SGY^{+}24] Baoli Sun, Yanjun Guo, Tiantian Yan, Xinchen Ye, Zhihui Wang, Haojie Li, and Zhiyong Wang. Digging into depth and color spaces: a mapping constraint network for depth super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):310:1–310:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3677123.

Song:2022:LHV

[SGYX22] Yaguang Song, Junyu Gao, Xiaoshan Yang, and Changsheng Xu. Learning hierarchical video graph networks for one-stop video delivery. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):10:1–10:23, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3466886.

Sharma:2021:TCO

[Sha21] Suraj Sharma. Table of contents: Online supplement volume 16, number 3s. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):117e– 1:117e–2, January 2021. ISSN 1551-6857 (print), 1551-6865 (electronic).

Shenoy:2013:MSR

[She13] Prashant Shenoy. Multimedia systems research: The first twenty years and lessons for the next twenty. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):38:1–38:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shorfuzzaman:2021:EDL

[SHE21]  Mohammad  Shorfuzzaman,

M. Shamim Hossain, and Abdulmotaleb El Saddik. An explainable deep learning ensemble model for robust diagnosis of diabetic retinopathy grading. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):113:1–113:24, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3469841.

Song:2024:CMC

[SHF^{+}24] Dan Song, Shumeng Huo, Xinwei Fu, Chumeng Zhang, Wenhui Li, and An-An Liu. Crossmodal contrastive learning with a style-mixed bridge for single image 3D shape retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (12):366:1–366:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3689645.

Shen:2015:ASM

[SHIE15] Siqi Shen, Shun-Yun Hu, Alexandru Iosup, and Dick Epema. Area of simulation: Mechanism and architecture for multi-avatar virtual environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1):8:1–8:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shirmohammadi:2012:ISS

[SHOG12] Shervin Shirmohammadi, Mohamed Hefeeda, Wei Tsang Ooi, and Romulus Grigoras. Introduction to special section on 3D mobile multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):41:1– 41:??, September 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shi:2024:DMC

[SHSR24] Piao Shi, Min Hu, Xuefeng Shi, and Fuji Ren. Deep modular co-attention shifting network for multimodal sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):109:1–109:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3634706.

Sun:2024:BSG

[SHT^{+}24] Shiqi Sun, Danlan Huang, Xiaoming Tao, Chengkang Pan, Guangyi Liu, and Changwen Chen. Boosting scene graph generation with contextual information. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):54:1–54:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3615868.

Sun:2019:ERF

[SHWC19] Jia Sun, Di Huang, Yunhong Wang, and Liming Chen. Expression robust 3D facial landmarking via progressive coarse-to-fine tuning. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):21:1– 21:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3282833.

Shen:2020:VRS

[SHZ^{+}20] Ling Shen, Richang Hong, Haoran Zhang, Xinmei Tian, and Meng Wang. Video retrieval with similarity-preserving deep temporal hashing. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–16, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3356316.

Singh:2021:DHC

[Sin21] A. K. Singh. Data hiding: Current trends, innovation and potential challenges. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):101:1–101:16, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3382772.

Shen:2019:MLS

[SJC^{+}19] Chen Shen, Zhongming Jin, Wenqing Chu, Rongxin Jiang, Yaowu Chen, Guo-Jun Qi, and Xian-Sheng Hua. Multi-level similarity perception network for person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):32:1–32:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3309881.

Song:2025:CAB

[SJC^{+}25] Yalin Song, Wenbin Jiang, Xiuli Chai, Zhihua Gan, Mengyuan Zhou, and Lei Chen. Cross-attention based twobranch networks for document image forgery localization in the Metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):55:1–55:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3686158.

Siekkinen:2023:NNA

[SK23] Matti Siekkinen and Teemu Kämäräinen. Neural network assisted depth map packing for compression using standard hardware video codecs. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):174:1– 174:??, October 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3588440.

Stephanie:2024:WBP

[SKA24] Veronika Stephanie, Ibrahim Khalil, and Mohammed Atiquzzaman. Weight-based privacypreserving asynchronous SplitFed for multimedia healthcare data. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (12):377:1–377:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3695876.

Singh:2024:ISI

[SKC24] Amit Kumar Singh, Deepa Kundur, and Mauro Conti. Introduction to the special issue on integrity of multimedia and multimodal data in Internet of Things. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(6):151:1–151:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3643040.

Siekkinen:2018:CYS

[SkFM18] Matti Siekkinen, Teemu kämäräinen, Leonardo Favario, and Enrico Masala. Can you see what I see? Quality-of-experience measurements of mobile live video broadcasting. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):34:1–34:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sivaram:2009:DMS

[SKR09] G. S. V. S. Sivaram, Mohan S. Kankanhalli, and K. R. Ramakrishnan. Design of multimedia surveillance systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3): 23:1–23:??, August 2009. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sakai:2013:PPC

[SKSZ13] Kazuya Sakai, Wei-Shinn Ku, Min-Te Sun, and Roger Zimmermann. Privacy preserving continuous multimedia streaming in MANETs. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4):23:1–23:??, August 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Skorin-Kapov:2018:GES

[SKVHC18a] Lea Skorin-Kapov, Martín Varela, Tobias Hoßfeld, and Kuan-Ta Chen. Guest editorial: Special issue on “QoE Management for Multimedia Services”. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (2s):28:1–28:??, May 2018. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Skorin-Kapov:2018:SEC

[SKVHC18b] Lea Skorin-Kapov, Martín Varela, Tobias Hoßfeld, and Kuan-Ta Chen. A survey of emerging concepts and challenges for QoE management of multimedia services. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):29:1– 29:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Schaber:2015:CAM

[SKW^{+}15] Philipp Schaber, Stephan Kopf, Sina Wetzel, Tyler Ballast, Christoph Wesch, and Wolfgang Effelsberg. CamMark: Analyzing, modeling, and simulating artifacts in camcorder copies. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2s):42:1–42:??, February 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shi:2022:HMU

[SL22] Wuzhen Shi and Shaohui Liu. Hiding message using a cycle generative adversarial network. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):143:1–143:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3495566.

Shen:2023:QFC

[SL23] Feihong Shen and Jun Liu. Quantum Fourier convolutional network. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):13:1–13:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514249.

Si:2020:MLT

[SLBS20] Wen Si, Cong Liu, Zhongqin Bi, and Meijing Shan. Modeling long-term dependencies from videos using deep multiplicative neural networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2s):63:1–63:19, July 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3357797.

Shi:2025:FRB

[SLC25] Zenan Shi, Wenyu Liu, and Haipeng Chen. Face reconstructionbased generalized deepfake detection model with residual outlook attention. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):109:1–109:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shen:2024:SAC

[SLG^{+}24] Chengji Shen, Zhenjiang Liu, Xin Gao, Zunlei Feng, and Mingli Song. Self-adaptive clothing mapping based virtual tryon. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):61:1–61:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3613453.

Singh:2021:ISI

[SLK21] Amit Kumar Singh, Zhihan Lv, and Hoon Ko. Introduction to the special issue on Recent Trends in Medical Data Security for e-Health Applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (2s):58:1–58:3, June 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3459601.

Spicer:2012:NAD

[SLKS12] Ryan Spicer, Yu-Ru Lin, Aisling Kelliher, and Hari Sundaram. NextSlidePlease: Authoring and delivering agile multimedia presentations. ACM Transactions on Multimedia Computing, Communications, and Applications, 8 (4):53:1–53:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Su:2021:HRP

[SLL^{+}21] Ge Su, Bo Lin, Wei Luo, Jianwei Yin, Shuiguang Deng, Honghao Gao, and Renjun Xu. Hypomimia recognition in Parkinson’s disease with semantic features. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):106:1–106:20, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3476778.

Sun:2024:MMD

[SLL^{+}24] Kuiyuan Sun, Xiaolong Liu, Xiaolong Li, Yao Zhao, and Wei Wang. Multi-modal driven pose-controllable talking head generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):375:1–375:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673901.

Song:2024:TFI

[SLMJ24] Jiajun Song, Zhuo Li, Weiqing Min, and Shuqiang Jiang. Towards food image retrieval via generalization-oriented sampling and loss function design. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (1):13:1–13:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3600095.

Su:2020:MVG

[SLNL20] Yu-Ting Su, Wen-Hui Li, Wei-Zhi Nie, and An-An Liu. Multiview graph matching for 3D model retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):77:1–77:20, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3387920.

Song:2015:SVT

[SLP15] Minseok Song, Yeongju Lee, and Jinhan Park. Scheduling a video transcoding server to save energy. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2s):45:1–45:??, February 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shirmohammadi:2011:IAM

[SLYE11] Shervin Shirmohammadi, Jiebo Luo, Jie Yang, and Abdulmotaleb El Saddik. Introduction to ACM Multimedia 2010 best paper candidates. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):20:1–20:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shao:2021:EBR

[SLZ^{+}21] Huiru Shao, Jing Li, Jia Zhang, Hui Yu, and Jiande Sun. Eye-based recognition for user identification on mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):117:1–117:19, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3399659.

Sun:2024:GIS

[SLZ24] Youqiang Sun, Jianyi Liu, and Ru Zhang. Generative image steganography based on guidance feature distribution. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (11):336:1–336:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3625297.

Sun:2025:DCM

[SLZ25] Bowen Sun, Guo Lu, and Shibao Zheng. DiFace: Crossmodal face recognition through controlled diffusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):161:1– 161:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

S:2024:ISM

[SMK^{+}24] Shitharth S., Hariprasath Manoharan, Alaa O. Khadidos, Achyut Shankar, Carsten Maple, Adil O. Khadidos, and Shahid Mumtaz. Improved security for multimedia data visualization using hierarchical clustering algorithm. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):335:1– 335:??, November 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3610296.

Shi:2022:OOS

[SMN^{+}22] Ran Shi, Jing Ma, King Ngi Ngan, Jian Xiong, and Tong Qiao. Objective object segmentation visual quality evaluation: Quality measure and pooling method. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):73:1–73:19, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3491229.

Salim:2022:PED

[SMTR22] Sara Salim, Nour Moustafa, Benjamin Turnbull, and Imran Razzak. Perturbationenabled deep federated learning for preserving Internet of Things-based social networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2s):120:1–120:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3537899.

Shi:2012:RTR

[SNC12] Shu Shi, Klara Nahrstedt, and Roy Campbell. A real-time remote rendering system for interactive mobile graphics. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):46:1– 46:??, September 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Siegfried:2022:RUG

[SO22] Rémy Siegfried and Jean-Marc Odobez. Robust unsupervised gaze calibration using conversation and manipulation attention priors. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):20:1–20:27, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472622.

Silva:2013:HPH

[SOC^{+}13] Juan M. Silva, Mauricio Orozco, Jongeun Cha, Abdulmotaleb El Saddik, and Emil M. Petriu. Human perception of haptic-to-video and haptic-to-audio skew in multimedia applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(2):9:1–9:??, May 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sahoo:2021:SAD

[SP21] Kshira Sagar Sahoo and Deepak Puthal. SDN-assisted DDoS defense framework for the Internet of Multimedia Things. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):98:1–98:18, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3394956.

Saxena:2022:PSU

[SR22] Nidhi Saxena and Balasubramanian Raman. Pansharpening scheme using bi-dimensional empirical mode decomposition and neural network. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):108:1– 108:22, November 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3506709.

Sun:2025:GDG

[SR25] Zekun Sun and Na Ruan. GANK: Dynamic geometric and appearance features for efficient and robust detection of face forgery. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):124:1–124:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Singh:2017:SCB

[SRAA17] Priyanka Singh, Balasubramanian Raman, Nishant Agarwal, and Pradeep K. Atrey. Secure cloud-based image tampering detection and localization using POB number system. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):23:1–23:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sikora:2018:SAS

[SRDJ18] Marjan Sikora, Mladen Russo, Jurica Derek, and Ante Jurcević. Soundscape of an archaeological site recreated with audio augmented reality. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):74:1– 74:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Silva:2016:MIB

[SRPH16] Bruno M. C. Silva, Joel J.

P. C. Rodrigues, Neeraj Kumar Mario L. Proença, Jr., and Guangjie Han. MobiCoop: an incentive-based cooperation solution for mobile applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (4):49:1–49:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sharrab:2017:MAP

[SS17] Yousef O. Sharrab and Nabil J.

Sarhan. Modeling and analysis of power consumption in live video streaming systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (4):54:1–54:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Srivastava:2020:DAI

[SS20] Gargi Srivastava and Rajeev Srivastava. Design, analysis, and implementation of efficient framework for image annotation. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):89:1–89:24, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3386249.

Singh:2022:TII

[SS22] Kedar Nath Singh and Amit Kumar Singh. Towards integrating image encryption with compression: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):89:1–89:21, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498342.

Sarma:2024:DID

[SS24] Sandipan Sarma and Arijit Sur. DiRaC-I: Identifying diverse and rare training classes for zero-shot learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):3:1– 3:??, January 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3603147.

Saini:2020:TEB

[SSBT20] Naveen Saini, Sriparna Saha, Pushpak Bhattacharyya, and Himanshu Tuteja. Textual entailment-based figure summarization for biomedical articles. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):35:1–35:24, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3357334.

Shamai:2020:SFP

[SSK20] Gil Shamai, Ron Slossberg, and Ron Kimmel. Synthesizing facial photometries and corresponding geometries using generative adversarial networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–24, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3337067.

Sahu:2021:LMP

[SSP21] Amiya Kumar Sahu, Suraj Sharma, and Deepak Puthal. Lightweight multi-party authentication and key agreement protocol in IoT-based ehealthcare service. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):64:1–64:20, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3398039.

Seufert:2024:COC

[SSP^{+}24] Michael Seufert, Marius Spangenberger, Fabian Poignée, Florian Wamser, Werner Robitza, Christian Timmerer, and Tobias Hoßfeld. COBIRAS: Offering a continuous bit rate slide to maximize DASH streaming bandwidth utilization. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):311:1– 311:??, October 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3677379.

Szeliski:2013:NWC

[SSS13] Richard Szeliski, Noah Snavely, and Steven M. Seitz. Navigating the worldwide community of photos. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):47:1–47:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Singh:2023:ESM

[SSSA23] KN Singh, OP Singh, Amit Kumar Singh, and Amrit Kumar Agrawal. EiMOL: a secure medical image encryption algorithm based on optimization and the Lorenz system. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):94:1– 94:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3561513.

Slivar:2018:GCD

[SSSK18] Ivan Slivar, Mirko Suznjevic, and Lea Skorin-Kapov. Game categorization for deriving QoE-driven video encoding configuration strategies for cloud gaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):56:1–56:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shacham:2007:UDP

[SSTK07] Ron Shacham, Henning Schulzrinne, Srisakul Thakolsri, and Wolfgang Kellerer. Ubiquitous device personalization and use: The next generation of IP multimedia communications. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(2):12:1– 12:??, May 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

She:2020:LDS

[SSY20] Dongyu She, Ming Sun, and Jufeng Yang. Learning discriminative sentiment representation from strongly- and weakly supervised CNNs. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1–19, January 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3326335.

Shao:2023:SIP

[SSZ^{+}23] Xuan Shao, Ying Shen, Lin Zhang, Shengjie Zhao, Dandan Zhu, and Yicong Zhou. SLAM for indoor parking: a comprehensive benchmark dataset and a tightly coupled semantic framework. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):1:1–1:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3510856.

Steinmetz:2010:OOD

[Ste10] Ralf Steinmetz. Obituary to our dear friend Professor Dr. Nicolas D. Georganas, PhD. ACM Transactions on Multimedia Computing, Communications, and Applications, 6 (4):23:1–23:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Steinmetz:2011:EN

[Ste11] Ralf Steinmetz. Editorial notice. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):13:1–13:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Steinmetz:2012:E

[Ste12a]  Ralf Steinmetz.  Editorial.

ACM Transactions on Multimedia Computing, Communications, and Applications, 8 (4):49:1–49:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Steinmetz:2012:EN

[Ste12b] Ralf Steinmetz. Editorial note. ACM Transactions on Multimedia Computing, Communications, and Applications, 8s (1):9:1–9:??, February 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Steinmetz:2012:ENC

[Ste12c] Ralf Steinmetz. Editorial note and call for nominations: Nicolas D. Georganas best paper award. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1):1:1–1:??, January 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Steinmetz:2013:EN

[Ste13a] Ralf Steinmetz. Editorial note.

ACM Transactions on Multimedia Computing, Communications, and Applications, 9 (1s):31:1–31:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Steinmetz:2013:ER

[Ste13b] Ralf Steinmetz. Editorial: Reviewers. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4):22:1–22:??, August 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Steinmetz:2014:EN

[Ste14] Ralf Steinmetz. Editorial note.

ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (1):1:1–1:??, August 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Song:2023:DSD

[STSW23] Yue Song, Hao Tang, Nicu Sebe, and Wei Wang. Disentangle saliency detection into cascaded detail modeling and body filling. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):7:1–7:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3513134.

Sundaram:2013:EMS

[Sun13] Hari Sundaram. Experiential media systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):49:1–49:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Silva:2021:UMC

[SVA^{+}21] Ellen P. Silva, Natália Vieira, Glauco Amorim, Renata Mousinho, Gustavo Guedes, Gheorghita Ghinea, and Joel A. F. Dos Santos. Using multisensory content to impact the quality of experience of reading digital books. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):124:1–124:18, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458676.

Snidaro:2012:FMV

[SVF12] Lauro Snidaro, Ingrid Visentini, and Gian Luca Foresti. Fusing multiple video sensors for surveillance. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(1):7:1–7:??, January 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Strezoski:2018:OLS

[SW18] Gjorgji Strezoski and Marcel Worring. OmniArt: a large-scale artistic benchmark. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (4):88:1–88:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Swaminathan:2013:WMV

[Swa13] Viswanathan Swaminathan.

Are we in the middle of a video streaming revolution? ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):40:1– 40:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Snoek:2006:LRS

[SWH06] Cees G. M. Snoek, Marcel Worring, and Alexander G. Hauptmann. Learning rich semantics from news video archives by style analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(2):91–108, May 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sun:2022:ESR

[SWK^{+}22] Guofei Sun, Yongkang Wong, Mohan S. Kankanhalli, Xiangdong Li, and Weidong Geng. Enhanced 3D shape reconstruction with knowledge graph of category concept. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):71:1–71:20, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491224.

Shi:2024:RSB

[SWM^{+}24] Hongjian Shi, Hao Wang, Ruhui Ma, Yang Hua, Tao Song, Honghao Gao, and Haibing Guan. Robust searchingbased gradient collaborative management in intelligent transportation system. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):34:1– 34:??, February 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3549939.

Shamsolmoali:2021:ISI

[SWS21] Pourya Shamsolmoali, Ruili Wang, and A. H. Sadka. Introduction to the special issue on Advanced Approaches for Multiple Instance Learning on Multimedia Applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):69:1– 69:2, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3459603.

Sun:2022:RGJ

[SWS^{+}22] Teng Sun, Chun Wang, Xuemeng Song, Fuli Feng, and Liqiang Nie. Response generation by jointly modeling personalized linguistic styles and emotions. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):52:1–52:20, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3475872.

Shi:2025:RNN

[SWZ25] Mengyu Shi, Miao Wang, and Yujun Zhang. RePC: a novel neural video quality enhancement system framework for ABR streaming of VBRencoded videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):137:1–137:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sang:2011:BCT

[SX11] Jitao Sang and Changsheng Xu. Browse by chunks: Topic mining and organizing on webscale social media. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):30:1–30:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sang:2013:SIA

[SX13] Jitao Sang and Changsheng Xu. Social influence analysis and application on multimedia sharing websites. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):53:1–53:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shao:2006:ASM

[SXM^{+}06] Xi Shao, Changsheng Xu, Namunu C. Maddage, Qi Tian, Mohan S. Kankanhalli, and Jesse S. Jin. Automatic summarization of music videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(2): 127–148, May 2006. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shi:2023:LVT

[SXY^{+}23] Yaya Shi, Haiyang Xu, Chunfeng Yuan, Bing Li, Weiming Hu, and Zheng-Jun Zha. Learning video-text aligned representations for video captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):63:1–63:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3546828.

Sauer:2009:MDC

[SY09] Danielle Sauer and Yee-Hong Yang. Music-driven character animation. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(4):27:1–27:??, October 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Song:2025:TCI

[SYM^{+}25] Tao Song, Kunlin Yang, Fan Meng, Xin Li, Handan Sun, and Chenglizhao Chen. Tropical cyclone image super-resolution via multimodality fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):131:1– 131:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sobhani:2017:VBA

[SYS17] Ashkan Sobhani, Abdulsalam Yassine, and Shervin Shirmohammadi. A video bitrate adaptation and prediction mechanism for HTTP adaptive streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(2):18:1–18:??, May 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Song:2023:SSC

[SYX23] Yaguang Song, Xiaoshan Yang, and Changsheng Xu. Selfsupervised calorie-aware heterogeneous graph networks for food recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):27:1– 27:??, February 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524618.

Shi:2024:BFS

[SYY^{+}24] Yanyan Shi, Shaowu Yang, Wenjing Yang, Dianxi Shi, and Xuehui Li. Boosting fewshot object detection with discriminative representation and class margin. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):75:1–75:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3608478.

Shen:2012:IFP

[SZ12] Zhijie Shen and Roger Zimmermann. ISP-friendly P2P live streaming: a roadmap to realization. ACM Transactions on Multimedia Computing, Communications, and Applications, 8s(1):11:1–11:??, February 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sun:2022:SSS

[SZB^{+}22] Ziyi Sun, Yunfeng Zhang, Fangxun Bao, Ping Wang, Xunxiang Yao, and Caiming Zhang. SADnet: Semisupervised single image dehazing method based on an attention mechanism. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):58:1–58:23, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3478457.

Shi:2024:DAT

[SZC24] Caijuan Shi, Yuanfan Zheng, and Zhen Chen. Domain adaptive thermal object detection with unbiased granularity alignment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):274:1–274:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665892.

Sharma:2021:ISI

[SZEST21] Suraj Sharma, Xuyun Zhang, Hesham El-Sayed, and Zhiyuan Tan. Introduction to the special issue on privacy and security in evolving Internet of Multimedia Things. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):93:1–93:3, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3423955.

Song:2025:LGV

[SZF^{+}25] Jianbo Song, Hong Zhang, Yachun Feng, Hanyang Liu, and Yifan Yang. Languageguided visual tracking: Comprehensive and effective multimodal information fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):290:1–290:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sun:2023:IRG

[SZFR23] Sophie C. C. Sun, Yongkang Zhao, Fang-Wei Fu, and Yawei Ren. Improved random gridbased cheating prevention visual cryptography using latin square. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):77:1–77:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3550275.

Shah:2019:PCB

[SZHY19] Mohsin Shah, Weiming Zhang, Honggang Hu, and Nenghai Yu. Paillier cryptosystem based mean value computation for encrypted domain image processing operations. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):76:1– 76:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3325194.

Shi:2022:SIN

[SZL^{+}22] Qinghongya Shi, Hong-Bo Zhang, Zhe Li, Ji-Xiang Du, Qing Lei, and Jing-Hua Liu. Shuffle-invariant network for action recognition in videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (3):69:1–69:18, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3485665.

Shi:2024:ICM

[SZL^{+}24] Dan Shi, Lei Zhu, Jingjing Li, Guohua Dong, and Huaxiang Zhang. Incomplete crossmodal retrieval with deep correlation transfer. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):127:1–127:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3637442.

Su:2017:DDP

[SZLL17] Zhuo Su, Kun Zeng, Hanhui Li, and Xiaonan Luo. A dualdomain perceptual framework for generating visual inconspicuous counterparts. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(2):22:1– 22:??, May 2017. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic).

Siarohin:2019:IIM

[SZM^{+}19] Aliaksandr Siarohin, Gloria Zen, Cveta Majtanovic, Xavier Alameda-Pineda, Elisa Ricci, and Nicu Sebe. Increasing image memorability with neural style transfer. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):42:1–42:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3311781.

Shen:2021:CDO

[SZM^{+}21] Xiangjun Shen, Jinghui Zhou, Zhongchen Ma, Bingkun Bao, and Zhengjun Zha. Crossdomain object representation via robust low-rank correlation analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):126:1–126:20, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3458825.

Sun:2016:SOR

[SZTL16] Shaoyan Sun, Wengang Zhou, Qi Tian, and Houqiang Li. Scalable object retrieval with compact image representation from generic object regions. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):29:1– 29:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Sun:2024:IRL

[SZW^{+}24] Yunda Sun, Lin Zhang, Zhong Wang, Yang Chen, Shengjie Zhao, and Yicong Zhou. I2P registration by learning the underlying alignment feature space from pixel-to-point similarities. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):388:1–388:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3697839.

Song:2023:SSI

[SZZ^{+}23] Dan Song, Chu-Meng Zhang, Xiao-Qian Zhao, Teng Wang, Wei-Zhi Nie, Xuan-Ya Li, and An-An Liu. Self-supervised image-based 3D model retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):70:1–70:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3548690.

Sun:2018:RPP

[SZZT18] Weiwei Sun, Jiantao Zhou, Shuyuan Zhu, and Yuan Yan Tang. Robust privacypreserving image sharing over online social networks (OSNs).

ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (1):14:1–14:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Shamsolmoali:2020:AAM

[SZZY20] Pourya Shamsolmoali, Masoumeh Zareapoor, Huiyu Zhou, and Jie Yang. AMIL: Adversarial multi-instance learning for human pose estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):23:1–23:23, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3355612.

Shao:2025:TRV

[SZZZ25] Xuan Shao, Lin Zhang, Tianjun Zhang, and Shengjie Zhao. Towards a robust visual-inertial-surround-view SLAM system for autonomous indoor parking. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):204:1–204:23, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Toni:2015:OSA

[TAPP^{+}15] Laura Toni, Ramon Aparicio-Pardo, Karine Pires, Gwendal Simon, Alberto Blanc, and Pascal Frossard. Optimal selection of adaptive streaming representations. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2s):43:1–43:??, February 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Thomee:2016:FSP

[TAS16] Bart Thomee, Ioannis Arapakis, and David A. Shamma. Finding social points of interest from georeferenced and oriented online photographs. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (2):36:1–36:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tasaka:2020:CSM

[Tas20] Shuji Tasaka. Causal structures of multidimensional QoE in haptic-audiovisual communications: Bayesian modeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):11:1–11:23, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3375922.

Tasaka:2022:EMC

[Tas22] Shuji Tasaka. An empirical method for causal inference of constructs for QoE in hapticaudiovisual communications. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1):17:1–17:24, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3473986.

Tasaka:2024:UQM

[Tas24] Shuji Tasaka. Usefulness of QoS in multidimensional QoE prediction for haptic-audiovisual communications. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):22:1–22:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3613246.

Timmerer:2025:HAS

[TAT^{+}25] Christian Timmerer, Hadi Amirpour, Farzad Tashtarian, Samira Afzal, Amr Rizk, Michael Zink, and Hermann Hellwagner. HTTP adaptive streaming: a review on current advances and future challenges. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):198:1–198:27, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Teodosio:2005:SS

[TB05] Laura Teodosio and Walter Bender. Salient stills. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(1):16–36, February 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Timmerer:2017:BPA

[TB17] Christian Timmerer and Ali C.

Begen. Best papers of the 2016 ACM Multimedia Systems (MMSys) Conference and Workshop on Network and Operating System Support for Digital Audio and Video (NOSSDAV) 2016. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):40:1– 40:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tan:2011:URS

[TBC^{+}11] Shulong Tan, Jiajun Bu, Chun Chen, Bin Xu, Can Wang, and Xiaofei He. Using rich social media information for music recommendation via hypergraph model. ACM Transactions on Multimedia Computing, Communications, and Applications, 7S(1):22:1–22:??, 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Thouin:2008:EAV

[TC08] Frederic Thouin and Mark Coates. Equipment allocation in video-on-demand network deployments. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1):5:1–5:??, October 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tjondronegoro:2008:SES

[TCJ08] Dian Tjondronegoro, Yi-Ping Phoebe Chen, and Adrien Joly. A scalable and extensible segmentevent-object-based sports video retrieval system. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2):13:1–13:??, May 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tang:2023:UDA

[TCL^{+}23] Zhenjun Tang, Zhiyuan Chen, Zhixin Li, Bineng Zhong, Xianquan Zhang, and Xinpeng Zhang. Unifying dual-attention and Siamese transformer network for full-reference image quality assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):205:1– 205:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3597434.

Filho:2020:DPV

[TCP^{+}20] Roberto Iraja Tavares Da Costa Filho, Marcelo Caggiani Luizelli, Stefano Petrangeli, Maria Torres Vega, Jeroen Van der Hooft, Tim Wauters, Filip De Turck, and Luciano Paschoal Gaspary. Dissecting the performance of VR video streaming through the VR-EXP experimentation platform. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–23, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3360286.

Takezoe:2025:CAA

[TCS^{+}25] Rinyoichi Takezoe, Hao Chen, Gang Shen, Xuefei Lv, Yaowei Wang, Shiliang Zhang, and Xiaoyu Wang. Context-assisted active learning for weakly supervised person search. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):138:1– 138:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tang:2013:TOH

[TCW^{+}13] Jinhui Tang, Qiang Chen, Meng Wang, Shuicheng Yan, Tat-Seng Chua, and Ramesh Jain. Towards optimizing human labeling for interactive image tagging. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4):29:1–29:??, August 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tang:2023:DUK

[TDW^{+}23] Hao Tang, Lei Ding, Songsong Wu, Bin Ren, Nicu Sebe, and Paolo Rota. Deep unsupervised key frame extraction for efficient video classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):119:1–119:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3571735.

Tyson:2016:MAM

[TESU16] Gareth Tyson, Yehia Elkhatib, Nishanth Sastry, and Steve Uhlig. Measurements and analysis of a major adult video portal. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):35:1–35:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Telili:2024:DBL

[TFHM24] Ahmed Telili, Sid Ahmed Fezza, Wassim Hamidouche, and Hanene F. Z. Brachemi Meftah. 2BiVQA: Double bi-LSTM-based video quality assessment of UGC videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):100:1–100:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3632178.

Tang:2025:QAO

[TFY^{+}25] Jianheng Tang, Kejia Fan, Wenjie Yin, Shihao Yang, Yajiang Huang, Anfeng Liu, Naixue Xiong, Mianxiong Dong, Tian Wang, and Shaobo Zhang. A quality-aware and obfuscation-based data collection scheme for cyber-physical metaverse systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):50:1– 50:??, February 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3659582.

Tang:2023:LSR

[TGC23] Geyu Tang, Xingyu Gao, and Zhenyu Chen. Learning semantic representation on visual attribute graph for person re-identification and beyond. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):206:1–206:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487044.

Tanveer:2021:PLT

[TGSF21] M. Tanveer, Tarun Gupta, Miten Shah, and For the Alzheimer’s Disease Neuroimaging Initiative. Pinball loss twin support vector clustering. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):63:1– 63:23, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3409264.

Tang:2022:HMB

[TH22] Zengming Tang and Jun Huang. Harmonious multibranch network for person re-identification with harder triplet loss. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):98:1–98:21, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3501405.

Telili:2025:CHP

[THA^{+}25] Ahmed Telili, Wassim Hamidouche, Hadi Amirpour, Sid Ahmed Fezza, Christian Timmerer, and Luce Morin. Convex hull prediction methods for bitrate ladder construction: Design, evaluation, and comparison. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (7):200:1–200:23, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tahir:2022:NAT

[THR^{+}22] Madiha Tahir, Zahid Halim, Atta Ur Rahman, Muhammad Waqas, Shanshan Tu, Sheng Chen, and Zhu Han. Non-acted text and keystrokes database and learning methods to recognize emotions. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):61:1–61:24, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480968.

Tan:2014:PVS

[TJN14] Song Tan, Yu-Gang Jiang, and Chong-Wah Ngo. Placing videos on a semantic hierarchy for search result navigation. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (4):37:1–37:??, June 2014. ISSN 1551-6857 (print), 1551-6865 (electronic).

Tu:2025:DFT

[TJZ^{+}25] Jiahang Tu, Wei Ji, Hanbin Zhao, Chao Zhang, Roger Zimmermann, and Hui Qian. DriveDiTFit: Fine-tuning diffusion transformers for autonomous driving data generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):85:1–85:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3712064.

Thirunarayanan:2017:CSE

[TKK^{+}17] Ishwarya Thirunarayanan, Khimya Khetarpal, Sanjeev Koppal, Olivier Le Meur, John Shea, and Eakta Jain. Creating segments and effects on comics by clustering gaze data. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):24:1– 24:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tanveer:2020:ISI

[TKPL20] M. Tanveer, P. Khanna, M. Prasad, and C. T. Lin. Introduction to the special issue on computational intelligence for biomedical data and imaging. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):29:1–29:4, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381919.

Tang:2022:LTS

[TLY^{+}22] Yansong Tang, Xingyu Liu, Xumin Yu, Danyang Zhang, Jiwen Lu, and Jie Zhou. Learning from temporal spatial cubism for cross-dataset skeletonbased action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):46:1– 46:24, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3472722.

Tong:2020:PND

[TLZ^{+}20] Chao Tong, Baoyu Liang, Mengze Zhang, Rongshan Chen, Arun Kumar Sangaiah, Zhigao Zheng, Tao Wan, Chenyang Yue, and Xinyi Yang. Pulmonary nodule detection based on ISODATA-improved faster RCNN and 3D-CNN with focal loss. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):36:1–36:9, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3365445.

Tang:2021:MCA

[TLZ^{+}21] Xiaochuan Tang, Mingzhe Liu, Hao Zhong, Yuanzhen Ju, Weile Li, and Qiang Xu. MILL: Channel attention-based deep multiple instance learning for landslide recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):76:1– 76:11, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3454009.

Tu:2024:RFI

[TLZL24] Hanyue Tu, Li Li, Wengang Zhou, and Houqiang Li. Reconstruction-free image compression for machine vision via knowledge transfer. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):313:1– 313:??, October 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3678471.

T:2025:TGS

[TM25] Anjali T. and Masilamani V. Text-guided synthesis of masked face images. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):8:1– 8:??, January 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3654667.

Tiotsop:2022:MIM

[TMB^{+}22] Lohic Fotio Tiotsop, Tomas Mizdos, Marcus Barkowsky, Peter Pocta, Antonio Servetti, and Enrico Masala. Mimicking individual media quality perception with neural network based artificial observers. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1):12:1–12:25, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3464393.

Tullimas:2008:MSU

[TNEcC08] Sunand Tullimas, Thinh Nguyen, Rich Edgecomb, and Sen ching Cheung. Multimedia streaming using multiple TCP connections. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2):12:1–12:??, May 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tran:2021:CQM

[TNH^{+}21] Huyen T. T. Tran, Nam Pham Ngoc, Tobias Hoßfeld, Michael Seufert, and Truong Cong Thang. Cumulative quality modeling for HTTP adaptive streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):22:1–22:24, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3423421.

Tulilaulu:2018:DM

[TNP^{+}18] Aurora Tulilaulu, Matti Nelimarkka, Joonas Paalasmaa, Daniel Johnson, Dan Ventura, Petri Myllys, and Hannu Toivonen. Data musicalization. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (2):47:1–47:??, May 2018. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tang:2025:FGS

[TNZ^{+}25] Jiajia Tang, Binbin Ni, Feiwei Zhou, Dongjun Liu, Yu Ding, Yong Peng, Andrzej Cichocki, Qibin Zhao, and Wanzeng Kong. Fine-grained semantic disentanglement network for multimodal sarcasm analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):162:1–162:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

TOMCCAP-STAFF:2012:TCO

[TOM12] TOMCCAP-STAFF. Table of contents: Online supplement volume 8, number 2s, online supplement volume 8, number 3s. ACM Transactions on Multimedia Computing, Communications, and Applications, 8 (4):48:1–48:??, November 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tiwari:2022:ODN

[TP22] Arti Tiwari and Millie Pant. Optimized deep-neural network for content-based medical image retrieval in a brownfield IoMT network. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):125:1–125:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3546194.

Tanveer:2020:MLT

[TRK^{+}20] M. Tanveer, B. Richhariya, R. U. Khan, A. H. Rashid, P. Khanna, M. Prasad, and C. T. Lin. Machine learning techniques for the diagnosis of Alzheimer’s disease: a review. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):30:1–30:35, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3344998.

Tanwar:2020:CPP

[TRRB20] Vishesh Kumar Tanwar, Balasubramanian Raman, Amitesh Singh Rajput, and Rama Bhargava. CryptoLesion: a privacypreserving model for lesion segmentation using whale optimization over cloud. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):50:1– 50:23, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3380743.

Tripathi:2020:CNC

[TS20] Suvidha Tripathi and Satish Kumar Singh. Cell nuclei classification in histopathological images using Hybrid O L ConvNet. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):32:1–32:22, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3345318.

Thong:2022:DSV

[TS22] William Thong and Cees G. M.

Snoek. Diversely-supervised visual product search. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):13:1– 13:22, January 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3461646.

Tu:2025:LFF

[TSH^{+}25] Zhewei Tu, Xiangbo Shu, Peng Huang, Rui Yan, Zhenxing Liu, and Jiachao Zhang. Leveraging frame- and feature-level progressive augmentation for semi-supervised action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):107:1–107:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tu:2005:ASP

[TSHP05] Yi-Cheng Tu, Jianzhong Sun, Mohamed Hefeeda, and Sunil Prabhakar. An analytical study of peer-to-peer media streaming systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(4):354–376, November 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tiotsop:2024:MID

[TSP^{+}24] Lohic Fotio Tiotsop, Antonio Servetti, Peter Pocta, Glenn Van Wallendael, Marcus Barkowsky, and Enrico Masala. Multiple image distortion DNN modeling individual subject quality assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (8):255:1–255:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3664198.

Tai:2023:MAS

[TSZ^{+}23] Yichun Tai, Hailin Shi, Dan Zeng, Hang Du, Yibo Hu, Zicheng Zhang, Zhijiang Zhang, and Tao Mei. Multi-agent semi-Siamese training for longtail and shallow face learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):196:1–196:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3594669.

Tian:2012:STL

[TTR12] Xinmei Tian, Dacheng Tao, and Yong Rui. Sparse transfer learning for interactive video search reranking. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3):26:1–26:??, July 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Turk:2013:TYE

[Tur13] Matthew Turk. Over twenty years of eigenfaces. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):45:1– 45:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Truong:2007:VAS

[TV07] Ba Tu Truong and Svetha Venkatesh. Video abstraction: a systematic review and classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(1):??, February 2007. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tiwari:2018:MMS

[TVK18] Akanksha Tiwari, Christian Von Der Weth, and Mohan S. Kankanhalli. Multimodal multiplatform social media event summarization. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):38:1–38:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Trabelsi:2019:UDS

[TVZ^{+}19] Rim Trabelsi, Jagannadan Varadarajan, Le Zhang, Issam Jabri, Yong Pei, Fethi Smach, Ammar Bouallegue, and Pierre Moulin. Understanding the dynamics of social interactions: a multi-modal multi-view approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):15:1– 15:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3300937.

Tan:2023:IVV

[TWB^{+}23] Chong Hong Tan, Koksheik Wong, Vishnu Monn Baskaran, Kiki Adhinugraha, and David Taniar. Is it violin or Viola? Classifying the instruments’ music pieces using descriptive statistics. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):93:1–93:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3563218.

Tan:2024:TBR

[TWFW24] Mingkui Tan, Zhiquan Wen, Leyuan Fang, and Qi Wu. Transformer-based relational inference network for complex visual relational reasoning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (1):10:1–10:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3605781.

Tian:2021:PIC

[TWKK21] Tao Tian, Hanli Wang, Sam Kwong, and C.-C. Jay Kuo. Perceptual image compression with block-level just noticeable difference prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):126:1– 126:15, January 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3408320.

Tang:2019:RVL

[TWL19] Pengjie Tang, Hanli Wang, and Qinyu Li. Rich visual and language representation with complementary semantics for video captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (2):31:1–31:??, June 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3303083.

Talhaoui:2025:VDI

[TWM^{+}25] Mohamed Zakariya Talhaoui, Zhelong Wang, Mohamed Amine Midoun, Abdelkarim Smaili, Djamel Eddine Mekkaoui, Mourad Lablack, and Ke Zhang. Vulnerability detection and improvements of an image cryptosystem for real-time visual protection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):75:1–75:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3708546.

Tan:2024:CLS

[TWY24] Jiawei Tan, Hongxing Wang, and Junsong Yuan. Characters link shots: Character attention network for movie scene segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):94:1–94:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3630257.

Tang:2025:GDC

[TXW^{+}25] Shengeng Tang, Feng Xue, Jingjing Wu, Shuo Wang, and Richang Hong. Gloss-driven conditional diffusion models for sign language production. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):105:1– 105:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tang:2024:MLF

[TXWK24] Zhangyong Tang, Tianyang Xu, Xiao-Jun Wu, and Josef Kittler. Multi-level fusion for robust RGBT tracking via enhanced thermal representation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):312:1–312:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3678176.

Tang:2023:QGP

[TY23] Yiming Tang and Yi Yu.

Query-guided prototype learning with decoder alignment and dynamic fusion in few-shot segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):84:1–84:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3555314.

Tang:2025:FEM

[TYL^{+}25] Long Tang, Dengpan Ye, Zhenhao Lu, Yunming Zhang, and Chuanxi Chen. Feature extraction matters more: an effective and efficient universal deepfake disruptor. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):46:1–46:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3653457.

Tan:2018:UCD

[TYY^{+}18] Min Tan, Jun Yu, Zhou Yu, Fei Gao, Yong Rui, and Dacheng Tao. User-click-data-based finegrained image recognition via weakly supervised metric learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):70:1–70:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Tan:2022:FGI

[TYY^{+}22] Min Tan, Fu Yuan, Jun Yu, Guijun Wang, and Xiaoling Gu. Fine-grained image classification via multi-scale selective hierarchical biquadratic pooling. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1s):31:1–31:23, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3492221.

Tang:2024:HED

[TZ24] Lv Tang and Xinfeng Zhang.

High efficiency deep-learning based video compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):244:1– 244:??, August 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3661311.

Tong:2021:IPP

[TZLZ21] Chao Tong, Mengze Zhang, Chao Lang, and Zhigao Zheng. An image privacy protection algorithm based on adversarial perturbation generative networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):43:1–43:14, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3381088.

Tang:2025:UUD

[TZZ25] Lv Tang, Xinfeng Zhang, and Li Zhang. UVC: a unified deep video compression framework. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (3):100:1–100:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3715144.

UlFazal:2021:EIC

[UFJ21] Muhammad Abu Ul Fazal, Sam Ferguson, and Andrew Johnston. Evaluation of information comprehension in concurrent speech-based designs. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):129:1– 129:19, January 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3409463.

Uddin:2022:NMM

[UJLS22] Md Azher Uddin, Joolekha Bibi Joolee, Young-Koo Lee, and Kyung-Ah Sohn. A novel multimodal network-based dynamic scene understanding. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):7:1– 7:19, January 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3462218.

Urruty:2007:DEF

[ULIS07] Thierry Urruty, Stanislas Lew, Nacim Ihadaddene, and Dan A. Simovici. Detecting eye fixations by projection clustering. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4): 5:1–5:20, December 2007. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

UlHaq:2024:MNA

[UMM24] Ijaz Ul Haq, Khalid Mahmood Malik, and Khan Muhammad.

Multimodal neurosymbolic approach for explainable deepfake detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):341:1–341:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624748.

Uddin:2025:ICA

[USM^{+}25] Shah Muhammad Imtiyaj Uddin, Rashedul Islam Sumon, Md Ariful Islam Mozumder, Md Kamran Hussin Chowdhury, Tagne Poupi Theodore Armand, and Hee Cheol Kim. Innovations and challenges of AI in film: a methodological framework for future exploration. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):199:1–199:55, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ursu:2008:ITN

[UTK^{+}08] Marian F. Ursu, Maureen Thomas, Ian Kegel, Doug Williams, Mika Tuomola, Inger Lindstedt, Terence Wright, Andra Leurdijk, Vilmos Zsombori, Julia Sussner, Ulf Myrestam, and Nina Hall. Interactive TV narratives: Opportunities, progress, and challenges. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(4): 25:1–25:??, October 2008. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Venkatagiri:2015:ALG

[VCO15] Seshadri Padmanabha Venkatagiri, Mun Choon Chan, and Wei Tsang Ooi. Automated link generation for sensorenriched Smartphone images. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (1s):13:1–13:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Vaz:2025:GPS

[VF25] Bruno Vaz and Álvaro Figueira. GANs in the panorama of synthetic data generation methods. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):3:1–3:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3657294.

Vu:2010:UOC

[VGNL10] Long Vu, Indranil Gupta, Klara Nahrstedt, and Jin Liang. Understanding overlay characteristics of a large-scale peerto-peer IPTV system. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(4):31:1– 31:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Vahdati:2025:MRW

[VLE25] Monireh (Monica) Vahdati, Fedwa Laamarti, and Abdulmotaleb El Saddik. Metareview of wearable devices for healthcare in the metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (7):203:1–203:36, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Valdes:2012:AEV

[VM12] Victor Valdes and Jose M. Martinez. Automatic evaluation of video summaries. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3):25:1–25:??, July 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Vellingiri:2020:SCB

[VMP20] Shanthi Vellingiri, Ryan P.

McMahan, and Balakrishnan Prabhakaran. SCeVE: a component-based framework to author mixed reality tours. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2):40:1–40:23, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3377353.

Verdugo:2011:IFC

[VNC^{+}11] Renato Verdugo, Miguel Nussbaum, Pablo Corro, Pablo Nuñnez, and Paula Navarrete. Interactive films and coconstruction. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(4):39:1–39:??, November 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Vyas:2024:ISI

[VNDB24] Ritesh Vyas, Michele Nappi, Alberto Del Bimbo, and Sambit Bakshi. Introduction to special issue on “Recent Trends in Multimedia Forensics”. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):330:1– 330:??, November 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3678473.

Villanueva:2013:HMB

[VPSS^{+}13] Arantxa Villanueva, Victoria Ponz, Laura Sesma-Sanchez, Mikel Ariz, Sonia Porta, and Rafael Cabeza. Hybrid method based on topography for robust detection of iris center and eye corners. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4):25:1–25:??, August 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

vanRensburg:2023:OWD

[vRPPP23] Bianca Jansen van Rensburg, Pauline Puteaux, William Puech, and Jean-Pierre Pedeboy. 3D object watermarking from data hiding in the homomorphic encrypted domain. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):175:1–175:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3588573.

Damme:2022:MLB

[VTD22] Sam Van Damme, Maria Torres Vega, and Filip De Turck. Machine learning based contentagnostic viewport prediction for 360-degree video. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):50:1– 50:24, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474833.

VanderHooft:2020:TBA

[VTP^{+}20] Jeroen Van der Hooft, Maria Torres Vega, Stefano Petrangeli, Tim Wauters, and Filip De Turck. Tile-based adaptive streaming for virtual reality video. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–24, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3362101.

VanLeuken:2011:SVO

[VV11] Reinier H. Van Leuken and Remco C. Veltkamp. Selecting vantage objects for similarity indexing. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(3):16:1–16:??, August 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Vandecasteele:2017:SSC

[VVSV17] Florian Vandecasteele, Karel Vandenbroucke, Dimitri Schuurman, and Steven Verstockt. Spott: On-the-spot ecommerce for television using deep learning-based video analysis techniques. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):38:1–38:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2025:WMA

[WAC^{+}25] Hao Wu, Maha Abdallah, Yuanfang Chi, Lehao Lin, and Wei Cai. Web3 multimedia applications: Under the impact of decentralization. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):191:1–191:38, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2018:ELV

[WAD^{+}18] Junjue Wang, Brandon Amos, Anupam Das, Padmanabhan Pillai, Norman Sadeh, and Mahadev Satyanarayanan. Enabling live video analytics with a scalable and privacy-aware framework. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):64:1–64:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2012:CPA

[WAK^{+}12] Wanmin Wu, Ahsan Arefin, Gregorij Kurillo, Pooja Agarwal, Klara Nahrstedt, and Ruzena Bajcsy. CZLoD: a psychophysical approach for 3D tele-immersive video. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3s):39:1– 39:??, September 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2021:SDM

[Wan21] Yang Wang. Survey on deep multi-modal data analytics: Collaboration, rivalry, and fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):10:1–10:25, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3408317.

Wu:2016:ERM

[WB16] Qiong Wu and Pierre Boulanger.

Enhanced reweighted MRFs for efficient fashion image parsing.

ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (3):42:1–42:??, June 2016. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wisniewski:2017:OAA

[WBB^{+}17] Piotr Wisniewski, Jordi Mongay Batalla, Andrzej Beben, Piotr Krawiec, and Andrzej Chydzinski. On optimizing adaptive algorithms based on rebuffering probability. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):43:1– 43:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2023:BHI

[WBH^{+}23] Shuo Wang, Huixia Ben, Yanbin Hao, Xiangnan He, and Meng Wang. Boosting hyperspectral image classification with dual hierarchical learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1):21:1–21:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3522713.

Wei:2009:CCM

[WBL09] Yong Wei, Suchendra M. Bhandarkar, and Kang Li. Clientcentered multimedia content adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3):22:1–22:??, August 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Willis:2025:EHC

[WBR25] Nina Willis, Abraham Bernstein, and Luca Rossetto. Effects of human cognitioninspired task presentation on interactive video retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):159:1–159:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2017:DAQ

[WBRZ17] Cong Wang, Divyashri Bhat, Amr Rizk, and Michael Zink. Design and analysis of QoEaware quality adaptation for DASH: a spectrum-based approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):45:1–45:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Weng:2012:CVR

[WC12] Ming-Fang Weng and Yung-Yu Chuang. Collaborative video reindexing via matrix factorization. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2):23:1–23:??, May 2012. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wei:2023:MLC

[WC23] Hao Wei and Rui Chen. A multi-level consistency network for high-fidelity virtual tryon. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):156:1–156:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580500.

Wang:2023:DCP

[WCCN23] Tianyi Wang, Harry Cheng, Kam Pui Chow, and Liqiang Nie. Deep convolutional pooling transformer for deepfake detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):179:1–179:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3588574.

Wang:2017:MCL

[WCF^{+}17] Shuai Wang, Yang Cong, Huijie Fan, Baojie Fan, Lianqing Liu, Yunsheng Yang, Yandong Tang, Huaici Zhao, and Haibin Yu. Multi-class latent concept pooling for computer-aided endoscopy diagnosis. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(2):15:1–15:??, May 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:DDT

[WCG^{+}25] Kun Wang, Jiuxin Cao, Jiawei Ge, Chang Liu, and Bo Liu. Dual-domain triple contrast for cross-dataset skeleton-based action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):115:1– 115:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2005:AFE

[WCK05] Huahui Wu, Mark Claypool, and Robert Kinicki. Adjusting forward error correction with temporal scaling for TCP-friendly streaming MPEG. ACM Transactions on Multimedia Computing, Communications, and Applications, 1(4):315–337, November 2005. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2022:BRJ

[WCL^{+}22] Xiaoqin Wang, Chen Chen, Rushi Lan, Licheng Liu, Zhenbing Liu, Huiyu Zhou, and Xiaonan Luo. Binary representation via jointly personalized sparse hashing. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):137:1–137:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3558769.

Wang:2023:DCB

[WCL23] Jinxia Wang, Rui Chen, and Zhihan Lv. DNA computingbased multi-source data storage model in digital twins. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3s):127:1–127:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3561823.

Wang:2018:SAM

[WCLC18] Anran Wang, Jianfei Cai, Jiwen Lu, and Tat-Jen Cham. Structure-aware multimodal feature fusion for RGB-D scene classification and beyond. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):39:1– 39:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2015:WMZ

[WCO15] Hui Wang, Mun Choon Chan, and Wei Tsang Ooi. Wireless multicast for zoomable video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1):5:1–5:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2014:BCM

[WCX^{+}14] Zhiyu Wang, Peng Cui, Lexing Xie, Wenwu Zhu, Yong Rui, and Shiqiang Yang. Bilateral correspondence model for words-and-pictures association in multimedia-rich microblogs. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (4):34:1–34:??, June 2014. ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2018:DAQ

[WCY^{+}18] Jiyan Wu, Bo Cheng, Yuan Yang, Ming Wang, and Junliang Chen. Delay-aware quality optimization in cloudassisted video streaming system. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):4:1–4:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2021:LDS

[WCZ^{+}21] Yang Wang, Yang Cao, Jing Zhang, Feng Wu, and Zheng-Jun Zha. Leveraging deep statistics for underwater image enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):116:1–116:20, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3489520.

Wang:2025:DAS

[WDC^{+}25] Yongxin Wang, Feng Dong, Zhen-Duo Chen, Xin Luo, and Xin-Shun Xu. Domain-aware semantic alignment hashing for large-scale zero-shot image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):171:1–171:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2007:EST

[WDCX07] Surong Wang, Manoranjan Dash, Liang-Tien Chia, and Min Xu. Efficient sampling of training set in large and noisy multimedia data. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(3):14:1–14:??, August 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:IMC

[WDE25] Haopeng Wang, Haiwei Dong, and Abdulmotaleb El Saddik. Immersive multimedia communication: State-of-the-art on extended reality streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):190:1–190:33, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2021:MIH

[WDJ^{+}21] Wenjie Wang, Ling-Yu Duan, Hao Jiang, Peiguang Jing, Xuemeng Song, and Liqiang Nie. Market2Dish: Health-aware food recommendation. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):33:1– 33:19, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3418211.

Wu:2023:DUD

[WDLW23] Dayan Wu, Qi Dai, Bo Li, and Weiping Wang. Deep uncoupled discrete hashing via similarity matrix decomposition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1):22:1–22:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3524021.

Wang:2023:CSA

[WDPX23] Kan Wang, Changxing Ding, Jianxin Pang, and Xiangmin Xu. Context sensing attention network for video-based person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):143:1–143:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3573203.

Wang:2024:CLK

[WDR^{+}24] Fei Wang, Liang Ding, Jun Rao, Ye Liu, Li Shen, and Changxing Ding. Can linguistic knowledge improve multimodal alignment in vision-language pretraining? ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):364:1–364:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3690640.

Wu:2021:SPP

[WDS21] Hongjiao Wu, Ashutosh Dhar Dwivedi, and Gautam Srivastava. Security and privacy of patient information in medical systems based on blockchain technology. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):60:1–60:17, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3408321.

Wang:2023:GAE

[WDX^{+}23] Yang Wang, Bo Dong, Ke Xu, Haiyin Piao, Yufei Ding, Baocai Yin, and Xin Yang. A geometrical approach to evaluate the adversarial robustness of deep neural networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):172:1– 172:??, October 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3587936.

Westerski:2025:SDO

[WF25] Adam Westerski and Wee Teck Fong. Synthetic data for object detection with neural networks: State-of-the-art survey of domain randomisation techniques. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):2:1–2:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3637064.

Wang:2025:TCT

[WFC^{+}25a] Yifan Wang, Liang Feng, Fenglin Cai, Lusi Li, Rui Wu, and Jie Li. TEC-CNN: Toward efficient compressing of convolutional neural nets with low-rank tensor decomposition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):60:1–60:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3702641.

Wu:2025:EEV

[WFC^{+}25b] Xiaotian Wu, Xinjie Feng, Bing Chen, Ching-Nung Yang, Qing-Yu Peng, and Weiqi Yan. EVCS-DAS: Evolving visual cryptography schemes for dynamic access structures. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):74:1– 74:??, March 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3708547.

Wang:2021:IBM

[WFZ^{+}21a] Yang Wang, Meng Fang, Joey Tianyi Zhou, Tingting Mu, and Dacheng Tao. Introduction to big multimodal multimedia data with deep analytics. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1s):1:1–1:3, March 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3447530.

Wang:2021:ISI

[WFZ^{+}21b] Yang Wang, Meng Fang, Joey Tianyi Zhou, Tingting Mu, and Dacheng Tao. Introduction to the special issue on fine-grained visual computing. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):11:1–11:3, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3447532.

Wang:2025:SDC

[WFZ25] Fan Wang, Zhangjie Fu, and Xiang Zhang. A self-defense copyright protection scheme for NFT image art based on information embedding. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):42:1– 42:??, February 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3652519.

Wu:2024:LCA

[WGHX24] Ziyue Wu, Junyu Gao, Shucheng Huang, and Changsheng Xu. Learning commonsense-aware moment-text alignment for fast video temporal grounding. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):318:1–318:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663368.

Wang:2023:LFR

[WGL^{+}23] Zhengxue Wang, Guangwei Gao, Juncheng Li, Hui Yan, Hao Zheng, and Huimin Lu. Lightweight feature deredundancy and self-calibration network for efficient image super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):110:1–110:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569900.

Wu:2025:PGE

[WGM25] Ji-Yan Wu, Kasun Gamlath, and Archan Misra. Pr-Ge-Ne: Efficient encoding of pervasive video sensing streams by pruned generative networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):68:1–68:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3706109.

Wu:2022:ALC

[WH22] Da-Chun Wu and Yu-Tsung Hsu. Authentication of LINE chat history files by information hiding. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):22:1–22:23, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3474225.

Wang:2018:DFI

[WHF^{+}18] Peisong Wang, Qinghao Hu, Zhiwei Fang, Chaoyang Zhao, and Jian Cheng. DeepSearch: a fast image search framework for mobile devices. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):6:1–6:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Whitman:2013:CSF

[Whi13] Brian Whitman. Care and scale: Fifteen years of music retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):46:1–46:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2020:PSE

[WHJ20] Shangfei Wang, Longfei Hao, and Qiang Ji. Posed and spontaneous expression distinction using latent regression Bayesian networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):80:1–80:18, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3391290.

Wang:2021:DNA

[WHL^{+}21] Jinwei Wang, Wei Huang, Xiangyang Luo, Yun-Qing Shi, and Sunil Kr. Jha. Detecting non-aligned double JPEG compression based on amplitudeangle feature. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):138:1–138:18, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3464388.

Wang:2023:JJD

[WHL^{+}23] Haidong Wang, Xuan He, Zhiyong Li, Jin Yuan, and Shutao Li. JDAN: Joint detection and association network for real-time online multiobject tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):45:1–45:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3533253.

Wu:2018:ICS

[WHW18] Jie Wu, Haifeng Hu, and Yi Wu. Image captioning via semantic guidance attention and consensus selection strategy. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (4):87:1–87:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2021:DSG

[WHW^{+}21] Junyi Wu, Yan Huang, Qiang Wu, Zhipeng Gao, Jianqiang Zhao, and Liqin Huang. Dualstream guided-learning via a priori optimization for person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):117:1–117:22, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447715.

Wang:2025:CAL

[WHX^{+}25] Miaohui Wang, Runnan Huang, Wuyuan Xie, Zhan Ma, and Siwei Ma. Compression approaches for LiDAR point clouds and beyond: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):188:1–188:31, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2018:ICA

[WHY18] Anqi Wang, Haifeng Hu, and Liang Yang. Image captioning with affective guiding and selective attention. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3):73:1–73:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2019:PAT

[WHY19] Jie Wu, Haifeng Hu, and Liang Yang. Pseudo-3D attention transfer network with content-aware strategy for image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):80:1–80:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3336495.

Wen:2024:VSI

[WHZ^{+}24] Wenying Wen, Minghui Huang, Yushu Zhang, Yuming Fang, and Yifan Zuo. Visual security index combining CNN and filter for perceptually encrypted light field images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):25:1–25:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3612924.

Wang:2015:INB

[WJ15a] Shuang Wang and Shuqiang Jiang. INSTRE: a new benchmark for instance-level object retrieval and recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(3):37:1– 37:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2015:CAM

[WJ15b] Ming-Ju Wu and Jyh-Shing R. Jang. Combining acoustic and multilevel visual features for music genre classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (1):10:1–10:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2022:IFD

[WJQ^{+}22a] Jingjing Wu, Jianguo Jiang, Meibin Qi, Cuiqun Chen, and Yimin Liu. Improving feature discrimination for object tracking by structural-similarity-based metric learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):90:1– 90:23, November 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3497746.

Wu:2022:EEH

[WJQ^{+}22b] Jingjing Wu, Jianguo Jiang, Meibin Qi, Cuiqun Chen, and Jingjing Zhang. An end-to-end heterogeneous restraint network for RGB-D cross-modal person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):109:1– 109:22, November 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3506708.

Wang:2021:OBR

[WJS^{+}21] Ting Wang, Xiangjun Ji, Aiguo Song, Kurosh Madani, Amine Chohra, Huimin Lu, and Ramon Monero. Output-bounded and RBFNN-based position tracking and adaptive force control for security tele-surgery. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):61:1–61:15, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3394920.

Wan:2024:ESI

[WJXN24] Shaohua Wan, Yi Jin, Guangdong Xu, and Michele Nappi. Editorial to special issue on multimedia cognitive computing for intelligent transportation system. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):32:1–32:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3604938.

Wang:2010:MBA

[WK10] Xiangyu Wang and Mohan Kankanhalli. MultiFusion: a boosting approach for multimedia fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(4):25:1–25:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wilk:2016:CAV

[WKE16] Stefan Wilk, Stephan Kopf, and Wolfgang Effelsberg. Collaborative annotation of videos relying on weak consistency. ACM Transactions on Multimedia Computing, Communications, and Applications, 12 (3):45:1–45:??, June 2016. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:TKP

[WKN25] Jiaqi Wang, Ricky Y.-K. Kwok, and Edith C. H. Ngai. Towards key point identification (KPI) for lecture videos: Approaches and performance evaluation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):201:1–201:23, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2023:REC

[WKS^{+}23] Jia Wang, Jingcheng Ke, Hong-Han Shuai, Yung-Hui Li, and Wen-Huang Cheng. Referring expression comprehension via enhanced cross-modal graph attention networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):65:1–65:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3548688.

Wang:2008:MST

[WKST08] Bing Wang, Jim Kurose, Prashant Shenoy, and Don Towsley. Multimedia streaming via TCP: an analytic performance study. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2):16:1–16:??, May 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:IFG

[WLB^{+}25] Shuo Wang, Jinda Lu, Huixia Ben, Yanbin Hao, Xingyu Gao, and Meng Wang. Interventional feature generation for few-shot learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):165:1–165:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2020:LLF

[WLC^{+}20] Xiao Wang, Wu Liu, Jun Chen, Xiaobo Wang, Chenggang Yan, and Tao Mei. Listen, look, and find the one: Robust person search with multimodality index. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):47:1–47:20, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380549.

Wang:2021:SCG

[WLCG21] Rui Wang, Dong Liang, Xiaochun Cao, and Yuanfang Guo. Semantic correspondence with geometric structure analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):83:1–83:21, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3441576.

Wu:2022:LTP

[WLCH22] Hanjie Wu, Yongtuo Liu, Hongmin Cai, and Shengfeng He. Learning transferable perturbations for image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2):57:1–57:18, May 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3478024.

Wang:2023:SSL

[WLFL23] Wenjing Wang, Lilang Lin, Zejia Fan, and Jiaying Liu. Semisupervised learning for Mars imagery classification and segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):144:1–144:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572916.

Wei:2021:ISS

[WLH^{+}21] Haiyang Wei, Zhixin Li, Feicheng Huang, Canlong Zhang, Huifang Ma, and Zhongzhi Shi. Integrating scene semantic knowledge into image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):52:1–52:22, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3439734.

Wang:2025:LLS

[WLH^{+}25] Jiapeng Wang, Zening Lin, Dayi Huang, Longfei Xiong, and Lianwen Jin. LiLTv2: Language-substitutable layoutimage transformer for visual information extraction. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):72:1– 72:??, March 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3708351.

Wang:2013:RAA

[WLHL13] Yichuan Wang, Ting-An Lin, Cheng-Hsin Hsu, and Xin Liu. Region- and action-aware virtual world clients. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1):6:1–6:??, February 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2019:EBD

[WLHT19] Zhangcheng Wang, Ya Li, Richang Hong, and Xinmei Tian. Eigenvector-based distance metric learning for image classification and retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):84:1–84:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3340262.

Wu:2019:HVO

[WLL^{+}19] Kan Wu, Guanbin Li, Haofeng Li, Jianjun Zhang, and Yizhou Yu. Harvesting visual objects from Internet images via deep-learning-based objectness assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):72:1– 72:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3318463.

Wang:2023:RVS

[WLL23] Jian Wang, Qiang Ling, and Peiyan Li. Robust video stabilization based on motion decomposition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):164:1–164:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580498.

Wu:2024:KVK

[WLL^{+}24] Zizhao Wu, Siyu Liu, Peioyan Lu, Ping Yang, Yongkang Wong, Xiaoling Gu, and Mohan S. Kankanhalli. KF-VTON: Keypoints-driven flow based virtual try-on network. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):293:1–293:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673903.

Wang:2025:BTA

[WLMZ25] Jingtian Wang, Xiaolong Li, Bin Ma, and Yao Zhao. Boosting transferability of adversarial examples with spatiotemporal context. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):297:1–297:22, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2012:ABP

[WLQL12] Chuan Wu, Zongpeng Li, Xuanjia Qiu, and Francis C. M. Lau. Auction-based P2P VoD streaming: Incentives and optimal scheduling. ACM Transactions on Multimedia Computing, Communications, and Applications, 8s(1):14:1–14:??, February 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2024:MMI

[WLS^{+}24a] Han Wang, Hui Li, Abla Smahi, Feng Zhao, Yao Yao, Ching Chuen Chan, Shiyu Wang, Wenyuan Yang, and Shuo-Yen Robert Li. MIS: a multi-identifier management and resolution system in the metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(7):191:1–191:??, July 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597641.

Wang:2024:CMM

[WLS^{+}24b] Yan Wang, Peize Li, Qingyi Si, Hanwen Zhang, Wenyu Zang, Zheng Lin, and Peng Fu. Cross-modality multiple relations learning for knowledgebased visual question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):63:1–63:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3618301.

Wang:2024:HRS

[WLSN24] Yuanyuan Wang, Meng Liu, Xuemeng Song, and Liqiang Nie. Harnessing representative spatial-temporal information for video question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):306:1–306:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3675399.

Wang:2024:WGO

[WLSX24] Jie Wang, Guoqiang Li, Jie Shi, and Jinwen Xi. Weighted guided optional fusion network for RGB-T salient object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):136:1–136:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3624984.

Wu:2022:ICG

[WLSZ22] Lei Wu, Hefei Ling, Yuxuan Shi, and Baiyan Zhang. Instance correlation graph for unsupervised domain adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1s):33:1–33:23, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3486251.

Wang:2023:ISR

[WLTL23] Li Wang, Ke Li, Jingjing Tang, and Yuying Liang. Image super-resolution via lightweight attention-directed feature aggregation network. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):60:1– 60:??, March 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3546076.

Wu:2023:RIE

[WLW^{+}23] Hua Wu, Xin Li, Gang Wang, Guang Cheng, and Xiaoyan Hu. Resolution identification of encrypted video streaming based on HTTP/2 features. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (2):73:1–73:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3551891.

Wu:2008:ELS

[WLZ08] Chuan Wu, Baochun Li, and Shuqiao Zhao. Exploring largescale peer-to-peer live streaming topologies. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(3):19:1–19:??, August 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2012:DNW

[WLZ12] Chuan Wu, Baochun Li, and Shuqiao Zhao. Diagnosing network-wide P2P live streaming inefficiencies. ACM Transactions on Multimedia Computing, Communications, and Applications, 8s(1):13:1–13:??, February 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2024:ETC

[WLZ^{+}24a] Yuanzhi Wang, Yong Li, Xiaoya Zhang, Xin Liu, Anbo Dai, Antoni B. Chan, and Zhen Cui. Edit temporalconsistent videos with image diffusion model. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):368:1–368:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3691344.

Wu:2024:LAC

[WLZ^{+}24b] Xu Wu, Zhihui Lai, Jie Zhou, Xianxu Hou, Witold Pedrycz, and Linlin Shen. Light-aware contrastive learning for lowlight image enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):276:1– 276:??, September 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3665498.

Wang:2022:MGB

[WLZF22] Quan Wang, Sheng Li, Xinpeng Zhang, and Guorui Feng. Multi-granularity brushstrokes network for universal style transfer. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):107:1–107:17, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3506710.

Wang:2022:LLS

[WMH^{+}22] Jing Wang, Weiqing Min, Sujuan Hou, Shengnan Ma, Yuanjie Zheng, and Shuqiang Jiang. LogoDet-3K: a large-scale image dataset for logo detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1):21:1–21:19, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3466780.

Wang:2023:AAM

[WMM^{+}23] Jingyao Wang, Luntian Mou, Lei Ma, Tiejun Huang, and Wen Gao. AMSA: Adaptive multimodal learning for sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):135:1–135:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572915.

Wang:2022:EEC

[WMW^{+}22] Yabin Wang, Zhiheng Ma, Xing Wei, Shuai Zheng, Yaowei Wang, and Xiaopeng Hong. ECCNAS: Efficient crowd counting neural architecture search. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):36:1–36:19, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3465455.

Wu:2018:TLD

[WPRC18] Hui-Yin Wu, Francesca Palù, Roberto Ranon, and Marc Christie. Thinking like a director: Film editing patterns for virtual cinematographic storytelling. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(4):81:1–81:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wen:2018:CEE

[WQL18] Longyin Wen, Honggang Qi, and Siwei Lyu. Contrast enhancement estimation for digital image forensics. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2):49:1– 49:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:MVP

[WQTX25] Weiyu Wang, Chunmei Qing, Junpeng Tan, and Xiangmin Xu. Multi-view panoramic image style transfer with multiscale attention and global sharing. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):208:1–208:19, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2014:FEM

[WRK14] Xiangyu Wang, Yong Rui, and Mohan Kankanhalli. Up-Fusion: an evolving multimedia fusion method. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1):6:1–6:??, August 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2024:LGR

[WSLC24] Jia Wang, Hong-Han Shuai, Yung-Hui Li, and Wen-Huang Cheng. Language-guided residual graph attention network and data augmentation for visual grounding. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):7:1–7:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3604557.

Wen:2018:VBR

[WSLM18] Jiqing Wen, James She, Xiaopeng Li, and Hui Mao. Visual background recommendation for dance performances using deep matrix factorization. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (1):11:1–11:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:AVA

[WSLZ25] Yu-Ao Wang, James She, Troy TianYu Lin, and Kang Zhang. AI visual art history: an art movement with expanded artistic horizon. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):185:1–185:16, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2024:GBC

[WSW^{+}24] Lilong Wang, Yunhui Shi, Jin Wang, Shujun Chen, Baocai Yin, and Nam Ling. Graph based cross-channel transform for color image compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):102:1–102:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3631710.

Wu:2023:APA

[WSX^{+}23] Shangxi Wu, Jitao Sang, Kaiyuan Xu, Jiaming Zhang, and Jian Yu. Attention, please! Adversarial defense via activation rectification and preservation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):142:1–142:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3572843.

Wu:2024:AAL

[WSX^{+}24] Shangxi Wu, Jitao Sang, Kaiyan Xu, Guanhua Zheng, and Changsheng Xu. Adaptive adversarial logits pairing. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (2):56:1–56:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3616375.

Wu:2010:ELT

[WT10] Junwen Wu and Mohan M. Trivedi. An eye localization, tracking and blink pattern recognition system: Algorithm and evaluation. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(2):8:1–8:??, March 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2021:UDE

[WTD^{+}21] Jie Wang, Kaibin Tian, Dayong Ding, Gang Yang, and Xirong Li. Unsupervised domain expansion for visual categorization. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):121:1–121:24, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3448108.

Wang:2024:HPV

[WTL24] Hongbin Wang, Rui Tang, and Fan Li. Hypercube pooling for visual semantic embedding. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (12):363:1–363:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3689637.

Wang:2025:SFD

[WTL^{+}25] Xin Wang, Ting Yu Tsai, Li Lin, Hui Guo, Shu Hu, Ming-Ching Chang, Pradeep K. Atrey, and Siwei Lyu. Spotting the fakes: a deep dive into GAN-generated face detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):193:1–193:24, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2020:IMP

[WTZ^{+}20] Xun Wang, Yan Tian, Xuran Zhao, Tao Yang, Judith Gelernter, Jialei Wang, Guohua Cheng, and Wei Hu. Improving multiperson pose estimation by mask-aware deep reinforcement learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):84:1–84:18, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3397340.

Wang:2024:MTA

[WWC^{+}24] Hanrui Wang, Shuo Wang, Cunjian Chen, Massimo Tistarelli, and Zhe Jin. A multi-task adversarial attack against face authentication. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):332:1–332:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665496.

Wang:2009:MLS

[WWGT09] Bing Wang, Wei Wei, Zheng Guo, and Don Towsley. Multipath live streaming via TCP: Scheme, performance and benefits. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(3):25:1–25:??, August 2009. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2017:IRS

[WWH17] Haojun Wu, Yong Wang, and Jiwu Huang. Identification of reconstructed speech. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):10:1– 10:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2014:EIP

[WWHW14] Xinxi Wang, Yi Wang, David Hsu, and Ye Wang. Exploration in interactive personalized music recommendation: a reinforcement learning approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (1):7:1–7:??, August 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2013:ECR

[WWL13] Bo Wang, Jinqiao Wang, and Hanqing Lu. Exploiting content relevance and social relevance for personalized ad recommendation on Internet TV. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(4): 26:1–26:??, August 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2020:UNC

[WWL20] Xueping Wang, Yunhong Wang, and Weixin Li. U-Net conditional GANs for photo-realistic and identity-preserving facial expression synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1– 23, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3355397.

Wu:2023:CPG

[WWL^{+}23] Huisi Wu, Zhaoze Wang, Zhuoying Li, Zhenkun Wen, and Jing Qin. Context prior guided semantic modeling for biomedical image segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):90:1–90:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3558520.

Weng:2024:HHC

[WWL^{+}24] Zejia Weng, Zuxuan Wu, Hengduo Li, Jingjing Chen, and Yu-Gang Jiang. HCMS: Hierarchical and conditional modality selection for efficient video recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):35:1–35:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572776.

Wang:2025:MSD

[WWQ^{+}25a] Shen Wang, Yu Wang, Renjie Qiao, Kejun Wu, Chia-Wen Lin, and Chengtao Cai. Multi-scale dynamic fusion for visible-infrared person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):101:1–101:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3715330.

Wen:2025:PBI

[WWQ^{+}25b] Chao Wen, Chen Wei, Yuhua Qian, Xiaodan Song, and Xuemei Xie. Prompt-based invertible mapping alignment for unsupervised domain adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):134:1–134:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2024:AGM

[WWS^{+}24] Ruimin Wang, Fasheng Wang, Yiming Su, Jing Sun, Fuming Sun, and Haojie Li. Attentionguided multi-modality interaction network for RGB-D salient object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):68:1–68:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3624747.

Wang:2022:WYE

[WWW^{+}22] Jiahao Wang, Yunhong Wang, Nina Weng, Tianrui Chai, Annan Li, Faxi Zhang, and Samsi Yu. Will you ever become popular? Learning to predict virality of dance clips. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):54:1– 54:24, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3477533.

Wu:2023:DGD

[WWWL23] Xintian Wu, Huanyu Wang, Yiming Wu, and Xi Li. D^{3}T-GAN: Data-dependent domain transfer GANs for image generation with limited data. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (4):146:1–146:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3576858.

Wei:2021:CNL

[WWX^{+}21] Yang Wei, Zhuzhu Wang, Bin Xiao, Ximeng Liu, Zheng Yan, and Jianfeng Ma. Controlling neural learning network with multiple scales for image splicing forgery detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):124:1– 124:22, January 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3408299.

Wu:2021:BBD

[WWY^{+}21] Zhenyu Wu, Zhaowen Wang, Ye Yuan, Jianming Zhang, Zhangyang Wang, and Hailin Jin. Black-box diagnosis and calibration on GAN intra-mode collapse: a pilot study. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):114:1– 114:18, October 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3472768.

Wang:2024:ACN

[WWY^{+}24] Haoran Wang, Yajie Wang, Baosheng Yu, Yibing Zhan, Chunfeng Yuan, and Wankou Yang. Attentional composition networks for long-tailed human action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):8:1– 8:??, January 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3603253.

Wang:2025:FFC

[WWY^{+}25] Hanzhang Wang, Haoran Wang, Zhongrui Yu, Mingming Sun, Junjun Jiang, Xianming Liu, and Deming Zhai. FAST: Flexibly controllable arbitrary style transfer via latent diffusion models. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):268:1–268:20, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2024:TRH

[WWZ24] Jian Wang, Xiao Wang, and Guosheng Zhao. Task recommendation via heterogeneous multi-modal features and decision fusion in mobile crowdsensing. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):78:1–78:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3626239.

Wang:2025:IAW

[WWZ^{+}25] Jinwei Wang, Haihua Wang, Jiawei Zhang, Hao Wu, Xiangyang Luo, and Bin Ma. Invisible adversarial watermarking: a novel security mechanism for enhancing copyright protection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):43:1–43:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3652608.

Wang:2024:KIM

[WXCW24] Depei Wang, Ruifeng Xu, Lianglun Cheng, and Zhuowei Wang. Knowledge-integrated multi-modal movie turning point identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):138:1– 138:??, May 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3638557.

Wang:2025:CDS

[WXH^{+}25] Yan Wang, Hong Xie, Jinyang He, Xiaoyu Shi, and Mingsheng Shang. Cross-domain semantic transfer for domain generalization. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):149:1–149:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:TFS

[WXJL25] Jiahui Wang, Qin Xu, Bo Jiang, and Bin Luo. Transductive few-shot learning via joint message passing and prototypebased soft-label propagation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (4):129:1–129:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2020:IMC

[WXQC20] Lingxiang Wu, Min Xu, Shengsheng Qian, and Jianwei Cui. Image to modern Chinese poetry creation via a constrained topic-aware model. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2):53:1–53:21, June 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381858.

Wang:2025:PGB

[WXS^{+}25] Jie Wang, Tingfa Xu, Liqiang Song, Lihe Ding, Hui Li, Peng Jiang, Yuqi Han, and Jianan Li. PAPooling: Graph-based position adaptive aggregation of local geometry in point clouds. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (4):130:1–130:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2022:MFF

[WXW^{+}22] Zhoutao Wang, Qian Xie, Mingqiang Wei, Kun Long, and Jun Wang. Multi-feature fusion VoteNet for 3D object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):6:1–6:17, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3462219.

Wang:2023:CSC

[WXZ^{+}23] Meng Wang, Jizheng Xu, Li Zhang, Junru Li, Kai Zhang, Shiqi Wang, and Siwei Ma. Compressed screen content image super resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):209:1– 209:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3589963.

Wu:2023:BBT

[WY23] Xiaotian Wu and Peng Yao.

Boolean-based two-in-one secret image sharing by adaptive pixel grouping. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):14:1–14:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3517140.

Wang:2024:EAE

[WYGC24] Dewang Wang, Gaobo Yang, Zhiqing Guo, and Jiyou Chen. Enhancing adversarial embedding based image steganography via clustering modification directions. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):20:1–20:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603377.

Weir:2012:IHV

[WYK12] Jonathan Weir, Weiqi Yan, and Mohan S. Kankanhalli. Image hatching for visual cryptography. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2S):32:1–32:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2007:UGP

[WYM07] Jun Wang, Lijun Yin, and Jason Moore. Using geometric properties of topographic manifold to detect and track eyes for human-computer interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(4):3:1–3:20, December 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2018:ICD

[WYM18] Cheng Wang, Haojin Yang, and Christoph Meinel. Image captioning with deep bidirectional LSTMs and multi-task learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):40:1–40:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wanyan:2023:DSG

[WYMX23] Yuyang Wanyan, Xiaoshan Yang, Xuan Ma, and Changsheng Xu. Dual scene graph convolutional network for motivation prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):137:1–137:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3572914.

Wen:2025:LGR

[WYY^{+}25] Wenying Wen, Yu Ye, Ziye Yuan, Baolin Qiu, and Dingli Hua. LFIZW-GRHFMR: Robust zero-watermarking with GRHFMR for light field image. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):126:1–126:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2020:DBD

[WZ20] Shui-Hua Wang and Yu-Dong Zhang. DenseNet-201-based deep neural network with composite learning factor and precomputation for multiple sclerosis classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):60:1– 60:19, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3341095.

Wang:2013:PBS

[WZC^{+}13] Zhi Wang, Wenwu Zhu, Xiangwen Chen, Lifeng Sun, Jiangchuan Liu, Minghua Chen, Peng Cui, and Shiqiang Yang. Propagation-based social-aware multimedia content distribution. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):52:1–52:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2020:RFS

[WZD^{+}20] Qingyong Wang, Yun Zhou, Weiping Ding, Zhiguo Zhang, Khan Muhammad, and Zehong Cao. Random forest with self-paced bootstrap learning in lung cancer prognosis. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):34:1–34:12, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3345314.

Wu:2024:ADS

[WZL^{+}24] Jingjing Wu, Xi Zhou, Xiaohong Li, Hao Liu, Meibin Qi, and Richang Hong. Asymmetric deformable spatio-temporal framework for infrared object tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):322:1–322:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3678882.

Wang:2014:HEK

[WZNM14] Feng Wang, Wan-Lei Zhao, Chong-Wah Ngo, and Bernard Merialdo. A Hamming embedding kernel with informative bag-of-visual words for video semantic indexing. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(3):26:1– 26:??, April 2014. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2025:MMD

[WZRD25] Yang Wang, Yixing Zhang, Xudie Ren, and Yuxin Deng. MoDA: Mixture of domain adapters for parameter-efficient generalizable person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (5):139:1–139:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wang:2019:DSS

[WZTL19] Min Wang, Wengang Zhou, Qi Tian, and Houqiang Li. Deep scalable supervised quantization by self-organizing map. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):81:1–81:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3328995.

Wu:2024:SMG

[WZW24] Tian Wu, Rongbo Zhu, and Shaohua Wan. Semantic map guided identity transfer GAN for person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (11):333:1–333:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631355.

Wang:2023:DPS

[WZWY23] Xiaohan Wang, Linchao Zhu, Fei Wu, and Yi Yang. A differentiable parallel sampler for efficient video classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):112:1–112:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3569584.

Wu:2022:WPM

[WZZ^{+}22] Yike Wu, Shiwan Zhao, Ying Zhang, Xiaojie Yuan, and Zhong Su. When pairs meet triplets: Improving lowresource captioning via multiobjective optimization. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):83:1– 83:20, August 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3492325.

Wu:2023:OBM

[WZZ^{+}23] Jun Wu, Tianliang Zhu, Jiahui Zhu, Tianyi Li, and Chunzhi Wang. A optimized BERT for multimodal sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):91:1–91:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3566126.

Wang:2024:MDE

[WZZ^{+}24] Hanzhang Wang, Deming Zhai, Xiong Zhou, Junjun Jiang, and Xianming Liu. Mix-DDPM: Enhancing diffusion models through fitting mixture noise with global stochastic offset. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):283:1–283:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672080.

Wei:2025:MMA

[WZZ25] Lisi Wei, Libo Zhao, and Xiaoli Zhang. MAINet: Modalityaware interaction network for medical image fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):166:1– 166:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Wu:2025:RRE

[WZZL25] Tong Wu, Jinhua Zhu, Wengang Zhou, and Houqiang Li. RESIST: Rationale-enhanced and reward model-based endto-end social influence dialogue system. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):170:1–170:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xu:2006:FAF

[XC06] Huaxin Xu and Tat-Seng Chua. Fusion of AV features and external information sources for event detection in team sports video. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(1):44–67, February 2006. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xie:2023:VPG

[XCC^{+}23] Jiayuan Xie, Jiali Chen, Yi Cai, Qingbao Huang, and Qing Li. Visual paraphrase generation with key information retained. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):184:1–184:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3585010.

Xu:2023:PAC

[XCM^{+}23] Cheng Xu, Zejun Chen, Jiajie Mai, Xuemiao Xu, and Shengfeng He. Pose- and attribute-consistent person image synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):81:1–81:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3554739.

Xu:2021:LIR

[XDL^{+}21] Qianli Xu, Ana Garcia Del Molino, Jie Lin, Fen Fang, Vigneshwaran Subbaraju, Liyuan Li, and Joo-Hwee Lim. Lifelog image retrieval based on semantic relevance mapping. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):92:1– 92:18, August 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3446209.

Xu:2021:TTF

[XFQ^{+}21] Xiaolong Xu, Zijie Fang, Lianyong Qi, Xuyun Zhang, Qiang He, and Xiaokang Zhou. TripRes: Traffic flow prediction driven resource reservation for multimedia IoV with edge computing. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):41:1–41:21, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3401979.

Xing:2020:ICC

[XFSZ20] Meng Xing, Zhiyong Feng, Yong Su, and Jianhai Zhang. An image cues coding approach for 3D human pose estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–20, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3368066.

Xie:2019:CAN

[XFZ^{+}19] Hongtao Xie, Shancheng Fang, Zheng-Jun Zha, Yating Yang, Yan Li, and Yongdong Zhang. Convolutional attention networks for scene text recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):3:1–3:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3231737.

Xie:2024:SGD

[XG24] Yurui Xie and Ling Guan. Sparsity-guided discriminative feature encoding for robust keypoint detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):88:1–88:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3628432.

Xiang:2025:LVS

[XGX^{+}25] Suncheng Xiang, Jingsheng Gao, Mingye Xie, Mengyuan Guan, Jiacheng Ruan, and Yuzhuo Fu. Learning visualsemantic embedding for generalizable person re-identification: a unified perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):278:1– 278:17, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xie:2024:RCA

[XGZ^{+}24] Tianxiu Xie, Keke Gai, Liehuang Zhu, Shuo Wang, and Zijian Zhang. RAC-Chain: an asynchronous consensus-based cross-chain approach to scalable blockchain for metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (7):187:1–187:??, July 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3586011.

Xin:2021:WEG

[XHL^{+}21] Qi Xin, Shaohao Hu, Shuaiqi Liu, Ling Zhao, and Shuihua Wang. WTRPNet: an explainable graph feature convolutional neural network for epileptic EEG classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):107:1–107:18, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460522.

Xiong:2024:ICI

[XHL^{+}24a] Mingfu Xiong, Kaikang Hu, Zhihan Lyu, Fei Fang, Zhongyuan Wang, Ruimin Hu, and Khan Muhammad. Intercamera identity discrimination for unsupervised person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):232:1–232:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3652858.

Xu:2024:IBC

[XHL^{+}24b] Zhenbo Xu, Hai-Miao Hu, Liu Liu, Dongping Zhang, Shifeng Zhang, and Wenming Tan. Instance-based continual learning: a real-world dataset and baseline for fresh recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):1:1–1:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3591209.

Xiong:2023:RRD

[XHYX23] Lizhi Xiong, Xiao Han, Ching-Nung Yang, and Zhihua Xia. RDH-DES: Reversible data hiding over distributed encrypted-image servers based on secret sharing. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):5:1–5:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3512797.

Xu:2021:LBO

[XHZ^{+}21] Xiaolong Xu, Qihe Huang, Yiwen Zhang, Shancang Li, Lianyong Qi, and Wanchun Dou. An LSH-based offloading method for IoMT services in integrated cloud-edge environment. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3s):94:1–94:19, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3408319.

Xu:2025:BPD

[XJB25] Liyong Xu, Yifan Jiao, and Bing-Kun Bao. Bool prompt with decomposition and enhancement: Zero-shot VQA based on PVLMs. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):260:1–260:21, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xia:2025:SIA

[XJFS25] Guiyu Xia, Zhedong Jin, Dongdong Fang, and Yubao Sun. Source informationassisted UV-space transformation network for person image generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):262:1–262:16, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xia:2022:FCS

[XJG^{+}22] Zhihua Xia, Qiuju Ji, Qi Gu, Chengsheng Yuan, and Fengjun Xiao. A format-compatible searchable encryption scheme for JPEG images using bag-of-words. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):85:1–85:18, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3492705.

Xu:2022:EAC

[XJW^{+}22] Haotian Xu, Xiaobo Jin, Qiufeng Wang, Amir Hussain, and Kaizhu Huang. Exploiting attention-consistency loss for spatial-temporal stream action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):119:1–119:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3538749.

Xiao:2025:PPI

[XLF^{+}25] Gui Xiao, Zhen Ling, Qunqun Fan, Xiangyu Xu, Wenjia Wu, Ding Ding, Chen Chen, and Xinwen Fu. Pivot: Panoramicimage-based VR user authentication against side-channel attacks. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):54:1–54:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3694975.

Xu:2023:CUH

[XLH^{+}23] Boqiang Xu, Jian Liang, Lingxiao He, Jinlin Wu, Chao Fan, and Zhenan Sun. Colorunrelated head-shoulder networks for fine-grained person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):210:1–210:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3599730.

Xiao:2024:RRD

[XLH^{+}24] Yi Xiao, Tong Liu, Yu Han, Yue Liu, and Yongtian Wang. Realtime recognition of dynamic hand gestures in practical applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):50:1–50:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561822.

Xu:2021:ANM

[XLN^{+}21] Mingliang Xu, Qingfeng Li, Jianwei Niu, Hao Su, Xiting Liu, Weiwei Xu, Pei Lv, Bing Zhou, and Yi Yang. ART-UP: a novel method for generating scanning-robust aesthetic QR codes. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):25:1–25:23, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3418214.

Xu:2025:MSC

[XLS^{+}25] Liming Xu, Hanqi Li, Jie Shao, Xianhua Zeng, and Weisheng Li. Multi-scale consistency deep lifelong cross-modal hashing. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):283:1–283:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xue:2023:HFF

[XLT^{+}23] Han Xue, Jun Ling, Anni Tang, Li Song, Rong Xie, and Wenjun Zhang. High-fidelity face reenactment via identitymatched correspondence learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):122:1–122:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3571857.

Xing:2023:PPI

[XLW23a] Kai Xing, Tao Li, and Xuanhan Wang. ProposalVLAD with proposal-intra exploring for temporal action proposal generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):118:1–118:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571747.

Xu:2023:CTC

[XLW^{+}23b] Kang Xu, Weixin Li, Xia Wang, Xiaoyan Hu, Ke Yan, Xiaojie Wang, and Xuan Dong. CUR transformer: a convolutional unbiased regional transformer for image denoising. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):104:1–104:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3566125.

Xu:2024:HHK

[XLX^{+}24] Yunhui Xu, Youru Li, Muhao Xu, Zhenfeng Zhu, and Yao Zhao. HKA: a hierarchical knowledge alignment framework for multimodal knowledge graph completion. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):256:1–256:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3664288.

Xu:2021:DSS

[XLZ^{+}21] Chunyan Xu, Rong Liu, Tong Zhang, Zhen Cui, Jian Yang, and Chunlong Hu. Dual-stream structured graph convolution network for skeleton-based action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):120:1–120:22, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450410.

Xu:2022:BIB

[XLZ^{+}22] Sheng Xu, Chang Liu, Baochang Zhang, Jinhu Lü, Guodong Guo, and David Doermann. BiRe-ID: Binary neural network for efficient person re-ID. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (1s):26:1–26:22, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3473340.

Xiao:2023:NRD

[XLZ^{+}23] Mengyao Xiao, Xiaolong Li, Yao Zhao, Bin Ma, and Guodong Guo. A novel reversible data hiding scheme based on pixel-residual histogram. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):46:1–46:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3534565.

Xu:2024:DIC

[XLZ^{+}24] Jing Xu, Bing Liu, Yong Zhou, Mingming Liu, Rui Yao, and Zhiwen Shao. Diverse image captioning via conditional variational autoencoder and dual contrastive learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):29:1– 29:??, January 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3614435.

Xu:2007:CAD

[XMST07] Changsheng Xu, Namunu C.

Maddage, Xi Shao, and Qi Tian. Content-adaptive digital music watermarking based on music structure analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(1): ??, February 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xie:2023:CFP

[XPP^{+}23] Jin Xie, Yanwei Pang, Jing Pan, Jing Nie, Jiale Cao, and Jungong Han. Complementary feature pyramid network for object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):178:1–178:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3584362.

Xiang:2023:LML

[XQG^{+}23] Suncheng Xiang, Dahong Qian, Mengyuan Guan, Binjie Yan, Ting Liu, Yuzhuo Fu, and Guanjie You. Less is more: Learning from synthetic data with fine-grained attributes for person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):173:1– 173:??, October 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3588441.

Xiang:2024:RPR

[XQG^{+}24] Suncheng Xiang, Dahong Qian, Jingsheng Gao, Zirui Zhang, Ting Liu, and Yuzhuo Fu. Rethinking person reidentification via semanticbased pretraining. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):90:1–90:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3628452.

Xu:2022:TCA

[XSD^{+}22] Yifan Xu, Kekai Sheng, Weiming Dong, Baoyuan Wu, Changsheng Xu, and Bao-Gang Hu. Towards corruption-agnostic robust domain adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (4):99:1–99:16, November 2022.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3501800.

Xu:2010:IBP

[XSEZ10] Changsheng Xu, Eckehard Steinbach, Abdulmotaleb El Saddik, and Michelle Zhou. Introduction to the best papers of ACM Multimedia 2009. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3): 12:1–12:??, August 2010. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xu:2023:SBS

[XSL^{+}23] Shunxin Xu, Ke Sun, Dong Liu, Zhiwei Xiong, and Zheng-Jun Zha. Synergy between semantic segmentation and image denoising via alternate boosting. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):69:1–69:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3548459.

Xu:2021:ZSC

[XTL^{+}21] Xing Xu, Jialin Tian, Kaiyi Lin, Huimin Lu, Jie Shao, and Heng Tao Shen. Zeroshot cross-modal retrieval by assembling AutoEncoder and generative adversarial network. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1s):3:1–3:17, March 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3424341.

Xu:2017:CSA

[XW17] Jingxi Xu and Benjamin W. Wah. Consistent synchronization of action order with least noticeable delays in fast-paced multiplayer online games. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):8:1– 8:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xiang:2025:JMD

[XW25] Linhua Xiang and Zengfu Wang. Joint mixing data augmentation for skeleton-based action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):108:1– 108:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xu:2023:AAS

[XWDC23] Yikun Xu, Xingxing Wei, Pengwen Dai, and Xiaochun Cao. A^{2}SC: Adversarial attacks on subspace clustering. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):191:1– 191:??, November 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3587097.

Xu:2021:CMH

[XWH^{+}21] Xing Xu, Yifan Wang, Yixuan He, Yang Yang, Alan Hanjalic, and Heng Tao Shen. Cross-modal hybrid feature fusion for image-sentence matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):127:1–127:23, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3458281.

Xiang:2025:PUR

[XWH^{+}25] Chong-Yang Xiang, Xiao Wu, Jun-Yan He, Zhaoquan Yuan, and Tingquan He. Person in uniforms re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):61:1–61:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3703839.

Xu:2025:SAC

[XWL^{+}25a] Kai Xu, Lichun Wang, Shuang Li, Tong Gao, and Baocai Yin. Scene adaptive context modeling and balanced relation prediction for scene graph generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):76:1–76:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3708350.

Xu:2025:MCS

[XWL^{+}25b] Ning Xu, Xiaowen Wang, Jing Liu, Lanjun Wang, Xuanya Li, Mengxiao Zhu, Yongdong Zhang, and An-An Liu. Model can be subtle: Two important mechanisms for social media popularity prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):71:1– 71:??, February 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3705319.

Xu:2021:EIE

[XWW^{+}21] Xin Xu, Shiqin Wang, Zheng Wang, Xiaolong Zhang, and Ruimin Hu. Exploring image enhancement for salient object detection in low light images. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1s):8:1–8:19, March 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3414839.

Xian:2025:DDT

[XWW^{+}25] Weizhi Xian, Junyi Wang, Xuekai Wei, Jielu Yan, Yueting Huang, Kunyin Guo, Weijia Jia, and Mingliang Zhou. DTSD: a dual teacher-student-based discrimination model for anomaly detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):178:1– 178:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xia:2021:SED

[XWY21] Bohui Xia, Xueting Wang, and Toshihiko Yamasaki. Semantic explanation for deep neural networks using feature interactions. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):115:1–115:19, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474557.

Xiong:2024:VLS

[XWZ24a] Yuan Xiong, Jingru Wang, and Zhong Zhou. Virtual-Loc: Large-scale visual localization using virtual images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (3):66:1–66:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3622788.

Xue:2024:SAA

[XWZ^{+}24b] Mingfu Xue, Yinghao Wu, Leo Yu Zhang, Dujuan Gu, Yushu Zhang, and Weiqiang Liu. SSAT: Active authorization control and User’s fingerprint tracking framework for DNN IP protection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):324:1– 324:??, October 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3679202.

Xu:2008:AKG

[XXD^{+}08] Min Xu, Changsheng Xu, Lingyu Duan, Jesse S. Jin, and Suhuai Luo. Audio keywords generation for sports video analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2):11:1–11:??, May 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xiao:2021:WSS

[XXG^{+}21] Junsheng Xiao, Huahu Xu, Honghao Gao, Minjie Bian, and Yang Li. A weakly supervised semantic segmentation network by aggregating seed cues: The multi-object proposal generation perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):15:1–15:19, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3419842.

Xu:2023:PTM

[XYC^{+}23] Yongzong Xu, Zhijing Yang, Tianshui Chen, Kai Li, and Chunmei Qing. Progressive transformer machine for natural character reenactment. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):92:1–92:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3559107.

Xia:2020:CDB

[XYJ^{+}20] Kaijian Xia, Hongsheng Yin, Yong Jin, Shi Qiu, and Hongru Zhao. Cross-domain brain CT image smart segmentation via shared hidden space transfer FCM clustering. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (2s):61:1–61:21, July 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3357233.

Xue:2023:LEE

[XYL^{+}23] Feng Xue, Tian Yang, Kang Liu, Zikun Hong, Mingwei Cao, Dan Guo, and Richang Hong. LCSNet: End-to-end lipreading with channel-aware feature selection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):28:1–28:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3524620.

Xu:2022:RRL

[XYW^{+}22] Xin Xu, Xin Yuan, Zheng Wang, Kai Zhang, and Ruimin Hu. Rank-in-rank loss for person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):130:1– 130:??, June 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3532866.

Xu:2025:RDD

[XZC^{+}25] Yuyu Xu, Pingping Zhang, Minghui Chen, Qiudan Zhang, Wenhui Wu, Yun Zhang, and Xu Wang. RGB-D data compression via bi-directional cross-modal prior transfer and enhanced entropy modeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):58:1–58:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3702997.

Xiang:2023:BPP

[XZCG23] Tao Xiang, Honghong Zeng, Biwen Chen, and Shangwei Guo. BMIF: Privacy-preserving blockchain-based medical image fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):36:1–36:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531016.

Xu:2021:SVM

[XZH^{+}21] Tong Xu, Peilun Zhou, Linkang Hu, Xiangnan He, Yao Hu, and Enhong Chen. Socializing the videos: a multimodal approach for social relation recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1):23:1–23:23, April 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3416493.

Xiao:2022:DAS

[XZL^{+}22a] Hui-Chu Xiao, Wan-Lei Zhao, Jie Lin, Yi-Geng Hong, and Chong-Wah Ngo. Deeply activated salient region for instance search. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):147:1–147:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3510004.

Xu:2022:DBJ

[XZL^{+}22b] Kexin Xu, Haijun Zhang, Keping Long, Jianquan Wang, and Lei Sun. DRL based joint affective services computing and resource allocation in ISTN. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):135:1–135:??, October 2022. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3561821.

Xu:2025:DDL

[XZL^{+}25] Liming Xu, Dengping Zhao, Hanqi Li, Xianhua Zeng, and Bochuan Zheng. Deep differential lifelong cross-modal hashing for stream medical data retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):284:1–284:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xu:2023:MMM

[XZLZ23] Liming Xu, Xianhua Zeng, Weisheng Li, and Bochuan Zheng. MFGAN: Multi-modal feature-fusion for CT metal artifact reduction using GANs. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1s):32:1–32:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3528172.

Xiao:2024:FDW

[XZY^{+}24a] Shuai Xiao, Zhuo Zhang, Jiachen Yang, Jiabao Wen, and Yang Li. Forgery detection by weighted complementarity between significant invariance and detail enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):346:1– 346:??, November 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3605893.

Xu:2024:CPC

[XZY^{+}24b] Yiling Xu, Yujie Zhang, Qi Yang, Xiaozhong Xu, and Shan Liu. Compressed point cloud quality index by combining global appearance and local details. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):291:1–291:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672567.

Xu:2025:MEA

[XZY^{+}25] Zishan Xu, Xiaofeng Zhang, Yuqing Yang, Wei Chen, Jueting Liu, Tingting Xu, Zehua Wang, and Abdulmotaleb El Saddik. MuralAgent: Enhancing ancient mural outpainting with RAG-based texts and multimodal integration. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):267:1– 267:17, September 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Xie:2023:TDW

[XZZL23] Chi Xie, Zikun Zhuang, Shengjie Zhao, and Shuang Liang. Temporal dropout for weakly supervised action localization. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):102:1–102:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3567827.

Yang:2010:DMD

[Yan10] Bo Yang. DSI: a model for distributed multimedia semantic indexing and content integration. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(1):3:1–3:??, February 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yan:2017:LCI

[Yan17] Zheng Yan. Learning from collective intelligence: Feature learning using social images and tags. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(1):1:1–1:??, January 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yeh:2014:PPR

[YBO14] Che-Hua Yeh, Brian A. Barsky, and Ming Ouhyoung. Personalized photograph ranking and selection system considering positive and negative user feedback. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(4):36:1–36:??, June 2014. ISSN 1551-6857 (print), 1551-6865 (electronic).

Yuan:2023:ATD

[YBTX23] Mengqi Yuan, Bing-Kun Bao, Zhiyi Tan, and Changsheng Xu. Adaptive text denoising network for image caption editing. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):41:1–41:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3532627.

Yan:2025:MMM

[YBW^{+}25] Kun Yan, Zied Bouraoui, Fangyun Wei, Chang Xu, Ping Wang, Shoaib Jameel, and Steven Schockaert. Modeling multi-modal cross-interaction for multi-label few-shot image classification based on local feature selection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):80:1–80:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3711867.

Yiu:2008:ODC

[YC08] Wai-Pun Ken Yiu and Shueng-Han Gary Chan. Offering data confidentiality for multimedia overlay multicast: Design and analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(2):13:1–13:??, November 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yu:2023:MVS

[YCFX23] Hang Yu, Chilam Cheang, Yanwei Fu, and Xiangyang Xue. Multi-view shape generation for a 3D human-like body. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1):11:1–11:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3514248.

Yuan:2014:UQE

[YCGM14] Zhenhui Yuan, Shengyang Chen, Gheorghita Ghinea, and Gabriel-Miro Muntean. User quality of experience of mulsemedia applications. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):15:1– 15:??, September 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2025:MHA

[YCL25] Yiping Yang, Baiyun Cui, and Yingming Li. A multimodal hierarchical attentional ordering network. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):78:1–78:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3711864.

Yuan:2022:LAS

[YCLH22] Di Yuan, Xiaojun Chang, Zhihui Li, and Zhenyu He. Learning adaptive spatial-temporal context-aware correlation filters for UAV tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):70:1–70:18, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3486678.

Yuan:2023:GAT

[YCZ^{+}23] Jin Yuan, Shikai Chen, Yao Zhang, Zhongchao Shi, Xin Geng, Jianping Fan, and Yong Rui. Graph attention transformer network for multi-label image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):150:1– 150:??, July 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3578518.

Yu:2024:RDH

[YCZ^{+}24] Chunqiang Yu, Shichao Cheng, Xianquan Zhang, Xinpeng Zhang, and Zhenjun Tang. Reversible data hiding in shared JPEG images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):372:1–372:??, December 2024. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3695463.

Yang:2023:NRQ

[YDCZ23] Yang Yang, Yingqiu Ding, Ming Cheng, and Weiming Zhang. No-reference quality assessment for contrastdistorted images based on gray and color-gray-difference space. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (2):64:1–64:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3555355.

Ye:2024:VSA

[YDJ24] Jin Ye, Meng Dan, and Wenchao Jiang. A visual sensitivity aware ABR algorithm for DASH via deep reinforcement learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):77:1–77:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3591108.

Yao:2022:SLM

[YF22] Peng Yao and Jieqing Feng. Sparse LIDAR measurement fusion with joint updating cost for fast stereo matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):1:1– 1:18, January 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3471870.

Yuan:2019:SSP

[YFLF19] Yuan Yuan, Jie Fang, Xiaoqiang Lu, and Yachuang Feng. Spatial structure preserving feature pyramid network for semantic image segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):73:1–73:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3321512.

Yang:2012:UCM

[YGHH12] Linjun Yang, Bo Geng, Alan Hanjalic, and Xian-Sheng Hua. A unified context model for web image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3):28:1–28:??, July 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2023:EMF

[YGL^{+}23] Dinghao Yang, Wei Gao, Ge Li, Hui Yuan, Junhui Hou, and Sam Kwong. Exploiting manifold feature representation for efficient classification of 3D point clouds. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):50:1–50:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3539611.

Yuan:2024:DCB

[YGMY24] Hang Yuan, Wei Gao, Siwei Ma, and Yiqiang Yan. Divideand-conquer-based RDO-free CU partitioning for 8K video compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):114:1–114:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3634705.

Yuan:2019:DLT

[YGNT19] Bo Yuan, Xinbo Gao, Zhenxing Niu, and Qi Tian. Discovering latent topics by Gaussian latent Dirichlet allocation and spectral clustering. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1):25:1–25:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3290047.

Yeh:2013:CAS

[YH13] Lo-Yao Yeh and Jiun-Long Huang. A conditional access system with efficient key distribution and revocation for mobile pay-TV systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(3):18:1– 18:??, June 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ye:2014:OBL

[YH14] Jun Ye and Kien A. Hua.

Octree-based 3D logic and computation of spatial relationships in live video query processing. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (2):28:1–28:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Huang:2021:PCN

[yHcCzH^{+}21] Chun ying Huang, Yun chen Cheng, Guan zhang Huang, Ching ling Fan, and Cheng hsin Hsu. On the performance comparisons of native and clientless real-time screensharing technologies. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):54:1– 54:26, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3437881.

Yang:2024:PAP

[YHG^{+}24] Kaihui Yang, Junwei Han, Guangyu Guo, Chaowei Fang, Yingzi Fan, Lechao Cheng, and Dingwen Zhang. Progressive adapting and pruning: Domain-incremental learning for saliency prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):243:1– 243:??, August 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3661312.

Yang:2022:ICD

[YHH^{+}22] Kai-Wei Yang, Yen-Yun Huang, Jen-Wei Huang, Ya-Rou Hsu, Chang-Lin Wan, Hong-Han Shuai, Li-Chun Wang, and Wen-Huang Cheng. Improving crowd density estimation by fusing aerial images and radio signals. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):84:1–84:23, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3492346.

Yang:2024:SSL

[YHH24] Yiming Yang, Weipeng Hu, and Haifeng Hu. Syncretic space learning network for NIR-VIS face recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):11:1–11:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3607143.

Ye:2025:RRA

[YHP25] Zhaoda Ye, Xiangteng He, and Yuxin Peng. RaT2IGen:

Relation-aware text-to-image generation via learnable prompt. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (5):151:1–151:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ye:2017:TOM

[YHQH17] Jun Ye, Hao Hu, Guo-Jun Qi, and Kien A. Hua. A temporal order modeling approach to human action recognition from multimodal sensor data. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (2):14:1–14:??, May 2017. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2023:MID

[YHX^{+}23] Taocun Yang, Yaping Huang, Yanlin Xie, Junbo Liu, and Shengchun Wang. MixOOD: Improving out-of-distribution detection with enhanced data mixup. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):155:1–155:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578935.

Yang:2020:CLR

[YHXL20] Liang Yang, Haifeng Hu, Songlong Xing, and Xinlong Lu. Constrained LSTM and residual attention for image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):75:1–75:18, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3386725.

Yue:2019:SRS

[YHZ19] Guanghui Yue, Chunping Hou, and Tianwei Zhou. Subtitle region selection of S3D images in consideration of visual discomfort and viewing habit. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):77:1– 77:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3325197.

Yu:2023:DAP

[YHZ23] Hongchuan Yu, Mengqing Huang, and Jian Jun Zhang. Domain adaptation problem in sketch based image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):106:1–106:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3565368.

Yang:2014:MDF

[YI14] Ying Yang and Ioannis Ivrissimtzis. Mesh discriminative features for 3D steganalysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (3):27:1–27:??, April 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yarnagula:2019:QMC

[YJM^{+}19] Hema Kumar Yarnagula, Parikshit Juluri, Sheyda Kiani Mehr, Venkatesh Tamarapalli, and Deep Medhi. QoE for mobile clients with Segmentaware Rate Adaptation Algorithm (SARA) for DASH video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(2):36:1–36:??, June 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3311749.

Yang:2024:SFS

[YJM^{+}24] Yuqing Yang, Boris Joukovsky, José Oramas Mogrovejo, Tinne Tuytelaars, and Nikos Deligiannis. SNIPPET: a framework for subjective evaluation of visual explanations applied to Deep-Fake detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):253:1–253:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665248.

Yu:2018:SPI

[YJTN18] Tuo Yu, Haiming Jin, Wai-Tian Tan, and Klara Nahrstedt.

SKEPRID: Pose and illumination change-resistant skeletonbased person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (4):82:1–82:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yan:2007:MSO

[YK07] Wei-Qi Yan and Mohan S. Kankanhalli. Multimedia simplification for optimized MMS synthesis. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(1):??, February 2007. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2021:CIR

[YKQ^{+}21] Peihao Yang, Linghe Kong, Meikang Qiu, Xue Liu, and Guihai Chen. Compressed imaging reconstruction with sparse random projection. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1):26:1–26:25, April 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3447431.

Yao:2022:RGA

[YKW^{+}22] Lingxiang Yao, Worapan Kusakunniran, Qiang Wu, Jingsong Xu, and Jian Zhang. Recognizing gaits across walking and running speeds. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):75:1–75:22, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3488715.

Yang:2018:JEA

[YLCC18] Huei-Fang Yang, Bo-Yao Lin, Kuang-Yu Chang, and Chu-Song Chen. Joint estimation of age and expression by combining scattering and convolutional networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(1):9:1–9:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2025:DVD

[YLD^{+}25] Rui Yang, Rushi Lan, Zhenrong Deng, Xiaonan Luo, and Xiyan Sun. Deepfake video detection using facial feature points and ch-transformer. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):49:1–49:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672566.

Yeh:2022:GVW

[YLHL22] Chih-Kuo Yeh, Thi-Ngoc-Hanh Le, Zhi-Ying Hou, and Tong-Yee Lee. Generating virtual wire sculptural art from 3D models. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):51:1–51:23, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3475798.

Yang:2020:LSS

[YLK^{+}20] Zhenguo Yang, Zehang Lin, Peipei Kang, Jianming Lv, Qing Li, and Wenyin Liu. Learning shared semantic space with correlation alignment for cross-modal event retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):9:1–9:22, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3374754.

Yang:2023:SCF

[YLL^{+}23] Song Yang, Qiang Li, Wenhui Li, Xuan-Ya Li, Ran Jin, Bo Lv, Rui Wang, and Anan Liu. Semantic completion and filtration for image-text retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):140:1–140:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3572844.

Yao:2024:CMS

[YLL^{+}24] Tao Yao, Yiru Li, Ying Li, Yingying Zhu, Gang Wang, and Jun Yue. Cross-modal semantically augmented network for image-text matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):99:1– 99:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631356.

Yan:2021:RSI

[YLLL21] Xuehu Yan, Lintao Liu, Longlong Li, and Yuliang Lu. Robust secret image sharing resistant to noise in shares. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):24:1– 24:22, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3419750.

Yang:2023:HHF

[YLLL23] Xin Yang, Hengrui Li, Xiaochuan Li, and Tao Li. HIFGAN: a high-frequency information-based generative adversarial network for image super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):159:1–159:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3578934.

Yang:2025:FAS

[YLLZ25] Shangrong Yang, Chunyu Lin, Kang Liao, and Yao Zhao. Fish-Former: Annulus slicing-based transformer for fisheye rectification. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):128:1–128:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yan:2023:FDP

[YLS^{+}23] Xuehu Yan, Longlong Li, Lei Sun, Jia Chen, and Shudong Wang. Fake and dishonest participant immune secret image sharing. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):139:1–139:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3572842.

Yuan:2024:UFL

[YLYB24] Bowen Yuan, Jiahao Lu, Sisi You, and Bing-Kun Bao. Unbiased feature learning with causal intervention for visible-infrared person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):298:1–298:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3674737.

Yin:2010:LEC

[YLZ^{+}10] Hao Yin, Xuening Liu, Tongyu Zhan, Vyas Sekar, Feng Qiu, Chuang Lin, Hui Zhang, and Bo Li. LiveSky: Enhancing CDN with P2P. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):16:1–16:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yan:2021:DID

[YLZ^{+}21] Chenggang Yan, Zhisheng Li, Yongbing Zhang, Yutao Liu, Xiangyang Ji, and Yongdong Zhang. Depth image denoising using nuclear norm and learning graph model. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):122:1–122:17, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3404374.

Yaqoob:2024:APT

[YM24] Abid Yaqoob and Gabriel-Miro Muntean. Advanced predictive tile selection using dynamic tiling for prioritized 360^{∘} video VR streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):6:1–6:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603146.

Yakubu:2017:SSN

[YMA17] Abukari M. Yakubu, Namunu C. Maddage, and Pradeep K. Atrey. Securing speech noise reduction in outsourced environment. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):51:1–51:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yu:2025:DHQ

[YMJ^{+}25] Dongjian Yu, Weiqing Min, Xin Jin, Qian Jiang, Ying Jin, and Shuqiang Jiang. Diverse and high-quality food image generation from only food names. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):153:1–153:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yan:2022:AIF

[YML^{+}22] Chenggang Yan, Lixuan Meng, Liang Li, Jiehua Zhang, Zhan Wang, Jian Yin, Jiyong Zhang, Yaoqi Sun, and Bolun Zheng. Age-invariant face recognition by multi-feature fusionand decomposition with selfattention. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):29:1–29:18, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3472810.

Yang:2024:LFR

[YMS^{+}24] Yancun Yang, Weiqing Min, Jingru Song, Guorui Sheng, Lili Wang, and Shuqiang Jiang. Lightweight food recognition via aggregation block and feature encoding. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):327:1–327:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3680285.

Yang:2016:AGV

[YMX^{+}16] Xuyong Yang, Tao Mei, Ying-Qing Xu, Yong Rui, and Shipeng Li. Automatic generation of visual-textual presentation layout. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):33:1–33:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2021:ACG

[YMY^{+}21] Xin Yang, Zongliang Ma, Letian Yu, Ying Cao, Baocai Yin, Xiaopeng Wei, Qiang Zhang, and Rynson W. H. Lau. Automatic comic generation with stylistic multi-page layouts and emotion-driven text balloon generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):55:1– 55:19, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3440053.

Yan:2018:GES

[YNC18] Yan Yan, Liqiang Nie, and Rita Cucchiara. Guest editorial: Special section on “Multimedia Understanding via Multimodal Analytics”. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):37:1–37:??, May 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yu:2022:DAM

[YNLZ22] Yang Yu, Rongrong Ni, Wenjie Li, and Yao Zhao. Detection of AI-Manipulated fake faces via mining generalized features. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (4):94:1–94:23, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3499026.

Yang:2025:PCD

[YNY25] Yuqing Yang, Anh Nguyen, and Zhisheng Yan. A patch can disrupt live video streaming: Physical adversarial attacks on deep learning compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):257:1–257:23, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

You:2015:UPD

[YP15] Shingchern D. You and Yi-Han Pu. Using paired distances of signal peaks in stereo channels as fingerprints for copy identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1):1:1–1:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Ye:2020:SCM

[YP20] Zhaoda Ye and Yuxin Peng.

Sequential cross-modal hashing learning via multi-scale correlation mining. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–20, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3356338.

Yang:2022:MNM

[YPSC22] Chun-Wei Yang, Thanh Hai Phung, Hong-Han Shuai, and Wen-Huang Cheng. Mask or non-mask? Robust face mask detector via triplet-consistency representation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):42:1– 42:20, February 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3472623.

Yang:2021:SSI

[YQC^{+}21] Xin Yang, Yu Qiao, Shaozhe Chen, Shengfeng He, Baocai Yin, Qiang Zhang, Xiaopeng Wei, and Rynson W. H.

Lau. Smart scribbles for image matting. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):121:1–121:21, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3408323.

Yang:2012:EMA

[YQH12] Rui Yang, Zhenhua Qu, and Jiwu Huang. Exposing MP3 audio forgeries using frame offsets. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2S):35:1–35:??, September 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2024:DPJ

[YQZP24] Yang Yang, Shuailong Qiu, Lanling Zeng, and Zhigeng Pan. Detail-preserving joint image upsampling. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):251:1–251:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665246.

Yadav:2020:EDA

[YRE^{+}20] Shweta Yadav, Pralay Ramteke, Asif Ekbal, Sriparna Saha, and Pushpak Bhattacharyya. Exploring disorder-aware attention for clinical event extraction. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):31:1–31:21, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372328.

Yu:2021:CLG

[YSC21] Yi Yu, Abhishek Srivastava, and Simon Canales. Conditional LSTM-GAN for melody generation from lyrics. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1):35:1– 35:20, April 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3424116.

Yang:2021:AWE

[YSF^{+}21] Xin Yang, Xuemeng Song, Fuli Feng, Haokun Wen, Ling-Yu Duan, and Liqiang Nie. Attribute-wise explainable fashion compatibility modeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (1):36:1–36:21, April 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3425636.

Yang:2006:SER

[YSG^{+}06] Guang Yang, Tony Sun, Mario Gerla, M. Y. Sanadidi, and Ling-Jyh Chen. Smooth and efficient real-time video transport in the presence of wireless errors. ACM Transactions on Multimedia Computing, Communications, and Applications, 2(2):109–126, May 2006. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2024:CCM

[YSSF24] Shih-Wei Yang, Li-Hsiang Shen, Hong-Han Shuai, and Kai-Ten Feng. CMAF: Crossmodal augmentation via fusion for underwater acoustic image recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):124:1–124:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3636427.

Yan:2016:UVR

[YSXH16] Ming Yan, Jitao Sang, Changsheng Xu, and M. Shamim Hossain. A unified video recommendation by cross-network user modeling. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):53:1–53:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yin:2022:MFL

[YSY^{+}22] Guanghao Yin, Shouqian Sun, Dian Yu, Dejian Li, and Kejun Zhang. A multimodal framework for large-scale emotion recognition by fusing music and electrodermal activity signals. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (3):78:1–78:23, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3490686.

Yin:2015:CVC

[YSZ15] Yifang Yin, Beomjoo Seo, and Roger Zimmermann. Content vs. context: Visual and geographic information use in video landmark retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(3):39:1– 39:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yin:2014:STT

[YSZZ14] Yifang Yin, Zhijie Shen, Luming Zhang, and Roger Zimmermann. Spatial-temporal tag mining for automatic geospatial video annotation. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(2):29:1–29:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yan:2021:PNR

[YTL^{+}21] Chenggang Yan, Tong Teng, Yutao Liu, Yongbing Zhang, Haoqian Wang, and Xiangyang Ji. Precise no-reference image quality evaluation based on distortion identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3s):110:1– 110:21, October 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3468872.

Yu:2024:RHD

[YTL^{+}24a] Mengzhu Yu, Zhenjun Tang, Xiaoping Liang, Xianquan Zhang, Zhixin Li, and Xinpeng Zhang. Robust hashing with deep features and Meixner moments for image copy detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):384:1–384:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3696669.

Yue:2024:MGR

[YTL^{+}24b] Shengbin Yue, Yunbin Tu, Liang Li, Shengxiang Gao, and Zhengtao Yu. Multigrained representation aggregating transformer with gating cycle for change captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):321:1–321:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3660346.

Yanagi:2022:IRR

[YTOH22] Rintaro Yanagi, Ren Togo, Takahiro Ogawa, and Miki Haseyama. Interactive reranking via object entropyguided question answering for cross-modal image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (3):68:1–68:17, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3485042.

Yan:2025:CAD

[YTR^{+}25] Jiebin Yan, Ziwen Tan, Jiale Rao, Lei Wu, Yifan Zuo, and Yuming Fang. Computational analysis of degradation modeling in blind panoramic image quality assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):140:1– 140:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yu:2019:DCM

[YTRC19] Yi Yu, Suhua Tang, Francisco Raposo, and Lei Chen. Deep cross-modal correlation learning for audio and lyrics in music retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1):20:1– 20:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3281746.

Ye:2024:UUR

[YTY^{+}24] Jiabo Ye, Junfeng Tian, Ming Yan, Haiyang Xu, Qinghao Ye, Yaya Shi, Xiaoshan Yang, Xuwu Wang, Ji Zhang, Liang He, and Xin Lin. UniQR-Net: Unifying referring expression grounding and segmentation with QRNet. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):246:1–246:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3660638.

Yadav:2023:DML

[YV23] Ashima Yadav and Dinesh Kumar Vishwakarma. A deep multi-level attentive network for multimodal sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):15:1–15:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3517139.

Yang:2023:ESI

[YW23] Bin-Cheng Yang and Gangshan Wu. Efficient single-image super-resolution using dual path connections with multiple scale learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):115:1–115:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570164.

Yang:2024:EIC

[YWC24] Ching-Nung Yang, Xiaotian Wu, and Min-Jung Chung. Enhancement of information carrying and decoding for visual cryptography with error correction. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):23:1–23:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3612927.

Yan:2025:VUB

[YWC^{+}25] Jiebin Yan, Kangcheng Wu, Junjie Chen, Ziwen Tan, Yuming Fang, and Weide Liu. Viewport-unaware blind omnidirectional image quality assessment: a flexible and effective paradigm. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):145:1–145:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2024:BBS

[YWF^{+}24] Wenyuan Yang, Shaocong Wu, Jianwei Fei, Xianwang Zeng, Yuemin Ding, and Zhihua Xia. A bitcoin-based secure outsourcing scheme for optimization problem in multimedia Internet of Things. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(6):152:1–152:??, June 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3637489.

Yang:2020:ASC

[YWG^{+}20] Liang Yang, Yuexue Wang, Junhua Gu, Xiaochun Cao, Xiao Wang, Di Jin, Guiguang Ding, Jungong Han, and Weixiong Zhang. Autonomous semantic community detection via adaptively weighted lowrank approximation. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1– 22, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3355393.

Yang:2017:EPR

[YWH^{+}17] Xun Yang, Meng Wang, Richang Hong, Qi Tian, and Yong Rui. Enhancing person re-identification in a selftrained subspace. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):27:1–27:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yi:2024:OVS

[YWL^{+}24] Peng Yi, Zhongyuan Wang, Laigan Luo, Kui Jiang, Zheng He, Junjun Jiang, Tao Lu, and Jiayi Ma. Omniscient video super-resolution with explicitimplicit alignment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):150:1– 150:??, May 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3640346.

Yao:2023:CRA

[YWLZ23] Xuanrong Yao, Xin Wang, Yue Liu, and Wenwu Zhu. Continual recognition with adaptive memory update. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):134:1–134:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3573202.

Yang:2010:EMP

[YWN^{+}10a] Zhenyu Yang, Wanmin Wu, Klara Nahrstedt, Gregorij Kurillo, and Ruzena Bajcsy. Enabling multi-party 3D teleimmersive environments with ViewCast. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(2):7:1–7:??, March 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2010:EMT

[YWN^{+}10b] Zhenyu Yang, Wanmin Wu, Klara Nahrstedt, Gregorij Kurillo, and Ruzena Bajcsy. Enabling multiparty 3D teleimmersive environments with ViewCast. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(4):29:1–29:??, November 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2015:RCI

[YWNW15] Hong-Ying Yang, Xiang-Yang Wang, Pan-Pan Niu, and Ai-Long Wang. Robust color image watermarking using geometric invariant quaternion polar harmonic transform. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(3):40:1– 40:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2024:AIG

[YWW^{+}24] Jifan Yang, Zhongyuan Wang, Guangcheng Wang, Baojin Huang, Yuhong Yang, and Weiping Tu. Auxiliary information guided self-attention for image quality assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (4):119:1–119:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3635716.

Yin:2023:FFM

[YWY^{+}23] Haibing Yin, Hongkui Wang, Li Yu, Junhui Liang, and Guangtao Zhai. Feedforward and feedback modulations based foveated JND estimation for images. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):154:1–154:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579094.

Yan:2024:TSP

[YXJ^{+}24] Yunyao Yan, Guoqing Xiang, Huizhu Jia, Jie Chen, Xiaofeng Huang, and Xiaodong Xie. Two-stage perceptual quality oriented rate control algorithm for HEVC. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):140:1–140:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3636510.

Yang:2025:HFI

[YXK^{+}25] Bing Yang, Xueqin Xiang, Wanzeng Kong, Jianhai Zhang, and Jinliang Yao. Hybrid feature integrated transformer for 3D hand reconstruction from a single RGB image. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):177:1–177:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2021:DCN

[YXYB21] Zhenzhen Yang, Pengfei Xu, Yongpeng Yang, and Bing-Kun Bao. A densely connected network based on U-Net for medical image segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (3):89:1–89:14, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3446618.

Yang:2022:DQN

[YXZQ22] Chenyi Yang, Xiaolong Xu, Xiaokang Zhou, and Lianyong Qi. Deep Q networkdriven task offloading for efficient multimedia data analysis in edge computing-assisted IoV. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):124:1–124:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3548687.

You:2024:MOT

[YYBX24] Sisi You, Hantao Yao, Bing-Kun Bao, and Changsheng Xu. Multi-object tracking with spatial-temporal tracklet association. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):129:1–129:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635155.

Yang:2013:ETT

[YYS13] Yang Yang, Yi Yang, and Heng Tao Shen. Effective transfer tagging from image to video. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(2): 14:1–14:??, May 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2024:LVC

[YYX^{+}24] Jiayu Yang, Chunhui Yang, Fei Xiong, Yongqi Zhai, and Ronggang Wang. Learned video compression with adaptive temporal prior and decoded motion-aided quality enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):238:1–238:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3661824.

Yu:2024:PGE

[YYY^{+}24] Jiaqi Yu, Jinhai Yang, Hua Yang, Renjie Pan, Pingrui Lai, and Guangtao Zhai. Psychology-guided environment aware network for discovering social interaction groups from videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):233:1–233:??, August 2024. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3657295.

Yuan:2023:SBD

[YZ23] Ye Yuan and Jiawan Zhang. Shot boundary detection using color clustering and attention mechanism. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):198:1–198:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3595923.

Yu:2025:TTG

[YZD^{+}25] Zhipeng Yu, Zimeng Zhao, Yanxi Du, Yuzhou Zheng, Binghui Zuo, and Yangang Wang. T2C: Text-guided 4D cloth generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):212:1–212:19, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yuan:2020:ICJ

[YZG^{+}20] Jin Yuan, Lei Zhang, Songrui Guo, Yi Xiao, and Zhiyong Li. Image captioning with a joint attention mechanism by visual concept samples. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):83:1–83:22, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3394955.

Yang:2025:APS

[YZJ^{+}25] Jiayu Yang, Yongqi Zhai, Wei Jiang, Chunhui Yang, Feng Gao, and Ronggang Wang. Adaptive prediction structure for learned video compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):59:1–59:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3703914.

Yuan:2014:MRB

[YZL^{+}14] Jin Yuan, Yi-Liang Zhao, Huanbo Luan, Meng Wang, and Tat-Seng Chua. Memory recall based video search: Finding videos you have seen before based on your memory. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (2):21:1–21:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yu:2024:MIQ

[YZL^{+}24] Encheng Yu, Jianer Zhou, Zhenyu Li, Gareth Tyson, Weichao Li, Xinyi Zhang, Zhiwei Xu, and Gaogang Xie. Mustang: Improving QoE for realtime video in cellular networks by masking jitter. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):285:1–285:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3672399.

Yang:2022:ISS

[YZRW22] Xun Yang, Liang Zheng, Elisa Ricci, and Meng Wang. Introduction to the special section on learning representations, similarity, and associations in dynamic multimedia environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2s):127:1–127:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569952.

Yang:2022:RIT

[YZS^{+}22] Dongbao Yang, Yu Zhou, Wei Shi, Dayan Wu, and Weiping Wang. RD-IOD: Twolevel residual-distillation-based triple-network for incremental object detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1):18:1–18:23, January 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472393.

Yan:2023:TIF

[YZS^{+}23] Han Yan, Haijun Zhang, Jianyang Shi, Jianghong Ma, and Xiaofei Xu. Toward intelligent fashion design: a texture and shape disentangled generative adversarial network. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):107:1–107:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3567596.

Yin:2024:EIL

[YZW^{+}24] Shukang Yin, Sirui Zhao, Hao Wang, Tong Xu, and Enhong Chen. Exploiting instancelevel relationships in weakly supervised text-to-video retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):316:1–316:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3663571.

Yang:2016:SFM

[YZX16] Xiaoshan Yang, Tianzhu Zhang, and Changsheng Xu. Semantic feature mining for video event understanding. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):55:1–55:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yang:2015:BML

[YZXY15] Xiaoshan Yang, Tianzhu Zhang, Changsheng Xu, and Ming-Hsuan Yang. Boosted multifeature learning for cross-domain transfer. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(3):35:1–35:??, January 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Yin:2025:JST

[YZY^{+}25] Qian Yin, Xinfeng Zhang, Ruoke Yan, Yuhuai Zhang, Shanshe Wang, and Siwei Ma. Joint structure-texture scanorder for point cloud attribute compression using affine transformation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):173:1–173:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

You:2024:IAV

[YZYX24] Sisi You, Yukun Zuo, Hantao Yao, and Changsheng Xu. Incremental audio-visual fusion for person recognition in earthquake scene. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):53:1–53:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3614434.

Yao:2025:CMT

[YZZZ25] Jiacheng Yao, Jing Zhang, Shuying Zhang, and Li Zhuo. Cross-modal tri-semantic correlation-CLIP for short video homogenization recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):300:1– 300:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2025:SPE

[ZBYZ25] Mohan Zhou, Yalong Bai, Qing Yang, and Tiejun Zhao. StyleInject: Parameter efficient tuning of text-to-image diffusion models. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):152:1–152:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhu:2012:JLS

[ZC12] Xinglei Zhu and Chang W.

Chen. A joint layered scheme for reliable and secure mobile JPEG-2000 streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3):30:1– 30:??, July 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zimmermann:2008:DMP

[ZCAP08] Roger Zimmermann, Elaine Chew, Sakire Arslan Ay, and Moses Pawar. Distributed musical performances: Architecture and stream management. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(2): 14:1–14:??, May 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2015:SDP

[ZCD15] Bo Zhang, Nicola Conci, and Francesco G. B. De Natale. Segmentation of discriminative patches in human activity video. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1):4:1–4:??, August 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2023:PSN

[ZCHD23] Yi Zhang, Fang-Yi Chao, Wassim Hamidouche, and Olivier Deforges. PAV-SOD: a new task towards panoramic audiovisual saliency detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3):101:1–101:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3565267.

Zhang:2012:PHL

[ZCL^{+}12] Tieying Zhang, Xueqi Cheng, Jianming Lv, Zhenhua Li, and Weisong Shi. Providing hierarchical lookup service for P2P– VoD systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 8s(1):15:1–15:??, February 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zeng:2022:MIL

[ZCL^{+}22] Yawen Zeng, Da Cao, Shaofei Lu, Hanling Zhang, Jiao Xu, and Zheng Qin. Moment is important: Language-based video moment retrieval via adversarial learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(2):56:1–56:21, May 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3478025.

Zhang:2024:ICS

[ZCL^{+}24] Liping Zhang, Shukai Chen, Fei Lin, Wei Ren, Kim-Kwang Raymond Choo, and Geyong Min. 1DIEN: Crosssession electrocardiogram authentication using 1D integrated EfficientNet. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):17:1– 17:??, January 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3609800.

Zhang:2023:DRI

[ZCQ^{+}23] Yushu Zhang, Nuo Chen, Shuren Qi, Mingfu Xue, and Zhongyun Hua. Detection of recolored image by texture features in chrominance components. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):121:1–121:??, May 2023.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3571076.

Zhu:2020:DSS

[ZCS^{+}20] Nengjun Zhu, Jian Cao, Kunwei Shen, Xiaosong Chen, and Siji Zhu. A decision support system with intelligent recommendation for multidisciplinary medical treatment. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):33:1–33:23, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3352573.

Zhou:2007:CCO

[ZCT^{+}07] Suiping Zhou, Wentong Cai, Stephen J. Turner, Bu-Sung Lee, and Junhu Wei. Critical causal order of events in distributed virtual environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 3(3): 15:1–15:??, August 2007. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2024:SSC

[ZCX^{+}24] Zi-Chao Zhang, Zhen-Duo Chen, Zhen-Yu Xie, Xin Luo, and Xin-Shun Xu. S3Mix: Same category same semantics mixing for augmenting finegrained images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):9:1–9:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3605892.

Zeng:2023:DTV

[ZCXL23] Xianhua Zeng, Saiyuan Chen, Yicai Xie, and Tianxing Liao. 3V3D: Three-view contextual cross-slice difference threedimensional medical image segmentation adversarial network. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):192:1–192:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3592614.

Zhao:2013:DPO

[ZCY^{+}13] Yi-Liang Zhao, Qiang Chen, Shuicheng Yan, Tat-Seng Chua, and Daqing Zhang. Detecting profilable and overlapping communities with user-generated multimedia contents in LBSNs. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (1):3:1–3:??, December 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhao:2019:DSM

[ZCY^{+}19] Liang Zhao, Zhikui Chen, Laurence T. Yang, M. Jamal Deen, and Z. Jane Wang. Deep semantic mapping for heterogeneous multimedia transfer learning using co-occurrence data. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):9:1–9:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3241055.

Zhang:2024:RWS

[ZCYL24] Yongheng Zhang, Yuanqiang Cai, Danfeng Yan, and Rongheng Lin. Real-world scene image enhancement with contrastive domain adaptation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (12):371:1–371:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3694973.

Zhang:2024:CAP

[ZDD^{+}24] Yichi Zhang, Gongchun Ding, Dandan Ding, Zhan Ma, and Zhu Li. On content-aware postprocessing: Adapting statistically learned models to dynamic content. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):28:1–28:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3612925.

Zhang:2016:SPS

[ZDE16] Longyu Zhang, Haiwei Dong, and Abdulmotaleb El Saddik. From 3D sensing to printing: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(2):27:1–27:??, March 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zheng:2023:PLN

[ZDQ^{+}23] Qi Zheng, Jianfeng Dong, Xiaoye Qu, Xun Yang, Yabing Wang, Pan Zhou, Baolong Liu, and Xun Wang. Progressive localization networks for language-based moment localization. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):55:1–55:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3543857.

Zhang:2022:OCC

[ZDZ^{+}22] Tianjun Zhang, Hao Deng, Lin Zhang, Shengjie Zhao, Xiao Liu, and Yicong Zhou. Online correction of camera poses for the surround-view system: a sparse direct approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):106:1– 106:24, November 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3505252.

Zhang:2023:LSS

[ZDZ^{+}23] Pengyi Zhang, Huanzhang Dou, Wenhu Zhang, Yuhan Zhao, Zequn Qin, Dongping Hu, Yi Fang, and Xi Li. A large-scale synthetic gait dataset towards in-the-wild simulation and comparison study. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):17:1–17:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3517199.

Zhuo:2024:UVE

[ZFC^{+}24] Linhai Zhuo, Yuqian Fu, Jingjing Chen, Yixin Cao, and Yu-Gang Jiang. Unified view empirical study for large pretrained model on cross-domain few-shot learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):294:1–294:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3673231.

Zhu:2024:IST

[ZFFX24] Shipeng Zhu, Jun Fang, Pengfei Fang, and Hui Xue. Improving scene text retrieval via stylized middle modality. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):379:1–379:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3696209.

Zhang:2021:BCR

[ZFSX21] Jianhai Zhang, Zhiyong Feng, Yong Su, and Meng Xing. Bayesian covariance representation with global informative prior for 3D action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):135:1–135:22, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460235.

Zou:2025:FED

[ZFY^{+}25] Kaifeng Zou, Sylvain Faisan, Boyang Yu, Sebastien Valette, and Hyewon Seo. 4D facial expression diffusion model. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(1):7:1–7:??, January 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3653455.

Zheng:2008:CVP

[ZG08] Qing-Fang Zheng and Wen Gao. Constructing visual phrases for effective and efficient object-based image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 5(1):7:1–7:??, October 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhao:2019:PER

[ZGD^{+}19] Sicheng Zhao, Amir Gholaminejad, Guiguang Ding, Yue Gao, Jungong Han, and Kurt Keutzer. Personalized emotion recognition by personalityaware high-order learning of physiological signals. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(1s):14:1– 14:??, February 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3233184.

Zhang:2021:ISI

[ZGD21] Yu-Dong Zhang, Juan Manuel Gorriz, and Zhengchao Dong. Introduction to the special issue on explainable deep learning for medical image computing. ACM Transactions on Multimedia Computing, Communications, and Applications, 17 (3s):99:1–99:2, October 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3485046.

Zhu:2018:MIV

[ZGL^{+}18] Yi Zhu, Sharath Chandra Guntuku, Weisi Lin, Gheorghita Ghinea, and Judith A. Redi. Measuring individual video QoE: a survey, and proposal for future directions using social media. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(2s):30:1–30:??, May 2018.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2023:IQA

[ZGL^{+}23] Xiaoyu Zhang, Wei Gao, Ge Li, Qiuping Jiang, and Runmin Cong. Image quality assessment-driven reinforcement learning for mixed distorted image restoration. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):42:1– 42:??, February 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3532625.

Zhang:2020:CSO

[ZGM^{+}20] Zhaoxin Zhang, Changyong Guo, Fanzhi Meng, Taizhong Xu, and Junkai Huang. Cov-Lets: a second-order descriptor for modeling multiple features. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):21:1–21:14, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3357525.

Zhang:2021:ROR

[ZGR21] Jing Zhang, Jiaqi Guo, and Yonggong Ren. Robust ordinal regression: User credit grading with triplet loss-based sampling. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):7:1–7:20, March 2021.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3408303.

Zhuang:2023:OPF

[ZGWZ23] Weiming Zhuang, Xin Gan, Yonggang Wen, and Shuai Zhang. Optimizing performance of federated person reidentification: Benchmarking and analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):38:1–38:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531013.

Zhang:2024:VLS

[ZGY^{+}24a] Jing Zhang, Dan Guo, Xun Yang, Peipei Song, and Meng Wang. Visual-linguistic-stylistic triple reward for crosslingual image captioning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):110:1– 110:??, April 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3634917.

Zhou:2024:GPI

[ZGY^{+}24b] Sheng Zhou, Dan Guo, Xun Yang, Jianfeng Dong, and Meng Wang. Graph pooling inference network for textbased VQA. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):112:1–112:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3634918.

Zhang:2022:HMM

[ZGZ^{+}22] La Zhang, Haiyun Guo, Kuan Zhu, Honglin Qiao, Gaopan Huang, Sen Zhang, Huichen Zhang, Jian Sun, and Jinqiao Wang. Hybrid modality metric learning for visible-infrared person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):25:1– 25:15, February 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3473341.

Zhang:2018:JHA

[ZH18] Junfeng Zhang and Haifeng Hu.

Joint head attribute classifier and domain-specific refinement networks for face alignment. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (4):79:1–79:??, November 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2019:EQA

[Zha19] Wei Zhang. Efficient QoEaware scheme for video quality switching operations in dynamic adaptive streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1):17:1–17:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3269494.

Zhou:2023:AAM

[ZHC^{+}23] Wei Zhou, Yanke Hou, Dihu Chen, Haifeng Hu, and Tao Su. Attention-augmented memory network for image multi-label classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):116:1–116:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3570166.

Zhang:2023:BLL

[ZHD^{+}23] Dengyong Zhang, Pu Huang, Xiangling Ding, Feng Li, Wenjie Zhu, Yun Song, and Gaobo Yang. L^{2}BEC^{2}: Local lightweight bidirectional encoding and channel attention cascade for video frame interpolation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):66:1–66:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3547660.

Zeng:2021:VRS

[ZHG^{+}21] Kun Zeng, Jiangchuan Hu, Yongyi Gong, Kanoksak Wattanachote, Runpeng Yu, and Xiaonan Luo. Vertical retargeting for stereoscopic images via stereo seam carving. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):125:1– 125:22, January 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3408295.

Zhou:2025:AAL

[ZHG^{+}25] Yang-Hao Zhou, Heyan Huang, Cunhan Guo, Rong-Cheng Tu, Zeyu Xiao, Bo Wang, and Xian-Ling Mao. ALOHA: Adapting local spatio-temporal context to enhance the audiovisual semantic segmentation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):176:1–176:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2019:MFA

[ZHL19] Junxuan Zhang, Haifeng Hu, and Xinlong Lu. Moving foreground-aware visual attention and key volume mining for human action recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):74:1–74:??, September 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3321511.

Zhou:2025:TBS

[ZHL^{+}25] Mingliang Zhou, Shuqi Han, Jun Luo, Xu Zhuang, Qin Mao, and Zhengguo Li. Transformerbased and structure-aware dual-stream network for lowlight image enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):293:1– 293:24, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2022:ISIb

[ZHLJ22] Yin Zhang, Iztok Humar, Jia Liu, and Alireza Jolfaei. Introduction to the special issue on affective services based on representation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):135:1– 135:??, October 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3567836.

Zhu:2011:NDK

[ZHLY11] Jianke Zhu, Steven C. H. Hoi, Michael R. Lyu, and Shuicheng Yan. Near-duplicate keyframe retrieval by semi-supervised learning and nonrigid image matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 7(1):4:1–4:??, January 2011. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2016:MDD

[Zho16] Liang Zhou. Mobile device-to-device video distribution: Theory and application. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(3):38:1– 38:??, June 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2020:JSH

[ZHS20] Junfeng Zhang, Haifeng Hu, and Guobin Shen. Joint stacked hourglass network and salient region attention refinement for robust face alignment. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1):10:1–10:18, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3374760.

Zhu:2025:IIR

[ZHS^{+}25] Hongyi Zhu, Jia-Hong Huang, Yixian Shen, Stevan Rudinac, and Evangelos Kanoulas. Interactive image retrieval meets query rewriting with large language and vision language models. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):286:1–286:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2025:SSR

[ZHY^{+}25] Jian Zhang, Kaihao He, Ting Yu, Jun Yu, and Zhenming Yuan. Semi-supervised RGB-D hand gesture recognition via mutual learning of selfsupervised models. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):104:1– 104:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2013:MAS

[ZI13] Qianni Zhang and Ebroul Izquierdo. Multifeature analysis and semantic context learning for image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(2): 12:1–12:??, May 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhu:2021:MAS

[ZJL^{+}21] Lin Zhu, Xiurong Jiang, Jianing Li, Yuanhong Hao, and Yonghong Tian. Motion-aware structured matrix factorization for foreground detection in complex scenes. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):123:1–123:23, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3407188.

Zhao:2020:ISI

[ZJSJ20] Sicheng Zhao, Dhiraj Joshi, Mohammad Soleymani, and Qiang Ji. Introduction to the special issue on affective computing for large-scale heterogeneous multimedia data.

ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–2, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3365845.

Zhao:2023:PMT

[ZJT^{+}23] Sirui Zhao, Hongyu Jiang, Hanqing Tao, Rui Zha, Kun Zhang, Tong Xu, and Enhong Chen. PEDM: a multi-task learning model for persona-aware emojiembedded dialogue generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (3s):132:1–132:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3571819.

Zhang:2020:MSS

[ZJZC20] Anran Zhang, Xiaolong Jiang, Baochang Zhang, and Xianbin Cao. Multi-scale supervised attentive encoder–decoder network for crowd counting. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):28:1– 28:20, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3356019.

Zhu:2023:DDB

[ZLC^{+}23] Xiaoke Zhu, Changlong Li, Xiaopan Chen, Xinyu Zhang, and Xiao-Yuan Jing. Distance and direction based deep discriminant metric learning for kinship verification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1s):37:1–37:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531014.

Zhang:2025:VFI

[ZLC^{+}25] Dengyong Zhang, Runqi Lou, Jiaxin Chen, Xiangling Ding, Xin Liao, and Gaobo Yang. Video frame interpolation via fast bidirectional 3D correlation volume. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):148:1–148:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2023:LLI

[ZLF^{+}23] Mingliang Zhou, Hongyue Leng, Bin Fang, Tao Xiang, Xuekai Wei, and Weijia Jia. Low-light image enhancement via a frequency-based model with structure and texture decomposition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):187:1–187:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3590965.

Zheng:2024:CAG

[ZLF^{+}24] Jianwei Zheng, Yu Liu, Yuchao Feng, Honghui Xu, and Meiyu Zhang. Contrastive attentionguided multi-level feature registration for reference-based super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):55:1–55:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616495.

Zhong:2016:FED

[ZLH16] Sheng-Hua Zhong, Yan Liu, and Kien A. Hua. Field effect deep networks for image recognition with incomplete data. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(4):52:1–52:??, August 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2025:MSF

[ZLJ^{+}25] Shunxiang Zhang, Jiajia Liu, Yixuan Jiao, Yulei Zhang, Lei Chen, and Kuanching Li. A multimodal semantic fusion network with cross-modal alignment for multimodal sentiment analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):280:1–280:22, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2019:SET

[ZLK^{+}19] Yun Zhang, Na Li, Sam Kwong, Gangyi Jiang, and Huanqiang Zeng. Statistical early termination and early skip models for fast mode decision in HEVC INTRA coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):70:1– 70:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3321510.

Zheng:2020:ULH

[ZLL20] Yunpeng Zheng, Xuelong Li, and Xiaoqiang Lu. Unsupervised learning of human action categories in still images with deep representations. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–20, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3362161.

Zhong:2022:DSA

[ZLL^{+}22] Sheng-Hua Zhong, Jingxu Lin, Jianglin Lu, Ahmed Fares, and Tongwei Ren. Deep semantic and attentive network for unsupervised video summarization. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2):55:1–55:21, May 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3477538.

Zhang:2024:ISS

[ZLL^{+}24a] Huijie Zhang, Pu Li, Xiaobai Liu, Xianfeng Yang, and Li An. An iterative semi-supervised approach with pixel-wise contrastive loss for road extraction in aerial images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):80:1–80:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3606374.

Zhao:2024:UOF

[ZLL^{+}24b] Jing Zhao, Bin Li, Jiahao Li, Ruiqin Xiong, and Yan Lu. A universal optimization framework for learning-based image codec. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):16:1–16:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580499.

Zhou:2024:MFG

[ZLL^{+}24c] Li Zhou, Zhenyu Liu, Yutong Li, Yuchi Duan, Huimin Yu, and Bin Hu. Multi fine-grained fusion network for depression detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):257:1–257:??, August 2024. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3665247.

Zhou:2024:PEP

[ZLL^{+}24d] Zhenyu Zhou, Qing Liao, Lei Luo, Xinwang Liu, and En Zhu. ProtoRefine: Enhancing prototypes with similar structure in few-shot learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(12):383:1–383:??, December 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3694686.

Zhang:2025:DEH

[ZLLH25] Donglin Zhang, Chang-Xing Li, Mengke Li, and Zhikai Hu. Discrete elective hashing with incomplete labels for efficient cross-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (6):160:1–160:??, June 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2013:SMV

[ZLLT13] Wengang Zhou, Houqiang Li, Yijuan Lu, and Qi Tian. SIFT match verification by geometric coding for large-scale partialduplicate web image search. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1): 4:1–4:??, February 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2016:SPR

[ZLN^{+}16] Luming Zhang, Xuelong Li, Liqiang Nie, Yan Yan, and Roger Zimmermann. Semantic photo retargeting under noisy image labels. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(3):37:1–37:??, June 2016. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2018:EOP

[ZLOL18] Chao Zhou, Zhenhua Li, Joe Osgood, and Yao Liu. On the effectiveness of offset projections for 360-degree video streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 14(3s):62:1–62:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhao:2014:SSS

[ZLP^{+}14] Xin Zhao, Xue Li, Chaoyi Pang, Quan Z. Sheng, Sen Wang, and Mao Ye. Structured streaming skeleton — a new feature for online human gesture recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):22:1–22:??, September 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhao:2025:GGS

[ZLS^{+}25] Chengxin Zhao, Hefei Ling, Jialie Shen, Han Fang, Sijing Xie, Yaokun Fang, Zongyi Li, and Ping Li. GSyncCode: Geometry synchronous hidden code for one-step photography decoding. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):65:1–65:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3706060.

Zhang:2025:TBA

[ZLT^{+}25a] Zhenqiang Zhang, Kun Li, Shengeng Tang, Yanyan Wei, Fei Wang, Jinxing Zhou, and Dan Guo. Temporal boundary awareness network for repetitive action counting. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):118:1– 118:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2025:SMS

[ZLT^{+}25b] Yuxuan Zhou, Mingyang Li, Jingze Tong, Linlin Li, and Zhiwei Yang. SD-Meta: The software-defined network of human-centric metaverse for multi-lead or multi-media data in spread spectrum communications. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(6):167:1–167:??, June 2025.

CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2017:CAC

[ZLW17] Cong Zhang, Jiangchuan Liu, and Haiyang Wang. Cloudassisted crowdsourced livecast. ACM Transactions on Multimedia Computing, Communications, and Applications, 13 (3s):46:1–46:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2021:GMM

[ZLW^{+}21] Yi Zhang, Miaomiao Li, Siwei Wang, Sisi Dai, Lei Luo, En Zhu, Huiying Xu, Xinzhong Zhu, Chaoyun Yao, and Haoran Zhou. Gaussian mixture model clustering with incomplete data. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):6:1–6:14, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3408318.

Zhang:2024:JAV

[ZLX24] Yibo Zhang, Weiguo Lin, and Junfeng Xu. Joint audiovisual attention with contrastive learning for more general deepfake detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):137:1– 137:??, May 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3625100.

Zhou:2025:SPH

[ZLX^{+}25] Yuanding Zhou, Xinran Li, Cheng Xiong, Heng Yao, and Chuan Qin. A survey of perceptual hashing for multimedia. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):195:1–195:28, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2024:LDI

[ZLY^{+}24] Jiehua Zhang, Liang Li, Chenggang Yan, Zhan Wang, Changliang Xu, Jiyong Zhang, and Chuqiao Chen. Learning domain invariant features for unsupervised indoor depth estimation adaptation. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):290:1– 290:??, September 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3672397.

Zhang:2022:ISIa

[ZLZ^{+}22a] Shiliang Zhang, Guorong Li, Weigang Zhang, Qingming Huang, Tiejun Huang, Mubarak Shah, and Nicu Sebe. Introduction to the special issue on finegrained visual recognition and re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):24:1–24:3, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505280.

Zhao:2022:BAP

[ZLZ22b] Jian Zhao, Xianhui Liu, and Weidong Zhao. Balanced and accurate pseudo-labels for semisupervised image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):145:1–145:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3506711.

Zhang:2023:ERI

[ZLZX23] Ling Zhang, Chengjiang Long, Xiaolong Zhang, and Chunxia Xiao. Exploiting residual and illumination with GANs for shadow detection and shadow removal. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3):120:1–120:??, May 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3571745.

Zhang:2020:PLB

[ZMH^{+}20] Rui-Xiao Zhang, Ming Ma, Tianchi Huang, Haitian Pang, Xin Yao, Chenglei Wu, and Lifeng Sun. A practical learning-based approach for viewer scheduling in the crowdsourced live streaming. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):67:1– 67:22, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3397226.

Zhang:2024:TEC

[ZML^{+}24] Tianyu Zhang, Weiqing Min, Tao Liu, Shuqiang Jiang, and Yong Rui. Toward egocentric compositional action anticipation with adaptive semantic debiasing. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):122:1–122:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3633333.

Zhang:2023:VNB

[ZMX^{+}23] Chenchi Zhang, Wenbo Ma, Jun Xiao, Hanwang Zhang, Jian Shao, Yueting Zhuang, and Long Chen. VL-NMS: Breaking proposal bottlenecks in two-stage visual-language matching. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5s):166:1–166:??, October 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3579095.

Zhu:2023:TVB

[ZMZ^{+}23] Yucheng Zhu, Xiongkuo Min, Dandan Zhu, Guangtao Zhai, Xiaokang Yang, Wenjun Zhang, Ke Gu, and Jiantao Zhou. Toward visual behavior and attention understanding for augmented 360 degree videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):99:1–99:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3565024.

Zhang:2025:DIP

[ZNNH25] Jian Zhang, Jiangqun Ni, Fan Nie, and Jiwu Huang. Domain-invariant and patchdiscriminative feature learning for general deepfake detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):47:1–47:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3657297.

Zhai:2025:FCP

[ZNYZ25] Rui Zhai, Rongrong Ni, Yang Yu, and Yao Zhao. FaceDefend: Copyright protection to prevent face embezzle. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(2):44:1–44:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3699718.

Zhao:2013:AAP

[ZO13] Zhen Wei Zhao and Wei Tsang Ooi. APRICOD: an accesspattern-driven distributed caching middleware for fast content discovery of noncontinuous media access. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(2):15:1–15:??, May 2013. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhu:2023:MLR

[ZPL^{+}23] Jie Zhu, Bo Peng, Wanqing Li, Haifeng Shen, Qingming Huang, and Jianjun Lei. Modeling long-range dependencies and epipolar geometry for multi-view stereo. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):200:1–200:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3596445.

Zhang:2025:ISI

[ZPR^{+}25] Yushu Zhang, William Puech, Anderson Rocha, Rongxing Lu, Stefano Cresci, and Roberto Di Pietro. Introduction to the special issue on security and privacy of avatar in metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):41:1–41:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3702485.

Zhang:2023:BSG

[ZPY^{+}23] Yong Zhang, Yingwei Pan, Ting Yao, Rui Huang, Tao Mei, and Chang-Wen Chen. Boosting scene graph generation with visual relation saliency. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):8:1–8:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514041.

Zhuang:2019:RCI

[ZQKH19] Naifan Zhuang, Guo-Jun Qi, The Duc Kieu, and Kien A. Hua. Rethinking the combined and individual orders of derivative of states for differential recurrent neural networks: Deep differential recurrent neural networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3):83:1– 83:??, September 2019. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3337928.

Zahran:2018:AAS

[ZQRS18] Ahmed H. Zahran, Jason J.

Quinlan, K. K. Ramakrishnan, and Cormac J. Sreenan. ASAP: Adaptive stall-aware pacing for improved DASH video experience in cellular networks.

ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (3s):61:1–61:??, August 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2013:IST

[ZR13] Lei Zhang and Yong Rui. Image search-from thousands to billions in 20 years. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1s):36:1–36:??, October 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2008:AEE

[ZRCH08] Cha Zhang, Yong Rui, Jim Crawford, and Li-Wei He. An automated end-to-end lecture capture and broadcasting system. ACM Transactions on Multimedia Computing, Communications, and Applications, 4(1):6:1–6:??, January 2008. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2021:DNP

[ZRZ^{+}21] Zhao Zhang, Jiahuan Ren, Haijun Zhang, Zheng Zhang, Guangcan Liu, and Shuicheng Yan. DLRF-Net: a progressive deep latent low-rank fusion network for hierarchical subspace discovery. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):5:1–5:24, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3402030.

Zeng:2023:JAC

[ZRZX23] Shaoning Zeng, Yunbo Rao, Bob Zhang, and Yong Xu. Joint augmented and compressed dictionaries for robust image classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):136:1–136:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3572910.

Zhai:2022:RML

[ZSJL22] Deming Zhai, Ruifeng Shi, Junjun Jiang, and Xianming Liu. Rectified meta-learning from noisy labels for robust image-based plant disease classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):30:1–30:17, February 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3472809.

Zhang:2023:ANM

[ZSLW23] Xianlin Zhang, Mengling Shen, Xueming Li, and Xiaojie Wang. AABLSTM: a novel multi-task based CNN-RNN deep model for fashion analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):19:1– 19:??, January 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3519029.

Zhai:2021:PQA

[ZSMZ21] Guangtao Zhai, Wei Sun, Xiongkuo Min, and Jiantao Zhou. Perceptual quality assessment of low-light image enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):130:1–130:24, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3457905.

Zhao:2013:MEU

[ZSO13] Zhen Wei Zhao, Sameer Samarth, and Wei Tsang Ooi. Modeling the effect of user interactions on mesh-based P2P VoD streaming systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(2):13:1– 13:??, May 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2020:KAN

[ZSS20] Dongyang Zhang, Jie Shao, and Heng Tao Shen. Kernel attention network for single image super-resolution. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):90:1–90:15, September 2020. CODEN ????

ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3398685.

Zheng:2023:EEA

[ZSS^{+}23] Na Zheng, Xuemeng Song, Tianyu Su, Weifeng Liu, Yan Yan, and Liqiang Nie. Egocentric early action prediction via adversarial knowledge distillation. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):59:1–59:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3544493.

Zhang:2025:BMD

[ZSX^{+}25] Zhichao Zhang, Wei Sun, Li Xinyue, Jun Jia, Xiongkuo Min, Zicheng Zhang, Chunyi Li, Zijian Chen, Wang Puyi, Sun Fengyu, Jui Shangling, and Guangtao Zhai. Benchmarking multi-dimensional AIGC video quality assessment: a dataset and unified model. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(9):269:1–269:24, September 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zeng:2023:CFG

[ZSYW23] Rongfei Zeng, Mai Su, Ruiyun Yu, and Xingwei Wang. CD^{2}: Fine-grained 3D mesh reconstruction with twice chamfer distance. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):181:1–181:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582694.

Zhao:2022:JGF

[ZSZ^{+}22] Zhongwei Zhao, Ran Song, Qian Zhang, Peng Duan, and Youmei Zhang. JoT-GAN: a framework for jointly training GAN and person reidentification model. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(1s):27:1– 27:18, February 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3491225.

Zhu:2023:NLA

[ZSZ^{+}23] Dandan Zhu, Xuan Shao, Qiangqiang Zhou, Xiongkuo Min, Guangtao Zhai, and Xiaokang Yang. A novel lightweight audio-visual saliency model for videos. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(4):147:1–147:??, July 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3576857.

Zhang:2024:SOQ

[ZSZ^{+}24a] Zicheng Zhang, Wei Sun, Yingjie Zhou, Jun Jia, Zhichao Zhang, Jing Liu, Xiongkuo Min, and Guangtao Zhai. Subjective and objective quality assessment for in-the-wild computer graphics images. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):96:1– 96:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631357.

Zheng:2024:SSJ

[ZSZ^{+}24b] Chengyu Zheng, Ning Song, Ruoyu Zhang, Lei Huang, Zhiqiang Wei, and Jie Nie. Scale-semantic joint decoupling network for image-text retrieval in remote sensing. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):4:1–4:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603628.

Zhao:2025:LGB

[ZSZ^{+}25] Enyuan Zhao, Ning Song, Ze Zhang, Jie Nie, Xinyue Liang, and Zhiqiang Wei. Language-guided bias generation contrastive strategy for visual question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):114:1– 114:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zeng:2022:CAS

[ZT22] Linghua Zeng and Xinmei Tian. CRAR: Accelerating stereo matching with cascaded residual regression and adaptive refinement. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3):74:1–74:19, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3488719.

Zink:2020:IBP

[ZTB20] Michael Zink, Laura Toni, and Ali C. Begen. Introduction to the best papers from the ACM Multimedia Systems (MMSys) 2019 and Co-Located Workshops. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):66:1–66:2, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3398384.

Zhang:2023:PBI

[ZTQX23] Yushu Zhang, Qing Tan, Shuren Qi, and Mingfu Xue. PRNU-based image forgery localization with deep multi-scale fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):67:1–67:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3548689.

Zhang:2022:AIA

[ZTR^{+}22] Yazhou Zhang, Prayag Tiwari, Lu Rong, Rui Chen, Nojoom A. Alnajem, and M. Shamim Hossain. Affective interaction: Attentive representation learning for multi-modal sentiment classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(3s):136:1–136:??, October 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3527175.

Zhao:2023:BTG

[ZTX^{+}23] Mengyi Zhao, Hao Tang, Pan Xie, Shuling Dai, Nicu Sebe, and Wei Wang. Bidirectional transformer GAN for longterm human motion prediction. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(5):163:1–163:??, September 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579359.

Zhuang:2022:MAD

[ZWC^{+}22] Wenlin Zhuang, Congyi Wang, Jinxiang Chai, Yangang Wang, Ming Shao, and Siyu Xia. Music2Dance: DanceNet for music-driven dance generation. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2):65:1–65:21, May 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3485664.

Zhu:2020:PSB

[ZWF^{+}20] Junjie Zhu, Yuxuan Wei, Yifan Feng, Xibin Zhao, and Yue Gao. Physiological signalsbased emotion recognition via high-order correlation learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (3s):1–18, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3332374.

Zeng:2023:LEI

[ZWH^{+}23] Donghuo Zeng, Jianming Wu, Gen Hattori, Rong Xu, and Yi Yu. Learning explicit and implicit dual common subspaces for audio-visual crossmodal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):97:1–97:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3564608.

Zou:2024:CLN

[ZWJ^{+}24] Cong Zou, Rui Wang, Cheng Jin, Sanyi Zhang, and Xin Wang. S^{2}CL-LeafNet: Recognizing leaf images like human botanists. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):30:1–30:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3615659.

Zhang:2015:PMC

[ZWL15] Yu Zhang, James Z. Wang, and Jia Li. Parallel massive clustering of discrete distributions. ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (4):49:1–49:??, April 2015. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2017:SDL

[ZWL^{+}17] Jun Zhang, Meng Wang, Liang Lin, Xun Yang, Jun Gao, and Yong Rui. Saliency detection on light field: a multi-cue approach. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3):32:1–32:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2021:RMR

[ZWL^{+}21] Chao Zhang, Xiaopei Wu, Jianchao Lu, Xi Zheng, Alireza Jolfaei, Quan Z. Sheng, and Dongjin Yu. RICA-MD: a refined ICA algorithm for motion detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(1s):17:1–17:17, March 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3416492.

Zhang:2024:RAR

[ZWL^{+}24] Zhihao Zhang, Jun Wang, Shengjie Li, Lei Jin, Hao Wu, Jian Zhao, and Bo Zhang. Review and analysis of RGBT single object tracking methods: a fusion perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):259:1– 259:??, August 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3651308.

Zhang:2024:DAS

[ZWLZ24] Shixiong Zhang, Wenmin Wang, Honglei Li, and Shenyong Zhang. E-detector: Asynchronous spatio-temporal for event-based object detection in intelligent transportation system. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):36:1–36:??, February 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3584361.

Zhang:2012:CPC

[ZWM12a] Xin Zhang, Tomás Ward, and Séamus Mcloone. Comparison of predictive contract mechanisms from an information theory perspective. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(2):18:1–18:??, May 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2012:IBD

[ZWM12b] Xin Zhang, Tomás E. Ward, and Séamus Mcloone. An information-based dynamic extrapolation model for networked virtual environments. ACM Transactions on Multimedia Computing, Communications, and Applications, 8(3): 27:1–27:??, July 2012. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhong:2020:SDM

[ZWR^{+}20] Sheng-Hua Zhong, Yuantian Wang, Tongwei Ren, Mingjie Zheng, Yan Liu, and Gangshan Wu. Steganographer detection via multi-scale embedding probability estimation. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(4):1–23, January 2020. ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3352691.

Zhao:2020:ACL

[ZWS^{+}20] Sicheng Zhao, Shangfei Wang, Mohammad Soleymani, Dhiraj Joshi, and Qiang Ji. Affective computing for large-scale heterogeneous multimedia data: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 15(3s):1–32, January 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3363560.

Zeng:2024:MPS

[ZWX24] Xianhua Zeng, Xinyu Wang, and Yicai Xie. Multiple pseudo-Siamese network with supervised contrast learning for medical multi-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (5):128:1–128:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3637441.

Zhang:2021:LCF

[ZWY21] Donglin Zhang, Xiao-Jun Wu, and Jun Yu. Label consistent flexible matrix factorization hashing for efficient crossmodal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):90:1–90:18, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446774.

Zhou:2020:RLV

[ZWYS20] Zhili Zhou, Q. M. Jonathan Wu, Yimin Yang, and Xingming Sun. Region-level visual consistency verification for large-scale partial-duplicate image search. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):54:1–54:25, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3383582.

Zhen:2023:TAO

[ZWZ^{+}23a] Peining Zhen, Shuqi Wang, Suming Zhang, Xiaotao Yan, Wei Wang, Zhigang Ji, and Hai-Bao Chen. Towards accurate oriented object detection in aerial images with adaptive multi-level feature fusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(1):6:1–6:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3513133.

Zhu:2023:AAM

[ZWZ^{+}23b] Hongguang Zhu, Yunchao Wei, Yao Zhao, Chunjie Zhang, and Shujuan Huang. AMC: Adaptive multi-expert collaborative network for text-guided image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(6):188:1–188:??, November 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3584703.

Zheng:2021:IIR

[ZWZL21] Hongdi Zheng, Junfeng Wang, Jianping Zhang, and Ruirui Li. IRTS: an intelligent and reliable transmission scheme for screen updates delivery in DaaS. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(3):82:1–82:24, August 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3440035.

Zhang:2022:DVS

[ZWZL22] Zheng Zhang, Jianning Wang, Lei Zhu, and Guangming Lu. Discriminative visual similarity search with semantically cycleconsistent hashing networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2s):114:1–114:??, June 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3532519.

Zhang:2014:CDM

[ZX14] Tianzhu Zhang and Changsheng Xu. Cross-domain multievent tracking via CO-PMHT. ACM Transactions on Multimedia Computing, Communications, and Applications, 10 (4):31:1–31:??, June 2014. ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2023:AIS

[ZXD^{+}23a] Wei Zhou, Zhiwu Xia, Peng Dou, Tao Su, and Haifeng Hu. Aligning image semantics and label concepts for image multi-label classification.

ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (2):75:1–75:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3550278.

Zhou:2023:DAB

[ZXD^{+}23b] Wei Zhou, Zhiwu Xia, Peng Dou, Tao Su, and Haifeng Hu. Double attention based on graph attention network for image multi-label classification. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1):18:1–18:??, January 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3519030.

Zhao:2024:DHO

[ZXS^{+}24] Xiaojia Zhao, Tingting Xu, Qiangqiang Shen, Youfa Liu, Yongyong Chen, and Jingyong Su. Double high-order correlation preserved robust multi-view ensemble clustering. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (1):21:1–21:??, January 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3612923.

Zou:2025:GHF

[ZXS^{+}25] Shihao Zou, Yuanlu Xu, Nikolaos Sarafianos, Federica Bogo, Tony Tung, Weixin Si, and Li Cheng. Generating highfidelity clothed human dynamics with temporal diffusion. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (3):88:1–88:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3712011.

Zhang:2025:WDW

[ZXT^{+}25] Xiaofeng Zhang, Zishan Xu, Hao Tang, Chaochen Gu, Wei Chen, and Abdulmotaleb El Saddik. Wakeup-darkness: When multimodal meets unsupervised low-light image enhancement. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):84:1–84:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3711929.

Zhang:2022:TIS

[ZXX22] Feifei Zhang, Mingliang Xu, and Changsheng Xu. Tell, imagine, and search: End-to-end learning for composing text and image to image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (2):59:1–59:23, May 2022. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3478642.

Zhuang:2020:MAR

[ZXY^{+}20] Yueting Zhuang, Dejing Xu, Xin Yan, Wenzhuo Cheng, Zhou Zhao, Shiliang Pu, and Jun Xiao. Multichannel attention refinement for video question answering. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):24:1– 24:23, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3366710.

Zhang:2024:PPA

[ZXZ^{+}24a] Qinnan Zhang, Zehui Xiong, Jianming Zhu, Sheng Gao, and Wanting Yang. A privacypreserving auction mechanism for learning model as an NFT in blockchain-driven metaverse. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (7):190:1–190:??, July 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3599971.

Zhou:2024:MML

[ZXZ^{+}24b] Yong Zhou, Zeming Xie, Jiaqi Zhao, Wenliang Du, Rui Yao, and Abdulmotaleb El Saddik. Multi-modal LiDAR point cloud semantic segmentation with salience refinement and boundary perception. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):305:1– 305:??, October 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3674979.

Zhang:2025:AVS

[ZXZ^{+}25] Nana Zhang, Min Xiong, Dandan Zhu, Kun Zhu, Guangtao Zhai, and Xiaokang Yang. Audio-visual saliency prediction model with implicit neural representation. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):117:1–117:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zareapoor:2021:EAN

[ZY21] Masoumeh Zareapoor and Jie Yang. Equivariant adversarial network for image-to-image translation. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2s):73:1–73:14, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3458280.

Zhou:2024:BQA

[ZYC^{+}24] Wei Zhou, Qi Yang, Wu Chen, Qiuping Jiang, Guangtao Zhai, and Weisi Lin. Blind quality assessment of dense 3D point clouds with structure guided resampling. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):247:1–247:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3664199.

Zhang:2024:CST

[ZYCD24] Yuantong Zhang, Daiqin Yang, Zhenzhong Chen, and Wenpeng Ding. Continuous space-time video super-resolution with multi-stage motion information reorganization. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):273:1–273:??, September 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3665646.

Zhao:2024:BTS

[ZYH^{+}24] Jing Zhao, Hongwei Yang, Hui He, Jie Peng, Weizhe Zhang, Jiangqun Ni, Arun Kumar Sangaiah, and Aniello Castiglione. Backdoor twostream video models on federated learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(11):338:1–338:??, November 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3651307.

Zhang:2025:MGA

[ZYJZ25] Qun Zhang, Chao Yang, Bin Jiang, and Bolin Zhang. Multigrained alignment with knowledge distillation for partially relevant video retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):285:1– 285:22, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2017:TDC

[ZYL^{+}17] Qingchen Zhang, Laurence T. Yang, Xingang Liu, Zhikui Chen, and Peng Li. A Tucker deep computation model for mobile multimedia feature learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(3s):39:1–39:??, August 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhao:2024:TEL

[ZYL^{+}24] Ruonan Zhao, Laurence T. Yang, Debin Liu, Wanli Lu, Chenlu Zhu, and Yiheng Ruan. Tensor-empowered LSTM for communication-efficient and privacy-enhanced cognitive federated learning in intelligent transportation systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(2):33:1– 33:??, February 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3575661.

Zhao:2021:GRC

[ZYLN21] Zhongying Zhao, Yonghao Yang, Chao Li, and Liqiang Nie. GuessUNeed: Recommending courses via neural attention network and course prerequisite relation embeddings. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):132:1–132:17, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410441.

Zha:2010:VQS

[ZYM^{+}10] Zheng-Jun Zha, Linjun Yang, Tao Mei, Meng Wang, Zengfu Wang, Tat-Seng Chua, and Xian-Sheng Hua. Visual query suggestion: Towards capturing user intent in Internet image search. ACM Transactions on Multimedia Computing, Communications, and Applications, 6(3):13:1–13:??, August 2010. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zeng:2020:DTN

[ZYO20] Donghuo Zeng, Yi Yu, and Keizo Oyama. Deep triplet neural networks with cluster-CCA for audio-visual crossmodal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(3):76:1–76:23, September 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3387164.

Zhang:2024:ECD

[ZYQM24] Yiheng Zhang, Ting Yao, Zhaofan Qiu, and Tao Mei. Explaining cross-domain recognition with interpretable deep classifier. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(3):67:1–67:??, March 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3623399.

Zhang:2023:VAC

[ZYT^{+}23] Jiwei Zhang, Yi Yu, Suhua Tang, Jianming Wu, and Wei Li. Variational autoencoder with CCA for audio-visual cross-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):130:1– 130:??, June 2023. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3575658.

Zhang:2024:SAR

[ZYW^{+}24] Yazhou Zhang, Yang Yu, Mengyao Wang, Min Huang, and M. Shamim Hossain. Selfadaptive representation learning model for multi-modal sentiment and sarcasm joint analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(5):125:1–125:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635311.

Zeng:2025:NTS

[ZYW^{+}25] Jiuzhen Zeng, Laurence T.

Yang, Chao Wang, Junjie Su, and Xianjun Deng. A new tensor summary statistic for real-time detection of stealthy anomaly in avatar interaction. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):52:1–52:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3689429.

Zhu:2020:PCA

[ZYY^{+}20] Suguo Zhu, Xiaoxian Yang, Jun Yu, Zhenying Fang, Meng Wang, and Qingming Huang. Proposal complementary action detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2s):64:1–64:12, July 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3361845.

Zheng:2013:GSD

[ZYZ^{+}13] Yan-Tao Zheng, Shuicheng Yan, Zheng-Jun Zha, Yiqun Li, Xiangdong Zhou, Tat-Seng Chua, and Ramesh Jain. GPSView: a scenic driving route planner. ACM Transactions on Multimedia Computing, Communications, and Applications, 9(1):3:1–3:??, February 2013. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2025:HOA

[ZYZ^{+}25a] Lu Zhang, Rui Yao, Yuhong Zhang, Yong Zhou, Fuyuan Hu, Jiaqi Zhao, and Zhiwen Shao. Historical object-aware prompt learning for universal hyperspectral object tracking. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):209:1–209:20, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhu:2025:ICC

[ZYZ^{+}25b] Hancheng Zhu, Li Yan, Yong Zhou, Rui Yao, Zhiwen Shao, Jiaqi Zhao, and Leida Li. Image cropping with content and composition attribute-aware global relation reasoning. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):125:1– 125:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2019:DLB

[ZYZE19a] Wei Zhang, Ting Yao, Shiai Zhu, and Abdulmotaleb El Saddik. Deep learning-based multimedia analytics: a review. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1s):2:1–2:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3279952.

Zhang:2019:ESI

[ZYZE19b] Wei Zhang, Ting Yao, Shiai Zhu, and Abdulmotaleb El Saddik. Editorial to special issue on deep learning for intelligent multimedia analytics. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1s):1:1–1:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3292059.

Zhang:2021:WTG

[ZZB^{+}21] Bo Zhang, Rui Zhang, Niccolo Bisagno, Nicola Conci, Francesco G. B. De Natale, and Hongbo Liu. Where are they going? Predicting human behaviors in crowded scenes. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):123:1–123:19, November 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3449359.

Zhang:2024:PCA

[ZZB24] Shupei Zhang, Chenqiu Zhao, and Anup Basu. Principal component approximation network for image compression. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (5):121:1–121:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3637490.

Zhu:2015:SSB

[ZZC^{+}15] Biao Zhu, Hongxin Zhang, Wei Chen, Feng Xia, and Ross Maciejewski. ShotVis: Smartphone-based visualization of OCR information from images. ACM Transactions on Multimedia Computing, Communications, and Applications, 12(1s):12:1–12:??, October 2015. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zeng:2023:TIA

[ZZC^{+}23] Zengri Zeng, Baokang Zhao, Han-Chieh Chao, Ilsun You, Kuo-Hui Yeh, and Weizhi Meng. Towards intelligent attack detection using DNA computing. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(3s):126:1–126:??, June 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3561057.

Zhao:2025:AVC

[ZZC^{+}25a] Hanqing Zhao, Wenbo Zhou, Dongdong Chen, Weiming Zhang, Ying Guo, Zhen Cheng, Pengfei Yan, and Nenghai Yu. Audio-visual contrastive pretrain for face forgery detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):45:1–45:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3651311.

Zhu:2025:MLT

[ZZC^{+}25b] Shaojun Zhu, Bincheng Zhu, Kaikai Chi, Jiefan Qiu, Hailong Shi, and Xingyu Gao. Maximizing long-term task completion ratio of UAV-enabled wirelessly powered MEC systems. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):97:1–97:??, March 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3712599.

Zhu:2021:PAP

[ZZCZ21] Anqi Zhu, Lin Zhang, Juntao Chen, and Yicong Zhou. Pedestrian-aware panoramic video stitching based on a structured camera array. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(4):136:1– 136:24, November 2021. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3460511.

Zheng:2020:DPC

[ZZG^{+}20] Zhedong Zheng, Liang Zheng, Michael Garrett, Yi Yang, Mingliang Xu, and Yi-Dong Shen.

Dual-path convolutional imagetext embeddings with instance loss. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(2):51:1–51:23, June 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3383184.

Zhang:2023:LCE

[ZZJ^{+}23] Yue Zhang, Fanghui Zhang, Yi Jin, Yigang Cen, Viacheslav Voronin, and Shaohua Wan. Local correlation ensemble with GCN based on attention features for cross-domain person re-ID. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2):56:1–56:??, March 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3542820.

Zhang:2017:SRB

[ZZL^{+}17] Hong-Bo Zhang, Bineng Zhong, Qing Lei, Ji-Xiang Du, Jialin Peng, Duansheng Chen, and Xiao Ke. Sparse representationbased semi-supervised regression for people counting. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):47:1– 47:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhao:2021:SAM

[ZZL21] Meiqi Zhao, Jianmin Zheng, and Elvis S. Liu. Server allocation for massively multiplayer online cloud games using evolutionary optimization. ACM Transactions on Multimedia Computing, Communications, and Applications, 17(2):51:1– 51:23, June 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3433027.

Zhu:2023:DLB

[ZZL^{+}23] Linwei Zhu, Yun Zhang, Na Li, Gangyi Jiang, and Sam Kwong. Deep learning-based intra mode derivation for versatile video coding. ACM Transactions on Multimedia Computing, Communications, and Applications, 19(2s):96:1–96:??, April 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3563699.

Zhang:2024:CWS

[ZZL^{+}24a] Chengyang Zhang, Yong Zhang, Bo Li, Xinglin Piao, and Baocai Yin. CrowdGraph: Weakly supervised crowd counting via pure graph neural network. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (5):135:1–135:??, May 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3638774.

Zhao:2024:EWZ

[ZZL^{+}24b] Yuli Zhao, Yin Zhang, Francis C. M. Lau, Hai Yu, Zhiliang Zhu, and Bin Zhang. Expanding-window zigzag decodable fountain codes for scalable multimedia transmission. ACM Transactions on Multimedia Computing, Communications, and Applications, 20 (8):248:1–248:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3664610.

Zhou:2024:RRT

[ZZL^{+}24c] Mingliang Zhou, Xinwen Zhao, Futing Luo, Jun Luo, Huayan Pu, and Tao Xiang. Robust RGB-T tracking via adaptive modality weight correlation filters and cross-modality learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(4):95:1–95:??, April 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3630100.

Zuo:2024:PEP

[ZZL^{+}24d] Ruifan Zuo, Chaoqun Zheng, Fengling Li, Lei Zhu, and Zheng Zhang. Privacyenhanced prototype-based federated cross-modal hashing for cross-modal retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(9):295:1– 295:??, September 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3674507.

Zhang:2025:SIL

[ZZL^{+}25a] Dengyong Zhang, Wenjie Zhu, Xin Liao, Feifan Qi, Gaobo Yang, and Xiangling Ding. Spatiotemporal inconsistency learning and interactive fusion for deepfake video detection. ACM Transactions on Multimedia Computing, Communications, and Applications, 21 (2):48:1–48:??, February 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3664654.

Zhang:2025:MPA

[ZZL^{+}25b] Zicheng Zhang, Yingjie Zhou, Chunyi Li, Wei Sun, Xiongkuo Min, Xiaohong Liu, and Guangtao Zhai. MM-PCQA+: Advancing multi-modal learning for point cloud quality assessment. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(4):119:1–119:??, April 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2025:QAE

[ZZL^{+}25c] Zicheng Zhang, Yingjie Zhou, Chunyi Li, Baixuan Zhao, Xiaohong Liu, and Guangtao Zhai. Quality assessment in the era of large models: a survey. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):189:1–189:31, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2025:PGS

[ZZL^{+}25d] Zhili Zhou, Wensheng Zhang, Zhengdao Li, Huilin Ge, Bin Qiu, Fengjun Xiao, and Yongfeng Huang. Progressive generative steganography via high-resolution image generation for covert communication. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(10):296:1–296:23, October 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhuang:2025:MMS

[ZZL25e] Xingjie Zhuang, Fengling Zhou, and Zhixin Li. Multi-modal sarcasm detection via knowledgeaware focused graph convolutional networks. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(5):143:1–143:??, May 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhou:2017:CCB

[ZZLL17] Mingliang Zhou, Yongfei Zhang, Bo Li, and Xupeng Lin. Complexity correlation-based CTUlevel rate control with direction selection for HEVC. ACM Transactions on Multimedia Computing, Communications, and Applications, 13(4):53:1– 53:??, October 2017. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhao:2025:DFF

[ZZLW25] Haifeng Zhao, Chi Zhang, Deyin Liu, and Lin Wu. Deformation field fusion for medical image registration. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(3):94:1– 94:??, March 2025. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3707462.

Zha:2014:ISI

[ZZMS14] Zheng-Jun Zha, Lei Zhang, Max Mühlhäuser, and Alan F. Smeaton. Introduction to the special issue best papers of ACM Multimedia 2013. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):18:1– 18:??, September 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhu:2021:LDA

[ZZMZ21] Yucheng Zhu, Guangtao Zhai, Xiongkuo Min, and Jiantao Zhou. Learning a deep agent to predict head movement in 360degree images. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(4):130:1–130:23, January 2021. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410455.

Zhang:2020:ABM

[ZZP^{+}20] Hongyi Zhang, Haoke Zhang, Sandeep Pirbhulal, Wanqing Wu, and Victor Hugo C. De Albuquerque. Active balancing mechanism for imbalanced medical data in deep learningbased classification models. ACM Transactions on Multimedia Computing, Communications, and Applications, 16(1s):39:1–39:15, April 2020. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3357253.

Zhang:2025:EID

[ZZS^{+}25] Xinjie Zhang, Tenggan Zhang, Lei Sun, Jinming Zhao, and Qin Jin. Exploring interpretability in deep learning for affective computing: a comprehensive review. ACM Transactions on Multimedia Computing, Communications, and Applications, 21(7):194:1–194:28, July 2025. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhang:2024:USD

[ZZT24] Gai Zhang, Xinfeng Zhang, and Lv Tang. Unified and scalable deep image compression framework for human and machine. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(10):314:1–314:??, October 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3678472.

Zhang:2024:BDG

[ZZTL24] Dejun Zhang, Mian Zhang, Xuefeng Tan, and Jun Liu. Bridging the domain gap in scene flow estimation via hierarchical smoothness refinement. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(8):236:1–236:??, August 2024. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3661823.

Zhao:2019:VCR

[ZZW^{+}19] Rui-Wei Zhao, Qi Zhang, Zuxuan Wu, Jianguo Li, and Yu-Gang Jiang. Visual content recognition by exploiting semantic feature map with attention and multi-task learning. ACM Transactions on Multimedia Computing, Communications, and Applications, 15 (1s):6:1–6:??, February 2019. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3231739.

Zhu:2022:SSR

[ZZW^{+}22] Xiaoguang Zhu, Ye Zhu, Haoyu Wang, Honglin Wen, Yan Yan, and Peilin Liu. Skeleton sequence and RGB frame based multi-modality feature fusion network for action recognition.

ACM Transactions on Multimedia Computing, Communications, and Applications, 18 (3):80:1–80:24, August 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3491228.

Zhang:2024:SSV

[ZZW^{+}24] Yuqing Zhang, Yong Zhang, Shaofan Wang, Yun Liang, and Baocai Yin. Semi-supervised video object segmentation via an edge attention gated graph convolutional network. ACM Transactions on Multimedia Computing, Communications, and Applications, 20(1):24:1– 24:??, January 2024. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3611389.

Zhang:2020:ISI

[ZZX^{+}20] Shengping Zhang, Huiyu Zhou, Dong Xu, M. Emre Celebi, and Thierry Bouwmans. Introduction to the special issue on multimodal machine learning for human behavior analysis. ACM Transactions on Multimedia Computing, Communications, and Applications, 16 (1s):19:1–19:2, April 2020. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3381917.

Zhang:2014:AAS

[ZZY^{+}14] Hanwang Zhang, Zheng-Jun Zha, Yang Yang, Shuicheng Yan, Yue Gao, and Tat-Seng Chua. Attribute-augmented semantic hierarchy: Towards a unified framework for contentbased image retrieval. ACM Transactions on Multimedia Computing, Communications, and Applications, 11(1s):21:1– 21:??, September 2014. CO-DEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zheng:2018:DLC

[ZZY18] Zhedong Zheng, Liang Zheng, and Yi Yang. A discriminatively learned CNN embedding for person reidentification. ACM Transactions on Multimedia Computing, Communications, and Applications, 14 (1):13:1–13:??, January 2018. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zhu:2023:CSA

[ZZY^{+}23] Guanyu Zhu, Yong Zhou, Rui Yao, Hancheng Zhu, and Jiaqi Zhao. Cyclic self-attention for point cloud recognition. ACM Transactions on Multimedia Computing, Communications, and Applications, 19 (1s):49:1–49:??, February 2023. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

URL https://dl.acm.org/
doi/10.1145/3538648.

Zhang:2014:AGS

[ZZZ14] Ying Zhang, Luming Zhang, and Roger Zimmermann. Aestheticsguided summarization from multiple user generated videos.

ACM Transactions on Multimedia Computing, Communications, and Applications, 11 (2):24:1–24:??, December 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Zheng:2022:CMS

[ZZZ^{+}22] Yi Zheng, Yong Zhou, Jiaqi Zhao, Ying Chen, Rui Yao, Bing Liu, and Abdulmotaleb El Saddik. Clustering matters: Sphere feature for fully unsupervised person re-identification. ACM Transactions on Multimedia Computing, Communications, and Applications, 18(4):97:1–97:18, November 2022. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3501404.